### v1

Abstract
Experiments deliver credible treatment effects estimates, but are often localized to specific sites, populations, or mechanisms. When such estimates are insufficient to extrapolate effects for broader policy questions as for external validity and general-equilibrium (GE) effects, researchers combine trials with external evidence from reduced-form or structural observational estimates, or prior experiments. We develop a unified framework for designing experiments in this setting: the researcher selects which parameters to identify experimentally from a feasible set (which treatment arms and/or individuals to include in the experiment), allocates sample size, and specifies how to weight experimental and observational estimators. Because observational inputs may be biased in ways unknown ex ante, we adopt a minimax proportional-regret objective that evaluates any candidate design relative to an oracle that knows the bias and jointly chooses the design and estimator. This yields a transparent bias-variance trade-off that requires no prespecified bias bound and depends only on information about the precision of the estimators and the estimand’s sensitivity to underlying parameters. We illustrate the framework by (i) designing small-scale cash-transfer experiments aimed at estimating GE effects and (ii) optimizing site selection for microfinance interventions.
1 Introduction
Randomized controlled trials have transformed empirical economics by delivering credible, internally valid estimates. However, feasibility constraints often confine trials to localized effects such as the effect in a specific site or subpopulations, or of a particular mechanism. These effects, although useful, are often not sufficient to answer broader questions about external validity, generalizability, or equilibrium effects. In response, a growing literature in development economics (e.g., Meghir et al., 2022; Gechter et al., 2024; Bassi et al., 2022; Attanasio et al., 2012), political economy (e.g., de Albuquerque et al., 2025), education (e.g., Allende et al., 2019; Larroucau et al., 2024), and labor economics (e.g., Chetty et al., 2016) complements localized experiments with external evidence from reduced-form or structural observational estimates, and/or trials in other contexts, with the goal of estimating complex counterfactuals that no single experiment can fully identify. For example, in our review of AEA journals (in 2015-2025), over 30% of experimental papers report, in addition to the experiment, at least one observational estimate – either reduced-form (e.g., via matching or instrumental variable regression), or structural (see Figure 12). This raises a design question: given constraints on what experiments can identify, which experiments should be run (and how) when they will be combined with external evidence to estimate those counterfactuals?
For an illustrative example, consider a government piloting a cash-transfer program in a small set of districts to increase children’s school attendance. The trial delivers a local average effect, but policy decisions often require counterfactuals at scale, allowing prices and wages to adjust (e.g. Todd and Wolpin, 2006; Egger et al., 2022). Researchers therefore combine experimental evidence with a supply-demand model that leverages observational data to map local impacts into economy-wide outcomes. The design question is twofold: (i) which experiment to run (i.e., which parameters/effects are the most valuable to learn experimentally) given feasibility constraints and (ii) how to allocate sample across sites and arms so that, once combined with external inputs, we precisely estimate the effect at scale.
To answer these questions, we develop a framework for designing experiments to be used alongside external evidence. By external evidence we mean reduced-form or structural estimates based on observational data, and/or results from experiments conducted in other times or places. Our main contribution is a joint procedure that selects which experiments to run subject to a budget (i.e., which treatment arms and/or sub-populations to include in the study), sets their precision via sample allocation, and prescribes how to combine experimental estimates with external evidence. A central consideration in our design is that external inputs may be biased in ways not known ex ante.
We consider a setting where the object of interest is a known function $\tau(\theta)$ with $\theta$ a vector of unknown parameters; $\tau$ may be univariate or multivariate and typically encodes a policy-relevant target. Researchers have access to external estimates of $\theta$ that may be biased—henceforth, observational estimates. They then run one or more experiments in the population of interest to learn a subset of parameters without bias, choosing a sample allocation across sites and arms subject to a budget. We parameterize the design so that some, but not necessarily all, components of $\theta$ can be learned experimentally. After the experiment, each experimental estimate is combined with its observational counterpart via parameter-specific shrinkage weights (chosen jointly with the design), while parameters not learned experimentally rely on the observational estimates alone.
Returning to the cash-transfer example, $\tau(\theta)$ is the effect on schooling when all poor households in a region receive the subsidy. The parameter vector $\theta$ bundles (i) the direct schooling response to a conditional cash transfer (CCT), (ii) the income effect of transfers in partial equilibrium, and (iii) wage/price adjustments that shift the returns to schooling. A large-scale trial that identifies GE effects is often infeasible; instead, one can run small, partial-equilibrium experiments and then combine these with prior evidence to recover $\tau(\theta)$. For instance, researchers may choose between two (small-scale) treatment arms such as a CCT to estimate the direct effect or an unconditional-cash arm to measure the income elasticity, and extrapolate the GE effects using price effects from other experimental or observational studies. When site selection is also part of the design, $\theta$ additionally collects site-specific effects and $\tau(\theta)$ may target the average effect across sites.
A natural starting point is to minimize the mean-squared error (MSE) for estimating $\tau(\theta)$, perhaps in a worst-case sense over bias in the observational estimates. In practice, however, the size and even the direction of bias are unknown (before the experiment is run). A worst-case approach would therefore be overly-conservative and disregard information about the variance. We instead adapt the definition of proportional (adaptation) regret previously studied when combining estimators generated by a fixed design (Armstrong et al., 2024; Tsybakov, 1998) and generalize it to the experimental design problem. Here, the regret is the ratio of the worst-case researcher’s mean-squared error relative to an oracle that knows the worst-case bound on the observational bias and chooses both the design and the (shrinkage) estimator. Taking the supremum of this ratio over any value of the bias yields a procedure that is robust to any level of misspecification without requiring priors about the bias.
Our main result gives an explicit, easy-to-compute characterization of proportional regret. For any design and shrinkage weights, the regret is the maximum of two normalized components. The first is a variance component: the estimator’s sampling variance under the chosen design and weights, divided by the smallest variance attainable by any feasible design. The second is a bias component: the worst-case squared bias induced by using external evidence, divided by the smallest feasible worst-case bias. The variance component depends on the (expected) variance-covariance matrix of the observational and experimental estimates implied by the design, the shrinkage weights, and the sensitivity of the policy parameter, i.e., the gradient of $\tau$ with respect to the parameters. This gradient is evaluated, under mild conditions, at the observational estimates. [^4] For this result to hold, we assume that $\tau(\theta)$ is a smooth and differentiable function in $\theta$. For non-linear $\tau(\theta)$ in $\theta$, we also require that the observational estimates bias is local to zero in the spirit of Andrews et al. (2020); Bonhomme and Weidner (2022), but consider a more general framework where its rate of convergence can be the same, faster, or even slower than the estimators’ standard error. Section 4.1 provides details. By contrast, the bias component depends only on the sensitivity vector and the shrinkage weights, both known ex-ante.
This characterization makes the bias-variance trade-off transparent. At the optimum, the two normalized components tend to be equalized: when the bias component dominates, weight shifts toward experimental evidence and the design prioritizes learning high-sensitivity coordinates; when variance dominates, the procedure invests sample where it most reduces variance and relies more (via shrinkage) on the most precise observational inputs. This yields a nested procedure: we (i) find (regret-optimal) shrinkage weights for each candidate design to balance the two ratios; (ii) allocate sample, subject to the budget, to reduce the variance; and (iii) select the experiment set, trading off designs that improve precision the most against those targeting parameters to which the policy estimand is the most sensitive. The resulting design and estimator only require knowledge of the expected variance-covariance matrix (standard in experimental design problems, Gerber and Green, 2012) and of the observational estimates. It can then be reported in a pre-analysis plan.
We illustrate the framework in the cash‐transfer application. As an illustrative example, we consider a researcher evaluating a conditional cash transfer (CCT) to increase schooling in rural Kenya. Budget and feasibility constraints permit only small-scale (partial-equilibrium) randomization, while the goal is to predict general-equilibrium effects. As preliminary (external) inputs, the researcher uses evidence from Mexico’s PROGRESA program (Todd and Wolpin, 2006). We are concerned that these preliminary estimates may lack external validity in Kenya. We compare three designs: (i) a CCT arm to estimate the direct schooling effect; (ii) an unconditional cash transfer (UCT) arm to estimate income effects; and (iii) a two-arm design that allocates sample across CCT and UCT under a common budget. Running a single arm captures settings with fixed costs per arm; allowing both arms with budgeted allocation captures cases with non-binding fixed costs but binding variable costs. When both arms are feasible, the regret-optimal allocation puts most participants in the CCT and a small fraction in the UCT, reflecting the target’s greater sensitivity to the direct effect despite the CCT’s higher variance. If only one arm can be run, the choice hinges on sample size: for very small $n$, the lower-variance of the UCT arm makes this preferred over a (too noisy) CCT arm; once $n\geq 500$, the more informative but noisier CCT yields lower regret. Relative to an oracle that knows the bias, the optimal design closely tracks oracle performance, and regret approaches one as sample size grows.
As a second application, we design where to run a microfinance experiment in rural India, integrating evidence from earlier nonrandomized microfinance introductions. We start from the observational estimates in Banerjee et al. (2024) and design an experiment that selects one or more areas for randomization. Because Banerjee et al. (2024) also implement a separate randomized expansion, we can calibrate performance of our design (that is agnostic of the bias) by comparing each candidate design’s MSE to that of an oracle that knows the bias, using the experimental estimates to calibrate the bias. The optimal design concentrates sample in high-sensitivity, high-variance areas and then spreads as the budget grows; compared to a benchmark that randomly chooses areas and splits sample equally, it reduces MSE by roughly 20% under the calibrated bias.
In summary, this paper links the experimental-design literature—which typically focuses on settings where all parameters are identified within the experiment and leaves aside questions of data combination—to recent work that integrates experimental and (reduced-form or structural) observational evidence to extrapolate effects in complex scenarios.
Recent advances for experimental design include balancing and variance-minimizing allocations (Tabord-Meehan, 2018; Bai, 2019; Cytrynbaum, 2021; Kallus, 2018; Bertsimas et al., 2015), adaptive designs for policy choice (Kasy and Sautmann, 2019; Russo et al., 2018; Kato et al., 2024; Cesa-Bianchi et al., 2025), and experimental design under correct model specification (Silvey, 2013; Chaudhuri and Mykland, 1993; Chaloner and Verdinelli, 1995; Higbee, 2024; Kiefer and Wolfowitz, 1959; Reeves et al., 2024; Viviano, 2020; Kasy, 2016). Traditional work on experimental design for robust-model estimation either focused on testing competing models (Atkinson and Fedorov, 1975; López-Fidalgo et al., 2007), or on using a-priori knowledge of (worst-case) bias for the design of an experiment (Box and Draper, 1959; Wiens, 1998; Tsirpitzi et al., 2023; Sacks and Ylvisaker, 1984). We complement this literature by introducing and studying the question of which experiment to design when combined with observational evidence (when not all parameters can be learned experimentally).
Our minimax regret criterion connects to a long-standing decision-theoretic tradition for experimental design. References include Manski and Tetenov (2016), Banerjee et al. (2020), Manski (2004), Dominitz and F. Manski (2017), Olea et al. (2024), Hu et al. (2024), Breza et al. (2025) among others. These focus on settings where researchers have only access to experimental variation. Here instead we study the problem of choosing which experiment to run (and how) when combined with observational evidence. We also complement literature that leverages correctly-specified models for site selection (Gechter et al., 2024; Abadie and Zhao, 2021) by allowing for misspecification in observational estimates.
A related line of work analyzes estimation under misspecification (Armstrong and Kolesár, 2018; Armstrong et al., 2024; Andrews et al., 2017, 2020; Bonhomme and Weidner, 2022; Christensen and Connault, 2023; James et al., 1961), and combining existing experiments with observational studies (Gechter, 2022; Athey et al., 2025, 2020; Kallus et al., 2018; Dutz et al., 2021; Rosenman et al., 2022; Bhattacharya, 2013; de Chaisemartin and D’Haultfœuille, 2020; Rambachan et al., 2024). We adapt the worst-case proportional regret idea from Armstrong et al. (2024) and Tsybakov (1998) here studied in the context of experimental design. Our definition of sensitivity of the estimand to each parameter directly links to sensitivity analysis in Andrews et al. (2020). Unlike both strands of this literature where the design is fixed and researchers optimize over the choice of the estimator only, here we optimize the design itself. Consequently, the objective changes: regret is computed against the best design-estimator combination, and the resulting estimator is learned jointly with the design.
Finally, we connect to a growing empirical literature that combines observational (or model-based) evidence with experimental variation to learn economic quantities – spanning general equilibrium effects (Egger et al., 2022; Attanasio et al., 2012; Meghir et al., 2022; Kreindler et al., 2023), external-validity counterfactuals (Gechter et al., 2024; Banerjee et al., 2024), large-scale information campaigns (de Albuquerque et al., 2025; Larroucau et al., 2024), and market-system experiments (Bergquist and Dinerstein, 2020; Allende et al., 2019). We provide a unifying framework to decide what to learn experimentally, how to split the sample across experiments, and how to integrate experimental and observational evidence.
2 Problem description
Consider a researcher interested in an arbitrary target estimand $\tau(\theta)\in\mathbb{R}$, indexed by a low-dimensional parameter vector $\theta\in\mathbb{R}^{p}$ and a known mapping $\tau$. The estimand may arise from a reduced-form (e.g., treatment effect in a given area) or a structural model. For simplicity, we assume that the estimand is univariate, while Section 4.3 shows how our framework directly extends to multivariate $\tau(\theta)\in\mathbb{R}^{q}$ for $q>1$.
The goal is to construct an estimator $\hat{\tau}$ that accurately approximates $\tau(\theta)$ by combining existing evidence (e.g., observational studies) with experimental variation designed by the researcher. Our question is how to design such experiments under feasibility constraints.
2.1 Setup
We assume that researchers have access to estimators (and their variance) of $\theta$ denoted as $\tilde{\theta}^{\mathrm{obs}}\in\mathbb{R}^{p}$. We impose no restrictions on how $\tilde{\theta}^{\mathrm{obs}}$ is formed: it can based on arbitrary exclusion restrictions implied by an economic or statistical model. However, $\tilde{\theta}^{\mathrm{obs}}$, that for exposition we define observational estimates, have unknown biases collected in a vector $b\in\mathbb{R}^{p}$. Examples of $\tilde{\theta}^{\mathrm{obs}}$ include a structural model estimate with confounding, an estimate from an instrumental variable regression that may fail the exclusion restriction, or an estimate from a country different from the one of interest that may lack external validity.
Researchers may also collect evidence for a subset of parameters that they know is unbiased. For instance, researchers may generate exogenous variation via an experiment (or more generally acquiring more data) that identifies one or some of the parameters of interest. We refer to these parameters as experimental estimates, adopting here the definition of experimental estimates as parameters that the researcher knows are unbiased. [^5] In some applications researchers may be confident that a few observational estimates have no bias. It is possible to extend our framework to these settings with a change of notation where those observational estimates assumed to be unbiased are defined as experimental estimates, and such experimental estimates are always contained in the set $\mathcal{E}$ defined below.
Setting 1 (Main setup).
For a subset $\mathcal{E}\subseteq\{1,\ldots,p\}$ and a known strictly positive-definite matrix $\Sigma(\mathcal{E})\in\mathcal{G}(\mathcal{E})\subset\mathbb{R}^{(p+|\mathcal{E}|)\times(p+|\mathcal{E}|)}$ with uniformly bounded entries, define $\tilde{\theta}^{\mathrm{obs}}\in\mathbb{R}^{p}$ an observational estimate and $\tilde{\theta}^{\mathrm{exp}}_{\mathcal{E},\Sigma}\in\mathbb{R}^{|\mathcal{E}|}$ an experimental estimate each satisfying
$\mathbb{E}\!\left[\tilde{\theta}^{\mathrm{obs}}\right]-\theta=b,\qquad\mathbb{E}\!\left[\tilde{\theta}^{\mathrm{exp}}_{\mathcal{E},\Sigma}\right]-\theta_{\mathcal{E}}=0,$
for an unknown bias vector $b\in\mathbb{R}^{p}$. Moreover, define its joint variance as
$\mathbb{V}\!\begin{pmatrix}\tilde{\theta}^{\mathrm{obs}}-\theta\\[2.0pt] \tilde{\theta}^{\mathrm{exp}}_{\mathcal{E},\Sigma}-\theta_{\mathcal{E}}\end{pmatrix}=\Sigma(\mathcal{E}).$ (1)
Given $(\mathcal{E},\Sigma)$, consider a class of linear plug-in estimators
$\hat{\tau}_{\gamma}\equiv\tau(\hat{\theta}(\gamma)),\qquad\hat{\theta}_{j}(\gamma)=\begin{cases}\gamma_{j}\,\tilde{\theta}_{j}^{\mathrm{exp}}+(1-\gamma_{j})\,\tilde{\theta}_{j}^{\mathrm{obs}},&j\in\mathcal{E},\\[2.0pt] \tilde{\theta}_{j}^{\mathrm{obs}},&j\notin\mathcal{E},\end{cases}$
where $\gamma=(\gamma_{j})_{j\in\mathcal{E}}\in\mathbb{R}^{|\mathcal{E}|}$ are shrinkage weights that can be an (implicit) function of $(\mathcal{E},\Sigma)$.
Here, $(\mathcal{E},\Sigma(\mathcal{E}))$ characterizes the experimental design (which parameters are learned and with what precision). When clear, we omit the subscript $\Sigma$ on $\tilde{\theta}^{\mathrm{exp}}$.
We assume $\Sigma(\mathcal{E})$ is known, which is standard in experimental planning (Gerber and Green, 2012) (in practice, any consistent estimator, e.g., from pilot studies would suffice). In addition, we assume that $\Sigma(\mathcal{E})$ is uniformly bounded, which implies that once we commit to learn a set of experimental parameters $\tilde{\theta}_{\mathcal{E}}^{\mathrm{exp}}$, their variance is finite. In addition $\Sigma$ is strictly positive definite, therefore assuming that the variance is bounded away from zero. [^6] For this latter condition to hold in an asymptotic framework with growing sample size, $\theta$ and $\tilde{\theta}$ can be defined as parameters of interest after appropriately rescaling by the square-root of the sample size; in this case $\Sigma$ denotes the asymptotic variance. See Section 4.1 for details.
We let $\mathcal{E}\in\mathcal{S},\Sigma(\mathcal{E})\in\mathcal{G}(\mathcal{E})$ for some constraint sets $(\mathcal{S},\mathcal{G}(\mathcal{E}))$ encoding feasibility or budget constraints. In our framework, restrictions on the experiments researchers can run $\mathcal{E}\in\mathcal{S}$ (and on their precision $\Sigma(\mathcal{E})\in\mathcal{G}(\mathcal{E})$) can take any desired form. In most applications, we think of such constraints as arising from fixed costs of running an experiment and/or standard lower bound constraints on power for experimental estimates, which are common in empirical practice (Duflo et al., 2007; Athey and Imbens, 2017; List et al., 2011)
We write compactly $(\mathcal{E},\Sigma,\gamma)\in\mathcal{D}$ indicating $\mathcal{D}$ the set of feasible experiments (e.g., sample size, etc.), with $\gamma\in\mathbb{R}^{|\mathcal{E}|}$ for a choice of experiments $\mathcal{E}$.
Our goal is to optimize both over the design and the weights $\gamma$. Before introducing our main research question, we impose the following assumption.
Assumption 1 (First-order estimation error).
For any admissible $(\mathcal{E},\Sigma,\gamma)\in\mathcal{D}$, assume
$\tau(\theta)-\tau\!\big(\hat{\theta}(\gamma)\big)=\sum_{j=1}^{p}\omega_{j}\,\big(\theta_{j}-\hat{\theta}_{j}(\gamma)\big),$ (2)
for known weights $\omega\in\mathbb{R}^{p}$, with $|\omega_{j}|>0$ for all $j$.
Assumption 1 holds exactly for linear $\tau(\cdot)$ and serves as a first-order approximation for smooth nonlinear $\tau(\cdot)$.
Remark 1 (Non-linear $\tau(\theta)$).
Section 4.1 (and Example 3 below) extends our framework to non-linear estimands and shows how Assumption 1 serves as an approximation via a Taylor expansion within a local asymptotic framework, where the bias is local to zero, but at a possibly faster, equal or even slower rate than the estimator’s standard error. For non-linear $\tau$, $\omega$ is replaced by the gradient of $\tau$ evaluated at the baseline observational estimates (i.e., $\omega=\frac{\partial\tau(\theta)}{\partial\theta}\Big|_{\theta=\tilde{\theta}^{\mathrm{obs}}}+o_{p}(1)$), which is typically observed before the experiment is conducted.
Remark 2 (When researchers can only identify linear combinations of the parameters).
In some applications, researchers may be able to identify via an experiment a linear combination of the parameters of interest (up-to a first order approximation), defined $\phi\equiv W^{\top}\theta$, where $W\in\mathbb{R}^{p\times s}$ is a known matrix. Suppose that $\widetilde{\phi}$ is an unbiased estimator of $\phi$, and the researcher can access a subset of its entries via an experiment.
Our framework extends to this setting as follows. Define the coefficient of the linear projection of $\omega$ onto the column space of $W$: $\pi=\left(W^{\top}W\right)^{-1}W^{\top}\omega$ (known ex-ante). Define the residual $v=\omega-W\pi$. Then,
$\tau=\left[W\pi+v\right]^{\top}\theta=\pi^{\top}\phi+v^{\top}\theta\implies\tau=\check{\omega}^{\top}\check{\theta},$
where $\check{\omega}\equiv(\beta^{\top},v^{\top})^{\top}$ and $\check{\theta}=(\phi^{\top},\theta^{\top})^{\top}$. This corresponds to our original setting where the researcher’s feasible set only allows access to unbiased estimates of a subset of the parameters, corresponding to the entries of $\phi$. ∎
2.2 Research questions
Our design problem can be described in three steps:
  1. 1.
    Preliminary step: weights. For each $(\mathcal{E},\Sigma(\mathcal{E}))\in\mathcal{D}$, choose $\gamma$ to combine experimental and observational estimates on the selected coordinates.
  2. 2.
    Middle step: precision. For each $\mathcal{E}$, choose $\Sigma(\mathcal{E})\in\mathcal{G}(\mathcal{E})$, where $\mathcal{G}(\mathcal{E})$ denotes an arbitrary set of positive definite covariances under feasibility constraints.
  3. 3.
    Outer step: experiment choice. Choose $\mathcal{E}\in\mathcal{S}$, the feasible set of parameters to learn experimentally for given arbitrary constraint set $\mathcal{S}$.
Choosing $\gamma$ (step 1) for a fixed design follows a long-standing tradition in econometrics and statistics (Armstrong et al., 2024; Andrews and Shapiro, 2021; Donoho, 1994; Athey et al., 2025). We study this question here in combination with the experimental design problem, step 2 and 3, which has not been studied by the references above; this, as we show, will change the underlying optimization problem (also for $\gamma$). The central challenge is that performance depends on the unknown bias $b$ in the observational estimates.
2.3 Mapping to a simple two parameters model
While we derive our results in general form, we will build intuition using a stylized description with a two-parameter model.
Setting 2 (Illustration with two-parameter model).
Consider two parameters $\theta=(\theta_{1},\theta_{2})^{\top}$, and mutually independent observational and experimental estimates
$\tilde{\theta}^{\mathrm{obs}}_{j}-\theta_{j}\sim\mathcal{N}(b_{j},\ \sigma_{j}^{2}),\qquad\tilde{\theta}^{\exp}_{j}-\theta_{j}\sim\mathcal{N}(0,\ v_{j}^{2}),\qquad j=1,2.$
We may run one experiment: $\mathcal{S}=\{\{1\},\{2\}\}$. If we pick $j$, we estimate
$\hat{\theta}_{j}=\gamma_{j}\,\tilde{\theta}^{\exp}_{j}+(1-\gamma_{j})\,\tilde{\theta}^{\mathrm{obs}}_{j},\qquad\hat{\theta}_{-j}=\tilde{\theta}^{\mathrm{obs}}_{-j},$
and, to first order, $\tau(\theta)-\tau(\hat{\theta})=\omega_{1}(\theta_{1}-\hat{\theta}_{1})+\omega_{2}(\theta_{2}-\hat{\theta}_{2}).$
Example 1 (Choosing the site for an experiment for external validity).
Gechter et al. (2024) study where to run an experiment. For illustration, consider two sites $j\in\{1,2\}$ with site-specific ATEs $\theta_{j}$ and target the cross-site average
$\tau(\theta)=\tfrac{1}{2}(\theta_{1}+\theta_{2})=\omega^{\top}\theta,\qquad\omega=\tfrac{1}{2}(1,1)^{\top}.$
Let $\tilde{\theta}^{\mathrm{obs}}=(\tilde{\theta}^{\mathrm{obs}}_{1},\tilde{\theta}^{\mathrm{obs}}_{2})^{\top}$ denote observational estimates obtained in Gechter et al. (2024) from a structural model and potentially biased due to misspecification: $\mathbb{E}[\tilde{\theta}^{\mathrm{obs}}]-\theta=(b_{1},b_{2})^{\top}.$ A budget constraint allows an experiment in only one site, so $\mathcal{S}=\{\{1\},\{2\}\}$. If site $j$ is chosen, we obtain an unbiased $\tilde{\theta}^{\mathrm{exp}}_{j}$; the other site $k\neq j$ remains observational. Given $\mathcal{E}\in\mathcal{S}$,
$\widehat{\tau}_{\gamma}=\tfrac{1}{2}\!\left[\gamma_{j}\,\tilde{\theta}^{\mathrm{exp}}_{j}+(1-\gamma_{j})\,\tilde{\theta}^{\mathrm{obs}}_{j}+\tilde{\theta}^{\mathrm{obs}}_{k}\right],\qquad k\neq j.$
Our question is how to choose the set $\{j\}$ where to conduct the experiment (jointly with $\gamma$).
Example 2 (Choosing which survey to conduct).
Egger et al. (2022) study the efficacy of cash-transfer programs on the marginal propensity to consume (MPC). Measuring the MPC requires capturing both short- and long-run effects. Because survey rounds are limited, the authors complement experimental data that lacks short-run effects with auxiliary information from prior studies that use short-run surveys (collected in other regions; see Egger et al. (2022)). This raises the question of which survey to conduct (and with which frequency).
In stylized form, suppose researchers can observe, for $s\in\{1,2\}$, potential outcomes $Y_{s}(t)$ denoting consumption in period $s$ when measured $t$ periods after the intervention. The authors have auxiliary estimates from previous studies,
$\alpha\;=\;\mathbb{E}\!\big[Y_{1}(t=1)-Y_{1}(t=\infty)\big],\qquad\beta\;=\;\mathbb{E}\!\big[Y_{2}(t=1)-Y_{2}(t=\infty)\big],$
and wish to estimate the total effect $\tau=\alpha+\beta$. Researchers consider two survey designs:
  • (i)
    Early survey design to estimate $\alpha$ precisely;
  • (ii)
    Later survey to estimate $\beta$ precisely.
Our goal is to study which survey design to implement. (More complex designs with additional time periods or mixed precision across rounds are also possible here.)
Example 3 (Supply or demand experiment with non linear target).
(Bergquist and Dinerstein, 2020) conduct demand and supply experiments in food markets in Kenya. Suppose here we are interested in similar applications in Uganda. For exposition, consider a basic linear demand and supply
$Q^{D}=a-\beta_{D}P+u_{D},\qquad Q^{S}=c+\beta_{S}P+u_{S},\qquad\theta=(\beta_{D},\beta_{S})^{\top},$
with $\beta_{D}\neq-\beta_{S}$. Several estimands may be of interest; one of such estimands is the effect of a tariff $t$ on prices
$\tau\equiv\frac{\beta_{S}}{\beta_{D}+\beta_{S}}\,t.$
Let $(\tilde{\beta}_{D}^{\mathrm{obs}},\tilde{\beta}_{S}^{\mathrm{obs}})^{\top}$ denote baseline estimates from Kenya, that may lack external validity.
Due to cost constraints, $\mathcal{S}=\{\{D\},\{S\}\},$ i.e., we can learn either demand (estimating $\tilde{\beta}_{D}^{\mathrm{exp}}$, with randomized price discounts) or supply (estimating $\tilde{\beta}_{S}^{\mathrm{exp}}$, by introducing regulatory cost shocks on firms). Let
$\big(\tilde{\theta}^{\mathrm{obs}}-\theta\big)\ \sim\ \mathcal{N}\big(b,\Sigma\big),\quad\tilde{\theta}^{\mathrm{obs}}\equiv\sqrt{n}(\tilde{\beta}_{D}^{\mathrm{obs}},\tilde{\beta}_{S}^{\mathrm{obs}})^{\top},\quad\theta\equiv\sqrt{n}(\beta_{D},\beta_{S})^{\top}$
with $\theta$ and $\tilde{\theta}$ denoting the parameter and estimator here rescaled by the square-root of the sample size, and $b$ capturing bias. Our goal is to choose whether to conduct a supply or demand experiment in Uganda to estimate $\tau$. For given estimators $\hat{\theta}$ that combine the estimates from Kenya with the chosen experimental estimate from Uganda, under local asymptotics and a first-order Taylor expansion described in Section 4.1, we can write [^7] Here, we are using a local asymptotic framework where the bias of the estimates in Kenya is non-negligible but small relative to $n$, i.e., $b/\sqrt{n}=o(1)$, so that $\omega$ can be consistently estimated using observational data.
$\tau-\widehat{\tau}=\underbrace{\frac{1}{\sqrt{n}}\check{\omega}(\mathrm{plim}(\tilde{\beta}^{\mathrm{obs}}))^{\top}(\theta-\widehat{\theta})}_{\text{main first order effect}}+\underbrace{o_{p}\left(\frac{b}{n^{1/2}}\right)}_{\text{small higher order effects}},\qquad\check{\omega}(\beta)=\frac{t}{(\beta_{D}+\beta_{S})^{2}}\begin{pmatrix}-\beta_{S}\\ \beta_{D}\end{pmatrix}$,
where we can replace $\check{\omega}$ with its consistent counterpart $\check{\omega}(\tilde{\beta}^{\mathrm{obs}})$. Our goal is to choose between a supply or demand experiment accounting for first-order bias (and variance) effect.
3 Robust experimental design
In this section we introduce an experimental design focusing on the mean-squared error (MSE) of the estimator $\hat{\tau}$, a common measure of precision. In Section 4.2 we extend the framework to minimize the length of the confidence intervals.
3.1 Adaptation regret for experimental design
For a given design $(\mathcal{E},\Sigma(\mathcal{E}))$ and shrinkage weights $\gamma$, define the MSE at a fixed observational-bias vector $b$ as
$\mathrm{MSE}_{b}(\mathcal{E},\Sigma,\gamma)\;=\;\mathbb{E}_{\mathcal{E},\Sigma,b}\!\left[(\widehat{\tau}_{\gamma}-\tau)^{2}\right],$
where $\mathbb{E}_{\mathcal{E},\Sigma,b}$ denotes expectation under the data-generating process implied by $(\mathcal{E},\Sigma(\mathcal{E}))$ and observational bias $b$. The MSE is a measure of precision of $\hat{\tau}$ and ideally, one would minimize $\mathrm{MSE}_{b}$ over $(\mathcal{E},\Sigma(\mathcal{E}),\gamma)\in\mathcal{D}$.
However, because $b$ is unknown, we first consider an uncertainty set given by an $\ell_{\infty}$-ball,
$\mathcal{B}(B)\equiv\{b:\ \|b\|_{\infty}\leq B\},$
with $B\geq 0$ an upper bound on the largest coordinate-wise bias. The $\ell_{\infty}$ choice has a desirable property relative to $\ell_{1}$ or $\ell_{2}$: it does not force a trade-off across coordinates (a large bias in one component need not be “offset” by a small bias elsewhere), which is attractive when biases may be positively correlated across parameters. The drawback is that worst-case solutions depend on the radius $B$ which may be unknown in practice.
If an oracle knew $B$, a natural choice would be to pick the minimax design
$\mathrm{MSE}^{*}(B)\;\equiv\;\inf_{(\mathcal{E},\Sigma,\gamma)\in\mathcal{D}}\ \sup_{b\in\mathcal{B}(B)}\ \mathrm{MSE}_{b}(\mathcal{E},\Sigma,\gamma).$
However, having to specify $B$ can pose a large burden on the researchers and make the choice of the experiment sensitive to $B$. Nevertheless, we could seek designs that perform as close as possible to the oracle that knows $B$, uniformly over the values of $B$.
This idea of performing as close as possible to the oracle is based on an extensive literature on regret minimization (Manski, 2004; Manski and Tetenov, 2007; Montiel Olea et al., 2023; Kitagawa and Tetenov, 2018; Manski and Tetenov, 2016). We minimize the worst-case proportional increase in MSE relative to the oracle across all bias radii:
$\mathcal{R}(\mathcal{E},\Sigma,\gamma)\;\equiv\;\sup_{B\geq 0}\ \frac{\displaystyle\sup_{b\in\mathcal{B}(B)}\mathrm{MSE}_{b}(\mathcal{E},\Sigma,\gamma)}{\displaystyle\mathrm{MSE}^{*}(B)},$
defined adaptation regret by Armstrong et al. (2024) (building in turn on Tsybakov, 1998).
Whereas the above references consider the choice of an estimator for fixed design, here, the regret is relative to an oracle that chooses both the estimator and the design. Our key contribution – different from the references above – is the study of the optimal experimental design informed by biased (observational) estimates. This leads to a different definition of regret optimization, both because we optimize over the class of designs and $\Sigma$, but also in $\gamma$ due to the different definition of oracle. The optimal solution consists of a pair of design and estimator that can be pre-specified.
3.2 Optimal experimental design: main result
Before stating the result, we introduce compact notation. For a given $(\mathcal{E},\gamma)$, with $\gamma\in\mathbb{R}^{|\mathcal{E}|}$ define
$\tilde{\omega}_{\text{obs},j}(\mathcal{E},\gamma)=\begin{cases}\omega_{j},&j\notin\mathcal{E},\\ \omega_{j}(1-\gamma_{j}),&j\in\mathcal{E},\end{cases}\qquad\tilde{\omega}_{\text{exp}}(\mathcal{E},\gamma)=\omega_{\mathcal{E}}\circ\gamma$
where $\circ$ denotes the component-wise product and $\omega_{\mathcal{E}}$ is the subvector of $\omega$ corresponding to the entries indexed by $\mathcal{E}$. Let
$\tilde{\omega}(\mathcal{E},\gamma)\equiv\begin{pmatrix}\tilde{\omega}_{\text{obs}}(\mathcal{E},\gamma)\\[2.0pt] \tilde{\omega}_{\text{exp}}(\mathcal{E},\gamma)\end{pmatrix}\in\mathbb{R}^{\,p+|\mathcal{E}|}.$ (3)
The estimator’s variance and worst–case squared bias divided by $B^{2}$ are
$\alpha(\mathcal{E},\Sigma,\gamma)\;\equiv\;\tilde{\omega}(\mathcal{E},\gamma)^{\top}\Sigma(\mathcal{E})\,\tilde{\omega}(\mathcal{E},\gamma),\qquad\beta(\mathcal{E},\gamma)\;\equiv\;\Big(\,\|\omega\|_{1}-\|\omega_{\,\mathcal{E}}\|_{1}+\,|1-\gamma|^{\top}|\omega_{\mathcal{E}}|\,\Big)^{\!2},$ (4)
with $|1-\gamma|$ indicating the absolute value of each entry of the vector $1-\gamma$.
Define the best achievable variance and worst-case bias divided by $B^{2}$ as
$\alpha^{\star}\;\equiv\;\min_{\mathcal{E}\in\mathcal{S}}\ \min_{\Sigma(\mathcal{E})\in\mathcal{G}(\mathcal{E})}\ \min_{\gamma\in\mathbb{R}^{|\mathcal{E}|}}\alpha(\mathcal{E},\Sigma,\gamma),\qquad\beta^{\star}\;\equiv\;\min_{\mathcal{E}\in\mathcal{S}}\ \min_{\gamma\in\mathbb{R}^{|\mathcal{E}|}}\beta(\mathcal{E},\gamma)=\min_{\mathcal{E}\in\mathcal{S}}\Big(||\omega||_{1}-\|\omega_{\mathcal{E}}\|_{1}\Big)^{2}.$ (5)
Here $\alpha^{\star}$ is the variance of the most precise feasible design, while $\beta^{\star}$ is the squared bias of the least biased design. [^8] The expression for $\beta^{\star}$ corresponds to the following experiment choice: choose $\mathcal{E}$ with the largest $\omega_{\mathcal{E}}||_{1}$ and set $\gamma_{\mathcal{E}}\equiv 1$ at those coordinates. Both quantities depend on the class of designs $\mathcal{S},\mathcal{G}(\mathcal{E})$ and are easy to compute: $\alpha^{\star}$ is a constrained quadratic minimization; $\beta^{\star}$ requires enumerating $\mathcal{E}\in\mathcal{S}$.
Ideally, we would like a design with variance equal to $\alpha^{\star}$ and bias equal to $\beta^{\star}$. Unfortunately, this may be infeasible as it might require different choices of experiments and estimators to achieve one or the other.
Instead, in our main theorem below we show that the regret-optimal pair of design and estimator trades-off the bias and variance and depends on how far each of these two components are from their smallest attainable value.
Theorem 1.
Consider Setting 1 and let Assumption 1 hold. Then, for any $(\mathcal{E},\Sigma,\gamma)$,
$\mathcal{R}(\mathcal{E},\Sigma,\gamma)\;=\;\max\!\left\{\frac{\alpha(\mathcal{E},\Sigma,\gamma)}{\alpha^{\star}},\ \ \frac{\beta(\mathcal{E},\gamma)}{\beta^{\star}}\right\}.$
Proof.
See Appendix B.1. ∎
Theorem 1 provides an explicit expression of the regret (which does not require researchers to specify $B$). To our knowledge, it is the first result to study adaptation regret for experimental design. Its derivation requires us to find precise binding patterns as a function of $B$, which accounts for its effect on optimizing the oracle solution over the design and estimator.
The adaptation regret balances the variance of a given design relative to the smallest possible variance $\alpha^{\star}$ against the worst-case bias $\beta$ relative to the smallest possible worst-case bias $\beta^{\star}$.
To gain insight into the optimal solution, provided that the class of designs $\mathcal{D}$ is sufficiently flexible (outside boundary solutions), we may expect the minimizer $(\mathcal{E}^{\star},\Sigma^{\star},\gamma^{\star})\in\arg\min\mathcal{R}(\mathcal{E},\Sigma,\gamma)$ to equalize
$\frac{\alpha(\mathcal{E}^{\star},\Sigma^{\star},\gamma^{\star})}{\alpha^{\star}}\;=\;\frac{\beta(\mathcal{E}^{\star},\gamma^{\star})}{\beta^{\star}}.$
That is, at the optimum the design equalizes the estimator’s variance $\mathbb{V}(\hat{\tau}_{\gamma})$ (over the design choice) divided by the smallest feasible variance $\alpha^{\star}$ to the worst-case squared bias (equal to $\beta B^{2}$) divided by the smallest achievable worst-case squared bias $\beta^{\star}B^{2}$.
[…]
Figure 1: Each feasible design $(\mathcal{E},\Sigma,\gamma)$ maps to a point $\big(\alpha/\alpha^{\star},\ \beta/\beta^{\star}\big)$: the xx–axis is the variance ratio and the yy–axis is the worst–case bias ratio. The blue curve depicts the attainable frontier as we vary shrinkage $\gamma$ and precision $\Sigma$. Level sets of the objective $\mathcal{R}=\max\{\alpha/\alpha^{\star},\ \beta/\beta^{\star}\}$ are axis–aligned squares (the dotted inverted “L” shows the smallest such square touching the frontier). The minimizer outside boundary solutions is where the frontier meets the $45^{\circ}$ line (red dot), i.e., where the two normalized components are equalized: $\alpha(\mathcal{E}^{\star},\Sigma^{\star},\gamma^{\star})/\alpha^{\star}=\beta(\mathcal{E}^{\star},\gamma^{\star})/\beta^{\star}$.
Theorem 1 provides us an immediate and simple to compute solution to the optimal design problem via backward induction. Specifically, we compute the estimator, the optimal variance and the experimental choice as follows:
$\gamma^{\star}(\mathcal{E},\Sigma)$ $\in\mathrm{arg}\min_{\gamma\in\mathbb{R}^{|\mathcal{E}|}}\,\max\!\left\{\frac{\alpha(\mathcal{E},\Sigma,\gamma)}{\alpha^{\star}},\ \ \frac{\beta(\mathcal{E},\gamma)}{\beta^{\star}}\right\}$
$\Sigma^{\star}(\mathcal{E})$ $\in\mathrm{arg}\min_{\Sigma\in\mathcal{G}(\mathcal{E})}\,\max\!\left\{\frac{\alpha(\mathcal{E},\Sigma,\gamma^{\star}(\mathcal{E},\Sigma))}{\alpha^{\star}},\ \ \frac{\beta(\mathcal{E},\gamma^{\star}(\mathcal{E},\Sigma))}{\beta^{\star}}\right\}$
$\mathcal{E}^{\star}$ $\in\mathrm{arg}\min_{\mathcal{E}\in\mathcal{S}}\,\max\!\left\{\frac{\alpha(\mathcal{E},\Sigma^{\star}(\mathcal{E}),\gamma^{\star}(\mathcal{E},\Sigma^{\star}(\mathcal{E})))}{\alpha^{\star}},\ \ \frac{\beta(\mathcal{E},\gamma^{\star}(\mathcal{E},\Sigma^{\star}(\mathcal{E})))}{\beta^{\star}}\right\}.$
Optimization proceeds as follows:
  • we first choose $\gamma^{\star}$ for each design $\mathcal{E},\Sigma(\mathcal{E})$. Importantly, the choice of $\gamma^{\star}$ depends on the class of designs available to researchers ($\mathcal{D}$), because $\mathcal{D}$ determines the oracle values $\alpha^{\star},\beta^{\star}$ in the denominators of the objective function (a larger class of designs improves the oracle solution);
  • the choice of the experimental variance $\Sigma$ (how precise each experiment is under budget constraints) also depends on the bias component $\beta$ because $\beta$ affects how we select the shrinkage weights $\gamma^{\star}$ in the previous step;
  • the optimal design $\mathcal{E}^{\star}$ then accounts for both choices of $\gamma^{\star}$ and $\Sigma^{\star}$.
3.3 Example with two parameters
Consider the two-parameters model of Setting 2 where for simplicity we fix the choice of the variances $\sigma^{2},v^{2}>0$. The smallest variance and worst–case squared bias (per unit $B^{2}$) are
$\alpha^{\star}=\min_{k\in\{1,2\}}\underbrace{\left\{\ \omega_{-k}^{2}\sigma_{-k}^{2}\;+\;\omega_{k}^{2}\frac{\sigma_{k}^{2}\,v_{k}^{2}}{\sigma_{k}^{2}+v_{k}^{2}}\ \right\}}_{\text{Variance with variance-optimal }\gamma},\qquad\beta^{\star}=\min_{k\in\{1,2\}}\underbrace{\omega_{k}^{2}}_{\text{Bias}^{2}/B^{2}\text{ with bias-optimal }\gamma}.$
Here, $\alpha^{\star}$ follows from the variance-only optimal shrinkage $\gamma_{k}=\sigma_{k}^{2}/(\sigma_{k}^{2}+v_{k}^{2})$ (i.e., the oracle choice when $B=0$); $\beta^{\star}$ follows from the fact that the oracle that minimizes the worst-case bias chooses the experiment $k$ with the largest $\omega_{k}$, sets $\gamma_{k}=1$ and incurs a bias proportional to $|\omega_{-k}|$, since no experiment is conducted for k-k.
The variance and worst-case squared bias (divided by $B^{2}$) incurred by an analyst choosing experiment $j$ and $\gamma_{j}$ are, respectively
$\alpha(j,\gamma_{j})$ $\equiv\underbrace{\omega_{-j}^{2}\sigma_{-j}^{2}}_{\text{var obs estimate}}+\underbrace{\omega_{j}^{2}\!\big[(1-\gamma_{j})^{2}\sigma_{j}^{2}+\gamma_{j}^{2}v_{j}^{2}\big]}_{\text{var exp and obs estimate}},\quad\beta(j,\gamma_{j})$ $\equiv\big(\underbrace{|\omega_{-j}|}_{\begin{subarray}{c}\text{bias obs}\\ \text{estimate}/B\end{subarray}}+\underbrace{(1-\gamma_{j})|\omega_{j}|}_{\begin{subarray}{c}\text{bias obs minus}\\ \text{ bias exp}/B\end{subarray}}\big)^{2}.$
By Theorem 1, the worst-case adaptation regret for choosing experiment $\{j\}$ and $\gamma_{j}$, equals
$\mathcal{R}(\{j\},\gamma_{j})=\max\Big\{\underbrace{\alpha(j,\gamma_{j})/\alpha^{\star}}_{\text{Variance/}\alpha^{\star}},\ \underbrace{\beta(j,\gamma_{j})/\beta^{\star}}_{\text{Bias}^{2}/(B^{2}\beta^{\star})}\Big\}.$
Both maximand components are evaluated with the same choice of $\gamma$.
3.3.1 Choosing $\gamma_{j}$ for given design $\{j\}$
The first step is to choose the shrinkage weight $\gamma$, noting that regret is computed relative to an oracle that optimizes both $\gamma$ and the experiment design $\{j\}$. Minimizing $\mathcal{R}$ over $\gamma_{j}$ yields a rule that either equalizes the two components or selects a boundary solution.
Corollary 1.
Consider Setting 2 and let Assumption 1 hold. Then
[…]
Proof.
See Appendix B.2. ∎
At the boundaries, if the relative variance $\alpha/\alpha^{\star}$ is always smaller than the relative bias $\beta/\beta^{\star}$, the optimal weight is $\gamma_{j}^{\star}=1$ (i.e., use only the experimental estimate). If instead the variance term always dominates the bias term, the optimal choice is the variance–minimizing weight $\gamma_{j}^{\star}=\sigma_{j}^{2}/(\sigma_{j}^{2}+v_{j}^{2})$. In most applications, the solution is interior: we choose $\gamma_{j}^{\star}$ strictly between the boundaries to equalize the two contributions, $\frac{\alpha(j,\gamma_{j}^{\star})}{\alpha^{\star}}=\frac{\beta(j,\gamma_{j}^{\star})}{\beta^{\star}},$ so that the variance ratio matches the worst–case squared bias ratio.
The optimal solution depends not only on the features of the experiment design $j$ but also on the features of the other feasible design j-j ($v_{-j}^{2},\sigma_{-j}^{2}$ and $\omega_{-j}$) through $\alpha^{\star}$ and $\beta^{\star}$. This differs from choosing the optimal $\gamma$ for a single fixed design, since the choice of $\gamma$ will then interact with the choice of the class of design the researcher (and oracle) can choose.
Illustration in Figure 2
Figure 2 illustrates three cases for $j=2$. In the top row, we plot $\alpha(2,\gamma_{2}^{\star})/\alpha^{\star}$ and $\beta(2,\gamma_{2}^{\star})/\beta^{\star}$ evaluated at the optimal weight while varying, column by column, $\omega_{2}$, $v_{2}$, and $\sigma_{2}$. In the bottom row, we plot the corresponding $\gamma_{2}^{\star}$.
Observation 1: small $\omega_{2}$ pushes $\gamma_{2}^{\star}$ toward one. The first (left) column sets $v_{1}=v_{2}/2$ (the first experiment is more precise) and $\omega_{1}=1$ (with $\sigma_{1}=\sigma_{2}=1$). For small $\omega_{2}$ (low sensitivity of the second coordinate to bias), the optimal choice is $\gamma_{2}^{\star}=1$. The reason is twofold. First, the bias ratio $\beta(2,\gamma)/\beta^{\star}=\big(|\omega_{1}|+|1-\gamma||\omega_{2}|\big)^{2}/\beta^{\star}$ is normalized by $\beta^{\star}=\min\{|\omega_{1}|,|\omega_{2}|\}^{2}=\omega_{2}^{2}$ when $\omega_{2}<\omega_{1}$; thus even a small observational component $|1-\gamma|>0$ makes the ratio increase substantially. Second, the variance ratio $\alpha(2,\gamma)/\alpha^{\star}$ is multiplied by $\omega_{2}^{2}$ in its $j=2$ contribution, so its dependence on $\gamma$ becomes negligible as $\omega_{2}\downarrow 0$. Together these forces make the bias ratio dominant and push $\gamma_{2}^{\star}$ to the boundary at $1$. The implication is that for parameters with low $\omega_{2}$, we typically shrink $\hat{\theta}_{2}$ more toward the experimental estimate $\tilde{\theta}_{2}^{\mathrm{exp}}$.
Observation 2: when $\omega_{2}$ is of the same order as $\omega_{1}$, an interior solution emerges. As $\omega_{2}$ increases toward $\omega_{1}$ (roughly $0.6\,\omega_{1}$ and above), the normalization ceases to penalize $j=2$ as harshly, and an interior solution appears: $\gamma_{2}^{\star}$ moves down from $1$ toward the variance–only weight, trading off bias and variance. The first kink in the bottom-left panel reflects this regime change, where $\gamma_{2}^{\star}$ is selected to equalize bias and variance.
Observation 3: for $\omega_{2}>1$, $\gamma_{2}^{\star}$ increases again. At $\omega_{2}=\omega_{1}$ we observe a second kink, since the worst–case bias normalizer switches to $\beta^{\star}=\omega_{1}^{2}$. This raises the bias cost of putting weight on the observational estimate; as $\omega_{2}$ increases above $1$, $\gamma_{2}^{\star}$ rises again. This pattern is driven by the oracle benchmark, which optimizes both the estimator and the design.
Observation 4: for $\omega_{2}\gtrsim 1.25$, $\gamma_{2}^{\star}$ increases slowly and stays interior. Around $\omega_{2}\approx 1.25$ we observe a third kink: the variance normalizer $\alpha^{\star}$ switches branches (from favoring $k=1$ to favoring $k=2$ in the variance–only comparison). The optimal $\gamma_{2}^{\star}$ remains interior but grows more slowly, because increases in $\gamma_{2}^{\star}$ now have a larger relative impact on $\alpha/\alpha^{\star}$. [^9] A further kink could occur if the solution transitioned into a pure variance–dominant regime (where the boundary weight is chosen), which does not arise for the range of $\omega_{2}$ shown here.
Observation 5: $\gamma_{2}^{\star}$ increases with experimental precision and decreases with observational precision The second and third columns vary $v_{2}$ and $\sigma_{2}$, respectively. In both, we fix $v_{1}=1$ and take $\omega_{1}=0.9\,\omega_{2}$ (so here $\omega_{1}=0.9$, $\omega_{2}=1$), with $\sigma_{1}=\sigma_{2}=1$ unless varied. Across $v_{2},\sigma_{2}\in[0.5,2]$, the solution is interior. As $v_{2}\downarrow 0.5$ or $\sigma_{2}\uparrow 2$, the optimal weight $\gamma_{2}^{\star}$ increases toward $1$, placing nearly all weight on the experimental estimate. Conversely, as $v_{2}$ becomes large or $\sigma_{2}$ becomes small, $\gamma_{2}^{\star}$ approaches its lower bound, the variance–minimizing weight $\sigma_{2}^{2}/(\sigma_{2}^{2}+v_{2}^{2})$. This is a typical (and desired) behavior of shrinkage estimators.
Table 1: Regimes for $\gamma_{2}^{\star}$ as a function of $\omega_{2}$ for example in Figure 2 with $\sigma_{1}^{2}=\sigma_{2}^{2}=1,v_{1}=1,v_{2}=0.5,\omega_{1}=1$. Each row summarizes each observation 1-4 for $\gamma_{2}^{\star}$.
Regime $\beta^{\star}$ $\alpha^{\star}$ Solution regime ($\gamma_{2}^{\star}$) Trend of $\gamma_{2}^{\star}$
$\boldsymbol{\omega_{2}\ll\omega_{1}}$ $|\omega_{2}|^{2}$ $\omega_{2}^{2}\sigma_{2}^{2}\;+\;\omega_{1}^{2}\dfrac{\sigma_{1}^{2}v_{1}^{2}}{\sigma_{1}^{2}+v_{1}^{2}}$ Bias–dominant Constant $\gamma_{2}^{\star}=1$
$\boldsymbol{\omega_{2}<\omega_{1}}$ $|\omega_{2}|^{2}$ $\omega_{2}^{2}\sigma_{2}^{2}\;+\;\omega_{1}^{2}\dfrac{\sigma_{1}^{2}v_{1}^{2}}{\sigma_{1}^{2}+v_{1}^{2}}$ Interior Decrease in $\omega_{2}$
$\boldsymbol{\omega_{2}>\omega_{1}}$ $|\omega_{1}|^{2}$ $\omega_{2}^{2}\sigma_{2}^{2}\;+\;\omega_{1}^{2}\dfrac{\sigma_{1}^{2}v_{1}^{2}}{\sigma_{1}^{2}+v_{1}^{2}}$ Interior Increase in $\omega_{2}$
$\boldsymbol{\omega_{2}\gg\omega_{1}}$ $|\omega_{1}|^{2}$ $\omega_{1}^{2}\sigma_{1}^{2}\;+\;\omega_{2}^{2}\dfrac{\sigma_{2}^{2}v_{2}^{2}}{\sigma_{2}^{2}+v_{2}^{2}}$ Interior Slow increase in $\omega_{2}$
Figure 2: Illustration for experiment $j=2$. Top row: relative bias $\beta/\beta^{\star}$ (dashed) and relative variance $\alpha/\alpha^{\star}$ (solid), both evaluated at the optimal weight $\gamma_{2}^{\star}$. Bottom row: the optimal weight $\gamma_{2}^{\star}$. Columns vary, respectively, (a) $\omega_{2}$ with $v_{1}=0.5$, $v_{2}=1$, $\sigma_{1}=\sigma_{2}=1$, $\omega_{1}=1$; (b) $v_{2}$ with $\omega=(0.9,1)$, $v_{1}=1$, $\sigma=(1,1)$; (c) $\sigma_{2}$ with $\omega=(0.9,1)$, $v=(1,1)$, $\sigma_{1}=1$.
3.3.2 Choosing the design $\{j\}$
After computing $\gamma_{j}^{\star}$, we then optimize over the design index $j$, by choosing
$j^{\star}\in\arg\min_{j\in\{1,2\}}\max\Big\{\alpha(j,\gamma_{j}^{\star})/\alpha^{\star},\ \beta(j,\gamma_{j}^{\star})/\beta^{\star}\Big\}.$ (6)
To build intuition, suppose $\gamma_{j}^{\star}$ is interior for both $j\in\{1,2\}$. Then $\alpha(j,\gamma_{j}^{\star})/\alpha^{\star}=\beta(j,\gamma_{j}^{\star})/\beta^{\star}$, which implies
$j^{\star}\in\arg\min_{j\in\{1,2\}}\underbrace{\left\{\omega_{-j}^{2}\sigma_{-j}^{2}+\omega_{j}^{2}\!\big[(1-\gamma_{j}^{\star})^{2}\sigma_{j}^{2}+(\gamma_{j}^{\star})^{2}v_{j}^{2}\big]\right\}}_{\text{overall variance at }\gamma_{j}^{\star}}=\arg\min_{j\in\{1,2\}}\underbrace{\left\{|\omega_{-j}|+|1-\gamma_{j}^{\star}||\omega_{j}|\right\}}_{\text{worst--case bias}/|B|\text{ at }\gamma_{j}^{\star}}.$
In this case, optimal $\gamma_{j}^{\star}$ balances variance and worst–case bias. Because $\alpha^{\star}$ and $\beta^{\star}$ do not depend on $j$, $j^{\star}$ minimizes either (and therefore both) criteria $\alpha(j,\gamma_{j}^{\star})$ and $\beta(j,\gamma_{j}^{\star})$.
A second case arises when one experiment (say $j=1$) is bias–dominant (i.e., $\gamma_{1}^{\star}=1$) or variance–dominant (i.e., $\gamma_{1}^{\star}=\sigma_{1}^{2}/(\sigma_{1}^{2}+v_{1}^{2})$), while the other admits an interior solution $\gamma_{2}^{\star}$ for which $\beta(2,\gamma_{2}^{\star})/\beta^{\star}=\alpha(2,\gamma_{2}^{\star})/\alpha^{\star}$. In this case, the optimal design selects the experiment that yields the smaller value of the dominating criterion (worst–case bias in the bias–dominant scenario, variance in the variance–dominant scenario).
Although we view these as the leading scenarios, cases at the extreme boundaries can occur. The first is when the relative variance $\alpha/\alpha^{\star}$ or the worst–case bias $\beta/\beta^{\star}$ uniformly dominates the other for both experiments. Then the optimal design minimizes the dominating term; alternatively, if for one experiment $j$ the variance ratio $\alpha/\alpha^{\star}$ is dominant while for the other experiment j-j the worst–case bias ratio $\beta/\beta^{\star}$ is dominant, the minimax regret is the smaller of the two dominating values. These are desiderable properties as the optimal design always prioritizes the dominating term. Table 2 summarizes the cases.
Table 2: Optimal shrinkage and experiment choice across different regimes. The first three regimes (row) corresponds to the case where the variance ratio $\alpha/\alpha^{\star}$ does not uniformly dominate $\beta/\beta^{\star}$ for all values of $\gamma_{j}$ and some $j\in\{1,2\}$ (which we view as leading cases). The other cases correspond to boundary solutions.
Case (condition) Optimal $\gamma_{j}^{\star}$ Optimal $j^{\star}$ solves Solution $(j^{\star},\gamma_{j}^{\star})$
Leading cases
Case 1 no bias/ variance dominance $\gamma_{j}^{\star}$ in the interior $\mathrm{arg}\min_{j}\left\{\omega_{-j}^{2}\sigma_{-j}^{2}+\omega_{j}^{2}\!\big[(1-\gamma_{j}^{\star})^{2}\sigma_{j}^{2}+(\gamma_{j}^{\star})^{2}v_{j}^{2}\big]\right\}$ $j^{\star}$ minimizes bias and variance
($\alpha(j,\gamma_{j})/\alpha^{\star}-\beta(j,\gamma_{j})/\beta^{\star}$ can flip sign) for $j\in\{1,2\}$ $\mathrm{arg}\min_{j}\left\{|\omega_{-j}|+|1-\gamma_{j}^{\star}||\omega_{j}|\right\}$ $\gamma_{j}^{\star}$ balances bias/variance
Case 2 $j=1$ is variance dominant $\gamma_{1}^{\star}=\frac{\sigma_{1}^{2}}{v_{1}^{2}+\sigma_{1}^{2}}$ $\mathrm{arg}\min_{j}\left\{\omega_{-j}^{2}\sigma_{-j}^{2}+\omega_{j}^{2}\!\big[(1-\gamma_{j}^{\star})^{2}\sigma_{j}^{2}+(\gamma_{j}^{\star})^{2}v_{j}^{2}\big]\right\}$ $j^{\star},\gamma_{1}^{\star}$ minimizes variance
$j=2$ has no bias/variance dominance $\gamma_{2}^{\star}$ is interior and $\gamma_{2}^{\star}$ balances bias/variance
Case 3 $j=1$ is bias dominant $\gamma_{1}^{\star}=1$ $\mathrm{arg}\min_{j}\left\{|\omega_{-j}|+|1-\gamma_{j}^{\star}||\omega_{j}|\right\}$ $j^{\star},\gamma_{1}^{\star}$ minimizes bias
$j=2$ has no bias/variance dominance $\gamma_{2}^{\star}$ is interior and $\gamma_{2}^{\star}$ balances bias/variance
Other (boundary) solutions
Case 4 Both $j\in\{1,2\}$ are variance-dominant $\gamma_{j}^{\star}=\frac{\sigma_{j}^{2}}{\sigma_{j}^{2}+v_{j}^{2}}$ $\mathrm{arg}\min_{j}\left\{\omega_{-j}^{2}\sigma_{-j}^{2}+\omega_{j}^{2}\frac{\sigma_{j}^{2}v_{j}^{2}}{\sigma_{j}^{2}+v_{j}^{2}}\right\}$ Minimizes variance
($\alpha/\alpha^{\star}>\beta/\beta^{\star}$ for all $\gamma$ and $j$) for $j\in\{1,2\}$
Case 5 Both $j\in\{1,2\}$ are bias-dominant $\gamma_{j}^{\star}=1$ $\mathrm{arg}\max_{j}|\omega_{j}|$ Minimizes bias
($\alpha/\alpha^{\star}<\beta/\beta^{\star}$ for all $\gamma$ and $j$) for $j\in\{1,2\}$
Case 6 bias dominant vs. variance dominant $\gamma_{1}^{\star}=1$ $\mathrm{arg}\min_{j}$ Minimizes bias $\beta/\beta^{\star}$ of $j=1$
($\alpha/\alpha^{\star}<\beta/\beta^{\star}$ for $j=1$ and vice versa for $j=2$) $\gamma_{2}^{\star}=\frac{\sigma_{2}^{2}}{\sigma_{2}^{2}+v_{2}^{2}}$ $\left\{\frac{|\omega_{2}|^{2}}{|\omega_{1}|^{2}}1\{j=1\}+(\omega_{1}^{2}\sigma_{1}^{2}+\omega_{2}^{2}\frac{\sigma_{2}^{2}v_{2}^{2}}{\sigma_{2}^{2}+v_{2}^{2}})1\{j=2\}\right\}$ vs. variance $\alpha/\alpha^{\star}$ of $j=2$
Illustration in Figure 3
For illustration, Figure 3 compares the maximum regret of choosing experiment $j=1$ versus $j=2$ in three scenarios. The vertical dashed line marks the value where the two curves intersect and the designer is indifferent between $j=1$ and $j=2$. In each column, we keep the same parameterizations as in Figure 2.
Observation 1: The regret for $j=2$ decreases rapidly in $\omega_{2}$ and then decreases more slowly. In the first column, we vary $\omega_{2}$ with the first experiment $j=1$ having a smaller experimental variance, $v_{1}=v_{2}/2$. The panel shows that a small $\omega_{2}$ makes the regret of choosing experiment $2$ larger than that of choosing experiment $1$. This is because a small $\omega_{2}$ corresponds to a small bias from not choosing $j=2$. As $\omega_{2}$ increases, the max–regret curve for $j=2$ declines until it reaches $\omega_{2}\approx 1.25$; after this point, the regret curve for $j=2$ rises slowly, since a further increase in $\omega_{2}$ is associated with an increase in estimator variance.
Observation 2: The regret for $j=1$ decreases slowly and then increases rapidly in $\omega_{2}$. The regret curve for choosing $j=1$ as we vary $\omega_{2}$ (fixing $\omega_{1}=1$) first decreases slowly and then increases. The reason is that for $\omega_{2}<\omega_{1}$, $\gamma_{1}^{\star}$ is an interior solution and the oracle squared bias is $\beta^{\star}=\omega_{2}^{2}$. Therefore, an increase in $\omega_{2}^{2}$ raises the oracle’s bias. When $\omega_{2}>\omega_{1}$, $\gamma_{1}^{\star}$ becomes a boundary solution and $\beta^{\star}=\omega_{1}^{2}$. In this case, a larger $\omega_{2}$ increases the variance of the estimator while the oracle bias $\beta^{\star}$ remains constant in $\omega_{2}$. As a result, it becomes more attractive for the analyst to run the experiment with $j=2$.
Observation 3: The regret for $j=2$ is monotonically increasing in $v_{2}^{2}$, and the opposite holds for $j=1$. The second plot shows that the regret from choosing $j=2$ increases monotonically with the experimental variance $v_{2}^{2}$, while the regret from choosing $j=1$ decreases correspondingly. The kinks in the curves are driven by regime shifts in the oracle solutions $(\beta^{\star},\alpha^{\star})$.
Observation 4: The regret for $j=2$ is monotonically decreasing in $\sigma_{2}^{2}$. The third plot shows that the regret for $j=2$ decreases monotonically with the observational variance $\sigma_{2}^{2}$. This is expected, since a larger $\sigma_{2}^{2}$ makes choosing $j=2$ more appealing relative to relying on observation.
Observation 5: The regret for $j=1$ first decreases slowly and then increases rapidly in $\sigma_{2}^{2}$. As we vary $\sigma_{2}^{2}$, the regret for $j=1$ is initially (very) slowly decreasing. The reason is that a larger $\sigma_{2}^{2}$ increases the oracle variance $\alpha^{\star}$ faster than the variance of the estimator for $j=1$ (with $\gamma_{1}^{\star}\approx 1$). However, this behavior does not affect the optimal solution: for smaller values of $\sigma_{2}^{2}$, choosing $j=1$ remains preferable to choosing $j=2$. When $\sigma_{2}^{2}$ is larger than a tipping point, however, the regret of choosing $j=1$ increases rapidly with $\sigma_{2}^{2}$, making the first experiment no longer preferable to the second. This aligns with the intuition that, as observational variance grows, we should favor conducting the experiment for $j=2$.
In sum, these patterns show how our framework disentangles the competing forces across signal strength and observational/experimental noise, guiding the analyst to transparent design choices even in complex regimes. Table 3 summarizes the discussion.
Table 3: Qualitative behavior of maximum regret as $\omega_{2}$, $v_{2}$, or $\sigma_{2}$ increase (rows) in Figure 3. Entries summarize the direction and relative speed of change in the maximum regret for choosing experiment $j=1$ or $j=2$.
Small regime ($\omega_{2}<\omega_{1}$) Large regime ($\omega_{2}>\omega_{1}$)
Trend regret $j=1$ regret $j=2$ regret $j=1$ regret $j=2$
$\uparrow~\omega_{2}$ slow decrease fast decrease fast increase slow increase
$\uparrow~v_{2}$ decrease increase decrease increase
$\uparrow~\sigma_{2}$ slow decrease decrease fast increase decrease
Figure 3: Regret comparisons. Top row: $\max\{\alpha(j,\gamma_{j}^{\star})/\alpha^{\star},\beta(j,\gamma_{j}^{\star})/\beta^{\star}\}$ for $j\in\{1,2\}$ as the x–axis parameter varies (columns: $\omega_{2}$, $v_{2}$, $\sigma_{2}$). The vertical dashed line marks indifference, where the two curves intersect; to its left/right, the optimal experiment is the one with lower max regret. Bottom row: optimal weights $\gamma_{j}^{\star}$ for $j=1,2$, which explain how sensitivity and precision interact to drive the design switch at $x^{\star}$. Columns vary, respectively, (a) $\omega_{2}$ with $v_{1}=0.5$, $v_{2}=1$, $\sigma_{1}=\sigma_{2}=1$, $\omega_{1}=1$; (b) $v_{2}$ with $\omega=(0.9,1)$, $v_{1}=1$, $\sigma=(1,1)$; (c) $\sigma_{2}$ with $\omega=(0.9,1)$, $v=(1,1)$, $\sigma_{1}=1$.
4 Extensions
4.1 Nonlinear estimand $\tau(\theta)$
This subsection extends the framework to smooth, potentially nonlinear $\tau(\theta)$, and consider a local asymptotic framework in the spirit of Andrews et al. (2020); Bonhomme and Weidner (2022).
Setup.
Take any given $(\mathcal{E},\Sigma)$ and $\gamma$. Let $\tilde{\theta}^{\mathrm{obs}}$ and $\tilde{\theta}^{\mathrm{exp}}$ denote the corresponding observational and experimental estimators of $\theta\in\mathbb{R}^{p}$. Assume a common $\sqrt{n}$ rate and the following local asymptotics:
$\sqrt{n}\!\left(\tilde{\theta}^{\mathrm{exp}}-\theta\right)$ $\;\overset{d}{\longrightarrow}\;\mathcal{N}(0,\Sigma^{\mathrm{exp}}),$
$\sqrt{n}\!\left(\tilde{\theta}^{\mathrm{obs}}-\theta-b_{n}\right)$ $\;\overset{d}{\longrightarrow}\;\mathcal{N}(0,\Sigma^{\mathrm{obs}}),\qquad b_{n}\to 0,$
where $b_{n}\in\mathbb{R}^{p}$ is an unknown drift capturing local misspecification, and $\Sigma^{\mathrm{exp}},\Sigma^{\mathrm{obs}}$ are finite positive–definite matrices. We take the experimental and observational estimators to be asymptotically independent with common rates; this is not necessary but simplifies exposition. Thus $\tilde{\theta}^{\mathrm{obs}}\stackrel{{\scriptstyle p}}{{\to}}\theta$ (consistency), while $\sqrt{n}\,b_{n}$ may be bounded or even diverge. To justify first order linearization when $\sqrt{n}\,b_{n}$ may grow, we impose $\sqrt{n}\,\|b_{n}\|^{2}\to 0$ (e.g., $\|b_{n}\|=n^{-\alpha}$ with $\alpha\in(1/4,1/2]$). This implies that even if $b_{n}$ converges to zero, the asymptotic bias is possibly comparable to, smaller or larger than the asymptotic variance, therefore imposing weak restrictions on the magnitude of $\sqrt{n}b_{n}$ relative to the asymptotic variance.
Linearization of $\tau$.
Let $\tau:\mathbb{R}^{p}\to\mathbb{R}$ be twice continuously differentiable in a neighborhood of $\theta$, with Lipschitz gradient and bounded Hessian. A Taylor expansion around $\theta$ yields
$\tau(\hat{\theta}_{\gamma})-\tau(\theta)=\omega(\theta)^{\top}(\hat{\theta}_{\gamma}-\theta)\;+\;\tfrac{1}{2}(\hat{\theta}_{\gamma}-\theta)^{\top}H_{\tau}(\bar{\theta})\,(\hat{\theta}_{\gamma}-\theta),$
[…]
$\sqrt{n}\bigl(\tau(\hat{\theta}_{\gamma})-\tau(\theta)\bigr)=\omega(\theta)^{\top}\sqrt{n}\,(\hat{\theta}_{\gamma}-\theta)\;+\;o_{p}(1).$ (7)
Using the fact that $\tilde{\theta}^{\mathrm{obs}}\rightarrow_{p}\theta$, we can consistently estimate the asymptotic MSE $n\mathbb{E}\left[(\tau(\hat{\theta}_{\gamma})-\tau(\theta))^{2}\right]$ by replacing $\omega(\theta)$ with $\omega(\tilde{\theta}^{\mathrm{obs}})\rightarrow_{p}\omega(\theta)$ and ignoring the $o_{p}(1)$ on the right-hand side. As a result all our results continue to hold under such linearization up-to a negligible $o_{p}(1)$.
4.2 Confidence interval length as a regret criterion
In this subsection we consider an alternative loss based on the length of confidence intervals (CIs). For a candidate design $(\mathcal{E},\Sigma(\mathcal{E}))$, shrinkage weights $\gamma$, and a bias vector $b\in\mathbb{R}^{p}$, let $\tilde{\omega}(\mathcal{E},\gamma)$ be as in (3). Define the two–sided, $(1-\eta)$ bias–aware lower and upper bounds for $\tau(\theta)$ by
$\ell_{b}(\mathcal{E},\Sigma,\gamma)$ $\equiv\hat{\tau}_{\gamma}\;-\;(\omega\circ(1-\gamma))^{\top}b\;-\;z_{1-\eta/2}\,\sqrt{\tilde{\omega}(\mathcal{E},\gamma)^{\top}\Sigma(\mathcal{E})\,\tilde{\omega}(\mathcal{E},\gamma)},$
$u_{b}(\mathcal{E},\Sigma,\gamma)$ $\equiv\hat{\tau}_{\gamma}\;-\;(\omega\circ(1-\gamma))^{\top}b\;+\;z_{1-\eta/2}\,\sqrt{\tilde{\omega}(\mathcal{E},\gamma)^{\top}\Sigma(\mathcal{E})\,\tilde{\omega}(\mathcal{E},\gamma)},$
where $z_{1-\eta/2}$ is the standard normal $(1-\eta/2)$ quantile.
An audience, indexed by a worst–case bias radius $B\geq 0$, forms the worst–case CI
$L_{B}(\mathcal{E},\Sigma,\gamma)~\equiv~\left[\inf_{\|b\|_{\infty}\leq B}\ell_{b}(\mathcal{E},\Sigma,\gamma),\;\sup_{\|b\|_{\infty}\leq B}u_{b}(\mathcal{E},\Sigma,\gamma)\right],$
where we write $|L_{B}(\mathcal{E},\Sigma,\gamma)|$ for its total length. The proportional regret of $(\mathcal{E},\Sigma,\gamma)$ is
$\tilde{\mathcal{R}}(\mathcal{E},\Sigma,\gamma)~\equiv~\sup_{B\geq 0}\;\frac{|L_{B}(\mathcal{E},\Sigma,\gamma)|}{\inf_{(\mathcal{E}^{\prime},\Sigma^{\prime},\gamma^{\prime})\in\mathcal{D}}|L_{B}(\mathcal{E}^{\prime},\Sigma^{\prime},\gamma^{\prime})|}\!.$
Theorem 2.
Consider Setting 1 and let Assumption 1 hold. Then, for any $(\mathcal{E},\Sigma,\gamma)$,
[…]
with $\alpha$ and $\beta$ as defined in Equation (4) and $\alpha^{\star},\beta^{\star}$ as defined in Equation (5).
Proof.
See Appendix C.1
Theorem 2 shows that the objective function is the same as the MSE-objective function up to a monotonic (square-root) transformation which does not affect the choice of the minimizer.
4.3 Vector valued estimands
Next, we extend our framework to vector value estimands of the form $\tau(\theta)\in\mathbb{R}^{q}$ for $q\geq 1$.
Setting 3 (Vector-valued estimands).
Consider a vector-valued target estimand with entries $\tau^{\ell}(\theta)\equiv(\omega^{\ell})^{\top}\theta$ for $\ell=1,\dots L$, and $\omega^{\ell},\theta\in\mathbb{R}^{k}$.
For a subset $\mathcal{E}\subseteq\{1,\ldots,p\}$ and a positive-definite matrix $\Sigma(\mathcal{E})\in\mathcal{G}(\mathcal{E})\subset\mathbb{R}^{(p+|\mathcal{E}|)\times(p+|\mathcal{E}|)}$, define $\tilde{\theta}^{\mathrm{obs}}\in\mathbb{R}^{p}$ an observational estimate and $\tilde{\theta}^{\mathrm{exp}}_{\mathcal{E},\Sigma}\in\mathbb{R}^{|\mathcal{E}|}$ an experimental estimate each satisfying
$\mathbb{E}\!\left[\tilde{\theta}^{\mathrm{obs}}\right]-\theta=b,\qquad\mathbb{E}\!\left[\tilde{\theta}^{\mathrm{exp}}_{\mathcal{E},\Sigma}\right]-\theta_{\mathcal{E}}=0,$
for an unknown bias vector $b\in\mathbb{R}^{p}$. Moreover, define its joint variance as
$\mathbb{V}\!\begin{pmatrix}\tilde{\theta}^{\mathrm{obs}}-\theta\\[2.0pt] \tilde{\theta}^{\mathrm{exp}}_{\mathcal{E},\Sigma}-\theta_{\mathcal{E}}\end{pmatrix}=\Sigma(\mathcal{E})$
assumed to be known, with uniformly bounded entries and strictly positive definite.
Given $(\mathcal{E},\Sigma)$, consider a class of linear plug-in estimators to allow the optimal shrinkage weights to depend on the specific entry of $\tau$.
$\hat{\tau}^{\ell}_{\gamma}\equiv\tau^{\ell}(\hat{\theta}^{\ell}(\gamma)),\qquad\hat{\theta}^{\ell}_{j}(\gamma)=\begin{cases}\gamma_{j}^{\ell}\,\tilde{\theta}_{j}^{\mathrm{exp}}+(1-\gamma_{j}^{\ell})\,\tilde{\theta}_{j}^{\mathrm{obs}},&j\in\mathcal{E},\\[2.0pt] \tilde{\theta}_{j}^{\mathrm{obs}},&j\notin\mathcal{E},\end{cases}$
where $\gamma=(\gamma^{\ell}_{j})_{\ell\in[L],j\in\mathcal{E}}\in\mathbb{R}^{L\times|\mathcal{E}|}$ are shrinkage weights that can be an (implicit) function of $(\mathcal{E},\Sigma)$. Let $\hat{\theta}(\gamma)\equiv(\hat{\theta}^{1}(\gamma),\dots,\hat{\theta}^{L}(\gamma))\in\mathbb{R}^{p\times L}$ and $\hat{\tau}_{\gamma}\equiv(\hat{\tau}^{1}_{\gamma}(\hat{\theta}^{1}(\gamma)),\dots,\hat{\tau}^{L}_{\gamma}(\hat{\theta}^{L}(\gamma)))^{\top}\in\mathbb{R}^{L}$.
Setting 3 is a flexible generalization of Setting 1, allowing each entry of $\tau$ to have its own vector of shrinkage weights.
Similar to the scalar case, we make the following assumption regarding the first-order estimation error.
Assumption 2 (First-order estimation error).
For any admissible $(\mathcal{E},\Sigma,\gamma)\in\mathcal{D}$, assume for all $\ell=1,\dots,L$
$\tau^{\ell}(\theta)-\tau^{\ell}\!\big(\hat{\theta}^{\ell}(\gamma)\big)=(\omega^{\ell})^{\top}\,\big(\theta-\hat{\theta}^{\ell}(\gamma)\big),$ (8)
for a known weighting vector $\omega^{\ell}\in\mathbb{R}^{p}$, with $|\omega^{\ell}_{i}|>0$, for all $i,\ell$.
We begin by generalizing the mean squared error to higher-dimensional parameters. For a given design $(\mathcal{E},\Sigma)$ and shrinkage weights $\gamma$, define the MSE at a fixed observational-bias vector $b$ as
$\mathrm{MSE}_{b}(\mathcal{E},\Sigma,\gamma)\;=\;\mathbb{E}_{\mathcal{E},\Sigma,b}\!\left[\|\widehat{\tau}_{\gamma}-\tau\|_{2}^{2}\right],$ (9)
where $\mathbb{E}_{\mathcal{E},\Sigma,b}$ denotes expectation under the data-generating process implied by $(\mathcal{E},\Sigma(\mathcal{E}))$ and observational bias $b$, and $\|\cdot\|_{2}$ denotes the $L^{2}$ Euclidean norm.
As before, we introduce more compact notation before proceeding to the main result. For a given $(\mathcal{E},\gamma)$ define
$\tilde{\omega}^{\ell}_{\text{obs},j}(\mathcal{E},\gamma)=\begin{cases}\omega^{\ell}_{j},&j\notin\mathcal{E},\\ \omega^{\ell}_{j}(1-\gamma^{\ell}_{j}),&j\in\mathcal{E},\end{cases}\qquad\tilde{\omega}^{\ell}_{\text{exp},j}(\mathcal{E},\gamma)=\begin{cases}0,&j\notin\mathcal{E},\\ \omega^{\ell}_{j}\gamma^{\ell}_{j},&j\in\mathcal{E},\end{cases}$
and let
$\tilde{\omega}^{\ell}(\mathcal{E},\gamma)\equiv\begin{pmatrix}\tilde{\omega}^{\ell}_{\text{obs}}(\mathcal{E},\gamma)\\[2.0pt] \tilde{\omega}^{\ell}_{\text{exp}}(\mathcal{E},\gamma)\end{pmatrix}\in\mathbb{R}^{\,p+|\mathcal{E}|}.$
Using this notation, we can write $\hat{\tau}^{\ell}_{\gamma}=\tilde{\omega}^{\ell}(\mathcal{E},\gamma)^{\top}\,\hat{\theta}$, where $\hat{\theta}=(\tilde{\theta}^{\text{obs}\top},\tilde{\theta}^{\text{exp}\top})^{\top}\in\mathbb{R}^{p+|\mathcal{E}|}$. Stacking across the $\ell$ index gives the estimator for the full vector $\tau$: $\hat{\tau}_{\gamma}=\tilde{\omega}(\mathcal{E},\gamma)\,\hat{\theta}$, where the $\ell^{th}$ row of $\tilde{\omega}(\mathcal{E},\gamma)\in\mathbb{R}^{L\times(p+|\mathcal{E}|)}$ is $\tilde{\omega}^{\ell}(\mathcal{E},\gamma)^{\top}$.
The estimator’s variance and worst–case squared bias divided by $B^{2}$ are
$\check{\alpha}(\mathcal{E},\Sigma,\gamma)$ $\equiv\text{Trace}\left(\tilde{\omega}(\mathcal{E},\gamma)\Sigma(\mathcal{E})\,\tilde{\omega}(\mathcal{E},\gamma)^{\top}\right),$ (10)
$\check{\beta}(\mathcal{E},\gamma)$ $\equiv\sum_{\ell=1}^{L}\Big(\,(\omega^{\ell})^{\top}\bar{v}(\gamma)-(\omega^{\ell}_{\,\mathcal{E}})^{\top}\bar{v}(\gamma)_{\,\mathcal{E}}+\,(1-\gamma^{\ell})^{\top}(\bar{v}(\gamma)_{\,\mathcal{E}}\circ\omega_{\,\mathcal{E}})\,\Big)^{\!2},$
where $\omega^{\ell}_{\mathcal{E}}$ is the subvector of $\omega^{\ell}$ corresponding to the entries indexed by the experiment $\mathcal{E}$, and
$\bar{v}(\gamma)\in\operatorname*{arg\,max}_{\{v\in\mathbb{R}^{p}:|v_{j}|\leq 1\forall j\}}\sum_{\ell=1}^{L}\Big(\,(\omega^{\ell})^{\top}v-(\omega^{\ell}_{\,\mathcal{E}})^{\top}v_{\,\mathcal{E}}+\,(1-\gamma^{\ell})^{\top}(v_{\,\mathcal{E}}\circ\omega_{\,\mathcal{E}})\,\Big)^{\!2},$
where the dependence of $\bar{v}(\gamma)$ on $\mathcal{E}$ is omitted for conciseness.
The smallest worst-case variance and bias are defined as
$\check{\alpha}^{\star}=\min_{(\mathcal{E},\Sigma,\gamma)\in\mathcal{D}}\alpha(\mathcal{E},\Sigma,\gamma),\quad\check{\beta}^{\star}=\min_{(\mathcal{E},\Sigma,\gamma)\in\mathcal{D}}\beta(\mathcal{E},\gamma).$
As before, we are interesting in minimizing
$\check{\mathcal{R}}(\mathcal{E},\Sigma,\gamma)\equiv\sup_{B\geq 0}\frac{\sup_{b\in\mathcal{B}(B)}\mathrm{MSE}_{b}(\mathcal{E},\Sigma,\gamma)}{\inf_{(\mathcal{E},\Sigma,\gamma)\in\mathcal{D}}\sup_{b\in\mathcal{B}(B)}\mathrm{MSE}_{b}(\mathcal{E},\Sigma,\gamma)}.$
The result below shows how our results directly extend to this case.
Theorem 3.
Consider Setting 3 and let Assumption 2 hold. Then, for any $(\mathcal{E},\Sigma,\gamma)$,
$\check{\mathcal{R}}(\mathcal{E},\Sigma,\gamma)\;=\;\max\!\left\{\frac{\check{\alpha}(\mathcal{E},\Sigma,\gamma)}{\check{\alpha}^{\star}},\ \ \frac{\check{\beta}(\mathcal{E},\gamma)}{\check{\beta}^{\star}}\right\}.$
Proof.
See Appendix C.2
5 Empirical applications
In this section we provide two empirical applications. In the first application we study the problem of choosing where to conduct an experiment using observational data from Banerjee et al. (2024). In the second application we illustrate the method for choosing treatment arms when estimating a structural model, calibrating our estimation to the PROGRESA experiment (Todd and Wolpin, 2006; Attanasio et al., 2012). We solve the optimization program by solving for each choice of experiments, both over $\gamma$ and the allocation of sample size via numerical optimizationand then enumerate the solutions across all possible choices.
5.1 Choosing experimental participants
As a first exercise, we show how observational evidence can inform whom to recruit into an experiment. We consider the setting in Banerjee et al. (2024), who study how the expansion of microfinance affects village social networks in Karnataka, India.
In their observational analysis, the authors assemble panel data from 75 villages in Karnataka, 43 of which were exposed to microfinance. Because program rollout was not randomized across villages, they estimate effects using difference-in-differences (DiD). While DiD is informative, the lack of randomized assignment may bias estimates in the presence of selection and lack of parallel-trends (e.g., Marx et al., 2024; Ghanem et al., 2022). Motivated by these limitations, the authors subsequently conducted an experimental evaluation in one metropolitan area, randomizing microfinance access across 104 urban neighborhoods. We only use observational variation for choosing the experimental design and estimator. We then use experimental variation generated by Banerjee et al. (2024) to validate our procedure.
Using preliminary observational estimates from Banerjee et al. (2024) in Karnataka, we ask: Which area(s) in Karnataka should be prioritized for an experimental evaluation, and how many villages should be enrolled? Concretely, we consider designs that randomize the introduction of microfinance in $n$ villages within one or more selected areas and study how the choice of areas and the allocation of $n$ vary with design constraints. Our objective is to design an experiment that yields an externally valid estimate of the effect of microfinance for rural Karnataka. We focus on the first outcome reported by Banerjee et al. (2024) corresponding to the density of the network, i.e., the percentage of connections a random household in a village has relative to the village size.
In practice, implementation costs often depend on how geographically dispersed the study sites are. Coordinating with microfinance partners, training and supervising field teams, and conducting surveys become more expensive when sites are far apart or span multiple administrative units. We therefore consider operational constraints that limit the dispersion of enrolled villages and number of villages enrolled in the area.
Figure 4: Observational villages and area-level DiD effects in Karnataka. The background heat map partitions the state into the four contiguous areas used to compute area-specific DiD estimates of microfinance’s impact on network density; the color scale encodes the DiD value. Points mark village locations: circles denote villages exposed to microfinance and triangles denote unexposed villages in the observational sample (authors’ survey). Major cities (Bengaluru, Mysuru, Tumakuru) are labeled for orientation.
Observational study
We use observational evidence from Banerjee et al. (2024) to construct area-level difference-in-differences (DiD) estimates for Karnataka. Specifically, we partition the observational sample into four geographically contiguous areas that group nearby villages; each area contains 11–12 villages that were exposed to microfinance during the study period. Figure 4 maps these four areas and reports the corresponding DiD point estimates. Three of the four areas display negative estimated effects, with meaningful variation in magnitudes across areas.
Table 4 summarizes, for each area, the number of treated and untreated villages, the pre-treatment sample variances, and the DiD point estimates together with their variances, computed assuming independence across villages. [^10] Alternative variance estimators (e.g., allowing within-area correlation) can be incorporated without changing the design logic.
These area-specific DiD estimates and their variances serve as our primary observational inputs to the experimental design problem. [^11] We use DiD estimates to mimic the estimator used by the authors. An analogous analysis can be conducted after empirical-Bayes shrinkage of area-level estimates, omitted for brevity.
Table 4: Area-level observational inputs (pre-prior) for Karnataka villages. The outcome is network density (corresponding to the average share of connection of an individual to other individuals in a village). For each area, we report the numbers of treated ($n_{1}$) and untreated ($n_{0}$) villages, the pre-intervention variance of density pooled across arms ($v_{\text{pre}}^{2}$), the area-specific DiD estimate $\hat{\mu}$, and its sampling variance $\widehat{\sigma}^{2}$.
Area $n_{1}$ $n_{0}$ $v_{\text{pre}}^{2}$ $\hat{\mu}$ $\widehat{\sigma}^{2}$
11 $11$ 66 0.000 4570.000\,457 0.0187-0.0187 0.000 045 30.000\,045\,3
22 $12$ 99 0.001 750.001\,75 0.0377-0.0377 0.000 1400.000\,140
33 $11$ $12$ 0.001 380.001\,38 0.003 90-0.003\,90 0.000 1870.000\,187
44 $11$ 66 0.000 9320.000\,932 $0.0148$ 0.000 1300.000\,130
Experimental design
We consider a family of designs that (i) select $E\in\{1,\ldots,4\}$ geographic areas in Karnataka from which to recruit experimental sites and (ii) assign $n_{1}$ villages to treatment and $n_{0}=n_{1}$ to control (so the total sample is $n=2n_{1}$). Here, $E=1$ corresponds to recruiting from a single area, while $E=4$ recruits from all four areas. We examine a grid of sample sizes that varies the number of participants from $10$ up to $104$ ($52$ treated units), with the latter corresponding to the size of the experiment in Banerjee et al. (2024). For variance calibration, we take $v_{\mathrm{pre},a}^{2},a\in\{1,\cdots,4\}$ to be the pre-intervention, area-level variance of network density in Table 4. We assume that the variance of a single treated–control difference in the experiment is $2v_{\mathrm{pre},a}^{2}$.
Results
We begin by studying which areas are selected and how sample is allocated when the total experimental sample is large ($n_{1}=52$) and the number of admissible areas may or may not be constrained. Figure 5 visualizes the resulting allocation across Karnataka’s four observational areas for $E\in\{1,2,3,4\}$. With $E=1$, the algorithm concentrates recruitment in the Tumakuru area—the location with the highest uncertainty in the observational estimates. As $E$ relaxes to $2$ and $3$, additional areas enter and the total sample splits across them in roughly—but not exactly—equal shares, reflecting a trade-off between (i) exploiting heterogeneity in the observational variance and (ii) keeping experimental variance low. When $E=4$, all areas are eligible and the allocation smooths further across space. In the rest of our discussion, we discuss properties when not all areas may be selected (E<4E<4). [^12] The solution for $E=4$ always favors minimizing the bias first, since $\beta^{\star}=0$.
Figure 6 tracks how area-level allocations evolve with the total number of villages. When the experiment is small (e.g., $10$ villages) and $E=1$, the algorithm favors an area with relatively low experimental variance—even if its observational estimate is less uncertain—because, at small $n$, experimental noise is first-order. As the total sample increases, the weight shifts toward areas with more uncertain observational estimates. Consistently, the area with the largest experimental variance (Area 2) is typically excluded unless all four areas can be chosen.
Figure 7 reports the optimal shrinkage $\gamma_{j}^{\star}$ by area as total villages vary. The solution is interior for all $E<4$: $\gamma_{j}^{\star}$ rises with sample size, approaching one from below as experimental noise diminishes. For example, at a moderate sample (around 15 villages), $\gamma^{\star}$ is about $0.65$ when $E\in\{2,3\}$; with $E=1$, concentrating the sample within one area lowers experimental variance and therefore yields a larger $\gamma^{\star}$.
Main MSE comparison
We compare three strategies: (i) our chosen design which chooses the area, sample size and $\gamma^{\star}$ as described in Section 3, (ii) a standard benchmark that selects areas uniformly at random, allocates villages evenly across the chosen areas, and sets $\gamma_{j}=1$ for experimental estimates, and (iii) an oracle that knows the bias vector $b$ and optimizes both design (areas and allocation) and $\gamma$.
Because $b$ is unknown in practice, for this exercise, we calibrate it using the difference between the Hyderabad RCT ATE obtained from the follow-up experiment of Banerjee et al. (2024) and the Karnataka DiD average effect across sites, treating $b$ as common across areas.
Figure 8 plots the ratio $\text{MSE}/\text{MSE}^{\star}$, where $\text{MSE}^{\star}$ is the oracle’s MSE. As the number of treated villages $n_{1}$ increases, all designs improve; because the denominator also falls with $n_{1}$, the ratio need not be monotone. Across $E\in\{1,2,3\}$, the random benchmark remains well above the oracle—by roughly 3–5 percentage points (pp) for $E=1$, about 10 pp for $E=2$, and around 20 pp for $E=3$ at the top of the grid—whereas our design tracks the oracle closely (within a few percentage points) even without knowledge of $b$. The gap is the largest when a larger number of areas $E$ can be included, because our procedure can better emulate the oracle by selecting (and differentially weighting) favorable areas. On the other hand, the random benchmark ignores information about observational and experimental variance heterogeneity. As sample size increases, the gap between the two stabilizes: with large $n_{1}$, optimal shrinkage approaches $\gamma^{\star}\!\approx\!1$ and an even allocation across the selected areas is nearly optimal once the area set is well chosen.
This empirical calibration highlights the benefit of the procedure that is able to track the oracle even in the absence of knowledge of the bias $b$, while significantly outperforming standard alternative benchmarks such as a random allocation strategy.
Figure 5: Optimal area selection and village allocation for a large experiment ($n_{1}=52$) under constraints on the number of areas $E$. Each panel corresponds to a value of $E$; shading indicates the total number of recruited villages in each area. With $E=1$ the design concentrates in the area with the noisiest observational estimate; as $E$ increases, allocation spreads across areas, reflecting a trade-off between observational uncertainty and experimental variance.
Figure 6: Area-level allocations as total treated villages $n_{1}$ varies (from ten to fifty-two villages), for each constraint on the number of eligible areas $E$. Lines show the number of villages assigned to each area.
Figure 7: Optimal shrinkage $\gamma_{j}^{\star}$ by area as a function of total treated villages $n_{1}$ (from ten to fifty-two villages), under different constraints $E$. For $E<4$, the solution is interior ($0<\gamma_{j}^{\star}<1$) and increases with $n_{1}$, approaching one as experimental noise falls.
Figure 8: Relative MSE across designs under an hypothetical scenario where the bias equal the difference between the experimental and observational ATE estimate from Banerjee et al. (2024) as $n_{1}$ varies (from ten to fifty-two villages). The figure reports $\mathrm{MSE}/\mathrm{MSE}^{\star}$ for (i) the proposed design (area targeting and optimal $\gamma^{\star}$), (ii) a random benchmark (uniform area selection and equal allocation), where $\mathrm{MSE}^{*}$ denotes the MSE of the oracle that knows the bias, as $n_{1}$ varies and for each $E$. The proposed design closely tracks the oracle across $n_{1}$ and $E$, while the random benchmark remains substantially above, especially when $E$ is small.
5.2 Choosing treatment arms for model estimation
Next, we illustrate how our framework selects among treatment arms when the goal is to estimate and deploy a structural model. Combining experimental or pre-program observational data with structural models is increasingly used for program evaluation (Todd and Wolpin, 2006; Attanasio et al., 2012; Meghir et al., 2022). Because large-scale experimentation may be infeasible due to cost or fairness constraints (Muralidharan and Niehaus, 2017), researchers are often interested in designing small-scale experiments whose information, combined with a model, is informative about general-equilibrium (GE) effects.
We consider a researcher evaluating a conditional cash-transfer (CCT) for sending children to school in rural Kenya where such a program is not in place. Due to budget and feasibility constraints, the researcher can only randomize at small scale (partial equilibrium), but ultimately wishes to predict large-scale GE effects. As preliminary observational estimates, we use information from the Mexican PROGRESA experiment.
5.2.1 Model
Individual choice model
Following Bonhomme and Weidner (2022) (and in turn Todd and Wolpin (2006)), let $S\in\{0,1\}$ denote school attendance, $C$ consumption, $Y$ (pre-transfer) household income, $W$ the child’s potential wage, and $t$ the stipend when enrolled. Abstracting from covariates for exposition (we will introduce covariates in the estimation), utility is
$U(C,S,t,\varepsilon)=\xi_{1}\,C+\xi_{2}\,CS+(\xi_{3}-\xi_{1}-\xi_{2})\,tS+\xi_{4}\,S+S\varepsilon,\qquad\varepsilon\sim\mathcal{N}(0,1),$ (11)
with budget C=Y+W(1S)+tSC=Y+W(1-S)+tS. The parametrization $(\xi_{3}-\xi_{1}-\xi_{2})$ is without loss and simplifies expressions below. Enrollment satisfies
$S=1\big\{U(Y+t,1,t,\varepsilon)>U(Y+W,0,0,0)\big\}.$
Letting $Z(Y,W,t)\equiv\xi_{1}W-\xi_{2}Y-\xi_{3}t-\xi_{4},$ we have $P(S=1\mid Y,W,t)=\Phi\!\big(-Z(Y,W,t)\big)$.
General equilibrium effects
We are interested in the effect of a small stipend $t$ to all eligible (poor) households in rural Kenya. GE feedback is allowed through income and wages: for functions $y(t),w(t)$ and mean-zero idiosyncratic income and wage shocks $\varepsilon_{Yi},\varepsilon_{Wi}$, we assume
$Y_{i}(t)=y(t)+\varepsilon_{Yi},\qquad W_{i}(t)=w(t)+\varepsilon_{Wi}.$
Let $y_{0}\equiv\partial_{t}y(t)|_{t=0}$ and $w_{0}\equiv\partial_{t}w(t)|_{t=0}$. Define $\phi_{0}\equiv\mathbb{E}\!\left[\phi\!\big(-Z\big(Y(0),W(0),0\big)\big)\right]$. Our estimand of interest is the marginal effect of an increase in a small transfer $t$ to all eligible households
$\left.\frac{\partial\,\mathbb{E}\!\left[\Pr(S{=}1\mid Y(t),W(t),t)\right]}{\partial t}\right|_{t=0}=\;\phi_{0}\cdot\big(\,\xi_{3}+\xi_{2}y_{0}-\xi_{1}w_{0}\,\big),$ (12)
which decomposes into the direct stipend effect $\phi_{0}\xi_{3}$ and the indirect GE effects via income and wages, $\phi_{0}\xi_{2}y_{0}$ and $\phi_{0}\xi_{1}w_{0}$.
Income and wage effects
We estimate $y_{0}=1.5$ using a local fiscal/income multiplier $M_{\text{tot}}\approx 2.5$ from Egger et al. (2022), which in turn implies $y_{0}=M_{\text{tot}}-1$. [^13] Egger et al. (2022) report a total income/consumption multiplier $M_{\text{tot}}=\partial_{t}\mathbb{E}\big[Y_{\text{pre}}(t)+t\big]|_{t=0}$; therefore the target derivative of pre-transfer income is $y_{0}=\partial_{t}\mathbb{E}\big[Y_{\text{pre}}(t)\big]|_{t=0}=M_{\text{tot}}-1$. We write $w_{0}$ as a function of the parameters: Under simple market clearing for labor supply and demand, we can show [^14] This follows from the following: set $L_{s}(W,t)=L_{d}(W)$, the labor supply equals the labor demand, and differentiate at $t=0$: $\frac{\partial L_{s}}{\partial W}w_{0}+\frac{\partial L_{s}}{\partial t}=\frac{\partial L_{d}}{\partial W}w_{0}$. Suppose we have a constant number of hours worked for working child, denoted as $H$, and any child not at school is working. From the probit index, $\partial_{t}\mathbb{E}[S\,|\,\cdot]=\phi(-Z)\,\xi_{3}$ and $\partial_{W}\mathbb{E}[S\,|\,\cdot]=-\phi(-Z)\,\xi_{1}$ at $t=0$. Hence $\partial L_{s}/\partial t=-\phi_{0}\xi_{3}H$ and $\partial L_{s}/\partial W=\phi_{0}\xi_{1}H$. Dividing the equation by $H$ we obtain the desired result.
$w_{0}=\;\frac{\phi_{0}\,\xi_{3}}{\phi_{0}\,\xi_{1}-d},$ (13)
where $d$ denotes the slope of the demand curve divided by total hours of work. We calibrate $d$ to $0.5\frac{1-S_{0}}{W_{0}}$, where $S_{0}$ and $W_{0}$ are baseline probability to go to school (assuming all children not in school go to work) and baseline wages calibrated to the pre-experimental data from PROGRESA and $0.5$ obtained form the labor demand elasticity in Espey and Thilmany (2000). For expositional convenience, we treat $y_{0},d$ as constant, while $\phi_{0}\xi_{3}$ and $\phi_{0}\xi_{1}$ are parameters of the model and $w_{0}$ is a (non-linear) function of such parameters (it is possible to treat also $y_{0},d$ as additional possibly misspecified parameters, omitted for brevity). [^15] We treat $d$ as a fixed constant given the extensive literature on estimating demand elasiticities, and the fact that $S_{0}$, $W_{0}$ can be calibrated to pre-experimental data (for simplicity here using the PROGRESA pre-experimental data), and $y_{0}$ because computed experimentally in Kenya. When we are concerned with misspecification of those, researchers can augment $\theta$ to also incorporate $y_{0},d$ without having to estimate those estimated in the experiment.
Map to the design parameters
We can now define the parameters of interest as
$\theta_{1}\;\equiv\;\phi_{0}\,\xi_{2},\qquad\theta_{2}\;\equiv\;\phi_{0}\,\xi_{3},\qquad\theta_{3}\;\equiv\;-\phi_{0}\,\xi_{1},$ (14)
with
$\tau(\theta)\;=\;\theta_{2}\;+\;y_{0}\,\theta_{1}\;+\;w_{0}(\theta)\,\theta_{3},\qquad w_{0}(\theta)\;=\;\frac{\theta_{2}}{-\theta_{3}-d},$ (15)
where the explicit form of $w_{0}(\theta)$ follows from (13). Given observational estimates $\tilde{\theta}^{\mathrm{obs}}=(\tilde{\theta}^{\mathrm{obs}}_{1},\tilde{\theta}^{\mathrm{obs}}_{2},\tilde{\theta}^{\mathrm{obs}}_{3})$ described below, and following Section 4.1, we use $\omega=\frac{\partial\tau}{\partial\theta}\Big|_{\theta=\tilde{\theta}^{\mathrm{obs}}}$. [^16] Because $w_{0}$ is a function of $\theta$, we have $\frac{\partial\tau}{\partial\theta_{1}}=y_{0},\frac{\partial\tau}{\partial\theta_{2}}=\frac{d}{\theta_{3}+d},\frac{\partial\tau}{\partial\theta_{3}}=-\,\frac{\theta_{2}\,d}{(-\theta_{3}-d)^{2}}.$
5.2.2 Observational study and design of the experiment
Using experimental data from Mexico’s PROGRESA program, we estimate $(\xi_{1},\xi_{2},\xi_{3},\xi_{4})$ via probit with standard controls (age, distance to school, eligibility, year, highest grade attained). Estimation is conducted separately by gender, focusing here on the effects on female students. We pool treated and control observations to estimate the schooling effect $\xi_{4}$, controlling for the individual-level subsidy. In turn, we obtain observational estimates $\big(\tilde{\theta}_{1}^{\mathrm{obs}},\tilde{\theta}_{2}^{\mathrm{obs}},\tilde{\theta}_{3}^{\mathrm{obs}}\big)$ that map to the model parametrization. Standard errors are constructed using the Delta method with clustering at the village level. Point estimates, estimated sensitivity $\omega$, and standard errors are reported below.
Parameter $\tilde{\theta}^{\mathrm{obs}}$ $\omega$ $\sigma$
[…]
Observational estimates for female students, corresponding sensitivity parameter $\omega$, and standard errors.
$\Sigma^{\mathrm{obs}}\;=\;\begin{bmatrix}4.31&-11.31&5.57\\ -11.31&973.11&-126.16\\ 5.57&-126.16&1038.56\end{bmatrix}\times e^{-9}$
Variance–covariance matrix of observational estimates for female students.
We observe heterogeneity in both the sensitivity weights $\omega$ and (more significantly) the precision of the observational estimates. Our goal is to trade off these features, together with prospective experimental variation, to design the experiment in Kenya. Because the setting differs in time and country, PROGRESA-based estimates provide informative but possibly biased baselines.
Candidate small-scale treatment-arms in Kenya (partial equilibrium)
We consider a researcher who can afford only small, partial-equilibrium experiments in Kenya, as they may not be able to randomize treatments across the full population of interest. We examine three possible experiments for a given sample size $n$:
  • $j=\{1\}$: Unconditional transfer (income shock). The researcher randomizes a small income shock to a small fraction of households (implying no general-equilibrium effects). Under a first-order (Taylor) approximation, the experiment identifies
    $\left.\frac{\partial\,\mathbb{E}[\Pr(S{=}1\mid Y,W,0)]}{\partial Y}\right|_{t=0}=\theta_{1},$
    with precision $v_{1}^{2}/n$. The researcher optimizes over $\gamma_{1}$, the weight used to combine the experimental and observational estimates of $\theta_{1}$, while $\theta_{2}$ and $\theta_{3}$ are calibrated to the PROGRESA study.
  • $j=\{2\}$: Conditional cash transfer (stipend). The researcher randomizes a small stipend $t$, conditional on attending school, to a small fraction of households (with prices held fixed). Under a first-order approximation, this identifies
    $\left.\frac{\partial\,\mathbb{E}[\Pr(S{=}1\mid Y,W,t)]}{\partial t}\right|_{t=0}=\theta_{2},$
    with precision $v_{2}^{2}/n$. The researcher optimizes over $\gamma_{2}$, the weight used to combine the experimental and observational estimates of $\theta_{2}$, while $\theta_{1}$ and $\theta_{3}$ are calibrated to the PROGRESA study.
  • $j=\{1,2\}$: Two-arm design. The researcher runs (j=1j{=}1) and (j=2j{=}2) on independent samples, identifying $(\theta_{1},\theta_{2})$ with precisions $(v_{1}^{2}/n_{1},\ v_{2}^{2}/n_{2})$ where $n_{1}+n_{2}=n$. They optimize over both $\gamma_{1}$ and $\gamma_{2}$, while $\theta_{3}$ is calibrated to the PROGRESA study.
We consider two scenarios. In the first, the researcher (and oracle) can choose one or both experiments. In the second, they can choose only one of the two (either $j=\{1\}$ or $j=\{2\}$). Regret is defined accordingly, conditional on the set of feasible designs.
We calibrate $v_{1}^{2}$ as $\sigma_{1}^{2}\times n^{\mathrm{obs}}$, where $\sigma_{1}^{2}$ is the variance of $\tilde{\theta}_{1}^{\mathrm{obs}}$ and $n^{\mathrm{obs}}$ is the observational-sample size ($n^{\mathrm{obs}}=1089$ for the female sample using data from eligible students in Todd and Wolpin (2006)). We calibrate $v_{2}^{2}$ analogously as $\sigma_{2}^{2}\times n^{\mathrm{obs}}$.
Experiment type Identified parameter Choice variables Experiment standard error Description
$j=\{1\}$ $\theta_{1}$ $\gamma_{1}$ $6.56e^{-5}/\sqrt{n}$ Unconditional transfer (income shock, UCT). Randomize an income shock to a small fraction of households.
$j=\{2\}$ $\theta_{2}$ $\gamma_{2}$ $9.86e^{-4}/\sqrt{n}$ Conditional cash transfer (education shock, CCT). Randomize a conditional transfer to a small fraction of households.
$j=\{1,2\}$ $(\theta_{1},\,\theta_{2})$ $\gamma_{1},\ \gamma_{2},\ n_{1}$ $6.56e^{-5}/\sqrt{n_{1}}$ (arm 1), $9.86e^{-4}/\sqrt{n-n_{1}}$ (arm 2) Two-arm design. Run (j=1j{=}1) and (j=2j{=}2) on independent samples with sample sizes respectively $n_{1},n-n_{1}$.
Table 5: Design options, identified parameters, choice variables, and per-unit variances.
5.2.3 Results
Figure 9: Regret–optimal sample allocation across treatment arms when both the unconditional cash transfer (UCT; income shock) and the conditional cash transfer (CCT; stipend) arms are available. The optimal solution is interior: the vast majority of participants (over 90%) are assigned to CCT, with a small but nonzero fraction assigned to UCT to hedge misspecification. This pattern reflects the higher payoff to learning about CCT in this application relative to UCT.
Figure 10: Optimal shrinkage weights and design choice as functions of the total sample size $n$. Left panel: When both arms are run, the weight on the UCT experimental estimator is essentially one across $n$, while the weight on the CCT experimental estimator, $\gamma^{\star}_{\text{CCT}}$, rises with precision (from about 0.5 at small $n$ to roughly 0.9 by $n=1000$), reflecting the increasing value of experimental evidence. Right panel: When restricted to a single arm, the optimal choice switches from UCT at small $n$ (variance dominates under CCT) to CCT once $n$ crosses a threshold (around $n\approx 500$), where bias considerations dominate and favor CCT. Notes: $\gamma_{j}^{\star}$ is the optimal weight on the experimental (vs. observational) estimator for parameter $j$; we set $\gamma_{j}^{\star}=0$ to indicate that arm $j$ is not chosen.
Figure 11: Worst–case regret by design. Left panel: Allowing both arms with an interior allocation yields regret that approaches one as $n$ increases, indicating near–oracle performance under our normalization. In contrast, single–arm designs exhibit nonvanishing regret because bias remains first–order even as variance shrinks. Right panel: Focusing on the single–arm problem, UCT dominates only at small $n$; beyond the same threshold ($\approx 500$), CCT delivers uniformly lower regret as variance becomes second–order relative to bias. Notes: Regret is the worst–case objective normalized so that a value of one corresponds to the oracle benchmark; UCT = unconditional cash transfer, CCT = conditional cash transfer.
We now summarize results as we vary both the set of experiments the researcher can run and the total sample size $n$. Specifically, Figure 9 reports the sample allocation implied by the regret-optimal design when both treatment arms are available. The optimal allocation is interior: most participants (over roughly $90\%$) are assigned to the CCT arm (education), with a small but nonzero fraction assigned to the UCT arm (income shock). Two forces rationalize this pattern. First, the CCT parameter is more misspecification-sensitive ($\omega_{2}\approx 1.98$ vs. $\omega_{1}\approx 1.5$), so generating experimental evidence on CCT has a larger payoff in bias reduction. Second, the income effect $\theta_{1}$’s variance is about an order of magnitude smaller than the stipend effect $\theta_{2}$’s variance in our PROGRESA calibration, lowering the marginal returns to learn about UCT relative to CCT. Hence, while the designer keeps some allocation on UCT, most of the sample is placed on CCT; our framework makes this bias–variance trade-off explicit.
Alongside the sample allocation, our method optimizes the shrinkage weights on the parameters of interest. Figure 10 (left panel) shows that when both arms can be run, we place essentially full weight on the experimental estimator for UCT ($\gamma^{\star}_{\text{UCT}}\approx 1$ across $n$), while for CCT we gradually shift toward the experimental estimator as precision improves: $\gamma^{\star}_{\text{CCT}}$ rises from about $0.5$ at small $n$ to approximately $0.9$ by $n=1000$. Intuitively, higher $n$ lowers experimental variance, making it optimal to rely more heavily on experimental evidence for the CCT parameter.
It is also informative to compare optimal choices when the researcher is restricted to one arm. Figure 10 (right panel) displays the selected arm as a function of $n$ (together with the corresponding $\gamma_{j}^{\star}$). We encode “not chosen” as $\gamma_{j}^{\star}=0$. For small samples ($n\lesssim 500$), the optimal choice is UCT: the experimental CCT estimator would be too noisy, so the resulting variance under CCT dominates, while the bias advantage of CCT over UCT is limited because $\omega_{1}$ and $\omega_{2}$ are of similar magnitude. Once $n$ crosses a critical threshold (around $n\approx 500$), CCT becomes dominant: by combining observational and experimental estimates for CCT, we can achieve comparable variance, and the worst-case bias is lower than under UCT.
Finally, Figure 11 reports worst-case regret for each design choice. The left panel shows that the regret of our method (green line, running both arms with an interior allocation) declines toward one as $n$ increases, reflecting near-oracle performance under our normalization. In contrast, single-arm designs exhibit nonvanishing regret because bias remains first order even as variance shrinks. The right panel focuses on the single-arm problem: both the oracle and the researcher can only choose a single arm. In this case, UCT is preferable only at small $n$; beyond the same threshold ($\approx 500$), CCT yields uniformly lower regret as variance becomes second order relative to bias. The regret vanishes relative to the oracle that can only choose a single arm.
These results illustrate how our approach systematically balances precision and misspecification risk to balance experimental variation with structural models.
6 Implications for practice
This paper studies experimental design in the presence of complementary observational evidence. Practitioners often have access to observational inputs—e.g., estimates from models trained on existing data (which may be misspecified), results from prior experiments that may not transport to the target context, or estimates from observational designs that may suffer from confounding. Our goal is to combine such evidence with new experimental data, which deliver unbiased estimators for specific parameters of interest. We ask which experiment to run, and how to run it, under budget constraints that limit the number of arms and/or impose per-unit (variable) costs on effective sample size.
A key challenge is that the bias of the observational estimators is unknown in practice. We adopt a minimax proportional regret criterion that compares the mean-squared error (MSE) of a candidate design to that of an oracle that knows the worst-case bias. This reveals a fundamental trade-off between precision and robustness. The optimal design balances the design’s variance normalized by the smallest achievable variance (variance gap) and its worst-case bias normalized by the smallest attainable bias (bias gap). We propose a procedure that jointly determines: (i) how to combine observational and experimental evidence; (ii) how to allocate precision across experiments given budget constraints; and (iii) which treatment arm and/or sub-population to include in the experiment with fixed experimental costs.
In practice, the workflow is:
  • Define the estimand(s) of interest. Specify $\tau(\theta)$ for a known mapping $\tau$ and unknown parameters $\theta\in\mathbb{R}^{p}$. For example, $\tau(\theta)$ may represent a counterfactual, a general-equilibrium effect, or an average impact across locations. Adopt a parametrization in which some (but not necessarily all) components of $\theta$ can be learned experimentally; this clarifies what the experiment can identify.
  • Assemble informative observational evidence. Collect observational estimates $\tilde{\theta}^{\rm obs}$ and their covariance $\tilde{\Sigma}^{\rm obs}$. These serve as informative (but potentially biased) baselines. Such evidence may come from observational designs, structural estimates with pre-experimental data, or prior experiments conducted in different contexts or periods.
  • Compute the sensitivity parameters. Using the observational baseline, compute the sensitivity weights $\omega\;=\;\frac{\partial\tau(\theta)}{\partial\theta}\big|_{\tilde{\theta}^{\rm obs}}$, which quantify how bias in each coordinate of $\theta$ propagates to $\tau(\theta)$. Large $|\omega_{j}|$ indicates greater payoff to learning the $j$th component.
  • Specify feasibility constraints and calibrate experimental variance. Enumerate the admissible design set $\mathcal{S}$. That is, $\mathcal{S}$ is the set containing combinations of parameters that can be learned jointly via experiments (which arms, sites, or mechanisms can be randomized). Define, for each combination of parameters $\mathcal{E}\in\mathcal{S}$, the corresponding precision set $\mathcal{G}(\mathcal{E})$ as the feasible set of experimental variances for a given experiment $\mathcal{E}$. This requires specifying how the sample size $n_{j}$ allocated to each arm maps to the variance of each experimental estimate. Calibrate per-unit experimental variances using pilot studies or historical data.
  • Run the method and decide. For each feasible set of experiments $\mathcal{E}\in\mathcal{S}$ and precisions $\Sigma\in\mathcal{G}(\mathcal{E})$, we first optimize over the shrinkage vector $\gamma$ (how to weight observational vs. experimental estimates). We then optimize over the precision/allocation $\Sigma$ (e.g., the sample size $n_{j}$ assigned to each arm). Finally, we optimize over the experiment choice $\mathcal{E}\in\mathcal{S}$. In practice, this involves enumerating feasible $\mathcal{E}$ and simple numerical optimization for $(\gamma,\Sigma)$. The output is a pre-analysis plan with the selected arm(s), sample sizes, and pre-specified combination rule ($\gamma$).
  • Reporting and diagnostics. In a pre-analysis plan, we recommend reporting (i) the estimands of interest; (ii) the chosen arm(s) and the final sample size allocation, (iii) the shrinkage weights $\gamma^{\star}$ by parameter between the observational and experimental study, and (iv) the two normalized components $\alpha/\alpha^{\star}$ (variance-gap) and $\beta/\beta^{\star}$ (bias-gap), that can help assess the quality of the allocation. Before committing to a single experiment choice, researchers may also want to explore sensitivity to a range of values for the experimental variances.
The applicability of our framework spans a wide range of settings in economics and beyond. Examples include estimating general equilibrium effects or structural models (Todd and Wolpin, 2006; Attanasio et al., 2012; Meghir et al., 2022; Kreindler et al., 2023; de Albuquerque et al., 2025); choosing among alternative treatment arms in factorial designs (Muralidharan et al., 2020; Bandiera et al., 2025); and deciding where to run the next experiment for external validity (Gechter et al., 2024; Olea et al., 2024). In industrial organization, applications include choices between demand-side and supply-side interventions (Bergquist and Dinerstein, 2020), the effects of information acquisition in markets (Allende et al., 2019; Larroucau et al., 2024), and decisions about which additional data source to acquire to improve statistical analysis. Beyond economics, medical applications include allocating sample size across subgroups and allocating doses across treatment arms (e.g., Porter et al., 2024; Morita et al., 2017; Manski, 2025). In all cases, the same recipe identifies which parameters to target, how much to learn about each, and how to blend experimental and observational evidence.
Several open questions remain for future research. These include settings where researchers have well-specified priors about bias (e.g., from meta-analysis), sequential or adaptive experimental choices, and environments where the exogenous source of variation yields only partial identification of some parameters of interest.
Appendix A Additional figures
Figure 12: Share of experimental papers published in AEA journals also presenting experimental results in combination with observational estimates (either from a reduced form, or from a structural model or from both).
Appendix B Proofs of main results
B.1 Proof of Theorem 1
Preliminary notation
It suffices to prove the theorem for any given feasible design and weights $(\mathcal{E},\Sigma(\mathcal{E}),\gamma)$. Throughout, we denote by $\mathcal{E}^{c}\equiv\{1,\dots,p\}\setminus\mathcal{E}$ the complement of $\mathcal{E}$.
Let $\tilde{\omega}(\mathcal{E},\gamma)=[\,\tilde{\omega}^{\mathrm{obs}}(\mathcal{E},\gamma)^{\top},\ \tilde{\omega}^{\mathrm{exp}}(\mathcal{E},\gamma)^{\top}\,]^{\top}$ as in (3), with
$\tilde{\omega}^{\mathrm{obs}}_{\mathcal{E}}(\mathcal{E},\gamma)=\omega_{\mathcal{E}}\circ(1-\gamma),\quad\tilde{\omega}^{\mathrm{obs}}_{\mathcal{E}^{c}}(\mathcal{E},\gamma)=\omega_{\mathcal{E}^{c}},\quad\tilde{\omega}^{\mathrm{exp}}(\mathcal{E},\gamma)=\omega_{\mathcal{E}}\circ\gamma,$
so that $\tilde{\omega}^{\mathrm{obs}}\in\mathbb{R}^{p}$ and $\tilde{\omega}^{\mathrm{exp}}\in\mathbb{R}^{|\mathcal{E}|}$. We let $\gamma\in\mathbb{R}^{|\mathcal{E}|}$ and denote $|1-\gamma|$ a vector of dimension $|\mathcal{E}|$ with entry $j$ equal to $|1-\gamma_{j}|$.
For an observational bias $b\in\mathbb{R}^{p}$ (i.e., $E[\tilde{\theta}^{\mathrm{obs}}]-\theta=b$), the mean–squared error is
$\mathrm{MSE}_{\mathcal{E},\Sigma,b}\!\left[\hat{\tau}_{\gamma}\right]~=~\tilde{\omega}(\mathcal{E},\gamma)^{\top}\Sigma(\mathcal{E})\,\tilde{\omega}(\mathcal{E},\gamma)\;+\;\bigl(\tilde{\omega}^{\mathrm{obs}}(\mathcal{E},\gamma)^{\top}b\bigr)^{2}.$
For an oracle that knows $\|b\|_{\infty}\leq\bar{B}$,
$R_{\bar{B}}(\mathcal{E},\Sigma,\gamma)~\equiv~\sup_{\|b\|_{\infty}\leq\bar{B}}\mathrm{MSE}_{\mathcal{E},\Sigma,b}\!\left[\hat{\tau}_{\gamma}\right]~=~\bar{B}^{2}\!\left(\|\omega_{\mathcal{E}^{c}}\|_{1}+|1-\gamma|_{\mathcal{E}}^{\top}|\omega_{\mathcal{E}}|\right)^{\!2}+\tilde{\omega}(\mathcal{E},\gamma)^{\top}\Sigma(\mathcal{E})\,\tilde{\omega}(\mathcal{E},\gamma),$
since $\sup_{\|b\|_{\infty}\leq\bar{B}}\tilde{\omega}^{\mathrm{obs}\,\top}b=\bar{B}\,\|\tilde{\omega}^{\mathrm{obs}}\|_{1}=\bar{B}\big(\|\omega_{\mathcal{E}^{c}}\|_{1}+|1-\gamma|_{\mathcal{E}}^{\top}|\omega_{\mathcal{E}}|\big)$.
B.1.1 The oracle problem
Given $(\mathcal{E},\Sigma)$, the oracle first chooses the shrinkage vector
$\gamma^{\star}_{o}(\bar{B},\mathcal{E},\Sigma)~\in~\arg\min_{\gamma\in\mathbb{R}^{|\mathcal{E}|}}R_{\bar{B}}(\mathcal{E},\Sigma,\gamma),$
and then (by backward induction) the design
$\big(\mathcal{E}^{\star}_{o}(\bar{B}),\ \Sigma^{\star}_{o}(\bar{B})\big)~\in~\arg\min_{\mathcal{E},\,\Sigma}R_{\bar{B}}\big(\mathcal{E},\Sigma,\gamma^{\star}_{o}(\bar{B},\mathcal{E},\Sigma)\big).$
Let $R^{\star}_{o}(\bar{B})\equiv\min_{\mathcal{E},\Sigma,\gamma}R_{\bar{B}}(\mathcal{E},\Sigma,\gamma)$ denote the oracle’s minimized worst–case MSE.
Lemma 1 (Convexity and strict convexity).
For fixed $(\mathcal{E},\Sigma)$ with $\Sigma(\mathcal{E})\succ 0$, the map $\gamma\mapsto R_{\bar{B}}(\mathcal{E},\Sigma,\gamma)$ is convex on $\gamma$. Moreover, if $|\omega_{j}|>0$ for all $j\in\mathcal{E}$, then $\gamma\mapsto R_{\bar{B}}(\mathcal{E},\Sigma,\gamma)$ is strictly convex on $\gamma$.
Proof.
Write $R_{\bar{B}}=\bar{B}^{2}g+h$ with
$g(\gamma)=\Big(\|\omega_{\mathcal{E}^{c}}\|_{1}+|1-\gamma|^{\top}|\omega_{\mathcal{E}}|\Big)^{2},\qquad h(\gamma)=\tilde{\omega}(\mathcal{E},\gamma)^{\top}\Sigma(\mathcal{E})\,\tilde{\omega}(\mathcal{E},\gamma).$
(i) Convexity of $g$. The map $\gamma\mapsto|1-\gamma|^{\top}|\omega_{\mathcal{E}}|$ is convex (sum of absolute values of affine functions). Adding the constant $\|\omega_{\mathcal{E}^{c}}\|_{1}$ preserves convexity, and squaring preserves convexity.
(ii) Strict convexity of $h$. Let $D\equiv\mathrm{diag}(\omega_{\mathcal{E}})$. Without loss, reorder the entries of $\theta$ so that we have $\theta=(\theta_{\mathcal{E}^{c}},\theta_{\mathcal{E}})$. Then
$\tilde{\omega}(\mathcal{E},\gamma)=\underbrace{\begin{bmatrix}\omega_{\mathcal{E}^{c}}\\ D\mathbf{1}\\ 0\end{bmatrix}}_{=:d}\;+\;\underbrace{\begin{bmatrix}0\\ -D\\ D\end{bmatrix}}_{=:A}\gamma,\quad\text{so}\quad h(\gamma)=(A\gamma+d)^{\top}\Sigma(\mathcal{E})\,(A\gamma+d).$
Hence $h(\gamma)=\gamma^{\top}(A^{\top}\Sigma A)\gamma+2\gamma^{\top}A^{\top}\Sigma d+d^{\top}\Sigma d$, and the Hessian is
$\nabla^{2}h(\gamma)=2\,A^{\top}\Sigma(\mathcal{E})\,A.$
Because $\Sigma(\mathcal{E})\succ 0$ and, when $|\omega_{j}|>0$ for all $j\in\mathcal{E}$, the columns of A=[ 0;D;D]A=[\,0;\,-D;\,D\,] are linearly independent, we have for any nonzero $u$,
$u^{\top}A^{\top}\Sigma A\,u=(Au)^{\top}\Sigma(Au)>0.$
Thus $A^{\top}\Sigma A\succ 0$ and $h$ is strictly convex (or weakly convex if $\omega_{j}=0$ for some $j$).
(iii) Conclusion. A sum of a convex function ($\bar{B}^{2}g$) and a strictly convex function ($h$) is strictly convex; when some $\omega_{j}=0$, the sum remains convex. ∎
Next, we show that as the worst-case bias becomes large enough, the oracle will put zero weight on the observational estimators.
Lemma 2.
Fix $(\mathcal{E},\Sigma)$ with $\Sigma(\mathcal{E})\succ 0$ and write $A\equiv\|\omega_{\mathcal{E}^{c}}\|_{1}$, with $|\omega_{j}|>0$ for all $j$. Let
$R_{\bar{B}}(\gamma)=\bar{B}^{2}\Big(A+t(\gamma)\Big)^{2}+\tilde{\omega}(\mathcal{E},\gamma)^{\top}\Sigma(\mathcal{E})\,\tilde{\omega}(\mathcal{E},\gamma),\qquad t(\gamma)\equiv|1-\gamma|^{\top}|\omega_{\mathcal{E}}|,$
for $\gamma\in\mathbb{R}^{|\mathcal{E}|}$. Then:
(i) If $A>0$ (equivalently, $\|\omega_{\mathcal{E}}\|_{1}<\|\omega\|_{1}$), there exists
$\bar{B}^{\max}(\mathcal{E},\Sigma,\omega)\;\leq\;\sqrt{\,2\,\|\Sigma(\mathcal{E})\|_{\infty}\cdot\max\!\Big\{2,\ \frac{\|\omega\|_{1}}{A}\Big\}\,}$
such that every minimizer satisfies $\gamma^{\star}_{o}(\bar{B},\mathcal{E},\Sigma)=\mathbf{1}$ for all $\bar{B}\geq\bar{B}^{\max}$.
(ii) If $A=0$ (equivalently, $\|\omega_{\mathcal{E}}\|_{1}=\|\omega\|_{1}$), then for any sequence $\bar{B}\to\infty$ every sequence of minimizers obeys $\gamma^{\star}_{o}(\bar{B},\mathcal{E},\Sigma)\to\mathbf{1}$.
Proof.
Let $x(\gamma)\equiv\tilde{\omega}(\mathcal{E},\gamma)$ and $x(1)\equiv\tilde{\omega}(\mathcal{E},\mathbf{1})$. Using $a^{\top}\Sigma a-b^{\top}\Sigma b=(a-b)^{\top}\Sigma(a+b)$ and the Holder’s bound $|u^{\top}\Sigma v|\leq\|\Sigma\|_{\infty}\|u\|_{1}\|v\|_{1}$, we get
$\big|x(\gamma)^{\top}\Sigma x(\gamma)-x(1)^{\top}\Sigma x(1)\big|\leq\|\Sigma\|_{\infty}\,\|x(\gamma)-x(1)\|_{1}\,\|x(\gamma)+x(1)\|_{1}.$
In addition,
$x(\gamma)-x(1)=\begin{bmatrix}\omega_{\mathcal{E}^{c}}-\omega_{\mathcal{E}^{c}}\\[1.0pt] \omega_{\mathcal{E}}\circ(1-\gamma)\\[1.0pt] -\,\omega_{\mathcal{E}}\!\circ\!(1-\gamma)\end{bmatrix}\Rightarrow\|x(\gamma)-x(1)\|_{1}=2\,t(\gamma).$
Also,
$x(\gamma)+x(1)=\begin{bmatrix}2\,\omega_{\mathcal{E}^{c}}\\[1.0pt] \omega_{\mathcal{E}}\!\circ\!(1-\gamma)\\[1.0pt] \omega_{\mathcal{E}}\!\circ\!(1+\gamma)\end{bmatrix}\Rightarrow\|x(\gamma)+x(1)\|_{1}\leq 2\|\omega_{\mathcal{E}^{c}}\|_{1}+t(\gamma)+\sum_{j\in\mathcal{E}}|\omega_{j}|\,|1+\gamma_{j}|.$
Using the triangular inequality for reals $u,v$, $|u|=|u-v+v|\leq|u-v|+|v|$, so that $|1+\gamma_{j}|\leq|1+\gamma_{j}-2|+2=|1-\gamma_{j}|+2$, we have $\sum_{j\in\mathcal{E}}|\omega_{j}||1+\gamma_{j}|\leq t(\gamma)+2\|\omega_{\mathcal{E}}\|_{1}.$ Therefore
$\|x(\gamma)+x(1)\|_{1}\leq 2\|\omega\|_{1}+2\,t(\gamma),$
and hence
$\big|x(\gamma)^{\top}\Sigma x(\gamma)-x(1)^{\top}\Sigma x(1)\big|\leq 4\,\|\Sigma\|_{\infty}\,t(\gamma)\,\big(\|\omega\|_{1}+t(\gamma)\big).$ (16)
(i) Case $A>0$. For any $\gamma$,
$R_{\bar{B}}(\gamma)-R_{\bar{B}}(\mathbf{1})$ $=\bar{B}^{2}\big[(A+t)^{2}-A^{2}\big]+\big[x(\gamma)^{\top}\Sigma x(\gamma)-x(1)^{\top}\Sigma x(1)\big]$
$\geq\bar{B}^{2}\big(2A\,t+t^{2}\big)-4\,\|\Sigma\|_{\infty}\,t\big(\|\omega\|_{1}+t\big)$
$=(\,\bar{B}^{2}-4\|\Sigma\|_{\infty}\,)t^{2}+(2A\bar{B}^{2}-4\|\Sigma\|_{\infty}\|\omega\|_{1}\,)t,$
where $t(\gamma)\geq 0$. If $\bar{B}^{2}\geq\max\{\,4\|\Sigma\|_{\infty},\ 2\,\|\Sigma\|_{\infty}\,\|\omega\|_{1}/A\,\},$ then both coefficients are nonnegative and the right-hand side is $\geq 0$, with equality iff $t=0$, i.e., $\gamma=\mathbf{1}$. Thus $\gamma^{\star}_{o}(\bar{B},\mathcal{E},\Sigma)=\mathbf{1}$ for all such $\bar{B}$.
(ii) Case $A=0$ (i.e., $\|\omega_{\mathcal{E}^{c}}\|_{1}=0$). Recall $t(\gamma)\equiv\sum_{j\in\mathcal{E}}|\omega_{j}|\,|1-\gamma_{j}|\geq 0$ and
$R_{\bar{B}}(\gamma)-R_{\bar{B}}(\mathbf{1})~\geq~\bar{B}^{2}\,t(\gamma)^{2}\;-\;4\,\|\Sigma(\mathcal{E})\|_{\infty}\,t(\gamma)\big(\|\omega\|_{1}+t(\gamma)\big),$
which follows from (16) with $A=0$. Let $\gamma^{\star}=\gamma^{\star}_{o}(\bar{B},\mathcal{E},\Sigma)$ be any minimizer, so $R_{\bar{B}}(\gamma^{\star})\leq R_{\bar{B}}(\mathbf{1})$ and hence
$0\;\geq\;R_{\bar{B}}(\gamma^{\star})-R_{\bar{B}}(\mathbf{1})\;\geq\;\bar{B}^{2}\,t_{\star}^{2}\;-\;4\,\|\Sigma(\mathcal{E})\|_{\infty}\,t_{\star}\big(\|\omega\|_{1}+t_{\star}\big),$
where $t_{\star}(\gamma^{\star})\geq 0$. Rearranging gives
$\big(\bar{B}^{2}-4\|\Sigma(\mathcal{E})\|_{\infty}\big)\,t_{\star}^{2}\;\leq\;4\,\|\Sigma(\mathcal{E})\|_{\infty}\,\|\omega\|_{1}\,t_{\star}.$
If $\bar{B}^{2}>4\|\Sigma(\mathcal{E})\|_{\infty}$, we can divide both sides by $B^{2}-4\|\Sigma(\mathcal{E})\|_{\infty}$ and obtain
$t_{\star}\;\leq\;\frac{\,4\,\|\Sigma(\mathcal{E})\|_{\infty}\,\|\omega\|_{1}\,}{\bar{B}^{2}-4\|\Sigma(\mathcal{E})\|_{\infty}}\;=\;O\!\left(\frac{1}{\bar{B}^{2}}\right).$
Hence $t(\gamma^{\star})\to 0$ as $\bar{B}\to\infty$.
Finally, since $t(\gamma^{\star})=\sum_{j\in\mathcal{E}}|\omega_{j}|\,|1-\gamma^{\star}_{j}|$ and each summand is nonnegative, for any coordinate with $|\omega_{j}|>0$ we have
$|1-\gamma^{\star}_{j}|\;\leq\;\frac{t(\gamma^{\star})}{|\omega_{j}|}\;\longrightarrow\;0\quad\text{as }\bar{B}\to\infty.$
Thus $\gamma^{\star}_{o}(\bar{B},\mathcal{E},\Sigma)\to\mathbf{1}$ as $\bar{B}\rightarrow\infty$. ∎
The lemma below characterizes the behavior of the implicitly-defined functions of $\bar{B}$: $\bar{B}\mapsto\beta\left(\mathcal{E},\gamma^{\star}\left(\bar{B},\mathcal{E},\Sigma\right)\right)$ and $\bar{B}\mapsto\alpha\left(\mathcal{E},\gamma^{\star}\left(\bar{B},\mathcal{E},\Sigma\right)\right)$.
Lemma 3.
Fix $(\mathcal{E},\Sigma)$ with $\Sigma(\mathcal{E})\succ 0$. Let
$\alpha(\mathcal{E},\Sigma,\gamma)\equiv\tilde{\omega}(\mathcal{E},\gamma)^{\top}\Sigma(\mathcal{E})\,\tilde{\omega}(\mathcal{E},\gamma),\qquad\beta(\mathcal{E},\gamma)\equiv\Big(\|\omega_{\mathcal{E}^{c}}\|_{1}+|1-\gamma|^{\top}|\omega_{\mathcal{E}}|\Big)^{2},$
and for $\bar{B}\geq 0$ define $R_{\bar{B}}(\gamma)\equiv\alpha(\mathcal{E},\Sigma,\gamma)+\bar{B}^{2}\,\beta(\mathcal{E},\gamma)$. Let $\gamma^{\star}(\bar{B})\in\arg\min_{\gamma}R_{\bar{B}}(\gamma)$. Then $\bar{B}\mapsto\beta\!\left(\mathcal{E},\gamma^{\star}(\bar{B})\right)$ is non-increasing and $\bar{B}\mapsto\alpha\!\left(\mathcal{E},\Sigma,\gamma^{\star}(\bar{B})\right)$ is non-decreasing.
Proof.
Fix $0\leq\bar{B}_{1}<\bar{B}_{2}$ and choose minimizers $\gamma_{1}\in\arg\min_{\gamma}R_{\bar{B}_{1}}(\gamma)$ and $\gamma_{2}\in\arg\min_{\gamma}R_{\bar{B}_{2}}(\gamma)$. By optimality,
$\alpha(\gamma_{1})+\bar{B}_{1}^{2}\,\beta(\gamma_{1})$ $\leq\alpha(\gamma_{2})+\bar{B}_{1}^{2}\,\beta(\gamma_{2}),$ (16)
$\alpha(\gamma_{2})+\bar{B}_{2}^{2}\,\beta(\gamma_{2})$ $\leq\alpha(\gamma_{1})+\bar{B}_{2}^{2}\,\beta(\gamma_{1}).$ (17)
Subtracting $\alpha(\gamma_{2})$ from (16) and $\alpha(\gamma_{1})$ from (17) yields
$\alpha(\gamma_{1})-\alpha(\gamma_{2})\;\leq\;\bar{B}_{1}^{2}\big(\beta(\gamma_{2})-\beta(\gamma_{1})\big),\qquad\alpha(\gamma_{1})-\alpha(\gamma_{2})\;\geq\;\bar{B}_{2}^{2}\big(\beta(\gamma_{2})-\beta(\gamma_{1})\big).$
Let $\Delta_{\beta}\equiv\beta(\gamma_{2})-\beta(\gamma_{1})$. Then
$\bar{B}_{2}^{2}\,\Delta_{\beta}\;\leq\;\alpha(\gamma_{1})-\alpha(\gamma_{2})\;\leq\;\bar{B}_{1}^{2}\,\Delta_{\beta}.$
If $\Delta_{\beta}>0$, the last display implies $\bar{B}_{2}^{2}\leq\bar{B}_{1}^{2}$, a contradiction. Hence $\Delta_{\beta}\leq 0$, i.e.
$\beta(\gamma_{2})\leq\beta(\gamma_{1}),$
so $\bar{B}\mapsto\beta(\gamma^{\star}(\bar{B}))$ is non-increasing. Plugging $\Delta_{\beta}\leq 0$ into the first inequality gives
$\alpha(\gamma_{1})-\alpha(\gamma_{2})\;\leq\;\bar{B}_{1}^{2}\,\Delta_{\beta}\;\leq\;0,$
so $\alpha(\gamma_{1})\leq\alpha(\gamma_{2})$, i.e. $\bar{B}\mapsto\alpha(\gamma^{\star}(\bar{B}))$ is non-decreasing. Because the choice of minimizer at each $\bar{B}$ was arbitrary, the monotonicity holds for any $\bar{B}$. ∎
Proposition 1.
Let $\Sigma(\mathcal{E})\succ 0$ with uniformly bounded entries and fix $\mathcal{E}$. Let $|\omega_{j}|>0$ for all $j$. Let
$\alpha(\mathcal{E},\Sigma,\gamma)\equiv\tilde{\omega}(\mathcal{E},\gamma)^{\top}\Sigma(\mathcal{E})\,\tilde{\omega}(\mathcal{E},\gamma),\qquad\beta(\mathcal{E},\gamma)\equiv\Big(\|\omega_{\mathcal{E}^{c}}\|_{1}+|1-\gamma|^{\top}|\omega_{\mathcal{E}}|\Big)^{2},$
and $\gamma^{\star}(\bar{B})\in\arg\min_{\gamma}\{\alpha(\mathcal{E},\Sigma,\gamma)+\bar{B}^{2}\beta(\mathcal{E},\gamma)\}$. Then, as $\bar{B}\to\infty$,
$\beta\big(\mathcal{E},\gamma^{\star}(\bar{B})\big)\,\downarrow\,\big(\|\omega\|_{1}-\|\omega_{\mathcal{E}}\|_{1}\big)^{2}\quad\text{and}\quad\alpha\big(\mathcal{E},\Sigma,\gamma^{\star}(\bar{B})\big)\,\uparrow\,\tilde{\omega}(\mathcal{E},\mathbf{1})^{\top}\Sigma(\mathcal{E})\,\tilde{\omega}(\mathcal{E},\mathbf{1}),$
both monotonically.
Proof.
Write $A\equiv\|\omega_{\mathcal{E}^{c}}\|_{1}$ and $t(\gamma)\equiv\sum_{j\in\mathcal{E}}|\omega_{j}|\,|1-\gamma_{j}|\geq 0$, so $\beta(\mathcal{E},\gamma)=(A+t(\gamma))^{2}$ and $\beta(\mathcal{E},\mathbf{1})=A^{2}=(\|\omega\|_{1}-\|\omega_{\mathcal{E}}\|_{1})^{2}$.
Step 1 (monotonicity). By Lemma 3, $\bar{B}\mapsto\beta(\mathcal{E},\gamma^{\star}(\bar{B}))$ is non-increasing and $\bar{B}\mapsto\alpha(\mathcal{E},\Sigma,\gamma^{\star}(\bar{B}))$ is non-decreasing.
Step 2 (limit of $\beta$). Since $t(\gamma)\geq 0$, we have $\beta(\mathcal{E},\gamma)\geq A^{2}$ for all $\gamma$, hence the non-increasing sequence $\beta(\mathcal{E},\gamma^{\star}(\bar{B}))$ is bounded below by $A^{2}$ and thus converges to some limit $\geq A^{2}$. By Lemma 2, $\gamma^{\star}(\bar{B})\to\mathbf{1}$ as $\bar{B}\to\infty$, hence $t(\gamma^{\star}(\bar{B}))\to 0$ and therefore
$\lim_{\bar{B}\to\infty}\beta\big(\mathcal{E},\gamma^{\star}(\bar{B})\big)=\lim_{\bar{B}\to\infty}\big(A+t(\gamma^{\star}(\bar{B}))\big)^{2}=A^{2}=(\|\omega\|_{1}-\|\omega_{\mathcal{E}}\|_{1})^{2}.$
By Step 1 the convergence is monotone (decreasing).
Step 3 (limit of $\alpha$). Feasibility of $\gamma=\mathbf{1}$ implies, for all $\bar{B}$,
$\alpha\big(\mathcal{E},\Sigma,\gamma^{\star}(\bar{B})\big)+\bar{B}^{2}\,\beta\big(\mathcal{E},\gamma^{\star}(\bar{B})\big)~\leq~\alpha\big(\mathcal{E},\Sigma,\mathbf{1}\big)+\bar{B}^{2}\,\beta\big(\mathcal{E},\mathbf{1}\big)=\alpha(\mathcal{E},\Sigma,\mathbf{1})+\bar{B}^{2}A^{2},$
and since $\beta(\mathcal{E},\gamma^{\star}(\bar{B}))\geq A^{2}$, this yields
$\alpha\big(\mathcal{E},\Sigma,\gamma^{\star}(\bar{B})\big)\ \leq\ \alpha\big(\mathcal{E},\Sigma,\mathbf{1}\big)\qquad\forall\,\bar{B}.$
Thus $\alpha(\mathcal{E},\Sigma,\gamma^{\star}(\bar{B}))$ is non-decreasing and bounded above by $\alpha(\mathcal{E},\Sigma,\mathbf{1})$, so it converges to some limit $\leq\alpha(\mathcal{E},\Sigma,\mathbf{1})$. To identify the limit, we use the bound in Lemma 2
$\big|\alpha(\mathcal{E},\Sigma,\gamma)-\alpha(\mathcal{E},\Sigma,\mathbf{1})\big|=\big|x(\gamma)^{\top}\Sigma x(\gamma)-x(\mathbf{1})^{\top}\Sigma x(\mathbf{1})\big|\leq 4\,\|\Sigma(\mathcal{E})\|_{\infty}\,t(\gamma)\big(\|\omega\|_{1}+t(\gamma)\big),$
where $x(\gamma)\equiv\tilde{\omega}(\mathcal{E},\gamma)$. Because $t(\gamma^{\star}(\bar{B}))\to 0$ from Step 2, the RHS tends to 0, so
$\lim_{\bar{B}\to\infty}\alpha\big(\mathcal{E},\Sigma,\gamma^{\star}(\bar{B})\big)=\alpha\big(\mathcal{E},\Sigma,\mathbf{1}\big)=\tilde{\omega}(\mathcal{E},\mathbf{1})^{\top}\Sigma(\mathcal{E})\,\tilde{\omega}(\mathcal{E},\mathbf{1}).$
By Step 1 the convergence is monotone (increasing). ∎
B.1.2 The researcher’s problem
To characterize the adaptation regret, it suffices to characterize
$\sup_{\bar{B}\geq 0}\,\sup_{\|b\|_{\infty}\leq\bar{B}}\frac{\text{MSE}_{b}\left[\,\widehat{\tau}\left(\mathcal{E},\gamma\right)\,\right]}{R^{\star}_{o}\left(\bar{B}\right)}=\sup_{\mathcal{E}^{\prime},\Sigma^{\prime}}\sup_{\bar{B}\geq 0}\frac{R_{\bar{B}}(\mathcal{E},\Sigma,\gamma)}{R_{\bar{B}}(\mathcal{E}^{\prime},\Sigma^{\prime},\gamma_{o}^{\star}(\bar{B},\mathcal{E}^{\prime},\Sigma^{\prime}))}.$
Write
$\rho(\gamma,\bar{B},\mathcal{E},\mathcal{E}^{\prime},\Sigma,\Sigma^{\prime})\equiv\frac{R_{\bar{B}}(\mathcal{E},\Sigma,\gamma)}{R_{\bar{B}}(\mathcal{E}^{\prime},\Sigma^{\prime},\gamma_{o}^{\star}(\bar{B},\mathcal{E}^{\prime},\Sigma^{\prime}))}.$
Lemma 4.
Let $\Sigma(\mathcal{E}),\Sigma^{\prime}(\mathcal{E}^{\prime})\succ 0$ have uniformly bounded entries and assume $|\omega_{j}|>0$ for all $j$. For any fixed $(\mathcal{E},\Sigma,\gamma)$ and $(\mathcal{E}^{\prime},\Sigma^{\prime})$, define
$\alpha\equiv\alpha(\mathcal{E},\Sigma,\gamma),\quad\beta\equiv\beta(\mathcal{E},\gamma),\qquad\alpha^{\star}(\bar{B})\equiv\alpha(\mathcal{E}^{\prime},\Sigma^{\prime},\gamma^{\star}_{o}(\bar{B},\mathcal{E}^{\prime},\Sigma^{\prime})),\quad\beta^{\star}(\bar{B})\equiv\beta(\mathcal{E}^{\prime},\gamma^{\star}_{o}(\bar{B},\mathcal{E}^{\prime},\Sigma^{\prime})).$
Let
$\rho(\bar{B})\;\equiv\;\frac{\bar{B}^{2}\,\beta+\alpha}{\bar{B}^{2}\,\beta^{\star}(\bar{B})+\alpha^{\star}(\bar{B})}.$
[…]
Proof.
By Proposition 1, along the oracle path $\bar{B}\mapsto\gamma^{\star}_{o}(\bar{B},\mathcal{E}^{\prime},\Sigma^{\prime})$ we have
$\beta^{\star}(\bar{B})\ \downarrow\ \big(\|\omega\|_{1}-\|\omega_{\mathcal{E}^{\prime}}\|_{1}\big)^{2},\qquad\alpha^{\star}(\bar{B})\ \uparrow\ \tilde{\omega}(\mathcal{E}^{\prime},\mathbf{1})^{\top}\Sigma^{\prime}(\mathcal{E}^{\prime})\,\tilde{\omega}(\mathcal{E}^{\prime},\mathbf{1}),$
hence the ratio
$\psi(\bar{B})\;\equiv\;\frac{\beta^{\star}(\bar{B})}{\alpha^{\star}(\bar{B})}$
is non-increasing in $\bar{B}$ (numerator $\downarrow$, denominator $\uparrow$, both nonnegative).
Fix $0\leq\bar{B}_{1}<\bar{B}_{2}$. Using optimality of the oracle at each $\bar{B}$,
$\alpha^{\star}(\bar{B}_{1})+\bar{B}_{1}^{2}\beta^{\star}(\bar{B}_{1})~\leq~\alpha^{\star}(\bar{B}_{2})+\bar{B}_{1}^{2}\beta^{\star}(\bar{B}_{2}),\qquad\alpha^{\star}(\bar{B}_{2})+\bar{B}_{2}^{2}\beta^{\star}(\bar{B}_{2})~\leq~\alpha^{\star}(\bar{B}_{1})+\bar{B}_{2}^{2}\beta^{\star}(\bar{B}_{1}).$
Therefore
$\rho(\bar{B}_{2})-\rho(\bar{B}_{1})$ $=\frac{(\bar{B}_{2}^{2}\beta+\alpha)\big(\bar{B}_{1}^{2}\beta^{\star}(\bar{B}_{1})+\alpha^{\star}(\bar{B}_{1})\big)-(\bar{B}_{1}^{2}\beta+\alpha)\big(\bar{B}_{2}^{2}\beta^{\star}(\bar{B}_{2})+\alpha^{\star}(\bar{B}_{2})\big)}{\big(\bar{B}_{2}^{2}\beta^{\star}(\bar{B}_{2})+\alpha^{\star}(\bar{B}_{2})\big)\big(\bar{B}_{1}^{2}\beta^{\star}(\bar{B}_{1})+\alpha^{\star}(\bar{B}_{1})\big)}$
$\leq\frac{(\bar{B}_{2}^{2}\beta+\alpha)\big(\bar{B}_{1}^{2}\beta^{\star}(\bar{B}_{2})+\alpha^{\star}(\bar{B}_{2})\big)-(\bar{B}_{1}^{2}\beta+\alpha)\big(\bar{B}_{2}^{2}\beta^{\star}(\bar{B}_{2})+\alpha^{\star}(\bar{B}_{2})\big)}{\big(\bar{B}_{2}^{2}\beta^{\star}(\bar{B}_{2})+\alpha^{\star}(\bar{B}_{2})\big)\big(\bar{B}_{1}^{2}\beta^{\star}(\bar{B}_{1})+\alpha^{\star}(\bar{B}_{1})\big)}$
$=\frac{(\bar{B}_{2}^{2}-\bar{B}_{1}^{2})\big[\alpha^{\star}(\bar{B}_{2})\,\beta-\beta^{\star}(\bar{B}_{2})\,\alpha\big]}{\big(\bar{B}_{2}^{2}\beta^{\star}(\bar{B}_{2})+\alpha^{\star}(\bar{B}_{2})\big)\big(\bar{B}_{1}^{2}\beta^{\star}(\bar{B}_{1})+\alpha^{\star}(\bar{B}_{1})\big)}$
$=\frac{(\bar{B}_{2}^{2}-\bar{B}_{1}^{2})\,\alpha^{\star}(\bar{B}_{2})\,\alpha}{\big(\bar{B}_{2}^{2}\beta^{\star}(\bar{B}_{2})+\alpha^{\star}(\bar{B}_{2})\big)\big(\bar{B}_{1}^{2}\beta^{\star}(\bar{B}_{1})+\alpha^{\star}(\bar{B}_{1})\big)}\left(\frac{\beta}{\alpha}-\frac{\beta^{\star}(\bar{B}_{2})}{\alpha^{\star}(\bar{B}_{2})}\right).$
Since all denominators are positive, the last display shows
$\rho(\bar{B}_{2})-\rho(\bar{B}_{1})\ \leq\ 0\quad\text{whenever}\quad\frac{\beta}{\alpha}\ \leq\ \psi(\bar{B}_{2}).$
A symmetric argument (now using $\alpha^{\star}(\bar{B}_{2})+\bar{B}_{2}^{2}\beta^{\star}(\bar{B}_{2})\leq\alpha^{\star}(\bar{B}_{1})+\bar{B}_{2}^{2}\beta^{\star}(\bar{B}_{1})$) yields
$\rho(\bar{B}_{2})-\rho(\bar{B}_{1})\ \geq\ 0\quad\text{whenever}\quad\frac{\beta}{\alpha}\ \geq\ \psi(\bar{B}_{1}).$
Define the threshold
$\tilde{B}\;\equiv\;\sup\Big\{B\geq 0:\ \psi(B)\geq\tfrac{\beta}{\alpha}\Big\}\ \in\ [0,\infty],$
[…]
Completion of the proof.
Fix $(\mathcal{E},\Sigma,\gamma)$ and write, for any $(\mathcal{E}^{\prime},\Sigma^{\prime})$,
$\rho(\bar{B})\equiv\rho(\gamma,\bar{B},\mathcal{E},\mathcal{E}^{\prime},\Sigma,\Sigma^{\prime})=\frac{\bar{B}^{2}\,\beta(\mathcal{E},\gamma)+\alpha(\mathcal{E},\Sigma,\gamma)}{\bar{B}^{2}\,\beta^{\star}(\bar{B};\mathcal{E}^{\prime},\Sigma^{\prime})+\alpha^{\star}(\bar{B};\mathcal{E}^{\prime},\Sigma^{\prime})},$
[…]
$\sup_{\bar{B}\geq 0}\rho(\bar{B})=\max\!\Big\{\rho(0),\ \lim_{\bar{B}\to\infty}\rho(\bar{B})\Big\}.$
Endpoint $\bar{B}=0$. Since $\rho(0)=\alpha(\mathcal{E},\Sigma,\gamma)\big/\alpha^{\star}(0;\mathcal{E}^{\prime},\Sigma^{\prime})$,
$\sup_{\mathcal{E}^{\prime},\Sigma^{\prime}}\rho(0)=\frac{\alpha(\mathcal{E},\Sigma,\gamma)}{\inf_{\mathcal{E}^{\prime},\Sigma^{\prime}}\alpha^{\star}(0;\mathcal{E}^{\prime},\Sigma^{\prime})}.$
[…]
$\sup_{\mathcal{E}^{\prime},\Sigma^{\prime}}\rho(0)=\frac{\tilde{\omega}(\mathcal{E},\gamma)^{\top}\Sigma(\mathcal{E})\,\tilde{\omega}(\mathcal{E},\gamma)}{\displaystyle\min_{\mathcal{E}^{\prime},\Sigma^{\prime}}\tilde{\omega}(\mathcal{E}^{\prime},\mathbf{0})^{\top}\Sigma^{\prime}(\mathcal{E}^{\prime})\,\tilde{\omega}(\mathcal{E}^{\prime},\mathbf{0})}.$
Endpoint $\bar{B}\to\infty$. By Proposition 1, $\beta^{\star}(\bar{B};\mathcal{E}^{\prime},\Sigma^{\prime})\downarrow\big(\|\omega\|_{1}-\|\omega_{\mathcal{E}^{\prime}}\|_{1}\big)^{2}$ and $\alpha^{\star}(\bar{B};\mathcal{E}^{\prime},\Sigma^{\prime})\uparrow\alpha(\mathcal{E}^{\prime},\Sigma^{\prime},\mathbf{1})$. Hence, after diving the numerators and denominator by $B^{2}$
$\lim_{\bar{B}\to\infty}\rho(\bar{B})=\frac{\beta(\mathcal{E},\gamma)}{\big(\|\omega\|_{1}-\|\omega_{\mathcal{E}^{\prime}}\|_{1}\big)^{2}}.$
Maximizing the ratio over $(\mathcal{E}^{\prime},\Sigma^{\prime})$ at this endpoint amounts to minimizing the denominator. This completes the proof. ∎
B.2 Two parameters examples: proof of Corollary 1
By Theorem 1, for fixed $(\mathcal{E},\Sigma)$ the adaptation regret as a function of the shrinkage vector equals
$\mathcal{R}(\gamma)\;=\;\max\!\Big\{\alpha(\mathcal{E},\Sigma,\gamma)/\alpha^{\star},\;\beta(\mathcal{E},\gamma)/\beta^{\star}\Big\},$
with $\alpha^{\star},\beta^{\star}$ constants.
Step 1. With independence and two parameters,
$\alpha(\mathcal{E},\Sigma,\gamma)=C+\omega_{j}^{2}\!\Big(\gamma_{j}^{2}v_{j}^{2}+(1-\gamma_{j})^{2}\sigma_{j}^{2}\Big),$
where $C$ does not depend on $\gamma_{j}$. Hence $\alpha(j,\gamma_{j})\equiv\alpha(\mathcal{E},\Sigma,\gamma)$ is a strictly convex quadratic in $\gamma_{j}$ with unique minimizer at
$\gamma_{j}^{\mathrm{var}}=\frac{\sigma_{j}^{2}}{\sigma_{j}^{2}+v_{j}^{2}}.$
Moreover, $\alpha(j,\gamma_{j})$ is strictly increasing on $\big(\gamma_{j}^{\mathrm{var}},1\big)$ and strictly decreasing on $\big[0,\gamma_{j}^{\mathrm{var}}\big)$.
Write the worst–case bias component holding the other coordinate fixed as
$\beta(j,\gamma_{j})\;=\;\Big(A_{j}+|\omega_{j}|\cdot|1-\gamma_{j}|\Big)^{2},$
where $A_{j}\geq 0$ collects all terms not involving $\gamma_{j}$ (including the contribution from the other coordinate). On $[0,1]$, $|1-\gamma_{j}|=1-\gamma_{j}$, so $\beta(j,\gamma_{j})$ is strictly decreasing in $\gamma_{j}$ and convex with minimum at $\gamma_{j}=1$. Therefore, any minimizer satisfies
$\gamma_{j}^{\star}\in\big[\gamma_{j}^{\mathrm{var}},\,1\big].$
Step 2. On $\big(\gamma_{j}^{\mathrm{var}},1\big)$ the map $\gamma_{j}\mapsto\alpha(j,\gamma_{j})/\alpha^{\star}$ is strictly increasing, while $\gamma_{j}\mapsto\beta(j,\gamma_{j})/\beta^{\star}$ is strictly decreasing. Hence the function
$f(\gamma_{j})\equiv\max\!\Big\{\alpha(j,\gamma_{j})/\alpha^{\star},\ \beta(j,\gamma_{j})/\beta^{\star}\Big\},\qquad\gamma_{j}\in\big[\gamma_{j}^{\mathrm{var}},1\big],$
is minimized either (i) at a boundary point, or (ii) at the unique interior point where the two arguments are equal (by strict monotonicity, at most one intersection exists). Therefore,
  • If $\alpha(j,\gamma_{j})/\alpha^{\star}<\beta(j,\gamma_{j})/\beta^{\star}$ for all $\gamma_{j}\in\big(\gamma_{j}^{\mathrm{var}},1\big)$, then $f(\gamma_{j})=\beta(j,\gamma_{j})/\beta^{\star}$ on that interval. Since this term is strictly decreasing, the minimizer is the right boundary $\gamma_{j}^{\star}=1$.
  • If $\alpha(j,\gamma_{j})/\alpha^{\star}>\beta(j,\gamma_{j})/\beta^{\star}$ for all $\gamma_{j}\in\big(\gamma_{j}^{\mathrm{var}},1\big)$, then $f(\gamma_{j})=\alpha(j,\gamma_{j})/\alpha^{\star}$ on that interval. Since this term is strictly increasing there, the minimizer is the left boundary $\gamma_{j}^{\star}=\gamma_{j}^{\mathrm{var}}=\sigma_{j}^{2}/(\sigma_{j}^{2}+v_{j}^{2})$.
  • Otherwise, by the intermediate value theorem and strict monotonicity of the two curves, there exists a unique $\gamma_{j}\in\big(\gamma_{j}^{\mathrm{var}},1\big)$ such that
    $\frac{\alpha(j,\gamma_{j})}{\alpha^{\star}}\;=\;\frac{\beta(j,\gamma_{j})}{\beta^{\star}}.$
Appendix C Proofs of the extensions
C.1 Proof of Theorem 2
The proof of Theorem 2 follows similarly the one of Theorem 1 with minor modifications.
Let
$\bar{\alpha}(\mathcal{E},\Sigma,\gamma)\equiv\sqrt{\tilde{\omega}(\mathcal{E},\gamma)^{\top}\Sigma(\mathcal{E})\,\tilde{\omega}(\mathcal{E},\gamma)},\qquad\bar{\beta}(\mathcal{E},\gamma)\equiv\|\tilde{\omega}(\mathcal{E},\gamma)\|_{1},$
so that
$|L_{B}(\mathcal{E},\Sigma,\gamma)|=2\{z_{1-\eta/2}\,\bar{\alpha}(\mathcal{E},\Sigma,\gamma)+B\,\bar{\beta}(\mathcal{E},\gamma)\}.$
For any comparison design $(\mathcal{E}^{\prime},\Sigma^{\prime})$, let the oracle choose
$\gamma^{\star}(B,\mathcal{E}^{\prime},\Sigma^{\prime})\in\arg\min_{\gamma^{\prime}\in\mathbb{R}^{|\mathcal{E}^{\prime}|}}\Big\{z_{1-\eta/2}\,\bar{\alpha}(\mathcal{E}^{\prime},\Sigma^{\prime},\gamma^{\prime})+B\,\bar{\beta}(\mathcal{E}^{\prime},\gamma^{\prime})\Big\},$
and denote the induced oracle path
$\bar{\alpha}^{\star}(B)\equiv\bar{\alpha}(\mathcal{E}^{\prime},\Sigma^{\prime},\gamma^{\star}(B,\mathcal{E}^{\prime},\Sigma^{\prime})),\qquad\bar{\beta}^{\star}(B)\equiv\bar{\beta}(\mathcal{E}^{\prime},\gamma^{\star}(B,\mathcal{E}^{\prime},\Sigma^{\prime})).$
Define the ratio (the factor $2$ cancels)
$\tilde{\rho}(B)\equiv\frac{z_{1-\eta/2}\,\bar{\alpha}(\mathcal{E},\Sigma,\gamma)+B\,\bar{\beta}(\mathcal{E},\gamma)}{z_{1-\eta/2}\,\bar{\alpha}^{\star}(B)+B\,\bar{\beta}^{\star}(B)}.$
Step 1: Endpoint reduction (CI analogue of Lemma 4).
Write $a\equiv z_{1-\eta/2}\,\bar{\alpha}(\mathcal{E},\Sigma,\gamma)$ and $b\equiv\bar{\beta}(\mathcal{E},\gamma)$, and $A(B)\equiv z_{1-\eta/2}\,\bar{\alpha}^{\star}(B)$, $B^{\star}(B)\equiv\bar{\beta}^{\star}(B)$. By Proposition 1 (applied with square roots), $\bar{\beta}^{\star}(B)$ is non-increasing in $B$, and $\bar{\alpha}^{\star}(B)$ is non-decreasing in $B$. Exactly as in Lemma 4 (replace $\alpha,\beta$ there with $\bar{\alpha},\bar{\beta}$ here, and use $x=B$ in place of $x=\bar{B}^{2}$), one verifies that
$\tilde{\rho}^{\prime}(B)\ \text{is non-decreasing in $B$}.$
Hence $\tilde{\rho}$ is first non-increasing and then non-decreasing, so its supremum is attained at the endpoints:
$\sup_{B\geq 0}\tilde{\rho}(B)=\max\Big\{\tilde{\rho}(0),\ \lim_{B\to\infty}\tilde{\rho}(B)\Big\}.$ (18)
Step 2: Evaluate the endpoints and optimize over comparison designs.
At $B=0$,
$\tilde{\rho}(0)=\frac{\bar{\alpha}(\mathcal{E},\Sigma,\gamma)}{\bar{\alpha}^{\star}(0)}.$
Since the oracle at $B=0$ minimizes standard error, $\bar{\alpha}^{\star}(0)=\min_{(\mathcal{E}^{\prime},\Sigma^{\prime},\gamma^{\prime})}\bar{\alpha}(\mathcal{E}^{\prime},\Sigma^{\prime},\gamma^{\prime})=\sqrt{\alpha^{\star}},$ where $\alpha^{\star}$ is as in (5). Thus
$\sup_{\mathcal{E}^{\prime},\Sigma^{\prime}}\tilde{\rho}(0)=\frac{\bar{\alpha}(\mathcal{E},\Sigma,\gamma)}{\sqrt{\alpha^{\star}}}=\left(\frac{\alpha(\mathcal{E},\Sigma,\gamma)}{\alpha^{\star}}\right)^{\!1/2}.$
As $B\to\infty$, the BB–term dominates. By Proposition 1 (again in square–root form), $\lim_{B\to\infty}\bar{\beta}^{\star}(B)=\min_{(\mathcal{E}^{\prime},\gamma^{\prime})}\bar{\beta}(\mathcal{E}^{\prime},\gamma^{\prime})=\sqrt{\beta^{\star}},$ with $\beta^{\star}$ from (5). Therefore
$\sup_{\mathcal{E}^{\prime},\Sigma^{\prime}}\ \lim_{B\to\infty}\tilde{\rho}(B)=\frac{\bar{\beta}(\mathcal{E},\gamma)}{\sqrt{\beta^{\star}}}=\left(\frac{\beta(\mathcal{E},\gamma)}{\beta^{\star}}\right)^{\!1/2}.$
Step 3: Combine.
Taking the supremum over comparison designs and using (18),
$\tilde{\mathcal{R}}(\mathcal{E},\Sigma,\gamma)=\sup_{B\geq 0}\ \sup_{(\mathcal{E}^{\prime},\Sigma^{\prime},\gamma^{\prime})\in\mathcal{D}}\frac{|L_{B}(\mathcal{E},\Sigma,\gamma)|}{|L_{B}(\mathcal{E}^{\prime},\Sigma^{\prime},\gamma^{\prime})|}=\max\!\left\{\left(\frac{\alpha(\mathcal{E},\Sigma,\gamma)}{\alpha^{\star}}\right)^{\!1/2},\ \left(\frac{\beta(\mathcal{E},\gamma)}{\beta^{\star}}\right)^{\!1/2}\right\},$
which is exactly the statement of Theorem 2. ∎
C.2 Proof of Theorem 3
Define the recursive solution to the oracle’s problem in (9):
$\gamma^{\star}_{o}\left(B,\mathcal{E},\Sigma\right)$ $=$ $\operatorname*{arg\,min}_{\gamma\in\mathbb{R}^{|\mathcal{E}|}}\sup_{b\in\mathcal{B}(B)}\ \mathrm{MSE}_{b}(\mathcal{E},\Sigma,\gamma)$
$\Sigma^{\star}_{o}\left(B,\mathcal{E}\right)$ $=$ $\operatorname*{arg\,min}_{\Sigma\in\mathcal{G}(\mathcal{E})}\sup_{b\in\mathcal{B}(B)}\ \mathrm{MSE}_{b}\left(\mathcal{E},\Sigma,\gamma^{\star}_{o}\left(B,\mathcal{E},\Sigma\right)\right)$
$\mathcal{E}^{\star}_{o}(B)$ $=$ $\operatorname*{arg\,min}_{\mathcal{E}\in\mathcal{S}}\sup_{b\in\mathcal{B}(B)}\ \mathrm{MSE}_{b}\left(\mathcal{E},\Sigma^{\star}_{o}\left(B,\mathcal{E}\right),\gamma^{\star}_{o}\left(B,\mathcal{E},\Sigma\right)\right),$
and let
$\check{R}_{B}\left(\mathcal{E},\Sigma,\gamma\right)\equiv\sup_{b\in\mathcal{B}(B)}\ \mathrm{MSE}_{b}(\mathcal{E},\Sigma,\gamma)=\check{\beta}(\mathcal{E},\gamma)B^{2}+\check{\alpha}(\mathcal{E},\Sigma,\gamma),$
with $\check{\beta},\check{\alpha}$ as defined in Equation (10). The proof mimics the proof of Theorem 1. We will refer to $\check{R}_{B}(\gamma)$ omitting its arguments $\mathcal{E},\Sigma$ whenever clear from the context.
Lemma 5 (Convexity and strict convexity).
Fix $(\mathcal{E},\Sigma)$ with $\Sigma(\mathcal{E})\succ 0$. Then the map
$\gamma~\longmapsto~\check{R}_{\bar{B}}(\mathcal{E},\Sigma,\gamma)~=~\check{\alpha}(\mathcal{E},\Sigma,\gamma)\;+\;\bar{B}^{2}\,\check{\beta}(\mathcal{E},\gamma)$
is convex in $\gamma$. Moreover, if for each $j\in\mathcal{E}$ there exists at least one $\ell\in\{1,\dots,L\}$ with $|\omega^{\ell}_{j}|>0$, then the map is strictly convex in $\gamma$.
Proof.
We show convexity of the two addends separately.
1) Convexity (and strict convexity) of $\check{\alpha}(\mathcal{E},\Sigma,\gamma)$.
For each $\ell$, define $D_{\ell}:=\mathrm{diag}(\omega^{\ell}_{\mathcal{E}})$ and
$d_{\ell}~:=~\begin{bmatrix}\omega_{\mathcal{E}^{c}}\\ \omega^{\ell}_{\mathcal{E}}\\ 0\end{bmatrix},\qquad A_{\ell}~:=~\begin{bmatrix}0\\ -D_{\ell}\\ D_{\ell}\end{bmatrix}\ \in\mathbb{R}^{(p+|\mathcal{E}|)\times|\mathcal{E}|}.$
Then, by the definition of $\tilde{\omega}^{\ell}(\mathcal{E},\gamma)$,
$\tilde{\omega}^{\ell}(\mathcal{E},\gamma)\;=\;d_{\ell}\;+\;A_{\ell}\,\gamma^{\ell},$
where $\gamma^{\ell}\in\mathbb{R}^{|\mathcal{E}|}$ is the row of $\gamma$ for target $\ell$. Using (10),
$\check{\alpha}(\mathcal{E},\Sigma,\gamma)=\mathrm{tr}\!\left(\tilde{\omega}\,\Sigma\,\tilde{\omega}^{\top}\right)=\sum_{\ell=1}^{L}(d_{\ell}+A_{\ell}\gamma^{\ell})^{\top}\Sigma(\mathcal{E})\,(d_{\ell}+A_{\ell}\gamma^{\ell}).$
Hence $\check{\alpha}$ is a (separable across $\ell$) quadratic form in $\gamma$ with Hessian
$\nabla^{2}_{\gamma}\check{\alpha}(\mathcal{E},\Sigma,\gamma)=2\,\mathrm{blkdiag}\!\big(A_{1}^{\top}\Sigma(\mathcal{E})A_{1},\ldots,A_{L}^{\top}\Sigma(\mathcal{E})A_{L}\big)\ \succeq\ 0,$
where $\mathrm{blkdiag}$ is a block-diagonal matrix with entries as above. It follows that $\check{\alpha}$ is convex.
For strict convexity: if $\Sigma(\mathcal{E})\succ 0$ and $A_{\ell}$ has full column rank for at least one $\ell$ in every coordinate direction, then $\sum_{\ell}(A_{\ell}^{\top}\Sigma A_{\ell})\succ 0$ and thus $\check{\alpha}$ is strictly convex. Since $A_{\ell}=[0;-D_{\ell};D_{\ell}]$, $A_{\ell}u=0$ iff $D_{\ell}u=0$. Thus $A_{\ell}$ has full column rank iff $D_{\ell}$ is invertible, i.e. iff $|\omega^{\ell}_{j}|>0$ for all $j\in\mathcal{E}$. A sufficient condition ensuring strict convexity is: for each $j\in\mathcal{E}$ there exists at least one $\ell$ with $|\omega^{\ell}_{j}|>0$; then, for any nonzero $u\in\mathbb{R}^{|\mathcal{E}|}$ we can pick an $\ell$ with $u_{j}\neq 0$ and $|\omega^{\ell}_{j}|>0$, whence
$u^{\top}\Big(\sum_{\ell}A_{\ell}^{\top}\Sigma A_{\ell}\Big)u=\sum_{\ell}(A_{\ell}u)^{\top}\Sigma(A_{\ell}u)\ \geq\ (A_{\ell}u)^{\top}\Sigma(A_{\ell}u)\ >\ 0.$
Therefore $\check{\alpha}$ is strictly convex under the stated condition.
2) Convexity of $\check{\beta}(\mathcal{E},\gamma)$.
From (10) and the definition of $\bar{v}(\gamma)$,
$\check{\beta}(\mathcal{E},\gamma)=\sup_{\|v\|_{\infty}\leq 1}\ \sum_{\ell=1}^{L}\Big((\omega^{\ell})^{\top}v-(\omega^{\ell}_{\mathcal{E}})^{\top}v_{\mathcal{E}}+(1-\gamma^{\ell})^{\top}(\,v_{\mathcal{E}}\circ\omega_{\mathcal{E}}\,)\Big)^{\!2}.$
For each fixed $v$, the inner expression is an affine function of $\gamma$:
$(\omega^{\ell})^{\top}v-(\omega^{\ell}_{\mathcal{E}})^{\top}v_{\mathcal{E}}+\underbrace{\mathbf{1}^{\top}(\,v_{\mathcal{E}}\circ\omega_{\mathcal{E}}\,)}_{\text{constant in }\gamma}\;-\;(\gamma^{\ell})^{\top}(\,v_{\mathcal{E}}\circ\omega_{\mathcal{E}}\,).$
The square of an affine function is convex; a finite sum of convex functions is convex; and the pointwise supremum of convex functions is convex. Therefore $\check{\beta}(\mathcal{E},\gamma)$ is convex in $\gamma$. This concludes the proof. ∎
Lemma 6.
Suppose $\Sigma(\mathcal{E})\succ 0$ with uniformly bounded entries and $|\omega^{\ell}_{j}|>0$ for all $j\in\{1,\dots,p\}$ and $\ell\in\{1,\dots,L\}$. Then, for fixed $(\mathcal{E},\Sigma)$,
$\gamma^{\star}_{o}(B,\mathcal{E},\Sigma)\ \longrightarrow\ \mathbf{1}\qquad\text{as }B\to\infty,$
where $\mathbf{1}$ denotes the $L\times|\mathcal{E}|$ matrix of ones.
Proof.
Recall that for fixed $(\mathcal{E},\Sigma)$ the oracle criterion is
$\check{R}_{B}(\mathcal{E},\Sigma,\gamma)=\sup_{\|b\|_{\infty}\leq B}\mathrm{MSE}_{b}(\mathcal{E},\Sigma,\gamma)=\check{\alpha}(\mathcal{E},\Sigma,\gamma)+B^{2}\,\check{\beta}(\mathcal{E},\gamma),$
with (see (10))
$\check{\alpha}(\mathcal{E},\Sigma,\gamma)=\mathrm{tr}\!\big(\tilde{\omega}(\mathcal{E},\gamma)\,\Sigma(\mathcal{E})\,\tilde{\omega}(\mathcal{E},\gamma)^{\top}\big),\qquad\check{\beta}(\mathcal{E},\gamma)=\sup_{\|v\|_{\infty}\leq 1}\ \sum_{\ell=1}^{L}\big(a_{\ell}(v)+\delta_{\ell}(\gamma,v)\big)^{2},$
where
$a_{\ell}(v):=\sum_{j\in\mathcal{E}^{c}}\omega^{\ell}_{j}v_{j},\qquad\delta_{\ell}(\gamma,v):=\sum_{j\in\mathcal{E}}(1-\gamma^{\ell}_{j})\,\omega^{\ell}_{j}v_{j}.$
Step 1: The bias term is uniquely minimized at $\gamma=\mathbf{1}$.
For each $j\in\mathcal{E}$, define the $L$-vector
$d_{j}(\gamma):=\big((1-\gamma_{j}^{1})\omega_{j}^{1},\ldots,(1-\gamma_{j}^{L})\omega_{j}^{L}\big)^{\top}\in\mathbb{R}^{L},\qquad S(\gamma):=\max_{j\in\mathcal{E}}\|d_{j}(\gamma)\|_{2}.$
Choose $v\in[-1,1]^{p}$ with $v_{\mathcal{E}^{c}}=0$, $v_{j}=1$ for some fixed $j\in\mathcal{E}$, and $v_{j^{\prime}}=0$ for $j^{\prime}\in\mathcal{E}\setminus\{j\}$. Then
$\sum_{\ell=1}^{L}\big(a_{\ell}(v)+\delta_{\ell}(\gamma,v)\big)^{2}=\sum_{\ell=1}^{L}\big(\delta_{\ell}(\gamma,v)\big)^{2}=\|d_{j}(\gamma)\|_{2}^{2}.$
Taking the supremum over $v$ yields
$\check{\beta}(\mathcal{E},\gamma)\ \geq\ \max_{j\in\mathcal{E}}\|d_{j}(\gamma)\|_{2}^{2}\;=\;S(\gamma)^{2}.$ (19)
In particular, $\check{\beta}(\mathcal{E},\gamma)>\check{\beta}(\mathcal{E},\mathbf{1})$ whenever $\gamma\neq\mathbf{1}$ (since then $S(\gamma)>0$ by $|\omega^{\ell}_{j}|>0$). Thus $\gamma\mapsto\check{\beta}$ is uniquely minimized at $\gamma=\mathbf{1}$, with a quadratic gap at least $S(\gamma)^{2}$.
Step 2: Variance bound.
[…]
$\big|\check{\alpha}(\mathcal{E},\Sigma,\gamma)-\check{\alpha}(\mathcal{E},\Sigma,\mathbf{1})\big|\leq\|\Sigma(\mathcal{E})\|_{\infty}\sum_{\ell=1}^{L}\|x_{\ell}(\gamma)-y_{\ell}\|_{1}\,\|x_{\ell}(\gamma)+y_{\ell}\|_{1}.$
A direct calculation gives
$\|x_{\ell}(\gamma)-y_{\ell}\|_{1}=2\sum_{j\in\mathcal{E}}|(1-\gamma^{\ell}_{j})\omega^{\ell}_{j}|=2\,t_{\ell}(\gamma),$
and, using $|1+\gamma^{\ell}_{j}|\leq|1-\gamma^{\ell}_{j}|+2$,
$\|x_{\ell}(\gamma)+y_{\ell}\|_{1}\leq 2\|\omega^{\ell}\|_{1}+2\,t_{\ell}(\gamma).$
Hence
$\big|\check{\alpha}(\mathcal{E},\Sigma,\gamma)-\check{\alpha}(\mathcal{E},\Sigma,\mathbf{1})\big|\leq 4\,\|\Sigma(\mathcal{E})\|_{\infty}\!\left(\sum_{\ell=1}^{L}t_{\ell}(\gamma)\,\|\omega^{\ell}\|_{1}+\sum_{\ell=1}^{L}t_{\ell}(\gamma)^{2}\right).$
Let $d(\gamma)\in\mathbb{R}^{L\times|\mathcal{E}|}$ stack the $d_{j}(\gamma)$ as columns; then $t_{\ell}(\gamma)=\|d(\gamma)_{\ell,\cdot}\|_{1}$ and
$\sum_{\ell=1}^{L}t_{\ell}(\gamma)=\sum_{j\in\mathcal{E}}\|d_{j}(\gamma)\|_{1}\leq|\mathcal{E}|\,\sqrt{L}\,S(\gamma),\qquad\sum_{\ell=1}^{L}t_{\ell}(\gamma)^{2}\leq\Big(\sum_{\ell=1}^{L}t_{\ell}(\gamma)\Big)^{2}\leq|\mathcal{E}|^{2}L\,S(\gamma)^{2}.$
Therefore there exist finite constants $C_{1},C_{2}$ (depending only on $\Sigma,\omega,L,|\mathcal{E}|$) such that
$\big|\check{\alpha}(\mathcal{E},\Sigma,\gamma)-\check{\alpha}(\mathcal{E},\Sigma,\mathbf{1})\big|\ \leq\ C_{1}\,S(\gamma)\;+\;C_{2}\,S(\gamma)^{2}.$ (20)
Step 3: convergence of the minimizer.
Combine (19)–(20) to get, for any $\gamma$,
$\check{R}_{B}(\mathcal{E},\Sigma,\gamma)-\check{R}_{B}(\mathcal{E},\Sigma,\mathbf{1})=\big[\check{\alpha}(\gamma)-\check{\alpha}(\mathbf{1})\big]+B^{2}\big[\check{\beta}(\gamma)-\check{\beta}(\mathbf{1})\big]\ \geq\ -\,C_{1}S(\gamma)+\big(B^{2}-C_{2}\big)S(\gamma)^{2}.$
Fix $\varepsilon>0$ and choose $B$ such that $(B^{2}-C_{2})\varepsilon^{2}>C_{1}\varepsilon$. Then the RHS is >0>0 for all $\gamma$ with $S(\gamma)\geq\varepsilon$. Hence any minimizer $\gamma^{\star}_{o}(B)$ must satisfy $S(\gamma^{\star}_{o}(B))<\varepsilon$. Because $\varepsilon>0$ is arbitrary and $B\to\infty$, we obtain $S(\gamma^{\star}_{o}(B))\to 0$.
Finally, $S(\gamma^{\star}_{o}(B))\to 0$ means $(1-\gamma^{\star,\ell}_{j}(B))\omega^{\ell}_{j}\to 0$ for every $j\in\mathcal{E}$ and $\ell\in\{1,\dots,L\}$. Since $|\omega^{\ell}_{j}|>0$, this implies $\gamma^{\star,\ell}_{j}(B)\to 1$ for all $j,\ell$, i.e. $\gamma^{\star}_{o}(B,\mathcal{E},\Sigma)\to\mathbf{1}$ componentwise. ∎
Lemma 7.
Fix $(\mathcal{E},\Sigma)$ with $\Sigma(\mathcal{E})\succ 0$. Let
$\check{\alpha}(\mathcal{E},\Sigma,\gamma)=\mathrm{tr}\!\big(\tilde{\omega}(\mathcal{E},\gamma)\,\Sigma(\mathcal{E})\,\tilde{\omega}(\mathcal{E},\gamma)^{\top}\big),\qquad\check{\beta}(\mathcal{E},\gamma)=\sup_{\|v\|_{\infty}\leq 1}\ \sum_{\ell=1}^{L}\big(a_{\ell}(v)+\delta_{\ell}(\gamma,v)\big)^{2},$
where
$a_{\ell}(v):=\sum_{j\in\mathcal{E}^{c}}\omega^{\ell}_{j}v_{j},\qquad\delta_{\ell}(\gamma,v):=\sum_{j\in\mathcal{E}}(1-\gamma^{\ell}_{j})\,\omega^{\ell}_{j}v_{j},$
and for $\bar{B}\geq 0$ define $R_{\bar{B}}(\gamma)\equiv\alpha(\mathcal{E},\Sigma,\gamma)+\bar{B}^{2}\,\beta(\mathcal{E},\gamma)$. Let $\gamma^{\star}(\bar{B})\in\arg\min_{\gamma}R_{\bar{B}}(\gamma)$. Then $\bar{B}\mapsto\beta\!\left(\mathcal{E},\gamma^{\star}(\bar{B})\right)$ is non-increasing and $\bar{B}\mapsto\alpha\!\left(\mathcal{E},\Sigma,\gamma^{\star}(\bar{B})\right)$ is non-decreasing.
Proof.
The proof is identical to that of Lemma 3, after replacing $\alpha$ and $\beta$ with $\check{\alpha}$ and $\check{\beta}$, respectively. ∎
Proposition 2.
Let $\Sigma(\mathcal{E})\succ 0$ with uniformly bounded entries and fix $\mathcal{E}$. Let $|\omega_{j}|>0$ for all $j$. Define $\check{\alpha}(\mathcal{E},\Sigma,\gamma)$ and $\check{\beta}(\mathcal{E},\gamma)$ as in Lemma. Then, as $\bar{B}\to\infty$,
$\check{\beta}\big(\mathcal{E},\gamma^{\star}(\bar{B})\big)\,\downarrow\,\big(\|\omega\|_{1}-\|\omega_{\mathcal{E}}\|_{1}\big)^{2}\quad\text{and}\quad\check{\alpha}\big(\mathcal{E},\Sigma,\gamma^{\star}(\bar{B})\big)\,\uparrow\,\mathrm{tr}\!\big(\tilde{\omega}(\mathcal{E},\mathbf{1})\,\Sigma(\mathcal{E})\,\tilde{\omega}(\mathcal{E},\mathbf{1})^{\top}\big),$
both monotonically.
Proof.
Write $A\equiv\|\omega_{\mathcal{E}^{c}}\|_{1}$ and, following the notation in Lemma 7, define $g(\mathcal{E},\gamma,v)=\sum_{\ell=1}^{L}\big(a_{\ell}(v)+\delta_{\ell}(\gamma,v)\big)^{2}$, so $\check{\beta}(\mathcal{E},\gamma)=\sup_{\|v\|_{\infty}\leq 1}g(\mathcal{E},\gamma,v)$.
Step 1 (monotonicity). By Lemma 7, $\bar{B}\mapsto\check{\beta}(\mathcal{E},\gamma^{\star}(\bar{B}))$ is non-increasing and $\bar{B}\mapsto\check{\alpha}(\mathcal{E},\Sigma,\gamma^{\star}(\bar{B}))$ is non-decreasing.
Step 2 (limit of $\beta$). Letting $\mathbf{1}_{\mathcal{E}}$ denote the $p\times 1$ vector with $j^{th}$ entry $1\{j\notin\mathcal{E}\}$, we have $\check{\beta}(\mathcal{E},\gamma^{\star}(\bar{B}))\geq g(\mathcal{E},\gamma^{\star}(\bar{B}),\mathbf{1}_{\mathcal{E}})=A^{2}$, hence the non-increasing sequence $\check{\beta}(\mathcal{E},\gamma^{\star}(\bar{B}))$ is bounded below by $A^{2}$ and thus converges to some limit $\geq A^{2}$.
Because $v\mapsto g(\mathcal{E},\gamma,v)$ is convex, $\check{\beta}(\mathcal{E},\gamma)=\max_{v\in V}g(\mathcal{E},\gamma,v)$, where $V=\{-1,1\}^{p}$. Additionally, Lemma 6 implies $\lim_{\bar{B}\rightarrow\infty}g(\mathcal{E},\gamma^{\star}(\bar{B}),v)=g(\mathcal{E},\mathbf{1},v)$ $\forall v\in V$, since $\gamma\mapsto g(\mathcal{E},\gamma,v)$ is continuous. Therefore, since $V$ is finite,
$|\max_{v\in V}g(\mathcal{E},\gamma^{\star}(\bar{B}),v)-\max_{v\in V}g(\mathcal{E},\mathbf{1},v)|\leq\max_{v\in V}|g(\mathcal{E},\gamma^{\star}(\bar{B}),v)-g(\mathcal{E},\mathbf{1},v)|\rightarrow 0\text{ as $\bar{B}\rightarrow\infty$.}$
We conclude that $\lim_{\bar{B}\rightarrow\infty}\check{\beta}\big(\mathcal{E},\gamma^{\star}(\bar{B})\big)=\max_{v\in V}g(\mathcal{E},\mathbf{1},v)=A^{2}$. By Step 1, the convergence is monotone (decreasing).
Step 3 (limit of $\alpha$). Feasibility of $\gamma=\mathbf{1}$ implies, for all $\bar{B}$,
$\check{\alpha}\big(\mathcal{E},\Sigma,\gamma^{\star}(\bar{B})\big)+\bar{B}^{2}\,\check{\beta}\big(\mathcal{E},\gamma^{\star}(\bar{B})\big)~\leq~\check{\alpha}\big(\mathcal{E},\Sigma,\mathbf{1}\big)+\bar{B}^{2}\,\check{\beta}\big(\mathcal{E},\mathbf{1}\big)=\check{\alpha}(\mathcal{E},\Sigma,\mathbf{1})+\bar{B}^{2}A^{2},$
and since $\check{\beta}(\mathcal{E},\gamma^{\star}(\bar{B}))\geq A^{2}$, this yields
$\check{\alpha}\big(\mathcal{E},\Sigma,\gamma^{\star}(\bar{B})\big)\ \leq\ \check{\alpha}\big(\mathcal{E},\Sigma,\mathbf{1}\big)\qquad\forall\,\bar{B}.$
Thus $\check{\alpha}(\mathcal{E},\Sigma,\gamma^{\star}(\bar{B}))$ is non-decreasing and bounded above by $\check{\alpha}(\mathcal{E},\Sigma,\mathbf{1})$, so it converges to some limit $\leq\check{\alpha}(\mathcal{E},\Sigma,\mathbf{1})$. To identify the limit, we use the bound in Lemma 6
$\big|\check{\alpha}(\mathcal{E},\Sigma,\gamma^{\star}(\bar{B}))-\check{\alpha}(\mathcal{E},\Sigma,\mathbf{1})\big|\leq C_{1}\,S(\gamma^{\star}(\bar{B}))+C_{2}\,S(\gamma^{\star}(\bar{B}))^{2}\rightarrow C_{1}\,S(\mathbf{1})+C_{2}\,S(\mathbf{1})^{2}=0,$
as $\bar{B}\rightarrow\infty$. Therefore,
$\lim_{\bar{B}\to\infty}\check{\alpha}\big(\mathcal{E},\Sigma,\gamma^{\star}(\bar{B})\big)=\check{\alpha}\big(\mathcal{E},\Sigma,\mathbf{1}\big)=\mathrm{tr}\!\big(\tilde{\omega}(\mathcal{E},\mathbf{1})\,\Sigma(\mathcal{E})\,\tilde{\omega}(\mathcal{E},\mathbf{1})^{\top}\big).$
By Step 1, the convergence is monotone (increasing). ∎
Lemma 8.
Let $\Sigma(\mathcal{E}),\Sigma^{\prime}(\mathcal{E}^{\prime})\succ 0$ have uniformly bounded entries and assume $|\omega_{j}^{\ell}|>0$ for all $j,\ell$. For any fixed $(\mathcal{E},\Sigma,\gamma)$ and $(\mathcal{E}^{\prime},\Sigma^{\prime})$, define
$\check{\alpha}\equiv\check{\alpha}(\mathcal{E},\Sigma,\gamma),\quad\check{\beta}\equiv\check{\beta}(\mathcal{E},\gamma),\qquad\check{\alpha}^{\star}(\bar{B})\equiv\check{\alpha}(\mathcal{E}^{\prime},\Sigma^{\prime},\gamma^{\star}_{o}(\bar{B},\mathcal{E}^{\prime},\Sigma^{\prime})),\quad\check{\beta}^{\star}(\bar{B})\equiv\check{\beta}(\mathcal{E}^{\prime},\gamma^{\star}_{o}(\bar{B},\mathcal{E}^{\prime},\Sigma^{\prime})).$
Let
$\check{\rho}(\bar{B})\;\equiv\;\frac{\bar{B}^{2}\,\check{\beta}+\check{\alpha}}{\bar{B}^{2}\,\check{\beta}^{\star}(\bar{B})+\check{\alpha}^{\star}(\bar{B})}.$
[…]
Proof.
By Proposition 2, along the oracle path $\bar{B}\mapsto\gamma^{\star}_{o}(\bar{B},\mathcal{E}^{\prime},\Sigma^{\prime})$ we have
$\check{\beta}^{\star}(\bar{B})\ \downarrow\ \big(\|\omega\|_{1}-\|\omega_{\mathcal{E}^{\prime}}\|_{1}\big)^{2},\qquad\check{\alpha}^{\star}(\bar{B})\ \uparrow\ \mathrm{tr}\!\big(\tilde{\omega}(\mathcal{E}^{\prime},\mathbf{1})\,\Sigma(\mathcal{E}^{\prime})\,\tilde{\omega}(\mathcal{E}^{\prime},\mathbf{1})^{\top}\big),$
hence the ratio
$\check{\psi}(\bar{B})\;\equiv\;\frac{\check{\beta}^{\star}(\bar{B})}{\check{\alpha}^{\star}(\bar{B})}$
is non-increasing in $\bar{B}$ (numerator $\downarrow$, denominator $\uparrow$, both nonnegative).
The rest of the proof is identical to that of Lemma 4, after replacing $\alpha$, $\beta$, $\alpha^{\star}$, $\beta^{\star}$, $\rho$, and $\psi$ with $\check{\alpha}$, $\check{\beta}$, $\check{\alpha}^{\star}$, $\check{\beta}^{\star}$, $\check{\rho}$, and $\check{\psi}$, respectively. ∎
The completion of the proof is a analogous to that of B.1, after replacing $\alpha$, $\beta$, $\alpha^{\star}$, $\beta^{\star}$, and $\rho$ with $\check{\alpha}$, $\check{\beta}$, $\check{\alpha}^{\star}$, $\check{\beta}^{\star}$, and $\check{\rho}$, respectively. ∎