### v2

Abstract
Experiments deliver credible treatment effects estimates, but are often localized to specific sites, populations, or mechanisms. When such estimates are insufficient to extrapolate effects for broader policy questions as for external validity and general-equilibrium (GE) effects, researchers combine trials with external evidence from reduced-form or structural observational estimates, or prior experiments. We develop a unified framework for designing experiments in this setting: the researcher selects which parameters (or moments) to identify experimentally from a feasible set of possible experiments, allocates sample size, and specifies how to weight experimental and observational estimators. Because observational inputs may be biased in ways unknown ex ante, we adopt a minimax proportional-regret objective that evaluates any candidate design relative to an oracle that knows the bias and jointly chooses the design and estimator. This yields a transparent bias-variance trade-off that requires no prespecified bias bound and depends only on information about the precision of the estimators and the estimand’s sensitivity to underlying parameters. We illustrate the framework by (i) designing small-scale cash-transfer experiments aimed at estimating GE effects and (ii) optimizing site selection for microfinance interventions.
1 Introduction
Randomized controlled trials (RCTs) have improved empirical economics by providing internally valid causal estimates. However, feasibility constraints often confine trials to localized effects such as the effect in a specific site or subpopulations, or of a particular mechanism. These effects, although useful, are often not sufficient to answer broader questions about external validity and (general) equilibrium effects of at-scale interventions. For instance, the median evaluation of experiments in top economics journals is “representative of a population of 10,885 units, studied a treatment delivered to 5,340 units, and randomized in clusters of 26 units per cluster” (Muralidharan and Niehaus, 2017).
In response, a growing literature in development economics (e.g., Meghir et al., 2022; Bassi et al., 2022; Attanasio et al., 2012; de Albuquerque et al., 2025), education (e.g., Allende et al., 2019; Larroucau et al., 2024), labor economics (e.g., Chetty et al., 2016), and industrial organization (e.g., Katz and Allcott, 2025; Allcott et al., 2025) combine experiments with external evidence—observational estimates and/or results from experiments in other contexts—to estimate complex counterfactuals that no single experiment can identify. In our review of AEA journals over the past decade, more than 30% of experimental papers complement RCT evidence with at least one observational input, either reduced-form or structural (Figure 12). This trend poses a design question: given constraints on feasible experiments, which experiments should be run (and how) to estimate such counterfactuals?
Our main contribution is a framework for designing experiments combined with external evidence. By external evidence we mean reduced-form or structural estimates based on observational data, and/or results from experiments conducted in other contexts. We provide a procedure that selects which experiments to run subject to a budget (e.g., which treatment arms and/or sub-populations to include), sets their precision via sample allocation, and prescribes how to combine experimental and external evidence. External input may be biased in ways not known ex ante (e.g., due to confounding or lack of external validity).
For an illustrative example, consider a government piloting a cash-transfer program in a small set of districts to increase children’s school attendance. The trial delivers a local average effect, but policy decisions often require counterfactuals at scale, allowing prices and wages to adjust (e.g. Todd and Wolpin, 2006; Egger et al., 2022). Researchers therefore combine experimental evidence with a supply-demand model that leverages observational data to map local impacts into economy-wide outcomes. The design question is twofold: (i) which experiment to run (i.e., which parameters/effects are the most valuable to learn experimentally) given feasibility constraints and (ii) how to allocate sample across sites and arms so that, once combined with external inputs, we precisely estimate the effect at scale.
The first step is to specify the estimand and the underlying parameters. Define $\tau(\theta)$ the (policy-relevant) object of interest, where $\tau$ is a known smooth function and $\theta$ is a vector of unknown parameters; $\tau$ may be scalar or vector-valued. Researchers also observe external baseline estimates of $\theta$—henceforth observational estimates—which may be biased. Each feasible experiment can identify a subset of parameters without bias. This parameterization makes explicit which components of $\theta$ are learned experimentally and which may rely on biased observational evidence; not all components need (or can) be learned experimentally.
Returning to the cash-transfer example, $\tau(\theta)$ is the effect on schooling when all poor households in rural Kenya receive the subsidy. The parameter vector $\theta$ bundles (i) the direct schooling response to a conditional cash transfer (CCT), (ii) the income effect of transfers in partial equilibrium , and (iii) wage adjustments that shift the returns to schooling. Due to cost constraints, researchers may only run small, partial-equilibrium experiments and then combine these with prior evidence to recover $\tau(\theta)$. They may choose between a CCT treatment arm to estimate the direct effect, or an unconditional-cash arm (UCT) to measure the income elasticity, and extrapolate the GE effects using evidence from a previous study in Mexico (Todd and Wolpin, 2006). (When site selection is also part of the design, $\theta$ additionally collects site-specific effects and $\tau(\theta)$ targets the average effect across sites.)
The second step is to design the experiment: (i) selecting which experiment to run; (ii) allocating sample sizes across arms; and (iii) choosing how to combine observational and experimental evidence. We allow flexibility in (iii): researchers may choose to either optimally combine both sources when available or rely solely on experimental estimates. Parameters not identified experimentally are based on observational estimates.
A natural starting point is to design the experiment to minimize the mean-squared error (MSE) for estimating $\tau(\theta)$, perhaps worst-case over the observational bias. In practice, however, the size and even the direction of bias are unknown (before the experiment). A worst-case approach would therefore be overly-conservative and disregard information about the variance. We instead adapt the definition of proportional (adaptation) regret previously studied when combining two estimators generated by a fixed design (Armstrong et al., 2024; Tsybakov, 1998) and generalize it to the experimental design problem (with multiple estimators). Here, the regret is the ratio of the worst-case researcher’s mean-squared error relative to an oracle that knows the worst-case bound on the observational bias and chooses both the design and the estimator. Taking its supremum over any value of the bias yields a procedure that is robust to any level of misspecification without requiring priors about the bias.
Our main result gives an explicit novel characterization of proportional regret. For any design and weights that combine the experimental and observational point estimates, the regret is the maximum of two normalized components. The first is a variance regret: the estimator’s sampling variance under the chosen design and weights, divided by the smallest attainable variance. The second is a bias regret: the worst-case squared bias induced by using external evidence, divided by the smallest feasible worst-case bias over the class of designs. The variance regret depends on the (expected) variance-covariance matrix of the observational and experimental estimates implied by the design, the shrinkage weights, and the sensitivity of the policy parameter, i.e., the gradient of $\tau$ with respect to the parameters. This gradient is evaluated, under mild conditions, at the observational estimates. [^4] For this result to hold, we assume that $\tau(\theta)$ is twice differentiable with bounded derivatives. For non-linear $\tau(\theta)$ in $\theta$, we also require that the observational estimates bias is local to zero in the spirit of Andrews et al. (2020); Armstrong and Kolesár (2021); Bonhomme and Weidner (2022), where its rate of convergence can be the same, faster, or even slower than the estimators’ standard error. Section 5.1 provides details. The bias regret depends only on the sensitivity vector and the shrinkage weights, known ex-ante.
This characterization makes the bias-variance trade-off transparent. At the optimum, the two normalized components tend to be equalized: when the bias component dominates, the design prioritizes high-sensitivity parameters; when variance dominates, the procedure invests sample where it most reduces variance and relies more on the most precise observational inputs. This yields a nested procedure that chooses jointly the estimator, the sample allocation and the experiment to run, trading off precision with sensitivity of the policy estimand. The resulting design and estimator only require knowledge of the expected variance-covariance matrix (standard in experimental design problems, Gerber and Green, 2012) and of the observational estimates. It can then be reported in a pre-analysis plan. That is, the inputs for the design are the same that would be required by standard power-analysis plans.
From a theoretical perspective, our result leverages a key insight. Because here bias arise from multiple parameters, the choice of the uncertainty set is pivotal. While we do not require researchers to specify such a set, regret is with respect to an oracle that knows the worst-case bias within an $\ell_{\infty}$-ball. Compared with $\ell_{1}$ or $\ell_{2}$, the $\ell_{\infty}$-ball prevents that a large bias in one component is “offset” by smaller biases elsewhere. This is attractive when biases may be positively correlated across parameters. We show that the $\ell_{\infty}$ choice combined with our design problem (with decisions made before experimental data is realized), yields a quasi-convex regret in the bias radius as the oracle optimizes over the designs. This result, worst-case over the radius, delivers a transparent and easy-to-compute variance-bias trade-off.
We then extend the model to general method-of-moments estimators that combine experimentally identified moments with observational moments. As in the baseline case, observational moments may incur unknown misspecification. We optimize (i) which moments to identify experimentally and (ii) how to combine experimental and observational moments to estimate the counterfactual of interest. Our main result continues to hold in this more general scenario provided that the moment estimators are (asymptotically) linear in the moments. Linearity occurs (under smoothness conditions) as long as the ratio of the bias to standard error is no faster than $n^{1/4}$, with $n$ denoting the observational sample size; standard local-asymptotics with the bias and standard error of the same order is a special case.
We illustrate the framework in the cash‐transfer application, when the researcher uses evidence from Mexico’s PROGRESA program (Todd and Wolpin, 2006; Attanasio et al., 2012). We are concerned that these preliminary estimates may lack external validity in Kenya. We compare three designs: (i) a CCT arm to estimate the direct schooling effect; (ii) an unconditional cash transfer (UCT) arm to estimate income effects; and (iii) a two-arm design that allocates sample across CCT and UCT under a common budget. Running a single arm captures settings with fixed costs per arm; allowing both arms with budgeted allocation captures cases with non-binding fixed costs but binding variable costs. When both arms are feasible, the regret-optimal allocation puts most participants in the CCT and a small fraction in the UCT, reflecting the target’s greater sensitivity to the direct effect despite the CCT’s higher variance. If only one arm can be run, the choice hinges on sample size: for very small $n$, the lower-variance of the UCT arm makes this preferred over a (too noisy) CCT arm; once $n\geq 500$, the more informative but noisier CCT yields lower regret. Relative to an oracle that knows the bias, the optimal design closely tracks the oracle. We provides a step-by-step guide for practitioners applicable in other contexts.
As a second application, we design where to run a microfinance experiment in rural India, integrating evidence from earlier nonrandomized microfinance introductions. We start from the observational estimates in Banerjee et al. (2024) and design an experiment that selects one or more areas for randomization. Because Banerjee et al. (2024) also implement a separate randomized expansion, we can calibrate performance of our design (that is agnostic of the bias) by comparing each candidate design’s MSE to that of an oracle that knows the bias, using the experimental estimates to calibrate the bias. The optimal design concentrates sample in high-sensitivity, high-variance areas and then spreads as the budget grows; compared to a benchmark that randomly chooses areas and splits sample equally, it reduces MSE by roughly 20% under the calibrated bias.
Related literature
This paper links the experimental-design literature—which mostly focuses on settings where all parameters are identified within the experiment and leaves aside questions of data combination—to recent work that integrates experimental and (reduced-form or structural) observational evidence to extrapolate effects in complex scenarios.
Recent advances for experimental design include balancing and variance-minimizing allocations (Tabord-Meehan, 2018; Bai, 2019; Cytrynbaum, 2021; Kallus, 2018; Bertsimas et al., 2015), adaptive designs for policy choice (Kasy and Sautmann, 2019; Russo et al., 2018; Cesa-Bianchi et al., 2025), and experimental design under correct model specification (Silvey, 2013; Chaudhuri and Mykland, 1993; Chaloner and Verdinelli, 1995; Higbee, 2024; Kiefer and Wolfowitz, 1959; Reeves et al., 2024; Viviano, 2020; Kasy, 2016). Traditional work on experimental design for robust-model estimation either focused on testing competing models (Atkinson and Fedorov, 1975; López-Fidalgo et al., 2007), or on using a-priori knowledge of (worst-case) bias for the design of an experiment (Box and Draper, 1959; Wiens, 1998; Tsirpitzi et al., 2023; Sacks and Ylvisaker, 1984). All these references leave aside questions about observational data combination. In the context of minimizing the variance of an experiment, Rosenman and Owen (2021) proposes to use observational studies to construct high-confidence bounds on the experimental variance, and then focus on variance-optimal stratification. This problem differs from our design problem and analysis, which instead studies the use of observational studies in combination with experiments for extrapolation.
In summary, we complement this literature by studying which experiment to run (and how) when not all parameters relevant for the target estimand can be learned experimentally.
Our minimax regret criterion connects to a long-standing decision-theoretic tradition for experimental design. References include Manski and Tetenov (2016), Banerjee et al. (2020), Manski (2004), Dominitz and F. Manski (2017), Olea et al. (2024), Hu et al. (2024), Breza et al. (2025) among others. These references focus on settings where researchers have only access to experimental variation, instead of combining it with observational evidence, motivating different designs and objective functions. In the context of site-selection, we complement literature that leverages correctly-specified models (Gechter et al., 2024; Abadie and Zhao, 2021) by allowing for misspecification in observational estimates.
A related line of work analyzes estimation under misspecification (Armstrong and Kolesár, 2018; Armstrong et al., 2024; Andrews et al., 2025; Andrews et al., 2020; Bonhomme and Weidner, 2022; Christensen and Connault, 2023; James et al., 1961), and combining existing experiments with observational studies (Gechter, 2022; Athey et al., 2025, 2020; Kallus et al., 2018; Dutz et al., 2021; Rosenman et al., 2022; Bhattacharya, 2013; de Chaisemartin and D’Haultfœuille, 2020; Rambachan et al., 2024). We adapt the worst-case proportional regret idea from Armstrong et al. (2024) and Tsybakov (1998) here studied in the context of experimental design. Our definition of sensitivity of the estimand to each parameter builds on sensitivity analysis in Andrews et al. (2020). Unlike both strands of this literature where the design is fixed and researchers optimize over the choice of the estimator only, here we optimize the design itself. The objective and analysis change: regret is computed against the best design-estimator combination, and the resulting estimator is learned jointly with the design.
2 Problem description
Consider a researcher interested in an arbitrary target estimand $\tau(\theta)\in\mathbb{R}$, indexed by a low-dimensional unknown parameter vector $\theta\in\mathbb{R}^{p}$ and a known mapping $\tau$. For simplicity, we assume that the estimand is univariate, while Section 5.3 shows how our framework directly extends to multivariate $\tau(\theta)\in\mathbb{R}^{q}$ for $q>1$.
The goal is to construct an estimator $\hat{\tau}$ that accurately approximates $\tau(\theta)$ by combining existing evidence (e.g., observational studies) with experimental variation designed by the researcher. Our question is how to design such experiments under feasibility constraints.
Basic setup
We assume that researchers have access to estimators (and their variance) of $\theta$ denoted as $\tilde{\theta}^{\mathrm{obs}}\in\mathbb{R}^{p}$. We impose no restrictions on how $\tilde{\theta}^{\mathrm{obs}}$ is formed: it can based on arbitrary exclusion restrictions implied by an economic or statistical model. However, $\tilde{\theta}^{\mathrm{obs}}$, defined as observational estimates, have unknown biases collected in a vector $b\in\mathbb{R}^{p}$. Examples include an estimate from an instrumental variable regression that may fail the exclusion restriction, or an estimate from a country different from the one of interest that may lack external validity. Researchers can also collect experimental evidence for a subset of parameters, producing estimates whose bias equals zero by design.
Setting 1.
For a subset $\mathcal{E}\subseteq\{1,\ldots,p\}$ and a known positive-definite matrix $\Sigma(\mathcal{E})\in\mathbb{R}^{(p+|\mathcal{E}|)\times(p+|\mathcal{E}|)}$, define $\tilde{\theta}^{\mathrm{obs}}\in\mathbb{R}^{p}$ an observational estimate and $\tilde{\theta}^{\mathrm{exp}}_{\mathcal{E},\Sigma}\in\mathbb{R}^{|\mathcal{E}|}$ an experimental estimate each satisfying
$\mathbb{E}\!\left[\tilde{\theta}^{\mathrm{obs}}\right]-\theta=b,\qquad\mathbb{E}\!\left[\tilde{\theta}^{\mathrm{exp}}_{\mathcal{E},\Sigma}\right]-\theta_{\mathcal{E}}=0,$
for an unknown bias vector $b\in\mathbb{R}^{p}$. Moreover, define its joint variance-covariance matrix as $\mathbb{V}\!\begin{pmatrix}\tilde{\theta}^{\mathrm{obs}}-\theta\\[2.0pt] \tilde{\theta}^{\mathrm{exp}}_{\mathcal{E},\Sigma}-\theta_{\mathcal{E}}\end{pmatrix}=\Sigma(\mathcal{E}).$ Given $(\mathcal{E},\Sigma)$, consider a class of linear plug-in estimators
$\hat{\tau}_{\gamma}\equiv\tau(\hat{\theta}(\gamma)),\qquad\hat{\theta}_{j}(\gamma)=\begin{cases}\gamma_{j}\,\tilde{\theta}_{j}^{\mathrm{exp}}+(1-\gamma_{j})\,\tilde{\theta}_{j}^{\mathrm{obs}},&j\in\mathcal{E},\\[2.0pt] \tilde{\theta}_{j}^{\mathrm{obs}},&j\notin\mathcal{E},\end{cases}$
where $\gamma=(\gamma_{j})_{j\in\mathcal{E}}\in\mathbb{R}^{|\mathcal{E}|}$ are shrinkage weights that can be an (implicit) function of $(\mathcal{E},\Sigma)$.
Here, $(\mathcal{E},\Sigma(\mathcal{E}))$ characterizes the experimental design (which parameters are learned and with what precision). When clear, we omit the subscript $\Sigma$ on $\tilde{\theta}^{\mathrm{exp}}$. We assume $\Sigma(\mathcal{E})$ is known, which is standard in experimental planning (Gerber and Green, 2012) (in practice, any consistent estimator, e.g., from pilot studies would suffice).
The class of experiments and estimators is assumed to be in a user-specific set $\mathcal{D}$.
Assumption 1 (Feasible experiments and estimators).
We write compactly $(\mathcal{E},\Sigma,\gamma)\in\mathcal{D}$ indicating $\mathcal{D}$ the set of feasible design (i.e., choice of $\mathcal{E}$ and $\Sigma(\mathcal{E})$) and estimators (i.e., $\gamma$). The set $\mathcal{D}$ can be arbitrary as long as (i) each feasible $\Sigma$ has uniformly bounded entries and is strictly positive definite; (ii) $\gamma_{j}=1$ for all $j\in\mathcal{E}$ is one admissible choice (among potentially other values).
We let $\mathcal{E},\Sigma(\mathcal{E})$ be within arbitrary constraint sets that may encode feasibility or budget constraints. In our framework, restrictions on the experiments researchers can run $\mathcal{E}$ and on their precision $\Sigma(\mathcal{E})$ can take any desired form. In most applications, we think of such constraints as arising from fixed costs of running an experiment and/or constraints on their power (Duflo et al., 2007; Athey and Imbens, 2017; List et al., 2011). We assume that $\Sigma(\mathcal{E})$ is uniformly bounded, which implies that once we commit to learn a set of experimental parameters $\tilde{\theta}_{\mathcal{E}}^{\mathrm{exp}}$, their variance is finite, and strictly positive definite, therefore assuming that the variance is bounded away from zero. [^5] For this latter condition to hold in an asymptotic framework with growing sample size, $\theta$ and $\tilde{\theta}$ can be defined without loss as parameters of interest after appropriately rescaling by the square-root of the sample size; in this case $\Sigma$ denotes the asymptotic variance. See Section 5.1 for details. The shrinkage estimator can be unrestricted (i.e., $\gamma$ can take any value on the real line) or restricted (when for example researchers aim to only use experimental evidence when available, setting $\gamma_{j}=1$ for $j\in\mathcal{E}$).
We extend our framework to more general methods of moments estimators in Section 4.
Remark 1 (Structure of $\Sigma$).
Because observational and experimental moments are typically computed from disjoint samples, in practice, it is natural to model the covariance as block diagonal, $\Sigma=\operatorname{diag}(\Sigma_{\text{obs}},\Sigma_{\text{exp}})$. The first block corresponds to observational estimates (fixed and not subject to design choices), while the second corresponds to experimental estimates, and it is a function of the sample size allocated to treatment arms. This is a special case of our framework which can be encoded in the constraint set $\mathcal{D}$. ∎
Research questions
Our design problem can be described in three steps:
  1. 1.
    Preliminary step: weights. For each feasible $\mathcal{E},\Sigma(\mathcal{E})$, choose $\gamma$ to combine experimental and observational estimates on the selected coordinates. (If researchers restrict their estimators to only use experimental evidence when available, set $\gamma=1$.)
  2. 2.
    Middle step: precision. For each feasible $\mathcal{E}$, choose $\Sigma(\mathcal{E})$ (i.e., sample allocation).
  3. 3.
    Outer step: experiment choice. Choose a feasible set of experiments $\mathcal{E}$.
Choosing $\gamma$ (step 1) for a fixed design follows a long-standing tradition in econometrics and statistics (Armstrong et al., 2024; Andrews and Shapiro, 2021; Donoho, 1994; Athey et al., 2025). We study this question here in combination with the experimental design problem, step 2 and 3; this, as we show, will change the underlying optimization problem (also for $\gamma$). The central challenge is that performance depends on the unknown bias $b$.
Mapping $\tau$
A key feature of our framework is that researchers explicitly parameterize $\theta$ and $\tau(\cdot)$, thereby pre-specifying and making transparent which biases drive the design problem; we see this as an advantage because it clarifies the role of the experiment in complementing existing evidence. The sensitivity of $\tau$ to $\theta$, captured by its gradient $\omega(\theta)=\partial\tau(\theta)/\partial\theta$, determines how the (first-order) bias propagates onto the final estimand. For simplicity, we impose exact linearity, $\tau(\theta)=\omega^{\top}\theta$; all results extend to smooth, nonlinear $\tau$ via a first-order expansion, as discussed below.
Assumption 2 (First-order estimation error).
For any $(\mathcal{E},\Sigma,\gamma)\in\mathcal{D}$, assume
$\tau(\theta)-\tau\!\big(\hat{\theta}(\gamma)\big)=\sum_{j=1}^{p}\omega_{j}\,\big(\theta_{j}-\hat{\theta}_{j}(\gamma)\big),$ (1)
for known weights $\omega\in\mathbb{R}^{p}$, with $|\omega_{j}|\in(0,\infty)$ for all $j$.
Assumption 2 holds exactly when $\tau(\cdot)$ is linear and serves as a first-order approximation for smooth non-linear $\tau(\cdot)$. Section 5.1 (and Example 3) extend the framework to non-linear estimands encompassing as special cases the local-misspecification setups in Andrews et al. (2020); Armstrong and Kolesár (2021). For non-linear $\tau$, we replace $\omega$ with the gradient of $\tau$ evaluated at a preliminary observational estimate, i.e., $\omega\;=\;\frac{\partial\tau(\theta)}{\partial\theta}\Big|_{\theta=\tilde{\theta}^{\mathrm{obs}}}\;+\;o_{p}(1),$ which is typically available before the experiment is conducted. That is, building on references above, $\omega$ captures the first-order effect of the observational bias $b$ on the bias of $\hat{\tau}$. [^6] To see how this assumption holds under a first-order approximation, define $\theta$ the parameter rescaled by $\sqrt{n}$ (and similarly its bias $b_{n}$, indexed by $n$), where $n^{-1/2}$ is the rate of convergence of the observational estimators. Then second-order effects are negligible (and Assumption 2 holds asymptotically) provided that $b_{n}^{2}/\sqrt{n}\rightarrow 0$. This implies that Assumption 2 holds when the asymptotic bias $b_{n}$ is proportional to, smaller or even much larger than the asymptotic variance, provided that $b_{n}$ does not grow at rate faster than $n^{1/4}$.
We conclude with three examples in simple two-parameter models.
Example 1 (Choosing the site for an experiment for external validity).
Gechter et al. (2024) study where to run an experiment. For illustration, consider two sites $j\in\{1,2\}$ with site-specific ATEs $\theta_{j}$ and target the cross-site average
$\tau(\theta)=\tfrac{1}{2}(\theta_{1}+\theta_{2})=\omega^{\top}\theta,\qquad\omega=\tfrac{1}{2}(1,1)^{\top}.$
Let $\tilde{\theta}^{\mathrm{obs}}=(\tilde{\theta}^{\mathrm{obs}}_{1},\tilde{\theta}^{\mathrm{obs}}_{2})^{\top}$ denote observational estimates obtained in Gechter et al. (2024) from a structural model and potentially biased due to misspecification: $\mathbb{E}[\tilde{\theta}^{\mathrm{obs}}]-\theta=(b_{1},b_{2})^{\top}.$ A budget constraint allows an experiment in only one site, so that $\mathcal{E}\in\{\{1\},\{2\}\}$. If site $j$ is chosen, we obtain an unbiased $\tilde{\theta}^{\mathrm{exp}}_{j}$; the other site $k\neq j$ remains observational. Given $\mathcal{E}$,
$\widehat{\tau}_{\gamma}=\tfrac{1}{2}\!\left[\gamma_{j}\,\tilde{\theta}^{\mathrm{exp}}_{j}+(1-\gamma_{j})\,\tilde{\theta}^{\mathrm{obs}}_{j}+\tilde{\theta}^{\mathrm{obs}}_{k}\right],\qquad k\neq j.$
Our question is how to choose the set $\{j\}$ where to conduct the experiment (jointly with $\gamma$).
Example 2 (Choosing which survey to conduct).
Egger et al. (2022) study the efficacy of cash-transfer programs on the marginal propensity to consume (MPC). Measuring the MPC requires capturing both short- and long-run effects. Because survey rounds are limited, the authors complement experimental data that lacks short-run effects with auxiliary information from prior studies that use short-run surveys (collected in other regions; see Egger et al. (2022)). This raises the question of which survey to conduct (and with which frequency).
In stylized form, suppose researchers can observe, for $s\in\{1,2\}$, potential outcomes $Y_{s}(t)$ denoting consumption in period $s$ when measured $t$ periods after the intervention. The authors have auxiliary estimates from previous studies,
$\alpha\;=\;\mathbb{E}\!\big[Y_{1}(t=1)-Y_{1}(t=\infty)\big],\qquad\beta\;=\;\mathbb{E}\!\big[Y_{2}(t=1)-Y_{2}(t=\infty)\big],$
and wish to estimate the total effect $\tau=\alpha+\beta$. Researchers consider two survey designs:
  • (i)
    Early survey design to estimate $\alpha$ precisely;
  • (ii)
    Later survey to estimate $\beta$ precisely.
Our goal is to study which survey design to implement. (More complex designs with additional time periods or mixed precision across rounds are also possible here.)
Example 3 (Supply or demand experiment with non linear target).
Bergquist and Dinerstein (2020) conduct demand and supply experiments in food markets in Kenya. Suppose here we are interested in similar applications in Uganda. For exposition, consider a basic linear demand and supply
$Q^{D}=a-\beta_{D}P+u_{D},\qquad Q^{S}=c+\beta_{S}P+u_{S},\qquad\theta=(\beta_{D},\beta_{S})^{\top},$
with $\beta_{D}\neq-\beta_{S}$. Several estimands may be of interest; one of such estimands is the effect of a tariff $t$ on prices
$\tau\equiv\frac{\beta_{S}}{\beta_{D}+\beta_{S}}\,t.$
Let $(\tilde{\beta}_{D}^{\mathrm{obs}},\tilde{\beta}_{S}^{\mathrm{obs}})^{\top}$ denote baseline estimates from Kenya, that may lack external validity.
Due to cost constraints, $\mathcal{E}\in\{\{D\},\{S\}\},$ i.e., we can learn either demand (estimating $\tilde{\beta}_{D}^{\mathrm{exp}}$, with randomized price discounts) or supply (estimating $\tilde{\beta}_{S}^{\mathrm{exp}}$, by introducing regulatory cost shocks on firms). Let
$\big(\tilde{\theta}^{\mathrm{obs}}-\theta\big)\ \sim\ \mathcal{N}\big(b,\Sigma^{\mathrm{obs}}\big),\quad\tilde{\theta}^{\mathrm{obs}}\equiv\sqrt{n}(\tilde{\beta}_{D}^{\mathrm{obs}},\tilde{\beta}_{S}^{\mathrm{obs}})^{\top},\quad\theta\equiv\sqrt{n}(\beta_{D},\beta_{S})^{\top}$
with $\theta$ and $\tilde{\theta}$ denoting the parameter and estimator here rescaled by the square-root of the sample size, and $b$ capturing bias. Our goal is to choose whether to conduct a supply or demand experiment in Uganda to estimate $\tau$. For given estimators $\hat{\theta}$ that combine the estimates from Kenya with the chosen experimental estimate from Uganda under local asymptotics described in Section 5.1
$\tau-\widehat{\tau}=\underbrace{\frac{1}{\sqrt{n}}\check{\omega}(\mathrm{plim}(\tilde{\beta}^{\mathrm{obs}}))^{\top}(\theta-\widehat{\theta})}_{\text{main first order effect}}+\underbrace{o_{p}\left(\frac{1}{n^{1/2}}\right)}_{\text{small higher order effects}},\qquad\check{\omega}(\beta)=\frac{t}{(\beta_{D}+\beta_{S})^{2}}\begin{pmatrix}-\beta_{S}\\ \beta_{D}\end{pmatrix}$,
where we can replace $\check{\omega}$ with its consistent counterpart $\check{\omega}(\tilde{\beta}^{\mathrm{obs}})$. Our goal is to choose between a supply or demand experiment accounting for first-order bias (and variance) effect.
3 Robust experimental design
Next, we introduce an experimental design focusing on the mean-squared error (MSE) of the estimator $\hat{\tau}$, defined, for given bias level $b$ and design $(\mathcal{E},\Sigma(\mathcal{E}))$ as
$\mathrm{MSE}_{b}(\mathcal{E},\Sigma,\gamma)\;=\;\mathbb{E}_{\mathcal{E},\Sigma,b}\!\left[(\widehat{\tau}_{\gamma}-\tau)^{2}\right],$
where $\mathbb{E}_{\mathcal{E},\Sigma,b}$ denotes expectation under the data-generating process implied by $(\mathcal{E},\Sigma(\mathcal{E}))$ and observational bias $b$. Section 5.2 shows that our results for MSE also extend when minimizing confidence intervals’ length.
Ideally, one would minimize $\mathrm{MSE}_{b}$. However, because $b$ is unknown, we consider an uncertainty set given by an $\ell_{\infty}$-ball,
$\mathcal{B}(B)\equiv\{b:\ \|b\|_{\infty}\leq B\},$
with $B\geq 0$ an upper bound on the largest coordinate-wise bias. In our framework, a key challenge is that biases may arise from multiple parameters and uncertainty sets defined by different norms may be possible. A key insight here is that the $\ell_{\infty}$ choice has a desirable property relative to $\ell_{1}$ or $\ell_{2}$: it does not force a trade-off across coordinates (a large bias in one component need not be “offset” by a small bias elsewhere), which is attractive when biases may be positively correlated across parameters. The drawback is that worst-case solutions depend on the radius $B$ which may be unknown in practice.
If an oracle knew $B$, a natural choice would be to pick the minimax design
$\mathrm{MSE}^{*}(B)\;\equiv\;\inf_{(\mathcal{E},\Sigma,\gamma)\in\mathcal{D}}\ \sup_{b\in\mathcal{B}(B)}\ \mathrm{MSE}_{b}(\mathcal{E},\Sigma,\gamma).$
However, having to specify $B$ can pose a large burden on the researchers and make the choice of the experiment sensitive to $B$. Therefore, we seek designs that perform as close as possible to the oracle that knows $B$, uniformly over the values of $B$, following a long-standing tradition in decision theory (e.g. Manski, 2004; Manski and Tetenov, 2007; Montiel Olea et al., 2023; Kitagawa and Tetenov, 2018). Specifically, we minimize the worst-case proportional regret
$\mathcal{R}(\mathcal{E},\Sigma,\gamma)\;\equiv\;\sup_{B\geq 0}\ \frac{\displaystyle\sup_{b\in\mathcal{B}(B)}\mathrm{MSE}_{b}(\mathcal{E},\Sigma,\gamma)}{\displaystyle\mathrm{MSE}^{*}(B)},$
defined adaptation regret by Armstrong et al. (2024) (building in turn on Tsybakov, 1998).
There are two key distinctions relative to such references using the adaptation regret. First (and importantly) we optimize over the design itself rather than fixing it ex ante; consequently, regret is defined relative to an oracle that jointly chooses both the estimator and the design. Second, we allow biases to arise across multiple parameters (rather than a single one), which necessitates a careful specification of the uncertainty set; we adopt an $\ell_{\infty}$-ball, which, as we show yields transparent and computationally-attractive solutions.
3.1 Optimal design
A natural question is whether the adaptation regret provides natural trade-offs between the variance and the bias. To do so, denote $\mathbb{V}_{\mathcal{E},\Sigma}(\cdot)$ the variance operator for given $(\mathcal{E},\Sigma)$.
Definition 1 (Variance regret).
Let
$\alpha(\mathcal{E},\Sigma,\gamma)\;\equiv\;\mathbb{V}_{\mathcal{E},\Sigma}(\hat{\tau}_{\gamma}),\qquad\alpha^{\star}\;\equiv\;\min_{\mathcal{E}\in\mathcal{S}}\ \min_{\Sigma(\mathcal{E})\in\mathcal{G}(\mathcal{E})}\ \min_{\gamma\in\mathbb{R}^{|\mathcal{E}|}}\alpha(\mathcal{E},\Sigma,\gamma),$ (2)
with $\alpha^{\star}$ denoting the smallest feasible variance. We will refer to $\alpha/\alpha^{\star}$ as the variance regret.
The variance regret denotes the ratio between the variance of a given design and estimator relative to the smallest achievable variance. We provide a similar definition for the bias. Let $|1-\gamma|$ define the absolute value of each entry of the vector $1-\gamma$.
Definition 2 (Bias regret).
Define
$\qquad\beta(\mathcal{E},\gamma)\;\equiv\;\Big(\,\|\omega\|_{1}-\|\omega_{\,\mathcal{E}}\|_{1}+\,|1-\gamma|^{\top}|\omega_{\mathcal{E}}|\,\Big)^{\!2},\quad\beta^{\star}\;\equiv\;\min_{\mathcal{E}\in\mathcal{S}}\Big(||\omega||_{1}-\|\omega_{\mathcal{E}}\|_{1}\Big)^{2},$
with $\beta/\beta^{\star}$ defined as the bias regret.
[…]
An ideal design, would set the variance equal to $\alpha^{\star}$ and bias equal to $\beta^{\star}$. Unfortunately, this may be infeasible as it might require different choices of experiments and estimators to achieve one or the other. This raises the question of how to trade-off these two components.
This trade-off is formally characterized in our main theorem below.
Theorem 1.
Consider Setting 1 and let Assumptions  1, 2 hold. Then, for any $(\mathcal{E},\Sigma,\gamma)\in\mathcal{D}$,
$\mathcal{R}(\mathcal{E},\Sigma,\gamma)\;=\;\max\!\left\{\frac{\alpha(\mathcal{E},\Sigma,\gamma)}{\alpha^{\star}},\ \ \frac{\beta(\mathcal{E},\gamma)}{\beta^{\star}}\right\}.$
Proof.
See Appendix A.1. ∎
Theorem 1 provides a simple expression of the regret which does not require researchers to specify $B$. The key insight is that the adaptation regret in our context, leads to a (strictly) quasi-convex objective function, whose worst-case solution is the maximum between the variance and the bias regret. This provides a simple and intuitive trade-off between the two.
Specifically, provided that the class of designs $\mathcal{D}$ is sufficiently flexible (outside boundary solutions), it follows that at the optimum we equalize the variance and bias regret as in Figure 1
$\frac{\alpha}{\alpha^{\star}}\;=\;\frac{\beta}{\beta^{\star}}.$
In practice, this amount to allocate sample size to noisier treatment arms when the variance dominates, and select treatment arms with largest sensitivity when the bias dominates.
[…]
Figure 1: Each feasible design $(\mathcal{E},\Sigma,\gamma)$ maps to a point $\big(\alpha/\alpha^{\star},\ \beta/\beta^{\star}\big)$: the xx–axis is the variance ratio and the yy–axis is the worst–case bias ratio. The blue curve depicts the attainable frontier as we vary shrinkage $\gamma$ and precision $\Sigma$. Level sets of the objective $\mathcal{R}=\max\{\alpha/\alpha^{\star},\ \beta/\beta^{\star}\}$ are axis–aligned squares (the dotted inverted “L” shows the smallest such square touching the frontier). The minimizer outside boundary solutions is where the frontier meets the $45^{\circ}$ line (red dot).
Theorem 1 provides us an immediate solution to the optimal design problem. We compute the estimator, the optimal variance and the experimental choice via backward induction:
$\gamma^{\star}(\mathcal{E},\Sigma)$ $\in\mathrm{arg}\min_{\gamma\in\mathbb{R}^{|\mathcal{E}|}}\,\max\!\left\{\frac{\alpha(\mathcal{E},\Sigma,\gamma)}{\alpha^{\star}},\ \ \frac{\beta(\mathcal{E},\gamma)}{\beta^{\star}}\right\}$
$\Sigma^{\star}(\mathcal{E})$ $\in\mathrm{arg}\min_{\Sigma\in\mathcal{G}(\mathcal{E})}\,\max\!\left\{\frac{\alpha(\mathcal{E},\Sigma,\gamma^{\star}(\mathcal{E},\Sigma))}{\alpha^{\star}},\ \ \frac{\beta(\mathcal{E},\gamma^{\star}(\mathcal{E},\Sigma))}{\beta^{\star}}\right\}$
$\mathcal{E}^{\star}$ $\in\mathrm{arg}\min_{\mathcal{E}\in\mathcal{S}}\,\max\!\left\{\frac{\alpha(\mathcal{E},\Sigma^{\star}(\mathcal{E}),\gamma^{\star}(\mathcal{E},\Sigma^{\star}(\mathcal{E})))}{\alpha^{\star}},\ \ \frac{\beta(\mathcal{E},\gamma^{\star}(\mathcal{E},\Sigma^{\star}(\mathcal{E})))}{\beta^{\star}}\right\}.$
That is, we first choose $\gamma^{\star}$ for each design $\mathcal{E},\Sigma(\mathcal{E})$. We then choose $\Sigma$, whose choice depends on $\beta$ through the shrinkage weights. We finally choose the set of experiments $\mathcal{E}$.
3.2 Intuition with two parameters
To build further intuition, it is useful to study the two-parameter model.
Setting 2 (Illustration with two-parameter model).
Let $\theta=(\theta_{1},\theta_{2})^{\top}$, and mutually independent observational and experimental estimates $\tilde{\theta}^{\mathrm{obs}}_{j}-\theta_{j}\sim\mathcal{N}(b_{j},\ \sigma_{j}^{2}),\quad\tilde{\theta}^{\exp}_{j}-\theta_{j}\sim\mathcal{N}(0,\ v_{j}^{2}),\quad j=1,2.$ For simplicity $\sigma_{j}^{2},v_{j}^{2}$ are fixed. We may run one experiment due to budget constraints: $\mathcal{E}\in\{\{1\},\{2\}\}$. If we pick $j$, we estimate $\hat{\theta}_{j}=\gamma_{j}\,\tilde{\theta}^{\exp}_{j}+(1-\gamma_{j})\,\tilde{\theta}^{\mathrm{obs}}_{j},\quad\hat{\theta}_{-j}=\tilde{\theta}^{\mathrm{obs}}_{-j},$ and, to first order, $\tau(\theta)-\tau(\hat{\theta})=\omega_{1}(\theta_{1}-\hat{\theta}_{1})+\omega_{2}(\theta_{2}-\hat{\theta}_{2}).$
How do the variance and bias regret look like in this easier scenario? We can write
$\alpha^{\star}=\min_{k\in\{1,2\}}\left\{\ \omega_{-k}^{2}\sigma_{-k}^{2}\;+\;\omega_{k}^{2}\frac{\sigma_{k}^{2}\,v_{k}^{2}}{\sigma_{k}^{2}+v_{k}^{2}}\ \right\},\qquad\beta^{\star}=\min_{k\in\{1,2\}}\omega_{k}^{2}.$
Here, $\alpha^{\star}$ follows from the variance-only optimal shrinkage $\gamma_{k}=\sigma_{k}^{2}/(\sigma_{k}^{2}+v_{k}^{2})$ (i.e., the oracle choice when $B=0$); $\beta^{\star}$ equals the smallest between the two sensitivity parameters $\omega$. Both depend on the choice of the estimator and on the class of experiments.
We cannot achieve both $\alpha^{\star}$ and $\beta^{\star}$, since these require different choices of $\gamma$. Instead,
$\alpha(j,\gamma_{j})$ $\equiv\underbrace{\omega_{-j}^{2}\sigma_{-j}^{2}}_{\text{var obs estimate}}+\underbrace{\omega_{j}^{2}\!\big[(1-\gamma_{j})^{2}\sigma_{j}^{2}+\gamma_{j}^{2}v_{j}^{2}\big]}_{\text{var exp and obs estimate}},\quad\beta(j,\gamma_{j})$ $\equiv\big(\underbrace{|\omega_{-j}|}_{\begin{subarray}{c}\text{bias obs}\\ \text{estimate}/B\end{subarray}}+\underbrace{(1-\gamma_{j})|\omega_{j}|}_{\begin{subarray}{c}\text{bias obs minus}\\ \text{ bias exp}/B\end{subarray}}\big)^{2}.$
The very first step is optimizing over the estimator (assuming here for simplicity $\gamma$ can take any value). The following characterizes the regret optimal $\gamma^{\star}$ (see Appendix C.1).
[…]
At the boundaries, if the variance regret $\alpha/\alpha^{\star}$ is always smaller than the bias regret $\beta/\beta^{\star}$, the optimal weight is $\gamma_{j}^{\star}=1$ (i.e., use only the experimental estimate). If instead the variance regret always dominates, the optimal choice is the variance-optimal $\gamma_{j}^{\star}=\sigma_{j}^{2}/(\sigma_{j}^{2}+v_{j}^{2})$.
When the solution is an interior solution for $\gamma$, the optimal design minimizes both the bias and variance regret (now equalized), so that
$j^{\star}\in\arg\min_{j\in\{1,2\}}\underbrace{\left\{\omega_{-j}^{2}\sigma_{-j}^{2}+\omega_{j}^{2}\!\big[(1-\gamma_{j}^{\star})^{2}\sigma_{j}^{2}+(\gamma_{j}^{\star})^{2}v_{j}^{2}\big]\right\}}_{\text{overall variance at }\gamma_{j}^{\star}}=\arg\min_{j\in\{1,2\}}\underbrace{\left\{|\omega_{-j}|+|1-\gamma_{j}^{\star}||\omega_{j}|\right\}}_{\text{worst--case bias}/|B|\text{ at }\gamma_{j}^{\star}}.$
However, cases at the extreme boundaries can also occur. For instance, if the variance regret $\alpha/\alpha^{\star}$ or the bias regret $\beta/\beta^{\star}$ uniformly dominates the other for one of the experiments, the optimal design minimizes the dominating term. These are desiderable properties as the optimal design always prioritizes the dominating term. Appendix Table 5 provides a complete overview of all possible scenarios. More complex scenarios where researchers may allocate sample sizes to each experiment are possible, see Section 6.
Figure 2: Example for $j=2$. Top row: relative bias $\beta/\beta^{\star}$ (dashed) and relative variance $\alpha/\alpha^{\star}$ (solid), both evaluated at the optimal weight $\gamma_{2}^{\star}$. Bottom row: the optimal weight $\gamma_{2}^{\star}$. Columns vary, respectively, (a) $\omega_{2}$ with $v_{1}=0.5$, $v_{2}=1$, $\sigma_{1}=\sigma_{2}=1$, $\omega_{1}=1$; (b) $v_{2}$ with $\omega=(0.9,1)$, $v_{1}=1$, $\sigma=(1,1)$; (c) $\sigma_{2}$ with $\omega=(0.9,1)$, $v=(1,1)$, $\sigma_{1}=1$.
Figure 3: Example for $j=2$ (regret). Different colors correspond to the regret for choosing either experiment. Figure reports $\max\{\alpha(j,\gamma_{j}^{\star})/\alpha^{\star},\beta(j,\gamma_{j}^{\star})/\beta^{\star}\}$ for $j\in\{1,2\}$ as the x–axis parameter varies (columns: $\omega_{2}$, $v_{2}$, $\sigma_{2}$). The vertical dashed line marks indifference, where the two curves intersect; to its left/right, the optimal experiment is the one with lower max regret. Columns vary, respectively, (a) $\omega_{2}$ with $v_{1}=0.5$, $v_{2}=1$, $\sigma_{1}=\sigma_{2}=1$, $\omega_{1}=1$; (b) $v_{2}$ with $\omega=(0.9,1)$, $v_{1}=1$, $\sigma=(1,1)$; (c) $\sigma_{2}$ with $\omega=(0.9,1)$, $v=(1,1)$, $\sigma_{1}=1$.
Example 4 (Numerical illustration).
For an illustrative example, consider choosing $\gamma_{2}^{\star}$ when experiment $j=2$ is twice as precise as $j=1$, with $\sigma_{1}^{2}=\sigma_{2}^{2}=1$, $v_{1}=1$, $v_{2}=0.5$, and $\omega_{1}=1$. The optimal $\gamma_{2}^{\star}$ depends not only on features of design $j$ but also on those of the alternative design j-j, $(v_{-j}^{2},\sigma_{-j}^{2},\omega_{-j})$, through the benchmarks $(\alpha^{\star},\beta^{\star})$. Figure 2 (top row) plots $\alpha(2,\gamma_{2}^{\star})/\alpha^{\star}$ and $\beta(2,\gamma_{2}^{\star})/\beta^{\star}$ at the optimal weight while varying, column by column, $\omega_{2}$, $v_{2}$, and $\sigma_{2}$. The bottom row reports the corresponding $\gamma_{2}^{\star}$. Varying $\omega_{2}$ yields two key effects: both very small and very large $\omega_{2}$ push $\gamma_{2}^{\star}$ toward one, but for different reasons—when $\omega_{2}$ is small the oracle bias decreases faster than the experimental variance (since $\beta^{\star}=\omega_{2}$), while when $\omega_{2}$ is large the oracle bias is $\beta^{\star}=\omega_{1}$ and $\omega_{2}$ has only on the feasible experiment’s bias. Consequently, both regret and the optimal estimator depend on the oracle’s choice of the experiment. In the interior region, the optimal weight first shifts toward the variance–optimal solution (as the oracle bias declines) and then rises again toward a bias–preferred solution. When we vary the experimental and observational variances, we observe the expected patterns: $\gamma_{2}^{\star}$ moves inversely with experimental variance and in the same direction as observational variance. These patterns translate into implications for the regret of the chosen experiment (Figure 3). For the second arm $j=2$, regret decreases sharply as $\omega_{2}$ increases and then decreases, with further decreases becoming gradual once the bias component is already small. By contrast, for the first arm, regret is at best slightly decreasing for small changes in $\omega_{2}$ (when variance effects dominate), but eventually increases with $\omega_{2}$ because the oracle increasingly favors arm $2$, so the opportunity cost of selecting arm $1$ increases. A summary appears in Table 4, with a complete discussion in Appendix C.2. ∎
4 Generalization to moments selection
In many scenarios, we may expect that experiments only identify relevant moments in the data. We now extend the framework to generalized method of moments (GMM) estimators.
Consider a vector of moment conditions $g(\theta)\in\mathbb{R}^{p}$ that stacks all moments (observational and experimental) potentially available for estimation. Let $\Lambda$ denote the Jacobian of the moment function with respect to $\theta$, evaluated at $\theta$ (i.e., $\Lambda:=\partial g(\theta)/\partial\theta^{\top}$). For simplicity, we will assume that $\Lambda$ is known (i.e., moments are linear in $\theta$). This simplifies exposition but it is not necessary in the local asymptotic framework we present in Section 5.1, where $\Lambda$ can be replaced by the Jacobian of $g(\theta)$ evaluated at the observational estimates. For a given covariance matrix $\Sigma$ of the sample moment vector, let $\bar{g}_{\Sigma}$ denote the sample analogue with
$\mathbb{E}[\bar{g}_{\Sigma}]=b,\qquad\mathbb{V}(\bar{g}_{\Sigma})=\Sigma,$
where the vector $b$ captures the biases in the moments. The researcher may decide to access only some of the entries of $\bar{g}$ by selecting a subset of experiments due to budget constraints.
Misspecification set.
Let $\mathcal{I}\subseteq\{1,\dots,p\}$ index all potential experimental moments researchers may select (unbiased by design) and $\mathcal{I}^{c}$ the observational ones (potentially misspecified). We posit bounded misspecification
$\mathcal{B}(B)\;=\;\Big\{b\in\mathbb{R}^{p}:\ b_{j}=0\ \text{for }j\in\mathcal{I},\ \ |b_{j}|\leq B\ \text{for }j\in\mathcal{I}^{c}\Big\}.$ (4)
This restriction states that experimental moments have no bias, whereas observational moments have bias bounded by $B$, potentially unknown to the researcher.
Estimator and design.
Given the empirical moments $\bar{g}_{\Sigma}$, we consider (asymptotically) linear estimators of the form
$\hat{\theta}(W,\Sigma)-\theta\;=\;\Gamma_{\Lambda}(W)\,\bar{g}_{\Sigma},\quad\Gamma_{\Lambda}(W)\equiv-(\Lambda^{\top}W\Lambda)^{-1}\Lambda^{\top}W$ (5)
where $W$ and $\Sigma$ are chosen by the researcher from a feasible set. The structure of the estimator follows a standard GMM asymptotic expansion (e.g. Andrews et al., 2020). [^7] The weighting matrix $W$ may be singular. The expansion $\hat{\theta}-\theta\;\dot{=}\;-(\Lambda^{\top}W\Lambda)^{-1}\Lambda^{\top}W\,\bar{g}_{\Sigma}$ holds whenever $\Lambda^{\top}W\Lambda$ is nonsingular, which can be forced through the set of constraints.
Here, $W$ may have columns of (effectively) zeros, which excludes the corresponding moments from the estimator. This occurs when researchers faces a trade-off in which experiment to run, leading to different moments $\bar{g}$ observed in the experiment.
The matrix $\Sigma$ depends on the experimental design (e.g., sample–size allocations across experimental moments). The constraints on $W$ and $\Sigma$ effectively limits the class of experiments, sample size allocations and potentially estimators the researcher may consider.
Assumption 3 (Constraint set).
Let $(W,\Sigma)\in\mathcal{D}^{\prime}$ for a set $\mathcal{D}^{\prime}$ so that $\Lambda^{\top}W\Lambda$ is invertible, $\Gamma_{\Lambda}(W)\Sigma\Gamma_{\Lambda}(W)^{\top}$ has uniformly bounded entries and is strictly positive definite.
Assumption 3 requires that the variance of the selected (reweighted) moments is uniformly bounded and strictly positive definite (even when $W$ does not select some moments). This ensures that any subset of moments the researcher chooses to use is nondegenerate.
Example 5 (GMM moment selection).
Let $g(\theta)=(g^{\text{obs}}(\theta),\,g^{\text{exp}}_{2}(\theta),\,g^{\text{exp}}_{3}(\theta))^{\top}\in\mathbb{R}^{3}$, where the first component is observational and the other two are experimental. Suppose at most one experimental moment can be used. Selection is encoded by the GMM weight $W\in\mathbb{R}^{3\times 3}$. A simple example of class weighting matrices is
$W^{(2)}=\begin{pmatrix}w_{\mathrm{obs}}&\rho&0\\ \rho&w_{2}&0\\ 0&0&0\end{pmatrix}\qquad W^{(3)}=\begin{pmatrix}w_{\mathrm{obs}}&0&\rho\\ 0&0&0\\ \rho&0&w_{3}\end{pmatrix}$,
provided that $\Lambda^{\top}W\Lambda$ is invertible. ∎
Target and MSE.
Let the estimand be a linear functional $\tau(\theta)=\omega^{\top}\theta$ with $\omega\in\mathbb{R}^{d}$. Then
$\mathrm{MSE}_{b}(W,\Sigma)$ $=\big(\omega^{\top}\Gamma_{\Lambda}(W)\,b\big)^{2}\;+\;\omega^{\top}\Gamma_{\Lambda}(W)\,\Sigma\,\Gamma_{\Lambda}(W)^{\top}\omega.$ (6)
The key distinction now is that the variance and bias regret depend on $\omega$ and $\Lambda$.
Definition 3 (Moment based variance and bias regret).
With a slight abuse of notation, let
$\alpha(W,\Sigma)\;\equiv\;\omega^{\top}\Gamma_{\Lambda}(W)\,\Sigma\,\Gamma_{\Lambda}(W)^{\top}\omega,\qquad\beta(W)\;\equiv\;\big\|\big(\Gamma_{\Lambda}(W)^{\top}\omega\big)_{\mathcal{I}^{c}}\big\|_{1}^{\,2}.$
and $\alpha^{\star},\beta^{\star}$ their corresponding minimizers over $(W,\Sigma)\in\mathcal{D}^{\prime}$. The moment-based variance and bias regret are defined as $\alpha/\alpha^{\star},\beta/\beta^{\star}$ respectively.
The bias now depends on the interaction between the sensitivity $\omega$, the moment Jacobian $\Lambda$ and the identity of which moments are observational and therefore potentially biased. The weighting matrix $W$ determines how these moments influence the downstream estimator $\hat{\theta}$.
As before, (and with a slight abuse of notation) we consider the adaptation regret of the following form:
$\mathcal{R}(W,\Sigma)\;\equiv\;\sup_{B\geq 0}\ \frac{\sup_{b\in\mathcal{B}(B)}\mathrm{MSE}_{b}(W,\Sigma)}{\inf_{(W^{\prime},\Sigma^{\prime})\in\mathcal{D}^{\prime}}\ \mathrm{MSE}_{b}(W^{\prime},\Sigma^{\prime})}.$ (7)
The next theorem generalizes Theorem 1 to this more general framework.
Theorem 2 (Regret characterization).
Let Assumptions 2, 3 hold, $\mathcal{R}$ be as defined in Equation (7),with $\alpha,\beta,\alpha^{\star},\beta^{\star}$ as in Definition 3. Consider an estimator as in Equation (5). Then
$\mathcal{R}(W,\Sigma)\;=\;\max\!\left\{\,\frac{\alpha(W,\Sigma)}{\alpha^{\star}}\,,\ \frac{\beta(W)}{\beta^{\star}}\,\right\}.$ (8)
Proof.
See Appendix B.2. ∎
Theorem 2 shows how our framework generalizes to two related problems (i) choosing the weighting matrix $W$ for estimation and (ii) moment selection. The key insight is to show that the regret is quasi-convex under (approximate) linearity in the moments.
5 Extensions
5.1 Local asymptotics in the presence of non-linearities
This subsection extends our framework for (i) nonlinear moment functions $g(\theta)$ used in GMM in Section 4 and (ii) a smooth, potentially nonlinear estimand $\tau(\theta)$.
Linearization of the moments
Assume the sample moments obey root–$n$ scaling, i.e., $\mathbb{V}(\bar{g}_{\Sigma})=O(n^{-1})$. Let $b=b_{n}$ be a sequence where retain the same misspecification structure as in Section 4: the moment vector has mean $b=b_{n}$ with $(b_{n})_{j}=0$ for $j\in\mathcal{I}$ (experiment-based moments) and $(b_{n})_{\mathcal{I}^{c}}$ potentially nonzero. Our key assumption is a local condition on the moments bias of the form
$\|b_{n}\|^{2}\sqrt{n}\;\to\;0.$ (9)
This condition builds on the local asymptotic frameworks in Andrews et al. (2020) and Armstrong and Kolesár (2021). In their context, $b_{n}||^{2}=1/n$ matching the rate of convergence of the variance. In our case, $b_{n}$ can grow faster, slower or at the same rate of the standard error, encompassing $b_{n}||^{2}=1/n$ as a special case.
Under standard smoothness conditions, a GMM estimator with weighting matrix $W$ satisfies
$\sqrt{n}\big(\hat{\theta}(W,\Sigma)-\theta\big)=-(\Lambda^{\top}W\Lambda)^{-1}\Lambda^{\top}W\,\sqrt{n}\big(\bar{g}_{\Sigma}-b_{n}\big)+o_{p}(1),$
so on the original (unscaled) scale the mean shift is $(\Lambda^{\top}W\Lambda)^{-1}\Lambda^{\top}W\,b_{n}$ and the remainder is $o_{p}(n^{-1/2})$. The condition $\|b_{n}\|^{2}\sqrt{n}\to 0$ ensures that higher order terms are negligible under sufficient smoothness conditions, justifying the use of the linear expansion (5) for our design analysis. Because $b_{n}\to 0$, the Jacobian $\Lambda$ can be consistently estimated by evaluating the derivative of $g(\theta)$ at a preliminary observational estimator $\tilde{\theta}^{\mathrm{obs}}\rightarrow_{p}\theta$, assuming such estimates are available from an observational study (Andrews et al., 2020).
Linearization of the estimand.
Let $\tau:\mathbb{R}^{p}\to\mathbb{R}$ be twice continuously differentiable near $\theta$, with gradient $\omega(\theta):=\partial\tau(\theta)/\partial\theta$ and bounded Hessian. A Taylor expansion around $\theta$ yields
$\tau(\hat{\theta})-\tau(\theta)=\omega(\theta)^{\top}(\hat{\theta}-\theta)\;+\;\tfrac{1}{2}\,(\hat{\theta}-\theta)^{\top}H_{\tau}(\bar{\theta})\,(\hat{\theta}-\theta),$
for some $\bar{\theta}$ between $\hat{\theta}$ and $\theta$, and $H_{\tau}$ is the Hessian. It follows that we can write
$\sqrt{n}\big(\tau(\hat{\theta})-\tau(\theta)\big)=\omega(\theta)^{\top}\,\Gamma_{\Lambda}(W)\,\sqrt{n}\big(\bar{g}_{\Sigma}-b_{n}\big)\;+\;o_{p}(1),$
because $\|\hat{\theta}-\theta\|=O_{p}(n^{-1/2}+\|b_{n}\|)$ and the local condition $\sqrt{n}\,\|b_{n}\|^{2}\to 0$ renders the quadratic remainder $o_{p}(1)$ after $\sqrt{n}$ scaling.
Consequently, the regret expressions derived for the linear case continue to apply up to $o_{p}(1)$, with $\omega$ interpreted as the gradient $\omega(\theta)$ (or any consistent plug-in $\omega(\tilde{\theta}^{\mathrm{obs}})$ from a preliminary estimator $\tilde{\theta}^{\mathrm{obs}}\to_{p}\theta$).
Remark 2 (Rescaling and change of variables linking Example 3).
An equivalent formulation uses $\sqrt{n}$–rescaled coordinates (as in Example 3). Define the rescaled moment vector $\tilde{g}:=\sqrt{n}\,\bar{g}$ and the rescaled estimation error $\tilde{\theta}\equiv\sqrt{n}\,(\hat{\theta}-\theta)$. Then
$\mathbb{V}(\tilde{g})=\Sigma=O(1),\qquad\mathbb{E}[\tilde{g}]=b_{n}\ \ \text{with}\ \ b_{n}\equiv\sqrt{n}\,b,$
so $b_{n}$ is the asymptotic bias and $\Sigma$ the asymptotic variance. The local misspecification condition in (9) is equivalent to $\frac{\|b_{n}\|^{2}}{\sqrt{n}}\;\to\;0,$ which coincides with Example 3. ∎
5.2 Confidence interval length as a regret criterion
In this subsection we consider an alternative loss based on the length of confidence intervals (CIs), focusing on Setting 1 for simplicity. For a candidate design $(\mathcal{E},\Sigma(\mathcal{E}))$, shrinkage weights $\gamma$, and a bias vector $b\in\mathbb{R}^{p}$, define the two–sided, $(1-\eta)$ lower and upper bounds for $\tau(\theta)$ by
$\ell_{b}(\mathcal{E},\Sigma,\gamma)$ $\equiv\hat{\tau}_{\gamma}\;-\;(\omega\circ(1-\gamma))^{\top}b\;-\;z_{1-\eta/2}\,\sqrt{\alpha(\mathcal{E},\Sigma,\gamma)},$
$u_{b}(\mathcal{E},\Sigma,\gamma)$ $\equiv\hat{\tau}_{\gamma}\;-\;(\omega\circ(1-\gamma))^{\top}b\;+\;z_{1-\eta/2}\,\sqrt{\alpha(\mathcal{E},\Sigma,\gamma)},$
where $z_{1-\eta/2}$ is the standard normal $(1-\eta/2)$ quantile.
An audience, indexed by a worst–case bias radius $B\geq 0$, forms the worst–case CI
$L_{B}(\mathcal{E},\Sigma,\gamma)~\equiv~\left[\inf_{\|b\|_{\infty}\leq B}\ell_{b}(\mathcal{E},\Sigma,\gamma),\;\sup_{\|b\|_{\infty}\leq B}u_{b}(\mathcal{E},\Sigma,\gamma)\right],$
where we write $|L_{B}(\mathcal{E},\Sigma,\gamma)|$ for its total length. The proportional regret of $(\mathcal{E},\Sigma,\gamma)$ is
$\tilde{\mathcal{R}}(\mathcal{E},\Sigma,\gamma)~\equiv~\sup_{B\geq 0}\;\frac{|L_{B}(\mathcal{E},\Sigma,\gamma)|}{\inf_{(\mathcal{E}^{\prime},\Sigma^{\prime},\gamma^{\prime})\in\mathcal{D}}|L_{B}(\mathcal{E}^{\prime},\Sigma^{\prime},\gamma^{\prime})|}\!.$
Theorem 3.
Consider Setting 1 and let Assumption 2 hold. Then, for any $(\mathcal{E},\Sigma,\gamma)$,
[…]
Proof.
See Appendix B.1
Theorem 3 shows that the objective function is the same as the MSE-objective function up to a monotonic (square-root) transformation which does not affect the minimizer.
5.3 Vector valued estimands
Next, we extend our framework to vector value estimands of the form $\tau(\theta)\in\mathbb{R}^{q}$ for $q\geq 1$.
Setting 3 (Vector-valued estimands).
Consider a vector-valued target estimand with entries $\tau^{\ell}(\theta)\equiv(\omega^{\ell})^{\top}\theta$ for $\ell=1,\dots q$, and $\omega^{\ell},\theta\in\mathbb{R}^{k}$.
Given $(\mathcal{E},\Sigma)$, and observational and experimental estimates $\tilde{\theta}^{\mathrm{obs}},\tilde{\theta}^{\mathrm{exp}}$, consider a class of linear plug-in estimators to allow the optimal shrinkage weights to depend on the specific entry of $\tau$.
$\hat{\tau}^{\ell}_{\gamma}\equiv\tau^{\ell}(\hat{\theta}^{\ell}(\gamma)),\qquad\hat{\theta}^{\ell}_{j}(\gamma)=\begin{cases}\gamma_{j}^{\ell}\,\tilde{\theta}_{j}^{\mathrm{exp}}+(1-\gamma_{j}^{\ell})\,\tilde{\theta}_{j}^{\mathrm{obs}},&j\in\mathcal{E},\\[2.0pt] \tilde{\theta}_{j}^{\mathrm{obs}},&j\notin\mathcal{E},\end{cases}$
where $\gamma=(\gamma^{\ell}_{j})_{\ell\in\{1,\cdots,q\},j\in\mathcal{E}}\in\mathbb{R}^{q\times|\mathcal{E}|}$ are shrinkage weights that can be an (implicit) function of $(\mathcal{E},\Sigma)$. Let $\hat{\theta}(\gamma)\equiv(\hat{\theta}^{1}(\gamma),\dots,\hat{\theta}^{q}(\gamma))\in\mathbb{R}^{p\times q}$ and $\hat{\tau}_{\gamma}\equiv(\hat{\tau}^{1}_{\gamma}(\hat{\theta}^{1}(\gamma)),\dots,\hat{\tau}^{q}_{\gamma}(\hat{\theta}^{q}(\gamma)))^{\top}\in\mathbb{R}^{q}$.
Setting 3 is a flexible generalization of Setting 1, allowing each entry of $\tau$ to have its own vector of shrinkage weights.
Assumption 4 (First-order estimation error).
For any admissible $(\mathcal{E},\Sigma,\gamma)\in\mathcal{D}$, assume for all $\ell=1,\dots,q$
$\tau^{\ell}(\theta)-\tau^{\ell}\!\big(\hat{\theta}^{\ell}(\gamma)\big)=(\omega^{\ell})^{\top}\,\big(\theta-\hat{\theta}^{\ell}(\gamma)\big),$ (10)
for a known weighting vector $\omega^{\ell}\in\mathbb{R}^{p}$, with $|\omega^{\ell}_{i}|\in(0,\infty)$, for all $i,\ell$.
We begin by generalizing the mean squared error to higher-dimensional parameters. For a given design $(\mathcal{E},\Sigma)$ and shrinkage weights $\gamma$, define the MSE at a fixed observational-bias vector $b$ as
$\mathrm{MSE}_{b}(\mathcal{E},\Sigma,\gamma)\;=\;\mathbb{E}_{\mathcal{E},\Sigma,b}\!\left[\|\widehat{\tau}_{\gamma}-\tau\|_{2}^{2}\right],$ (11)
where $\mathbb{E}_{\mathcal{E},\Sigma,b}$ denotes expectation under the data-generating process implied by $(\mathcal{E},\Sigma(\mathcal{E}))$ and observational bias $b$, and $\|\cdot\|_{2}$ denotes the $L^{2}$ Euclidean norm.
We introduce more compact notation before proceeding to the main result. For a given $(\mathcal{E},\gamma)$ define
$\tilde{\omega}^{\ell}_{\text{obs},j}(\mathcal{E},\gamma)=\begin{cases}\omega^{\ell}_{j},&j\notin\mathcal{E},\\ \omega^{\ell}_{j}(1-\gamma^{\ell}_{j}),&j\in\mathcal{E},\end{cases}\qquad\tilde{\omega}^{\ell}_{\text{exp},j}(\mathcal{E},\gamma)=\begin{cases}0,&j\notin\mathcal{E},\\ \omega^{\ell}_{j}\gamma^{\ell}_{j},&j\in\mathcal{E},\end{cases}$
and let $\tilde{\omega}^{\ell}(\mathcal{E},\gamma)\equiv\begin{pmatrix}\tilde{\omega}^{\ell}_{\text{obs}}(\mathcal{E},\gamma)\\[2.0pt] \tilde{\omega}^{\ell}_{\text{exp}}(\mathcal{E},\gamma)\end{pmatrix}\in\mathbb{R}^{\,p+|\mathcal{E}|}.$ We can write $\hat{\tau}^{\ell}_{\gamma}=\tilde{\omega}^{\ell}(\mathcal{E},\gamma)^{\top}\,\hat{\theta}$, where $\hat{\theta}=(\tilde{\theta}^{\text{obs}\top},\tilde{\theta}^{\text{exp}\top})^{\top}\in\mathbb{R}^{p+|\mathcal{E}|}$. Stacking across the $\ell$ index gives the estimator for the full vector $\tau$: $\hat{\tau}_{\gamma}=\tilde{\omega}(\mathcal{E},\gamma)\,\hat{\theta}$, where the $\ell^{th}$ row of $\tilde{\omega}(\mathcal{E},\gamma)\in\mathbb{R}^{q\times(p+|\mathcal{E}|)}$ is $\tilde{\omega}^{\ell}(\mathcal{E},\gamma)^{\top}$.
The estimator’s variance and worst–case squared bias divided by $B^{2}$ are
$\check{\alpha}(\mathcal{E},\Sigma,\gamma)$ $\equiv\text{Trace}\left(\tilde{\omega}(\mathcal{E},\gamma)\Sigma(\mathcal{E})\,\tilde{\omega}(\mathcal{E},\gamma)^{\top}\right),$ (12)
$\check{\beta}(\mathcal{E},\gamma)$ $\equiv\sum_{\ell=1}^{L}\Big(\,(\omega^{\ell})^{\top}\bar{v}(\gamma)-(\omega^{\ell}_{\,\mathcal{E}})^{\top}\bar{v}(\gamma)_{\,\mathcal{E}}+\,(1-\gamma^{\ell})^{\top}(\bar{v}(\gamma)_{\,\mathcal{E}}\circ\omega_{\,\mathcal{E}})\,\Big)^{\!2},$
where $\omega^{\ell}_{\mathcal{E}}$ is the subvector of $\omega^{\ell}$ corresponding to the entries indexed by the experiment $\mathcal{E}$, and
$\bar{v}(\gamma)\in\operatorname*{arg\,max}_{\{v\in\mathbb{R}^{p}:|v_{j}|\leq 1\forall j\}}\sum_{\ell=1}^{L}\Big(\,(\omega^{\ell})^{\top}v-(\omega^{\ell}_{\,\mathcal{E}})^{\top}v_{\,\mathcal{E}}+\,(1-\gamma^{\ell})^{\top}(v_{\,\mathcal{E}}\circ\omega_{\,\mathcal{E}})\,\Big)^{\!2},$
where the dependence of $\bar{v}(\gamma)$ on $\mathcal{E}$ is omitted for conciseness and $\circ$ denotes the component-wise multiplication. The smallest worst-case variance and bias are defined as
$\check{\alpha}^{\star}=\min_{(\mathcal{E},\Sigma,\gamma)\in\mathcal{D}}\alpha(\mathcal{E},\Sigma,\gamma),\quad\check{\beta}^{\star}=\min_{(\mathcal{E},\Sigma,\gamma)\in\mathcal{D}}\beta(\mathcal{E},\gamma).$
As before, we are interesting in minimizing
$\check{\mathcal{R}}(\mathcal{E},\Sigma,\gamma)\equiv\sup_{B\geq 0}\frac{\sup_{b\in\mathcal{B}(B)}\mathrm{MSE}_{b}(\mathcal{E},\Sigma,\gamma)}{\inf_{(\mathcal{E},\Sigma,\gamma)\in\mathcal{D}}\sup_{b\in\mathcal{B}(B)}\mathrm{MSE}_{b}(\mathcal{E},\Sigma,\gamma)}.$
The result below shows how our results directly extend to this case.
Theorem 4.
Consider Setting 3 and let Assumption 4 hold. Then, for any $(\mathcal{E},\Sigma,\gamma)$,
$\check{\mathcal{R}}(\mathcal{E},\Sigma,\gamma)\;=\;\max\!\left\{\frac{\check{\alpha}(\mathcal{E},\Sigma,\gamma)}{\check{\alpha}^{\star}},\ \ \frac{\check{\beta}(\mathcal{E},\gamma)}{\check{\beta}^{\star}}\right\}.$
Proof.
See Appendix B.2
6 Empirical applications
In this section we provide two empirical applications, one for measuring GE effect and the second one for externally valid site selection.
6.1 An implementation guide for measuring GE effects
Combining experimental or pre-program observational data with structural models is increasingly used for program evaluation (Todd and Wolpin, 2006; Attanasio et al., 2012; Meghir et al., 2022; Allcott et al., 2025). [^8] Across all AEA articles in 2015-2025, about $10\%$ of the experimental papers combine experimental results with structural models (i.e., one-third of the papers that use observational in combination with experimental estimates). Because large-scale experimentation are often infeasible due to cost constraints (Muralidharan and Niehaus, 2017) these models are meant to extrapolate effects from small-scale experiments to learn general-equilibrium (GE) effects. This extrapolation typically relies on other experiments or observational data.
We present a (step-by-step) illustration below when these models are calibrated using previous studies conducted in a different country.
Question
Consider a researcher evaluating a cash transfers for sending children to school in rural Kenya where such a program is not in place. Due to budget constraints, the researcher can only randomize at a small scale (partial equilibrium), but ultimately wishes to predict GE effects. Researchers have access to preliminary estimates from the Mexican PROGRESA experiment (Todd and Wolpin, 2006), which we may be concerned, lack external validity.
6.1.1 Step 1: model description
The first step for a researcher is to describe the modeling choices. We present a stylized model of school choice from Bonhomme and Weidner (2022) and Todd and Wolpin (2006).
Individual choices
Let $S\in\{0,1\}$ denote school attendance, $C$ consumption, $Y$ (pre-transfer) household income, $W$ the child’s potential wage, and $t$ the stipend when enrolled. Abstracting from covariates (we will introduce covariates in the estimation), utility is
$U(C,S,t,\varepsilon)=\xi_{1}\,C+\xi_{2}\,CS+(\xi_{3}-\xi_{1}-\xi_{2})\,tS+\xi_{4}\,S+S\varepsilon,\qquad\varepsilon\sim\mathcal{N}(0,1),$ (13)
with budget C=Y+W(1S)+tSC=Y+W(1-S)+tS. The parametrization $(\xi_{3}-\xi_{1}-\xi_{2})$ is without loss and simplifies expressions below. Enrollment satisfies $S=1\big\{U(Y+t,1,t,\varepsilon)>U(Y+W,0,0,0)\big\}.$ Letting $Z(Y,W,t)\equiv\xi_{1}W-\xi_{2}Y-\xi_{3}t-\xi_{4},$ we have $P(S=1\mid Y,W,t)=\Phi\!\big(-Z(Y,W,t)\big)$, with $\Phi(\cdot)$ denoting the standard Gaussian CDF.
General equilibrium effects
We are interested in the effect of a small stipend $t$ to all eligible (poor) households in rural Kenya. GE feedback is allowed through income and wages: for functions $y(t),w(t)$ and mean-zero idiosyncratic income and wage shocks $\varepsilon_{Yi},\varepsilon_{Wi}$, let
$Y_{i}(t)=y(t)+\varepsilon_{Yi},\qquad W_{i}(t)=w(t)+\varepsilon_{Wi}.$
Let $y_{0}\equiv\partial_{t}y(t)|_{t=0}$ and $w_{0}\equiv\partial_{t}w(t)|_{t=0}$. Define $\phi_{0}\equiv\mathbb{E}\!\left[\phi\!\big(-Z\big(Y(0),W(0),0\big)\big)\right]$. Our estimand of interest is the marginal effect of an increase in a small transfer $t$ to all eligible households
$\left.\frac{\partial\,\mathbb{E}\!\left[\Pr(S{=}1\mid Y(t),W(t),t)\right]}{\partial t}\right|_{t=0}=\;\phi_{0}\cdot\big(\,\xi_{3}+\xi_{2}y_{0}-\xi_{1}w_{0}\,\big),$ (14)
which decomposes into the direct stipend effect $\phi_{0}\xi_{3}$ and the indirect GE effects via income and wages, $\phi_{0}\xi_{2}y_{0}$ and $\phi_{0}\xi_{1}w_{0}$. Under simple market clearing for labor supply and demand, we can show [^9] This follows from the following: set $L_{s}(W,t)=L_{d}(W)$, the labor supply equals the labor demand, and differentiate at $t=0$: $\frac{\partial L_{s}}{\partial W}w_{0}+\frac{\partial L_{s}}{\partial t}=\frac{\partial L_{d}}{\partial W}w_{0}$. Suppose we have a constant number of hours worked for working child, denoted as $H$, and any child not at school is working. From the probit index, $\partial_{t}\mathbb{E}[S\,|\,\cdot]=\phi(-Z)\,\xi_{3}$ and $\partial_{W}\mathbb{E}[S\,|\,\cdot]=-\phi(-Z)\,\xi_{1}$ at $t=0$. Hence $\partial L_{s}/\partial t=-\phi_{0}\xi_{3}H$ and $\partial L_{s}/\partial W=\phi_{0}\xi_{1}H$. Dividing the equation by $H$ we obtain the desired result.
$w_{0}=\;\frac{\phi_{0}\,\xi_{3}}{\phi_{0}\,\xi_{1}-d},$ (15)
where $d$ denotes the slope of the demand curve divided by total hours of work.
6.1.2 Step 2: parametrization and transparency on biases
Given the model, the following step for researchers is to define the parameters of interest. Here, we define the parameters of interest as
$\theta_{1}\;\equiv\;\phi_{0}\,\xi_{2},\qquad\theta_{2}\;\equiv\;\phi_{0}\,\xi_{3},\qquad\theta_{3}\;\equiv\;-\phi_{0}\,\xi_{1},$ (16)
with corresponding estimand
$\tau(\theta)\;=\;\underbrace{\theta_{2}}_{\text{Direct effect}}\;+\;\underbrace{y_{0}}_{\text{Income multiplier}}\,\underbrace{\theta_{1}}_{\text{Income effect}}\;+\;\underbrace{w_{0}(\theta)\,\theta_{3}}_{\text{Wage effect}},\qquad w_{0}(\theta)\;=\;\frac{\theta_{2}}{-\theta_{3}-d},$ (17)
where the explicit form of $w_{0}(\theta)$ follows from (15).
We treat $d$ as known for simplicity, using meta-analyses on demand elasiticities, and $y_{0}=1.5$ because computed experimentally in Kenya by Egger et al. (2022). [^10] Egger et al. (2022) report a total income/consumption multiplier $M_{\text{tot}}=\partial_{t}\mathbb{E}\big[Y_{\text{pre}}(t)+t\big]|_{t=0}=2.5$; therefore the target derivative of pre-transfer income is $y_{0}=\partial_{t}\mathbb{E}\big[Y_{\text{pre}}(t)\big]|_{t=0}=M_{\text{tot}}-1$. Specifically, $d=0.5\frac{1-S_{0}}{W_{0}}$, where $S_{0}$ and $W_{0}$ are baseline probability to go to school (assuming all children not in school go to work) and baseline wages calibrated to the pre-experimental data from PROGRESA and $0.5$ obtained from meta-studies in Espey and Thilmany (2000). [^11] One could also calibrate $S_{0},W_{0}$ to pre-experimental data in Kenya directly when available.
In this example, the primary concern is bias in the observational estimator $\tilde{\theta}^{\mathrm{obs}}$ constructed as described below, abstracting from potential biases in $y_{0},d$ which, for simplicity, we assume to be second-order. When these biases are considered first-order, researchers should let $y_{0},d$ be additional parameters in $\theta$. This difference illustrates a feature of our framework: the chosen parameterization ensures transparency about which biases’ concerns drive the design; these considerations can be collected in a pre-analysis plan.
6.1.3 Step 3: observational study estimates
Using experimental data from Mexico’s PROGRESA program, we estimate $(\xi_{1},\xi_{2},\xi_{3},\xi_{4})$ via probit with standard controls (age, distance to school, eligibility, year, highest grade attained). Estimation is conducted separately by gender, and we will be focusing here on the effects on female students.
We pool treated and control observations to estimate the schooling effect $\xi_{4}$, controlling for the individual-level subsidy. In turn, we obtain observational estimates $\big(\tilde{\theta}_{1}^{\mathrm{obs}},\tilde{\theta}_{2}^{\mathrm{obs}},\tilde{\theta}_{3}^{\mathrm{obs}}\big)$ that map to the model parametrization. Standard errors are constructed using the Delta method with clustering at the village level. Following Section 5.1, we then construct $\omega=\frac{\partial\tau}{\partial\theta}\Big|_{\theta=\tilde{\theta}^{\mathrm{obs}}}$. Point estimates, estimated sensitivity $\omega$, and standard errors are reported below.
Parameter $\tilde{\theta}^{\mathrm{obs}}$ $\omega$ $\sigma$
[…]
Observational estimates for female students, corresponding sensitivity parameter $\omega$, and standard errors.
$\Sigma^{\mathrm{obs}}\;=\;\begin{bmatrix}4.31&-11.31&5.57\\ -11.31&973.11&-126.16\\ 5.57&-126.16&1038.56\end{bmatrix}\times e^{-9}$
Variance–covariance matrix of observational estimates for female students.
We observe heterogeneity in both the sensitivity weights $\omega$ the precision of the observational estimates.
6.1.4 Step 4: describing the candidate experiments
The following step is to list the feasible experiments, subject to budget constraints.
In this example, we consider a researcher who can afford only small, partial-equilibrium experiments in Kenya, and examine three possible experiments for a given sample size $n$:
  • $j=\{1\}$: Unconditional transfer (income shock). The researcher randomizes a small income shock to a small fraction of households (implying no general-equilibrium effects). Under a first-order (Taylor) approximation, the experiment identifies
    $\left.\frac{\partial\,\mathbb{E}[\Pr(S{=}1\mid Y,W,0)]}{\partial Y}\right|_{t=0}=\theta_{1},$
    with precision $v_{1}^{2}/n$. The researcher optimizes over $\gamma_{1}$, while $\theta_{2}$ and $\theta_{3}$ are calibrated to the PROGRESA study.
  • $j=\{2\}$: Conditional cash transfer (stipend). The researcher randomizes a small stipend $t$, conditional on attending school, to a small fraction of households (with prices held fixed). Under a first-order approximation, this identifies
    $\left.\frac{\partial\,\mathbb{E}[\Pr(S{=}1\mid Y,W,t)]}{\partial t}\right|_{t=0}=\theta_{2},$
    with precision $v_{2}^{2}/n$. The researcher optimizes over $\gamma_{2}$, while $\theta_{1}$ and $\theta_{3}$ are calibrated to the PROGRESA study.
  • $j=\{1,2\}$: Two-arm design. The researcher runs (j=1j{=}1) and (j=2j{=}2) on independent samples, identifying $(\theta_{1},\theta_{2})$ with precisions $(v_{1}^{2}/n_{1},\ v_{2}^{2}/n_{2})$ where $n_{1}+n_{2}=n$. They optimize over both $\gamma_{1}$ and $\gamma_{2}$, while $\theta_{3}$ is calibrated to the PROGRESA study.
We consider two scenarios. In the first, the researcher (and oracle) can choose one or both experiments. In the second, they can choose only one of the two (either $j=\{1\}$ or $j=\{2\}$). Regret is defined accordingly, conditional on the set of feasible designs.
For simplicity, we calibrate $v_{1}^{2}$ as $\sigma_{1}^{2}\times n^{\mathrm{obs}}$, where $\sigma_{1}^{2}$ is the variance of $\tilde{\theta}_{1}^{\mathrm{obs}}$ and $n^{\mathrm{obs}}$ is the observational-sample size ($n^{\mathrm{obs}}=1089$ for the female sample using data from eligible students in Todd and Wolpin (2006)). We calibrate $v_{2}^{2}$ analogously as $\sigma_{2}^{2}\times n^{\mathrm{obs}}$. In practice, researchers may use pilot studies to calibrate the experimental variances when available.
Experiment type Identified parameter Choice Variables Calibrated standard error Description
$j=\{1\}$ $\theta_{1}$ $\gamma_{1}$ $6.56e^{-5}/\sqrt{n}$ Unconditional transfer (income shock, UCT). Randomize an income shock to a small fraction of households.
$j=\{2\}$ $\theta_{2}$ $\gamma_{2}$ $9.86e^{-4}/\sqrt{n}$ Conditional cash transfer (education shock, CCT). Randomize a conditional transfer to a small fraction of households.
$j=\{1,2\}$ $(\theta_{1},\,\theta_{2})$ $\gamma_{1},\ \gamma_{2},\ n_{1}$ $6.56e^{-5}/\sqrt{n_{1}}$ (arm 1), $9.86e^{-4}/\sqrt{n-n_{1}}$ (arm 2) Two-arm design. Run (j=1j{=}1) and (j=2j{=}2) on independent samples with sample sizes respectively $n_{1},n-n_{1}$.
Table 1: Design options, identified parameters, choice variables, and per-unit variances.
6.1.5 Step 5: experimental design
Figure 4: Regret–optimal sample allocation across treatment arms when both the unconditional cash transfer (UCT; income shock) and the conditional cash transfer (CCT; stipend) arms are feasible. The optimal solution is interior: the vast majority of participants (over 90%) are assigned to CCT, with a small but nonzero fraction assigned to UCT to hedge misspecification. This pattern reflects the higher payoff to learning about CCT in this application relative to UCT.
Figure 5: Optimal shrinkage weights and design choice as functions of the total sample size $n$. Left panel: When both arms are run, the weight on the UCT experimental estimator is essentially one across $n$, while the weight on the CCT experimental estimator, $\gamma^{\star}_{\text{CCT}}$, rises with precision (from about 0.5 at small $n$ to roughly 0.9 by $n=1000$), reflecting the increasing value of experimental evidence. Right panel: When restricted to a single arm, the optimal choice switches from UCT at small $n$ (variance dominates under CCT) to CCT once $n$ crosses a threshold (around $n\approx 500$), where bias considerations dominate and favor CCT. Notes: $\gamma_{j}^{\star}$ is the optimal weight on the experimental (vs. observational) estimator for parameter $j$; we set $\gamma_{j}^{\star}=0$ to indicate that arm $j$ is not chosen.
Figure 6: Worst–case regret by design. Left panel: Allowing both arms with an interior allocation yields regret that approaches one as $n$ increases, indicating near–oracle performance under our normalization. In contrast, single–arm designs exhibit nonvanishing regret because bias remains first–order even as variance shrinks. Right panel: Focusing on the single–arm problem, UCT dominates only at small $n$; beyond the same threshold ($\approx 500$), CCT delivers uniformly lower regret as variance becomes second–order relative to bias. Notes: Regret is the worst–case objective normalized so that a value of one corresponds to the oracle benchmark; UCT = unconditional cash transfer, CCT = conditional cash transfer.
The final step is to design the experiment and report it in a pre-analysis plan.
Choice of the experiment when only one treatment arm is available
Figure 5 (right panel) displays the selected arm as a function of $n$ (together with the corresponding $\gamma_{j}^{\star}$). We encode “not chosen” as $\gamma_{j}^{\star}=0$. For small samples ($n\lesssim 500$), the optimal choice is UCT: the experimental CCT estimator would be too noisy, so the resulting variance under CCT dominates, while the bias advantage of CCT over UCT is limited because $\omega_{1}$ and $\omega_{2}$ are of similar magnitude. Once $n$ crosses $n\approx 500$, CCT becomes dominant: by combining observational and experimental estimates for CCT, we can achieve comparable variance, and the worst-case bias is lower than under UCT.
Choice of the estimator
In addition to sample allocation, our method optimizes the shrinkage weights of the parameters of interest. Figure 5 (left panel) shows that when both arms can be run, we place essentially full weight on the experimental estimator for UCT ($\gamma^{\star}_{\text{UCT}}\approx 1$ across $n$), while for CCT we gradually shift toward the experimental estimator as precision improves: $\gamma^{\star}_{\text{CCT}}$ rises from about $0.5$ at small $n$ to approximately $0.9$ by $n=1000$. Intuitively, higher $n$ lowers experimental variance, making it optimal to rely more heavily on experimental evidence for the CCT parameter.
Sample size allocation when both arms are feasible
Figure 4 reports the sample allocation implied by the regret-optimal design when both treatment arms are feasible. The optimal allocation is interior: most participants (over roughly $90\%$) are assigned to the CCT arm (education), with a small but nonzero fraction assigned to the UCT arm (income shock). Two forces rationalize this pattern. First, the CCT parameter is more misspecification-sensitive ($\omega_{2}\approx 1.98$ vs. $\omega_{1}\approx 1.5$), so generating experimental evidence on CCT has a larger payoff in bias reduction. Second, the income effect $\theta_{1}$’s variance is about an order of magnitude smaller than the stipend effect $\theta_{2}$’s variance in our PROGRESA calibration, lowering the marginal returns to learn about UCT relative to CCT. Hence, while the designer keeps some allocation on UCT, most of the sample is placed on CCT; our framework makes this bias–variance trade-off explicit.
Regret comparisons
Figure 6 reports worst-case regret for each design choice. The left panel shows that the regret of our method (green line, running both arms with an interior allocation) declines toward one as $n$ increases, reflecting near-oracle performance under our normalization. In contrast, single-arm designs exhibit nonvanishing regret because bias remains first order even as variance shrinks. The right panel focuses on the single-arm problem: both the oracle and the researcher can only choose a single arm. In this case, UCT is preferable only for small $n$; beyond the same threshold ($\approx 500$), CCT yields uniformly lower regret as variance becomes second order relative to bias. The regret vanishes relative to the oracle that can only choose a single arm.
Implications
When only one arm is feasible, CCT is preferable only once the total sample size is sufficiently large; otherwise UCT is preferred. This pattern reflects the interaction between bias and variance: for small samples, CCT’s bias advantage is not large enough to compensate for its higher variance relative to UCT, whereas at larger samples the bias reduction becomes dominant. Existing designs based on variance considerations would fail to provide solutions that balance bias and variance comparisons.
When feasible, it is instead optimal to run two arms with highly unbalanced sample, reflecting the relative contribution of the variance for each treatment arm.
Following these steps, researchers can similarly report their design in a pre-analysis plan.
6.2 Measuring the performance of our method for site selection
Next, we show how observational evidence can inform whom to recruit into an experiment. We consider the setting in Banerjee et al. (2024), who study how the expansion of microfinance affects village social networks in Karnataka, India. We focus on the first outcome reported by Banerjee et al. (2024) corresponding to the density of the network, i.e., the percentage of connections a random household in a village has relative to the village size.
Background
In their observational analysis, the authors assemble panel data from 75 villages in Karnataka, 43 of which were exposed to microfinance. Because program rollout was not randomized across villages, they estimate effects using Difference-in-Differences (DiD). While DiD is informative, the lack of randomized assignment may bias estimates in the presence of selection (e.g., Ghanem et al., 2022). Motivated by these limitations, the authors subsequently conducted an experimental evaluation in one metropolitan area, randomizing microfinance access across 104 urban neighborhoods. We only use observational variation for choosing the experimental design and estimator. We then use experimental variation generated by Banerjee et al. (2024) to validate our procedure.
Research question
In practice, implementation costs often depend on how geographically dispersed the study sites are due to coordination and survey costs. Using preliminary observational estimates from Banerjee et al. (2024) in Karnataka, we ask: Which area(s) in Karnataka should be prioritized for an experimental evaluation, and how many villages should be enrolled?
Figure 7: Observational villages and area-level DiD effects in Karnataka. The background heat map partitions the state into the four contiguous areas used to compute area-specific DiD estimates of microfinance’s impact on network density; the color scale encodes the DiD value. Points mark village locations: circles denote villages exposed to microfinance and triangles denote unexposed villages in the observational sample (authors’ survey). Major cities (Bengaluru, Mysuru, Tumakuru) are labeled for orientation.
Observational study
We partition the observational sample into four geographically contiguous areas that group nearby villages; each area contains 11–12 villages that were exposed to microfinance during the study period. Figure 7 maps these four areas and reports the corresponding DiD point estimates. Three of the four areas display negative estimated effects, with meaningful variation in magnitudes across areas. Table 2 summarizes, for each area, samples sizes and variances, clustered at the village level. These area-specific DiD estimates and their variances serve as our primary observational inputs for experimental design. [^12] We use DiD estimates to mimic the estimator used by the authors. An analogous analysis can be conducted after empirical-Bayes shrinkage of area-level estimates, omitted for brevity.
Table 2: Area-level observational inputs for Karnataka villages. The outcome is network density (corresponding to the average share of connection of an individual to other individuals in a village). For each area, we report the numbers of treated ($n_{1}$) and untreated ($n_{0}$) villages, the pre-intervention variance of density pooled across arms ($v_{\text{pre}}^{2}$), the area-specific DiD estimate $\hat{\mu}$, and its sampling variance $\widehat{\sigma}^{2}$.
Area $n_{1}$ $n_{0}$ $v_{\text{pre}}^{2}$ $\hat{\mu}$ $\widehat{\sigma}^{2}$
11 $11$ 66 0.000 4570.000\,457 0.0187-0.0187 0.000 045 30.000\,045\,3
22 $12$ 99 0.001 750.001\,75 0.0377-0.0377 0.000 1400.000\,140
33 $11$ $12$ 0.001 380.001\,38 0.003 90-0.003\,90 0.000 1870.000\,187
44 $11$ 66 0.000 9320.000\,932 $0.0148$ 0.000 1300.000\,130
Experimental design
We consider a family of designs that (i) select $E\in\{1,\ldots,4\}$ geographic areas in Karnataka from which to recruit experimental sites and (ii) assign $n_{1}$ villages to treatment and $n_{0}=n_{1}$ to control (so the total sample is $n=2n_{1}$). Here, $E=1$ corresponds to recruiting from a single area, while $E=4$ recruits from all four areas. We examine a grid of sample sizes that varies the number of participants from $10$ up to $104$ ($52$ treated units), with the latter corresponding to the size of the experiment in Banerjee et al. (2024). For variance calibration, we take $v_{\mathrm{pre},a}^{2},a\in\{1,\cdots,4\}$ to be the pre-intervention, area-level variance of network density in Table 2. We assume that the variance of a single treated–control difference in the experiment is $2v_{\mathrm{pre},a}^{2}$.
Figure 8: Optimal area selection and village allocation for a large experiment ($n_{1}=52$) under constraints on the number of areas $E$. Each panel corresponds to a value of $E$; shading indicates the total number of recruited villages in each area. With $E=1$ the design concentrates in the area with the noisiest observational estimate; as $E$ increases, allocation spreads across areas, reflecting a trade-off between observational uncertainty and experimental variance.
Figure 9: Area-level allocations as total treated villages $n_{1}$ varies (from ten to fifty-two villages), for each constraint on the number of eligible areas $E$. Lines show the number of villages assigned to each area.
Figure 10: Relative MSE across designs under an hypothetical scenario where the bias equal the difference between the experimental and observational ATE estimate from Banerjee et al. (2024) as $n_{1}$ varies (from ten to fifty-two villages). The figure reports $\mathrm{MSE}/\mathrm{MSE}^{\star}$ for (i) the proposed design (area targeting and optimal $\gamma^{\star}$), (ii) a random benchmark (uniform area selection and equal allocation), where $\mathrm{MSE}^{*}$ denotes the MSE of the oracle that knows the bias, as $n_{1}$ varies and for each $E$. The proposed design closely tracks the oracle across $n_{1}$ and $E$, while the random benchmark remains substantially above, especially when $E$ is small.
Experimental design
We begin by studying which areas are selected and how sample is allocated when the total experimental sample is large ($n_{1}=52$) and the number of admissible areas may or may not be constrained. Figure 8 visualizes the resulting allocation across Karnataka’s four observational areas for $E\in\{1,2,3,4\}$. With $E=1$, the algorithm concentrates recruitment in the Tumakuru area—the location with the highest uncertainty in the observational estimates. As $E$ relaxes to $2$ and $3$, the total sample splits across them in roughly (but not exactly) equal shares, reflecting a trade-off between (i) exploiting heterogeneity in the observational variance and (ii) keeping experimental variance low. When $E=4$, all areas are eligible and the allocation smooths further across space. In the rest of our discussion, we discuss properties when not all areas may be selected ($E<4$).
Figure 9 tracks how area-level allocations evolve with the total number of villages. When the experiment is small (e.g., $10$ villages) and $E=1$, the algorithm favors an area with relatively low experimental variance, even if its observational estimate is less uncertain, because, at small $n$, experimental noise is first-order. As the total sample increases, the weight shifts toward areas with more uncertain observational estimates. Consistently, the area with the largest experimental variance (Area 2) is typically excluded.
Appendix Figure 13 reports the optimal shrinkage $\gamma_{j}^{\star}$ by area as total villages vary. The solution is interior for all $E<4$: $\gamma_{j}^{\star}$ rises with sample size, approaching one from below as experimental noise diminishes. For example, at a moderate sample (around 15 villages), $\gamma^{\star}$ is about $0.65$ when $E\in\{2,3\}$; with $E=1$, concentrating the sample within one area lowers experimental variance and therefore yields a larger $\gamma^{\star}$.
Main MSE comparisons
We compare three strategies: (i) our chosen design which chooses the area, sample size and $\gamma^{\star}$ as described in Section 3, (ii) a standard benchmark that selects areas uniformly at random, allocates villages evenly across the chosen areas, and sets $\gamma_{j}=1$ for experimental estimates, and (iii) an oracle that knows the bias vector $b$ and optimizes both design (areas and allocation) and $\gamma$.
Because $b$ is unknown in practice, for this exercise, we calibrate it using the difference between the Hyderabad RCT ATE obtained from the follow-up experiment of Banerjee et al. (2024) and the Karnataka DiD average effect across sites, treating $b$ as common across areas.
Figure 10 plots the ratio $\text{MSE}/\text{MSE}^{\star}$, where $\text{MSE}^{\star}$ is the oracle’s MSE. As the number of treated villages $n_{1}$ increases, all designs improve; because the denominator also falls with $n_{1}$, the ratio need not be monotone. Across $E\in\{1,2,3\}$, the random benchmark remains well above the oracle—by roughly 3–5 percentage points (pp) for $E=1$, about 10 pp for $E=2$, and around 20 pp for E=3E=3—whereas our design tracks the oracle closely (within a few percentage points) even without knowledge of $b$. The gap is the largest when a larger number of areas $E$ can be included, because our procedure can better emulate the oracle by selecting (and differentially weighting) favorable areas.
This result illustrates the merit of the method that improves MSE up to $20$ percentage points compared to a standard benchmark and leads to a near-oracle solution.
7 Implications for practice
This paper studies experimental design in the presence of observational evidence. A key challenge is that the bias of the observational estimators is unknown in practice. We adopt a minimax proportional regret criterion that compares the mean-squared error (MSE) of a candidate design to that of an oracle that knows the worst-case bias. This reveals a fundamental trade-off between precision and robustness. The optimal design balances the design’s variance normalized by the smallest achievable variance (variance gap) and its worst-case bias normalized by the smallest attainable bias (bias gap). We propose a procedure that jointly determines: (i) how to combine observational and experimental evidence; (ii) how to allocate precision across experiments given budget constraints; and (iii) which treatment arm and/or sub-population to include in the experiment with fixed experimental costs.
In practice, the workflow is:
  • Define the estimand(s) of interest. Specify $\tau(\theta)$ for a known mapping $\tau$ and unknown parameters $\theta\in\mathbb{R}^{p}$. For example, $\tau(\theta)$ may represent a counterfactual, a general-equilibrium effect, or an average impact across locations. Adopt a parametrization in which some (but not necessarily all) components of $\theta$ can be learned experimentally; this clarifies what the experiment can identify. When such parametrization is not readily available, Section 4 shows how our procedure applies to $\tau$ identified through arbitrary moment restrictions.
  • Assemble informative observational evidence. Collect observational estimates $\tilde{\theta}^{\rm obs}$ and their covariance $\tilde{\Sigma}^{\rm obs}$. These serve as informative (but potentially biased) baselines. Such evidence may come from observational designs, structural estimates with pre-experimental data, or prior experiments conducted in different contexts.
  • Compute the sensitivity parameters. Using the observational baseline, compute the sensitivity weights $\omega\;=\;\frac{\partial\tau(\theta)}{\partial\theta}\big|_{\tilde{\theta}^{\rm obs}}$, which quantify how bias in each coordinate of $\theta$ propagates to $\tau(\theta)$. Large $|\omega_{j}|$ indicates greater payoff to learning the $j$th component.
  • Specify feasibility constraints and calibrate experimental variance. Enumerate the admissible design set, and the budget on sample size allocation. Calibrate per-unit experimental variances using pilot studies or historical data to form the set of feasible designs $\mathcal{D}$.
  • Run the method and decide. Optimize over the design and corresponding estimators. The output is a pre-analysis plan with the selected arm(s), sample sizes, and pre-specified combination rule ($\gamma$).
The applicability of our framework spans a wide range of settings in economics and beyond. Examples include estimating general equilibrium effects or structural models (Todd and Wolpin, 2006; Attanasio et al., 2012; Meghir et al., 2022; Kreindler et al., 2023; de Albuquerque et al., 2025); choosing among alternative treatment arms in factorial designs (Muralidharan et al., 2020; Bandiera et al., 2025); and deciding where to run the next experiment for external validity (Gechter et al., 2024; Olea et al., 2024). In industrial organization, applications include choices between demand-side and supply-side interventions (Bergquist and Dinerstein, 2020), the effects of information acquisition in markets (Allende et al., 2019; Larroucau et al., 2024), and decisions about which additional data source to acquire to improve statistical analysis (Allcott et al., 2025). Beyond economics, medical applications include allocating sample size across subgroups and allocating doses across treatment arms (e.g., Porter et al., 2024; Morita et al., 2017; Manski, 2025).
Several open questions remain for future research. These include settings where researchers have well-specified priors about bias such as from Bayesian models (Gechter et al., 2024), or other modeling choices (Olea et al., 2024). This may be potentially incorporated through additional moment conditions as in Section 4. Additional future extensions include sequential or adaptive experimental choices or settings where the target estimand is only be partially identified (e.g. Cesa-Bianchi et al., 2025).
Appendix A Proofs of main results
A.1 Proof of Theorem 1 and Theorem 2
Theorem 1 is a special case of Theorem 2 proved below.
Step 1 (worst-case MSE). By Hölder’s inequality,
$\sup_{b\in\mathcal{B}(B)}\big(\omega^{\top}\Gamma_{\Lambda}(W)\,b\big)^{2}\leq\Big(B\,\big\|\big(\Gamma_{\Lambda}(W)^{\top}\omega\big)_{\mathcal{I}^{c}}\big\|_{1}\Big)^{2}=B^{2}\,\beta(W).$
In addition, this inequality is attained by $b_{j}=B\,\mathrm{sign}\!\big((\Gamma_{\Lambda}(W)^{\top}\omega)_{j}\big)$ for $j\in\mathcal{I}^{c}$ (and $b_{j}=0$ for $j\in\mathcal{I}$). Hence $\sup_{b\in\mathcal{B}(B)}\mathrm{MSE}_{b}(W,\Sigma)=\alpha(W,\Sigma)+B^{2}\beta(W).$
Step 2 (oracle envelope and basic properties). Let $\delta(t):=\inf_{W,\Sigma}\{\alpha(W,\Sigma)+t\beta(W)\}$. Recall that $\beta(W)$ and $\alpha(W,\Sigma)$ are bounded from above and $\alpha(W,\Sigma)>0$ by Assumption 3. Being the pointwise infimum of affine functions of $t$, $\delta$ is concave, nondecreasing, and finite for $t\geq 0$. Moreover,
$\delta(0)=\alpha^{\star},\qquad\lim_{t\to\infty}\frac{\delta(t)}{t}=\beta^{\star}.$
Step 3 (quasi-convexity and boundary maximization). Fix $(\alpha,\beta)$. Consider
$\phi(t):=\frac{\alpha+t\beta}{\delta(t)},\qquad t\geq 0.$
Because the numerator is affine and the denominator is positive and concave in $t$, $\phi$ is quasi-convex on $0,\infty)$. Hence, on any compact interval $[0,T]$, $\max_{t\in[0,T]}\phi(t)$ is attained at the boundary $\{0,T\}$. Letting $T\to\infty$ and using the limits in Step 2,
$\sup_{t\geq 0}\phi(t)=\max\!\left\{\frac{\alpha}{\delta(0)},\ \lim_{t\to\infty}\frac{\alpha+t\beta}{\delta(t)}\right\}=\max\!\left\{\frac{\alpha}{\alpha^{\star}},\ \frac{\beta}{\beta^{\star}}\right\},$
with the stated conventions $0/0=0$.
Appendix B Proofs of the extensions
B.1 Proof of Theorem 3
The proof of Theorem 3 follows similarly the one of Theorem 2 with minor modifications. It is easy to show that
$|L_{B}(\mathcal{E},\Sigma,\gamma)|=2\{z_{1-\eta/2}\,\alpha(\mathcal{E},\Sigma,\gamma)+B\,\sqrt{\beta(\mathcal{E},\gamma)}\}.$
We can now follow verbatim the proof of Theorem 2, with $\delta(t):=\inf_{\mathcal{E},\Sigma,\gamma}2\{z_{1-\eta/2}\sqrt{\alpha(\mathcal{E},\Sigma,\gamma)}+t\sqrt{\beta(\mathcal{E},\gamma)}\}$ with limits
$\delta(0)=2z_{1-\eta/2}\sqrt{\alpha^{\star}},\qquad\lim_{t\to\infty}\frac{\delta(t)}{t}=2\sqrt{\beta^{\star}}.$
Taking the ratios and limits as in Step 3 of the proof of Theorem 2 the proof completes.
B.2 Proof of Theorem 4
It is easy to show that
$\check{R}_{B}\left(\mathcal{E},\Sigma,\gamma\right)\equiv\sup_{b\in\mathcal{B}(B)}\ \mathrm{MSE}_{b}(\mathcal{E},\Sigma,\gamma)=\check{\beta}(\mathcal{E},\gamma)B^{2}+\check{\alpha}(\mathcal{E},\Sigma,\gamma),$
with $\check{\beta},\check{\alpha}$ as defined in Equation (12), where $\check{\beta}\geq 0$ by construction. In addition $\check{\alpha}>0$ and bounded away from zero by Assumption 1. We can therefore follow verbatim the proof of Theorem 1.
Appendix C Examples
C.1 Two parameters examples in Equation (3)
By Theorem 1, for fixed $(\mathcal{E},\Sigma)$ the adaptation regret as a function of the shrinkage vector equals
$\mathcal{R}(\gamma)\;=\;\max\!\Big\{\alpha(\mathcal{E},\Sigma,\gamma)/\alpha^{\star},\;\beta(\mathcal{E},\gamma)/\beta^{\star}\Big\},$
with $\alpha^{\star},\beta^{\star}$ constants.
Step 1. With independence and two parameters,
$\alpha(\mathcal{E},\Sigma,\gamma)=C+\omega_{j}^{2}\!\Big(\gamma_{j}^{2}v_{j}^{2}+(1-\gamma_{j})^{2}\sigma_{j}^{2}\Big),$
where $C$ does not depend on $\gamma_{j}$. Hence $\alpha(j,\gamma_{j})\equiv\alpha(\mathcal{E},\Sigma,\gamma)$ is a strictly convex quadratic in $\gamma_{j}$ with unique minimizer at
$\gamma_{j}^{\mathrm{var}}=\frac{\sigma_{j}^{2}}{\sigma_{j}^{2}+v_{j}^{2}}.$
Moreover, $\alpha(j,\gamma_{j})$ is strictly increasing on $\big(\gamma_{j}^{\mathrm{var}},1\big)$ and strictly decreasing on $\big[0,\gamma_{j}^{\mathrm{var}}\big)$.
Write the worst–case bias component holding the other coordinate fixed as
$\beta(j,\gamma_{j})\;=\;\Big(A_{j}+|\omega_{j}|\cdot|1-\gamma_{j}|\Big)^{2},$
where $A_{j}\geq 0$ collects all terms not involving $\gamma_{j}$ (including the contribution from the other coordinate). On $[0,1]$, $|1-\gamma_{j}|=1-\gamma_{j}$, so $\beta(j,\gamma_{j})$ is strictly decreasing in $\gamma_{j}$ and convex with minimum at $\gamma_{j}=1$. Therefore, any minimizer satisfies
$\gamma_{j}^{\star}\in\big[\gamma_{j}^{\mathrm{var}},\,1\big].$
Step 2. On $\big(\gamma_{j}^{\mathrm{var}},1\big)$ the map $\gamma_{j}\mapsto\alpha(j,\gamma_{j})/\alpha^{\star}$ is strictly increasing, while $\gamma_{j}\mapsto\beta(j,\gamma_{j})/\beta^{\star}$ is strictly decreasing. Hence the function
$f(\gamma_{j})\equiv\max\!\Big\{\alpha(j,\gamma_{j})/\alpha^{\star},\ \beta(j,\gamma_{j})/\beta^{\star}\Big\},\qquad\gamma_{j}\in\big[\gamma_{j}^{\mathrm{var}},1\big],$
is minimized either (i) at a boundary point, or (ii) at the unique interior point where the two arguments are equal (by strict monotonicity, at most one intersection exists). Therefore,
  • If $\alpha(j,\gamma_{j})/\alpha^{\star}<\beta(j,\gamma_{j})/\beta^{\star}$ for all $\gamma_{j}\in\big(\gamma_{j}^{\mathrm{var}},1\big)$, then $f(\gamma_{j})=\beta(j,\gamma_{j})/\beta^{\star}$ on that interval. Since this term is strictly decreasing, the minimizer is the right boundary $\gamma_{j}^{\star}=1$.
  • If $\alpha(j,\gamma_{j})/\alpha^{\star}>\beta(j,\gamma_{j})/\beta^{\star}$ for all $\gamma_{j}\in\big(\gamma_{j}^{\mathrm{var}},1\big)$, then $f(\gamma_{j})=\alpha(j,\gamma_{j})/\alpha^{\star}$ on that interval. Since this term is strictly increasing there, the minimizer is the left boundary $\gamma_{j}^{\star}=\gamma_{j}^{\mathrm{var}}=\sigma_{j}^{2}/(\sigma_{j}^{2}+v_{j}^{2})$.
  • Otherwise, by the intermediate value theorem and strict monotonicity of the two curves, there exists a unique $\gamma_{j}\in\big(\gamma_{j}^{\mathrm{var}},1\big)$ such that $\frac{\alpha(j,\gamma_{j})}{\alpha^{\star}}\;=\;\frac{\beta(j,\gamma_{j})}{\beta^{\star}}.$
C.2 Numerical example
Here, we present a numerical example in Setting 2.
Choice of $\gamma^{\star}$
Figure 2 (top row) reports $\alpha(2,\gamma_{2}^{\star})/\alpha^{\star}$ and $\beta(2,\gamma_{2}^{\star})/\beta^{\star}$ evaluated at the optimal weight while varying, column by column, $\omega_{2}$, $v_{2}$, and $\sigma_{2}$. In the bottom row, we plot the corresponding $\gamma_{2}^{\star}$.
Observation 1: small $\omega_{2}$ pushes $\gamma_{2}^{\star}$ toward one. The first (left) column sets $v_{1}=v_{2}/2$ (the first experiment is more precise) and $\omega_{1}=1$ (with $\sigma_{1}=\sigma_{2}=1$). For small $\omega_{2}$ (low sensitivity of the second coordinate to bias), the optimal choice is $\gamma_{2}^{\star}=1$. The reason is twofold. First, the bias ratio $\beta(2,\gamma)/\beta^{\star}=\big(|\omega_{1}|+|1-\gamma||\omega_{2}|\big)^{2}/\beta^{\star}$ is normalized by $\beta^{\star}=\min\{|\omega_{1}|,|\omega_{2}|\}^{2}=\omega_{2}^{2}$ when $\omega_{2}<\omega_{1}$; thus even a small observational component $|1-\gamma|>0$ makes the ratio increase substantially. Second, the variance ratio $\alpha(2,\gamma)/\alpha^{\star}$ is multiplied by $\omega_{2}^{2}$ in its $j=2$ contribution, so its dependence on $\gamma$ becomes negligible as $\omega_{2}\downarrow 0$. Together these forces make the bias ratio dominant and push $\gamma_{2}^{\star}$ to the boundary at $1$. The implication is that for parameters with low $\omega_{2}$, we typically shrink $\hat{\theta}_{2}$ more toward the experimental estimate $\tilde{\theta}_{2}^{\mathrm{exp}}$.
Observation 2: when $\omega_{2}$ is of the same order as $\omega_{1}$, an interior solution emerges. As $\omega_{2}$ increases toward $\omega_{1}$ (roughly $0.6\,\omega_{1}$ and above), the normalization ceases to penalize $j=2$ as harshly, and an interior solution appears: $\gamma_{2}^{\star}$ moves down from $1$ toward the variance–only weight, trading off bias and variance. The first kink in the bottom-left panel reflects this regime change, where $\gamma_{2}^{\star}$ is selected to equalize bias and variance.
Observation 3: for $\omega_{2}>1$, $\gamma_{2}^{\star}$ increases again. At $\omega_{2}=\omega_{1}$ we observe a second kink, since the worst–case bias normalizer switches to $\beta^{\star}=\omega_{1}^{2}$. This raises the bias cost of putting weight on the observational estimate; as $\omega_{2}$ increases above $1$, $\gamma_{2}^{\star}$ rises again. This pattern is driven by the oracle benchmark, which optimizes both the estimator and the design.
Observation 4: for $\omega_{2}\gtrsim 1.25$, $\gamma_{2}^{\star}$ increases slowly and stays interior. Around $\omega_{2}\approx 1.25$ we observe a third kink: the variance normalizer $\alpha^{\star}$ switches branches (from favoring $k=1$ to favoring $k=2$ in the variance–only comparison). The optimal $\gamma_{2}^{\star}$ remains interior but grows more slowly, because increases in $\gamma_{2}^{\star}$ now have a larger relative impact on $\alpha/\alpha^{\star}$. [^13] A further kink could occur if the solution transitioned into a pure variance–dominant regime (where the boundary weight is chosen), which does not arise for the range of $\omega_{2}$ shown here.
Observation 5: $\gamma_{2}^{\star}$ increases with experimental precision and decreases with observational precision The second and third columns vary $v_{2}$ and $\sigma_{2}$, respectively. In both, we fix $v_{1}=1$ and take $\omega_{1}=0.9\,\omega_{2}$ (so here $\omega_{1}=0.9$, $\omega_{2}=1$), with $\sigma_{1}=\sigma_{2}=1$ unless varied. Across $v_{2},\sigma_{2}\in[0.5,2]$, the solution is interior. As $v_{2}\downarrow 0.5$ or $\sigma_{2}\uparrow 2$, the optimal weight $\gamma_{2}^{\star}$ increases toward $1$, placing nearly all weight on the experimental estimate. Conversely, as $v_{2}$ becomes large or $\sigma_{2}$ becomes small, $\gamma_{2}^{\star}$ approaches its lower bound, the variance–minimizing weight $\sigma_{2}^{2}/(\sigma_{2}^{2}+v_{2}^{2})$. This is a typical (and desired) behavior of shrinkage estimators.
Table 3: Example: Regimes for $\gamma_{2}^{\star}$ as a function of $\omega_{2}$ for example in Appendix Figure 2 with $\sigma_{1}^{2}=\sigma_{2}^{2}=1,v_{1}=1,v_{2}=0.5,\omega_{1}=1$.
Regime $\beta^{\star}$ $\alpha^{\star}$ Solution regime ($\gamma_{2}^{\star}$) Trend of $\gamma_{2}^{\star}$
$\boldsymbol{\omega_{2}\ll\omega_{1}}$ $|\omega_{2}|^{2}$ $\omega_{2}^{2}\sigma_{2}^{2}\;+\;\omega_{1}^{2}\dfrac{\sigma_{1}^{2}v_{1}^{2}}{\sigma_{1}^{2}+v_{1}^{2}}$ Bias–dominant Constant $\gamma_{2}^{\star}=1$
$\boldsymbol{\omega_{2}<\omega_{1}}$ $|\omega_{2}|^{2}$ $\omega_{2}^{2}\sigma_{2}^{2}\;+\;\omega_{1}^{2}\dfrac{\sigma_{1}^{2}v_{1}^{2}}{\sigma_{1}^{2}+v_{1}^{2}}$ Interior Decrease in $\omega_{2}$
$\boldsymbol{\omega_{2}>\omega_{1}}$ $|\omega_{1}|^{2}$ $\omega_{2}^{2}\sigma_{2}^{2}\;+\;\omega_{1}^{2}\dfrac{\sigma_{1}^{2}v_{1}^{2}}{\sigma_{1}^{2}+v_{1}^{2}}$ Interior Increase in $\omega_{2}$
$\boldsymbol{\omega_{2}\gg\omega_{1}}$ $|\omega_{1}|^{2}$ $\omega_{1}^{2}\sigma_{1}^{2}\;+\;\omega_{2}^{2}\dfrac{\sigma_{2}^{2}v_{2}^{2}}{\sigma_{2}^{2}+v_{2}^{2}}$ Interior Slow increase in $\omega_{2}$
Choice of the design
Figure 11 compares the maximum regret of choosing experiment $j=1$ versus $j=2$ in three scenarios. The vertical dashed line marks the value where the two curves intersect and the designer is indifferent between $j=1$ and $j=2$. In each column, we keep the same parameterizations as in Figure 2.
Observation 1: The regret for $j=2$ decreases rapidly in $\omega_{2}$ and then decreases more slowly. In the first column, we vary $\omega_{2}$ with the first experiment $j=1$ having a smaller experimental variance, $v_{1}=v_{2}/2$. The panel shows that a small $\omega_{2}$ makes the regret of choosing experiment $2$ larger than that of choosing experiment $1$. This is because a small $\omega_{2}$ corresponds to a small bias from not choosing $j=2$. As $\omega_{2}$ increases, the max–regret curve for $j=2$ declines until it reaches $\omega_{2}\approx 1.25$; after this point, the regret curve for $j=2$ rises slowly, since a further increase in $\omega_{2}$ is associated with an increase in estimator variance.
Observation 2: The regret for $j=1$ decreases slowly and then increases rapidly in $\omega_{2}$. The regret curve for choosing $j=1$ as we vary $\omega_{2}$ (fixing $\omega_{1}=1$) first decreases slowly and then increases. The reason is that for $\omega_{2}<\omega_{1}$, $\gamma_{1}^{\star}$ is an interior solution and the oracle squared bias is $\beta^{\star}=\omega_{2}^{2}$. Therefore, an increase in $\omega_{2}^{2}$ raises the oracle’s bias. When $\omega_{2}>\omega_{1}$, $\gamma_{1}^{\star}$ becomes a boundary solution and $\beta^{\star}=\omega_{1}^{2}$. In this case, a larger $\omega_{2}$ increases the variance of the estimator while the oracle bias $\beta^{\star}$ remains constant in $\omega_{2}$. As a result, it becomes more attractive for the analyst to run the experiment with $j=2$.
Observation 3: The regret for $j=2$ is monotonically increasing in $v_{2}^{2}$, and the opposite holds for $j=1$. The second plot shows that the regret from choosing $j=2$ increases monotonically with the experimental variance $v_{2}^{2}$, while the regret from choosing $j=1$ decreases correspondingly. The kinks in the curves are driven by regime shifts in the oracle solutions $(\beta^{\star},\alpha^{\star})$.
Observation 4: The regret for $j=2$ is monotonically decreasing in $\sigma_{2}^{2}$. The third plot shows that the regret for $j=2$ decreases monotonically with the observational variance $\sigma_{2}^{2}$. This is expected, since a larger $\sigma_{2}^{2}$ makes choosing $j=2$ more appealing relative to relying on observation.
Observation 5: The regret for $j=1$ first decreases slowly and then increases rapidly in $\sigma_{2}^{2}$. As we vary $\sigma_{2}^{2}$, the regret for $j=1$ is initially (very) slowly decreasing. The reason is that a larger $\sigma_{2}^{2}$ increases the oracle variance $\alpha^{\star}$ faster than the variance of the estimator for $j=1$ (with $\gamma_{1}^{\star}\approx 1$). However, this behavior does not affect the optimal solution: for smaller values of $\sigma_{2}^{2}$, choosing $j=1$ remains preferable to choosing $j=2$. When $\sigma_{2}^{2}$ is larger than a tipping point, however, the regret of choosing $j=1$ increases rapidly with $\sigma_{2}^{2}$, making the first experiment no longer preferable to the second. This aligns with the intuition that, as observational variance grows, we should favor conducting the experiment for $j=2$.
In sum, these patterns show how our framework disentangles the competing forces across signal strength and observational/experimental noise, guiding the analyst to transparent design choices even in complex regimes. Table 4 summarizes the discussion.
Table 4: Example: Qualitative behavior of maximum regret as $\omega_{2}$, $v_{2}$, or $\sigma_{2}$ increase (rows) in Figure 11. Entries summarize the direction and relative speed of change in the maximum regret for choosing experiment $j=1$ or $j=2$.
Small regime ($\omega_{2}<\omega_{1}$) Large regime ($\omega_{2}>\omega_{1}$)
Trend regret $j=1$ regret $j=2$ regret $j=1$ regret $j=2$
$\uparrow~\omega_{2}$ slow decrease fast decrease fast increase slow increase
$\uparrow~v_{2}$ decrease increase decrease increase
$\uparrow~\sigma_{2}$ slow decrease decrease fast increase decrease
Figure 11: Regret comparisons. Top row: $\max\{\alpha(j,\gamma_{j}^{\star})/\alpha^{\star},\beta(j,\gamma_{j}^{\star})/\beta^{\star}\}$ for $j\in\{1,2\}$ as the x–axis parameter varies (columns: $\omega_{2}$, $v_{2}$, $\sigma_{2}$). The vertical dashed line marks indifference, where the two curves intersect; to its left/right, the optimal experiment is the one with lower max regret. Bottom row: optimal weights $\gamma_{j}^{\star}$ for $j=1,2$, which explain how sensitivity and precision interact to drive the design switch at $x^{\star}$. Columns vary, respectively, (a) $\omega_{2}$ with $v_{1}=0.5$, $v_{2}=1$, $\sigma_{1}=\sigma_{2}=1$, $\omega_{1}=1$; (b) $v_{2}$ with $\omega=(0.9,1)$, $v_{1}=1$, $\sigma=(1,1)$; (c) $\sigma_{2}$ with $\omega=(0.9,1)$, $v=(1,1)$, $\sigma_{1}=1$.
Appendix D Additional figures and tables
Figure 12: Share of experimental papers published in AEA journals also presenting experimental results in combination with observational estimates (either from a reduced form, or from a structural model or from both).
Table 5: Optimal shrinkage and experiment choice across different regimes in the presence of two parameters (Setting 2) The first three regimes (row) corresponds to the case where the variance ratio $\alpha/\alpha^{\star}$ does not uniformly dominate $\beta/\beta^{\star}$ for all values of $\gamma_{j}$ and some $j\in\{1,2\}$ (which we view as leading cases). The other cases correspond to boundary solutions.
Case (condition) Optimal $\gamma_{j}^{\star}$ Optimal $j^{\star}$ solves Solution $(j^{\star},\gamma_{j}^{\star})$
Leading cases
Case 1 no bias/ variance dominance $\gamma_{j}^{\star}$ in the interior $\mathrm{arg}\min_{j}\left\{\omega_{-j}^{2}\sigma_{-j}^{2}+\omega_{j}^{2}\!\big[(1-\gamma_{j}^{\star})^{2}\sigma_{j}^{2}+(\gamma_{j}^{\star})^{2}v_{j}^{2}\big]\right\}$ $j^{\star}$ minimizes bias and variance
($\alpha(j,\gamma_{j})/\alpha^{\star}-\beta(j,\gamma_{j})/\beta^{\star}$ can flip sign) for $j\in\{1,2\}$ $\mathrm{arg}\min_{j}\left\{|\omega_{-j}|+|1-\gamma_{j}^{\star}||\omega_{j}|\right\}$ $\gamma_{j}^{\star}$ balances bias/variance
Case 2 $j=1$ is variance dominant $\gamma_{1}^{\star}=\frac{\sigma_{1}^{2}}{v_{1}^{2}+\sigma_{1}^{2}}$ $\mathrm{arg}\min_{j}\left\{\omega_{-j}^{2}\sigma_{-j}^{2}+\omega_{j}^{2}\!\big[(1-\gamma_{j}^{\star})^{2}\sigma_{j}^{2}+(\gamma_{j}^{\star})^{2}v_{j}^{2}\big]\right\}$ $j^{\star},\gamma_{1}^{\star}$ minimizes variance
$j=2$ has no bias/variance dominance $\gamma_{2}^{\star}$ is interior and $\gamma_{2}^{\star}$ balances bias/variance
Case 3 $j=1$ is bias dominant $\gamma_{1}^{\star}=1$ $\mathrm{arg}\min_{j}\left\{|\omega_{-j}|+|1-\gamma_{j}^{\star}||\omega_{j}|\right\}$ $j^{\star},\gamma_{1}^{\star}$ minimizes bias
$j=2$ has no bias/variance dominance $\gamma_{2}^{\star}$ is interior and $\gamma_{2}^{\star}$ balances bias/variance
Other (boundary) solutions
Case 4 Both $j\in\{1,2\}$ are variance-dominant $\gamma_{j}^{\star}=\frac{\sigma_{j}^{2}}{\sigma_{j}^{2}+v_{j}^{2}}$ $\mathrm{arg}\min_{j}\left\{\omega_{-j}^{2}\sigma_{-j}^{2}+\omega_{j}^{2}\frac{\sigma_{j}^{2}v_{j}^{2}}{\sigma_{j}^{2}+v_{j}^{2}}\right\}$ Minimizes variance
($\alpha/\alpha^{\star}>\beta/\beta^{\star}$ for all $\gamma$ and $j$) for $j\in\{1,2\}$
Case 5 Both $j\in\{1,2\}$ are bias-dominant $\gamma_{j}^{\star}=1$ $\mathrm{arg}\max_{j}|\omega_{j}|$ Minimizes bias
($\alpha/\alpha^{\star}<\beta/\beta^{\star}$ for all $\gamma$ and $j$) for $j\in\{1,2\}$
Case 6 bias dominant vs. variance dominant $\gamma_{1}^{\star}=1$ $\mathrm{arg}\min_{j}$ Minimizes bias $\beta/\beta^{\star}$ of $j=1$
($\alpha/\alpha^{\star}<\beta/\beta^{\star}$ for $j=1$ and vice versa for $j=2$) $\gamma_{2}^{\star}=\frac{\sigma_{2}^{2}}{\sigma_{2}^{2}+v_{2}^{2}}$ $\left\{\frac{|\omega_{2}|^{2}}{|\omega_{1}|^{2}}1\{j=1\}+(\omega_{1}^{2}\sigma_{1}^{2}+\omega_{2}^{2}\frac{\sigma_{2}^{2}v_{2}^{2}}{\sigma_{2}^{2}+v_{2}^{2}})1\{j=2\}\right\}$ vs. variance $\alpha/\alpha^{\star}$ of $j=2$
Figure 13: Optimal shrinkage $\gamma_{j}^{\star}$ by area as a function of total treated villages $n_{1}$ (from ten to fifty-two villages), under different constraints $E$. For $E<4$, the solution is interior ($0<\gamma_{j}^{\star}<1$) and increases with $n_{1}$, approaching one as experimental noise falls.