### v3

Abstract
Experiments deliver credible treatment-effect estimates but, because they are costly, are often restricted to specific sites, small populations, or particular mechanisms. A common practice across several fields is therefore to combine experimental estimates with reduced-form or structural external (observational) evidence to answer broader policy questions such as those involving general equilibrium effects or external validity. We develop a unified framework for the design of experiments when combined with external evidence, i.e., choosing which experiment(s) to run and how to allocate sample size under arbitrary budget constraints. Because observational evidence may suffer bias unknown ex-ante, we evaluate designs using a minimax proportional-regret criterion that compares any candidate design to an oracle that knows the observational study bias and jointly chooses the design and estimator. This yields a transparent bias-variance trade-off that does not require the researcher to specify a bias bound and relies only on information already needed for conventional power calculations. We illustrate the framework by (i) designing cash-transfer experiments aimed at estimating general equilibrium effects and (ii) optimizing site selection for microfinance interventions.
1 Introduction
Randomized controlled trials (RCTs) have improved empirical economics by providing internally valid estimates. However, in practice, feasibility constraints often confine trials to localized effects such as the effect in a specific site or subpopulations, or of a particular mechanism, as Muralidharan and Niehaus (2017) document. [^4] Experiments in top economics journals are in median “representative of a population of 10,885 units […] and randomized in clusters of 26 units per cluster” (Muralidharan and Niehaus, 2017).These effects, although useful, are often not sufficient to answer relevant policy questions about external validity and general equilibrium (GE) effects of at-scale interventions.
In response, a large literature in development (e.g., Meghir et al., 2022; Todd and Wolpin, 2006; Attanasio et al., 2012; de Albuquerque et al., 2025), education and labor economics (e.g., Allende et al., 2019; Larroucau et al., 2024; Chetty et al., 2016), and more recently industrial organization (e.g., Katz and Allcott, 2025; Allcott et al., 2025) combines experiments with external evidence—observational and/or results from experiments in other settings—to extrapolate policy-relevant counterfactuals that no single experiment can identify. In our review of AEA journals over the past decade, about one-third of experimental papers complement experimental with observational inputs, either reduced-form or structural (Figure 11). Since budget constraints often bind, such common practice raises the question of how to choose the most informative experiment(s) and design for policy-relevant estimands from a (typical large) menu of possible experiments (e.g. Ravallion, 2012; Niederle, 2025).
Our main contribution is a framework for choosing which experiment to run and how to design it in this common setting of combining experimental with external evidence. By external evidence we mean reduced-form or structural estimates based on observational data, and/or results from experiments in other contexts. External evidence may be biased in ways not known ex ante (e.g., due to confounding or lack of external validity). We provide a procedure that (1) helps researchers disentangle the role of each experiment in this setting, (2) selects which experiments to run subject to a budget (e.g., which treatment arms and/or sub-populations), (3) optimally allocates sample size, and (4) prescribes how to combine experimental and external evidence for estimation. We allow for arbitrary (user-specific) constraints in each step, for both the choice of the design and of the estimator. [^5] For (4), researchers may constrain the estimator to either only use experimental evidence when available or optimally combine experimental with (biased) observational evidence for estimating certain parameters.
For an illustrative example, consider a government piloting a cash-transfer program in a small set of districts to increase children’s school attendance. The trial delivers a local average effect, but policy decisions often require counterfactuals at scale, allowing prices and wages to adjust (e.g. Todd and Wolpin, 2006; Egger et al., 2022). Researchers therefore combine experimental evidence with a supply-demand model that leverages observational data to map local impacts into economy-wide outcomes. The design question is (i) which mechanisms are the most valuable to learn experimentally given constraints on the size of the experiment and (ii) how to allocate sample size across treatment arms and sites.
The first step is to specify the estimand and the underlying parameters. Define $\tau(\theta)$ the (policy-relevant) object of interest, where $\tau$ is a known smooth function and $\theta$ is a vector of unknown parameters; $\tau$ may be scalar or vector-valued. Researchers also observe external baseline estimates of $\theta$–henceforth observational estimates–which may be biased. Each feasible experiment can identify a subset of parameters without bias. This parameterization makes explicit which components of $\theta$ are learned experimentally and which may rely on (biased) observational evidence; not all components need or can be learned experimentally.
Returning to the cash-transfer example, $\tau(\theta)$ is the effect on schooling when all poor households in rural Kenya receive the subsidy. The entries of $\theta$ consist of (i) the direct schooling response to a conditional cash transfer (CCT), (ii) the income effect of transfers in partial equilibrium, and (iii) wage adjustments that shift the returns to schooling. Due to cost constraints, researchers may only run small, partial-equilibrium experiments and then combine these with prior evidence to recover $\tau(\theta)$. They may choose between a CCT arm to estimate the direct effect, or an unconditional-cash arm (UCT) to measure the income elasticity, and extrapolate the GE effects using evidence from a previous study in Mexico (Todd and Wolpin, 2006). (When site selection is also part of the design, $\theta$ additionally collects site-specific effects and $\tau(\theta)$ targets the average effect across sites.)
The second step is to select the experiment(s), sample size and estimator. A natural starting point would be to design the experiment to minimize the mean-squared error (MSE) for $\tau(\theta)$. However, in practice, the bias is unknown prior to the experiment. A worst-case approach would be overly-conservative and disregard information about the variance. We instead adapt the definition of adaptation regret, previously studied for combining two estimators under a fixed design (Armstrong et al., 2024; Tsybakov, 1998), to our experimental design problem (with multiple parameters). Here, the regret is the ratio of the researcher’s worst-case mean-squared error relative to an oracle that knows the worst-case bound on the observational bias and chooses both the design and the estimator. Taking its supremum over the worst-case bias yields a robust procedure without requiring a prior about the bias or on its bound. However, our task of also choosing the experimental design and the lack of an unbiased estimate for $\tau(\theta)$ when not all parameters can be learned experimentally make existing solutions not applicable.
Our main result provides an explicit and novel characterization of adaptation regret for (asymptotically) linear estimators in experimental and observational moments. These include generalized methods of moments where we may also optimize over the weighting matrix. For any design and weighting matrix that combine the experimental and observational point estimates, the regret is the maximum of two normalized components. The first is a variance regret: the estimator’s sampling variance under the chosen design and weights, divided by the smallest attainable variance. The second is a bias regret: the worst-case squared bias induced by using external evidence, divided by the smallest feasible worst-case bias over the class of designs. The variance regret depends on the (expected) variance-covariance matrix of the observational and experimental estimates implied by the design, the weighting matrix, and the sensitivity of the policy parameter, i.e., the gradient of $\tau$ with respect to the parameters evaluated, under mild conditions, at the observational estimates. [^6] For this result to hold, we assume that $\tau(\theta)$ is twice differentiable with bounded derivatives. For non-linear $\tau(\theta)$ in $\theta$, we also require that the observational estimates bias is local to zero in the spirit of Andrews et al. (2020); Armstrong and Kolesár (2021); Bonhomme and Weidner (2022), where its rate of convergence can be the same, faster, or even slower than the estimators’ standard error. Section 5.3 provides details. The bias regret depends only on the sensitivity vector and the weighting matrix, known ex-ante.
Our characterization makes the bias-variance trade-off intuitive. At the optimum, the two normalized components tend to be equalized: when the bias dominates, the design prioritizes high-sensitivity parameters; when variance dominates, the procedure invests sample where it most reduces variance. This yields a nested procedure that jointly selects the estimator, the sample allocation and the experiment(s) to run, trading off precision with bias sensitivity. The resulting design and estimator only require knowledge of the expected variance-covariance matrix and observational estimates. It can be reported in a pre-analysis plan. That is, for given observational estimates, the inputs are the same as those required by standard power-analysis plans for $\tau(\theta)$ (e.g. Gerber and Green, 2012).
We then generalize our framework to general moment selection problems that combine experimentally identified moments with observational (misspecified) moments. With parameters defined by non-linear but smooth moments our results directly apply under standard local asymptotics, where the bias is local to zero at same or possibly different rate than the standard error (e.g. Andrews et al., 2017, 2020). We also generalize to partially identified $\tau$ and settings where researchers have priors (confidence sets) about the bias.
From a theoretical (and computational) perspective, our key insight is that, as we restrict the class of estimators to be (asymptotically) linear in the moments, the worst-case adaptation regret is quasi-convex in the radius of the bias, for any norm characterizing the bias’ uncertainty set. This result yields a transparent variance-bias trade-off and differs from existing solutions under the fixed-design optimal estimators as in Armstrong et al. (2024), Tsybakov (1998); there, quasi-convexity fails due to non-linearity of the estimator, and optimization corresponds to a minimax program over worst-case priors. In our joint design and estimation problem, quasi-convexity substantially simplifies optimization which is quadratic for the choice of the estimator, with a closed form expression for sample allocation and a mixed-integer formulation for the experiment choice only (typically from a finite set). This program is solved via off-the-shelf mixed-integer quadratic programming.
We illustrate the framework in a cash‐transfer application in Kenya, where the researcher combines new experimental evidence with external evidence from Mexico’s PROGRESA program (Todd and Wolpin, 2006; Attanasio et al., 2012), where these preliminary estimates may lack external validity. We compare three designs and their combinations: (i) a CCT arm to estimate direct schooling effects; (ii) a UCT arm to estimate income effects; and (iii) an employment program that identifies individual wage elasticities. Experimentation costs are binding, as cash transfers and employment programs require distinct infrastructures and implementation teams. We show that the wage experiment is the most informative for sufficiently large sample sizes, while for smaller samples a combination of CCT and UCT—allocating most of the sample to the CCT arm—is optimal. With one thousand participants, relative to the variance-minimizing (Neyman) design, our optimal two-arm design delivers more than a 400% reduction in bias (and adaptation regret) with variance no larger than $\sim$30% of the Neyman variance; for a single-arm design, our procedure yields a 140% reduction in bias at the cost of only about a 12% increase in variance. These comparisons illustrate the benefit of our method in balancing robustness and precision.
As a second application, we study where to run a microfinance experiment in India, integrating evidence from earlier nonrandomized microfinance introductions. We target the average effect in the entire region, but constraint the researcher to choose one or two smaller areas for randomization due to coordination (fixed) costs constraints. We use observational estimates in Banerjee et al. (2024) to select one or more areas for randomization. Because Banerjee et al. (2024) also implement a separate randomized expansion, we can calibrate performance of our design (agnostic of the bias) by comparing each candidate design’s MSE to that of an oracle that knows the bias, using the experimental estimates to calibrate the bias. Compared to the variance-optimal benchmark, we reduce MSE by more than 250%.
Related literature
This paper links the experimental-design literature—which mostly focuses on settings where all parameters are identified within the experiment and leaves aside questions of data combination—to recent work that integrates experimental and (reduced-form or structural) observational evidence to extrapolate effects in complex scenarios.
Recent advances for experimental design include balancing and variance-minimizing allocations (Tabord-Meehan, 2018; Bai, 2019; Cytrynbaum, 2021; Kallus, 2018; Bertsimas et al., 2015), adaptive designs for policy choice (Kasy and Sautmann, 2019; Russo et al., 2018; Cesa-Bianchi et al., 2025), and experimental design under correct model specification (Silvey, 2013; Chaudhuri and Mykland, 1993; Chaloner and Verdinelli, 1995; Higbee, 2024; Kiefer and Wolfowitz, 1959; Reeves et al., 2024; Viviano, 2020; Kasy, 2016). Traditional work on experimental design for robust-model estimation either focused on testing competing models (Atkinson and Fedorov, 1975; López-Fidalgo et al., 2007), or on using a-priori knowledge of (worst-case) bias for the design of an experiment (Box and Draper, 1959; Wiens, 1998; Tsirpitzi et al., 2023; Sacks and Ylvisaker, 1984). All these references leave aside questions about observational data combination. In the context of minimizing the variance of an experiment, Rosenman and Owen (2021) proposes to use observational studies to construct high-confidence bounds on the experimental variance, and then focus on variance-optimal stratification. This problem differs from our design problem and analysis, which instead studies the use of observational studies in combination with experiments for extrapolation.
In summary, we complement this literature by studying which experiment to run (and how) when not all parameters are learned experimentally, a common problem in economics.
Our minimax regret criterion connects to a long-standing decision-theoretic tradition for experimental design. References include Manski and Tetenov (2016), Banerjee et al. (2020), Manski (2004), Dominitz and F. Manski (2017), Olea et al. (2024), Hu et al. (2024), Breza et al. (2025) among others. These references focus on settings where researchers only have access to experimental variation, instead of combining it with external (observational) evidence, motivating different designs and objective functions. We also complement literature that leverages correctly-specified models for site-selection (Gechter et al., 2024; Abadie and Zhao, 2021) by allowing for misspecification in observational estimates.
A related line of work analyzes estimation under misspecification (Armstrong et al., 2024; Armstrong and Kolesár, 2018; Andrews et al., 2025; Andrews et al., 2020; Bonhomme and Weidner, 2022; Donoho, 1994; Kwon and Sun, 2025), and combining existing experiments with observational studies (Gechter, 2022; Athey et al., 2025, 2020; Kallus et al., 2018; Dutz et al., 2021; Rosenman et al., 2022; de Chaisemartin and D’Haultfœuille, 2020; Rambachan et al., 2024). We provide a novel characterization of adaptation regret as a quasi-convex function in the bias bound (and therefore as the maximum of the variance and bias regret) by leveraging linearity of the estimator. Such characterization shows how, in our problem, misspecification is captured through the first-order sensitivity of the estimand to each parameter, an important measure for sensitivity analysis in Andrews et al. (2020); it provides a decision-theoretic foundation for simultaneously protecting against a correctly specified and misspecified model (e.g. Bickel, 1984; Kempthorne, 1988), here through the maximum of a normalized variance and bias components. Unlike the references above, where the design is fixed and researchers optimize over the choice of the estimator, our main contribution is to optimize the design itself. This changes the optimization and regret, computed against the best design-estimator.
2 Baseline scenario: problem description
We will focus first on a simple shrinkage estimator and a univariate estimand. Our results generalize directly to general methods of moments and multivariate targets in Section 4.
Consider a researcher interested in an arbitrary target estimand $\tau(\theta)\in\mathbb{R}$, indexed by a low-dimensional unknown parameter vector $\theta\in\mathbb{R}^{p}$ and a known mapping $\tau$. The goal is to construct an estimator $\hat{\tau}$ that accurately approximates $\tau(\theta)$ by combining existing evidence (e.g., observational studies) with experimental variation designed by the researcher. Our question is how to design such experiments under feasibility constraints.
Basic setup
We assume that researchers have access to estimators (and their variance) of $\theta$ denoted as $\tilde{\theta}^{\mathrm{obs}}\in\mathbb{R}^{p}$. We impose no restrictions on how $\tilde{\theta}^{\mathrm{obs}}$ is formed: it can based on arbitrary exclusion restrictions implied by an economic or statistical model. However, $\tilde{\theta}^{\mathrm{obs}}$, defined as observational estimates, have biases collected in a vector $b\in\mathbb{R}^{p}$ unknown at the experimental design stage (see Remark 2). Examples include an estimate from an instrumental variable regression that may fail the exclusion restriction, or an experimental estimate from a country different from the one of interest that may lack external validity. Researchers can then collect experimental evidence for a subset of parameters, producing estimates which are unbiased by design.
Setting 1.
For a subset $\mathcal{E}\subseteq\{1,\ldots,p\}$ and a known positive-definite matrix $\Sigma(\mathcal{E})\in\mathbb{R}^{(p+|\mathcal{E}|)\times(p+|\mathcal{E}|)}$, define $\tilde{\theta}^{\mathrm{obs}}\in\mathbb{R}^{p}$ an observational estimate and $\tilde{\theta}^{\mathrm{exp}}_{\mathcal{E},\Sigma}\in\mathbb{R}^{|\mathcal{E}|}$ an experimental estimate each satisfying
$\mathbb{E}\!\left[\tilde{\theta}^{\mathrm{obs}}\right]-\theta=b,\qquad\mathbb{E}\!\left[\tilde{\theta}^{\mathrm{exp}}_{\mathcal{E},\Sigma}\right]-\theta_{\mathcal{E}}=0,$
for an unknown bias vector $b\in\mathbb{R}^{p}$. Moreover, define its joint variance-covariance matrix as $\mathbb{V}\!\begin{pmatrix}\tilde{\theta}^{\mathrm{obs}}-\theta\\[2.0pt] \tilde{\theta}^{\mathrm{exp}}_{\mathcal{E},\Sigma}-\theta_{\mathcal{E}}\end{pmatrix}=\Sigma(\mathcal{E}).$ Given $(\mathcal{E},\Sigma)$, consider a class of linear plug-in estimators
$\hat{\tau}_{\gamma}\equiv\tau(\hat{\theta}(\gamma)),\qquad\hat{\theta}_{j}(\gamma)=\begin{cases}\gamma_{j}\,\tilde{\theta}_{j}^{\mathrm{exp}}+(1-\gamma_{j})\,\tilde{\theta}_{j}^{\mathrm{obs}},&j\in\mathcal{E},\\[2.0pt] \tilde{\theta}_{j}^{\mathrm{obs}},&j\notin\mathcal{E},\end{cases}$
where $\gamma=(\gamma_{j})_{j\in\mathcal{E}}\in\mathbb{R}^{|\mathcal{E}|}$ are shrinkage weights that can be an (implicit) function of $(\mathcal{E},\Sigma)$.
Here, $(\mathcal{E},\Sigma(\mathcal{E}))$ characterizes the experimental design, encoding which parameters are learned and with what precision. When clear, we omit the subscript $\Sigma$ on $\tilde{\theta}^{\mathrm{exp}}$. We assume $\Sigma(\mathcal{E})$ is known, which is standard in experimental planning (Gerber and Green, 2012); when $\Sigma$ is unknown, we can interpret $\Sigma$ as the expected variance-covariance matrix. [^7] In this case, we interpret the MSE in Equation (2) as the expected MSE under a common expectation over the experimental variance-covariance matrix. It is also possible to extend our framework by allowing $\Sigma$ to be a function of $\tilde{\theta}^{\mathrm{obs}}$ provided that $\Sigma(\tilde{\theta}^{\mathrm{obs}})\rightarrow_{p}\Sigma(\mathrm{plim}(\tilde{\theta}^{\mathrm{obs}}))$, with $\Sigma(\mathrm{plim}(\tilde{\theta}^{\mathrm{obs}}))$ denoting the common expectation of the variance-covariance experimental matrix. In this case our analysis should be interpreted in an asymptotic regime similar to what discussed in Section 5.3.
The class of experiments and estimators is assumed to be in a user-specific set $\mathcal{D}$.
Assumption 1 (Feasible experiments and estimators).
We write $(\mathcal{E},\Sigma,\gamma)\in\mathcal{D}$ indicating $\mathcal{D}$ the set of feasible designs (i.e., choice of $\mathcal{E}$ and $\Sigma(\mathcal{E})$) and estimators (i.e., $\gamma$). The set $\mathcal{D}$ can be arbitrary as long as (i) each feasible $\Sigma$ has uniformly bounded entries and is strictly positive definite; (ii) $\gamma_{j}=1$ for all $j\in\mathcal{E}$ is one admissible choice (among potentially others).
We let $(\mathcal{E},\Sigma(\mathcal{E}))$ be within arbitrary constraint sets that may encode feasibility or budget constraints. In our framework, restrictions on the experiments researchers can run $\mathcal{E}$ and on their precision $\Sigma(\mathcal{E})$ can take any desired form. In most applications, we think of such constraints as arising from fixed costs of running an experiment and/or constraints on their power (Duflo et al., 2007; Athey and Imbens, 2017; List et al., 2011). We assume that $\Sigma(\mathcal{E})$ is uniformly bounded, which implies that once we commit to learn a set of experimental parameters $\tilde{\theta}_{\mathcal{E}}^{\mathrm{exp}}$, their variance is finite, and strictly positive definite, therefore assuming that the variance is bounded away from zero. [^8] For this latter condition to hold in an asymptotic framework with growing sample size, $\theta$ and $\tilde{\theta}$ can be defined without loss as parameters of interest after appropriately rescaling by the square-root of the sample size; in this case $\Sigma$ denotes the asymptotic variance. See Section 5.3 for details. The shrinkage estimator can be unrestricted (i.e., $\gamma$ can take any value on the real line) or restricted (when for example researchers aim to only use experimental evidence when available, setting $\gamma_{j}=1$ for $j\in\mathcal{E}$).
Example 1 (Structure of $\Sigma$).
Because observational and experimental moments are typically computed from disjoint samples, in practice, it is natural to model the covariance as block diagonal, $\Sigma=\operatorname{diag}(\Sigma_{\text{obs}},\Sigma_{\text{exp}})$. The first block corresponds to observational estimates (fixed and not subject to design choices), while the second corresponds to experimental estimates, and is a function of the sample size allocated to treatment arms (that we may optimize). This is a special case of our framework which can be encoded in the constraint set $\mathcal{D}$. ∎
Remark 1 (Some unbiased observational estimates).
In some applications, researchers may assume that some of the observational study estimates are not biased ($b_{j}=0$ for some $j$). This is possible to accommodate as a special case of our general framework in Section 4. ∎
Main design problem
Our design problem can be described in three steps:
  1. 1.
    Preliminary step: weights. For each feasible $\mathcal{E},\Sigma(\mathcal{E})$, choose $\gamma$ to combine experimental and observational estimates on the selected coordinates. (If researchers restrict their estimators to only use experimental evidence when available, set $\gamma=1$.)
  2. 2.
    Middle step: precision. For each feasible $\mathcal{E}$, choose $\Sigma(\mathcal{E})$ (i.e., sample allocation).
  3. 3.
    Outer step: experiment choice. Choose a feasible set of experiments $\mathcal{E}$.
Choosing $\gamma$ (step 1) for a fixed design follows a long-standing tradition in econometrics and statistics (Armstrong et al., 2024; Andrews and Shapiro, 2021; Donoho, 1994; Athey et al., 2025). We study this question here in combination with the experimental design problem, step 2 and 3, which, as we show will change the decision problem. The challenge is that performance depends on the bias $b$, unknown when designing the experiment.
Properties of $\tau$
A key feature of our framework is that researchers explicitly parameterize $\tau(\cdot)$, thereby pre-specifying and making transparent which biases drive the design problem; this feature clarifies the role of the experiment in complementing existing evidence. The sensitivity of $\tau$ to $\theta$, captured by its gradient $\omega(\theta)=\partial\tau(\theta)/\partial\theta$, determines how (to first order) the bias propagates onto the final estimand. For simplicity, we impose exact linearity, $\tau(\theta)=\omega^{\top}\theta$; all results extend to smooth, nonlinear $\tau$ via a first-order expansion. When $\tau$ is unknown, Section 5.1 studies settings where researchers only partially identify $\tau$.
Assumption 2 (First-order estimation error).
For any $(\mathcal{E},\Sigma,\gamma)\in\mathcal{D}$, assume
$\tau(\theta)-\tau\!\big(\hat{\theta}(\gamma)\big)=\sum_{j=1}^{p}\omega_{j}\,\big(\theta_{j}-\hat{\theta}_{j}(\gamma)\big),$ (1)
for known weights $\omega\in\mathbb{R}^{p}$, with $|\omega_{j}|\in(0,\infty)$ for all $j$.
[…]
We conclude with three examples in simple two-parameter models.
Example 2 (Choosing the site for an experiment for external validity).
Gechter et al. (2024) study where to run an experiment. For illustration, consider two sites $j\in\{1,2\}$ with site-specific ATEs $\theta_{j}$ and target the cross-site average $\tau(\theta)=\omega_{1}\theta_{1}+\omega_{2}\theta_{2},$ with $\omega$ denoting the population share in each site. Let $\tilde{\theta}^{\mathrm{obs}}=(\tilde{\theta}^{\mathrm{obs}}_{1},\tilde{\theta}^{\mathrm{obs}}_{2})^{\top}$ denote observational estimates obtained in Gechter et al. (2024) from a structural model and potentially biased due to misspecification: $\mathbb{E}[\tilde{\theta}^{\mathrm{obs}}]-\theta=(b_{1},b_{2})^{\top}.$ A budget constraint allows an experiment in only one site, so that $\mathcal{E}\in\{\{1\},\{2\}\}$. If site $j$ is chosen, we obtain an unbiased $\tilde{\theta}^{\mathrm{exp}}_{j}$; the other site $k\neq j$ remains observational. Given $\mathcal{E}$,
$\widehat{\tau}_{\gamma}=\omega_{j}\gamma_{j}\,\tilde{\theta}^{\mathrm{exp}}_{j}+\omega_{j}(1-\gamma_{j})\,\tilde{\theta}^{\mathrm{obs}}_{j}+\omega_{k}\tilde{\theta}^{\mathrm{obs}}_{k},\qquad k\neq j.$
Our question is how to choose the set $\{j\}$ where to conduct the experiment (jointly with $\gamma$).
Example 3 (Choosing which survey to conduct).
Egger et al. (2022) study the efficacy of cash-transfer programs on the marginal propensity to consume (MPC). Measuring the MPC requires capturing both short- and long-run effects. Because survey rounds are limited, the authors complement experimental data that lacks short-run effects with auxiliary information from prior studies that use short-run surveys (collected in other regions; see Egger et al. (2022)). This raises the question of which survey to conduct (and with which frequency).
In stylized form, suppose researchers can observe, for $s\in\{1,2\}$, potential outcomes $Y_{s}(t)$ denoting consumption in period $s$ when measured $t$ periods after the intervention. The authors have auxiliary estimates from previous studies,
$\alpha\;=\;\mathbb{E}\!\big[Y_{1}(t=1)-Y_{1}(t=\infty)\big],\qquad\beta\;=\;\mathbb{E}\!\big[Y_{2}(t=1)-Y_{2}(t=\infty)\big],$
and wish to estimate the total effect $\tau=\alpha+\beta$. Researchers consider two survey designs:
  • (i)
    Early survey design to estimate $\alpha$ precisely;
  • (ii)
    Later survey to estimate $\beta$ precisely.
Our goal is to study which survey design to implement. (More complex designs with additional time periods or mixed precision across rounds are also possible here.)
Example 4 (Experiment with non linear target and equilibrium model).
Bergquist and Dinerstein (2020) conduct demand and supply experiments in food markets in Kenya. Suppose here we are interested in similar applications in Uganda. For exposition, consider a stylized linear demand and supply model (similar reasoning applies to more complex models)
$Q^{D}=a-\beta_{D}P+u_{D},\qquad Q^{S}=c+\beta_{S}P+u_{S},\qquad\theta=(\beta_{D},\beta_{S})^{\top},$
with $\beta_{D}\neq-\beta_{S}$. Several estimands may be of interest; one of such estimands is the effect of a tariff $t$ on prices $\tau\equiv\frac{\beta_{S}}{\beta_{D}+\beta_{S}}\,t.$ Let $(\tilde{\beta}_{D}^{\mathrm{obs}},\tilde{\beta}_{S}^{\mathrm{obs}})^{\top}$ denote baseline estimates from Kenya, that may lack external validity.
Due to fixed costs of each experiment, $\mathcal{E}\in\{\{D\},\{S\}\},$ i.e., we can learn either demand (estimating $\tilde{\beta}_{D}^{\mathrm{exp}}$, with randomized price discounts) or supply (estimating $\tilde{\beta}_{S}^{\mathrm{exp}}$, by introducing regulatory cost shocks on firms). Let
$\big(\tilde{\theta}^{\mathrm{obs}}-\theta\big)\ \sim\ \mathcal{N}\big(b,\Sigma^{\mathrm{obs}}\big),\quad\tilde{\theta}^{\mathrm{obs}}\equiv\sqrt{n}(\tilde{\beta}_{D}^{\mathrm{obs}},\tilde{\beta}_{S}^{\mathrm{obs}})^{\top},\quad\theta\equiv\sqrt{n}(\beta_{D},\beta_{S})^{\top}$
with $\theta$ and $\tilde{\theta}$ denoting the parameter and estimator here rescaled by the square-root of the sample size, and $b$ capturing bias. Our goal is to choose whether to conduct a supply or demand experiment in Uganda to estimate $\tau$. For given estimators $\hat{\theta}$ that combine the estimates from Kenya with the chosen experimental estimate from Uganda under local asymptotics described in Section 5.3 and Remark 6
$\tau-\widehat{\tau}=\underbrace{\frac{1}{\sqrt{n}}\omega^{\top}(\theta-\widehat{\theta})}_{\text{main first order effect}}+\underbrace{o_{p}\left(\frac{1}{n^{1/2}}\right)}_{\text{small higher order effects}},\hskip 18.49988pt\omega=\frac{t}{(\mathrm{plim}(\tilde{\beta}_{D}^{\mathrm{obs}}+\tilde{\beta}_{S}^{\mathrm{obs}}))^{2}}\begin{pmatrix}-\mathrm{plim}(\tilde{\beta}_{S}^{\mathrm{obs}})\\ \mathrm{plim}(\tilde{\beta}_{D}^{\mathrm{obs}})\end{pmatrix}$,
where we can replace $\mathrm{plim}(\tilde{\beta}^{\mathrm{obs}})$ with its consistent counterpart $\tilde{\beta}^{\mathrm{obs}}$.
3 Robust experimental design
Next, we introduce an experimental design focusing on the mean-squared error (MSE) of the estimator $\hat{\tau}$. We consider confidence interval length in Section 5.1. The MSE is defined, for given bias level $b$, design $(\mathcal{E},\Sigma(\mathcal{E}))$, and choice of weights $\gamma$ as
$\mathrm{MSE}_{b}(\mathcal{E},\Sigma,\gamma)\;=\;\mathbb{E}_{\mathcal{E},\Sigma,b}\!\left[(\widehat{\tau}_{\gamma}-\tau)^{2}\right],$ (2)
where $\mathbb{E}_{\mathcal{E},\Sigma,b}$ denotes expectation under the design $(\mathcal{E},\Sigma(\mathcal{E}))$ and observational bias $b$.
Ideally, one would minimize $\mathrm{MSE}_{b}$. However, because $b$ is unknown, we consider an uncertainty set given by an $\ell_{\infty}$-ball,
$\mathcal{B}(B)\equiv\{b:\ \|b\|_{\infty}\leq B\},$ (3)
with $B\geq 0$ an upper bound on the largest coordinate-wise bias. In our framework, a key challenge is that biases may arise from multiple parameters. Here, for exposition we first focus on the $\ell_{\infty}$-ball given its interpretability, while Section 4 generalizes the framework to arbitrary norms. The $\ell_{\infty}$ norm imposes symmetric restrictions on the biases and it does not force a trade-off across coordinates (a large bias in one component need not be “offset” by a small bias elsewhere), which is attractive when biases may be positively correlated across parameters. The drawback is that worst-case solutions depend on the radius $B$ which may be unknown in practice.
If an oracle knew $B$, a natural choice would be to pick the minimax design
$\mathrm{MSE}^{*}(B)\;\equiv\;\inf_{(\mathcal{E},\Sigma,\gamma)\in\mathcal{D}}\ \sup_{b\in\mathcal{B}(B)}\ \mathrm{MSE}_{b}(\mathcal{E},\Sigma,\gamma).$
However, having to specify $B$ can pose a large burden on the researchers and make the choice of the experiment sensitive to $B$. Therefore, we seek designs that perform as close as possible to the oracle that knows $B$, uniformly over the values of $B$, following a long-standing tradition in decision theory (e.g. Manski, 2004; Manski and Tetenov, 2007; Montiel Olea et al., 2023; Kitagawa and Tetenov, 2018). Specifically, we minimize the worst-case proportional regret
$\mathcal{R}(\mathcal{E},\Sigma,\gamma)\;\equiv\;\sup_{B\geq 0}\ \frac{\displaystyle\sup_{b\in\mathcal{B}(B)}\mathrm{MSE}_{b}(\mathcal{E},\Sigma,\gamma)}{\displaystyle\mathrm{MSE}^{*}(B)},$ (4)
defined adaptation regret by Armstrong et al. (2024) and Tsybakov (1998) for estimation under fixed designs (see Remark 3 for a comparison with our experimental design problem).
Remark 2 (Adaptation to $B$).
We consider worst-case adaptation regret over $B$ because researchers typically do not know $B$ before running an experiment and learning $B$ would require running pilots for each design in the menu of feasible experiments. Also, Armstrong and Kolesár (2018) show that pointwise adaptation to unknown misspecification levels is impossible in general, motivating the adaptation regret. Note however that if researchers know that the misspecification level $B$ lies in a bounded interval $B\in[0,\bar{B}]$, our characterization extends by simply restricting the worst-case over $B$ to this interval; see Section 5.2.∎
3.1 Optimal design
Our main results show that the adaptation regret provides natural trade-offs between the variance and the bias. To do so, denote $\mathbb{V}_{\mathcal{E},\Sigma}(\cdot)$ the variance operator for given $(\mathcal{E},\Sigma)$.
Definition 1 (Variance regret).
Let
$\alpha(\mathcal{E},\Sigma,\gamma)\;\equiv\;\mathbb{V}_{\mathcal{E},\Sigma}(\hat{\tau}_{\gamma}),\qquad\alpha^{\star}\;\equiv\;\min_{(\mathcal{E},\Sigma,\gamma)\in\mathcal{D}}\alpha(\mathcal{E},\Sigma,\gamma),$ (5)
with $\alpha^{\star}$ denoting the smallest feasible variance. We will refer to $\alpha/\alpha^{\star}$ as the variance regret.
The variance regret denotes the ratio between the variance of a given design and estimator relative to the smallest achievable variance. We provide a similar definition for the bias.
Definition 2 (Bias regret).
Define
$\qquad\beta(\mathcal{E},\gamma)\;\equiv\;\Big(\,\|\omega\|_{1}-\|\omega_{\,\mathcal{E}}\|_{1}+\,|1-\gamma|^{\top}|\omega_{\mathcal{E}}|\,\Big)^{\!2},\quad\beta^{\star}\;\equiv\;\min_{\mathcal{E}\in\mathcal{S}}\Big(\|\omega\|_{1}-\|\omega_{\mathcal{E}}\|_{1}\Big)^{2},$
with $\beta/\beta^{\star}$ defined as the bias regret.
To gain insight, note that $B^{2}\beta=B^{2}\Big(\|\omega\|_{1}-\|\omega_{\mathcal{E}}\|_{1}\Big)^{2}$ denotes the worst-case bias for a given design, value $B^{2}$ and shrinkage estimator. $B^{2}\beta^{\star}$ denotes its minimizer.
An ideal design would set the variance equal to $\alpha^{\star}$ and bias equal to $\beta^{\star}$. Unfortunately, this may be infeasible as it might require different choices of experiments and estimators to achieve one or the other. This raises the question of how to trade-off these two components.
This trade-off is formally characterized in our main theorem below. We will use the convention throughout that 0/0 = 1.
Theorem 1.
Consider Setting 1 and let Assumptions  1, 2 hold. Then, for any $(\mathcal{E},\Sigma,\gamma)\in\mathcal{D}$,
$\mathcal{R}(\mathcal{E},\Sigma,\gamma)\;=\;\max\!\left\{\frac{\alpha(\mathcal{E},\Sigma,\gamma)}{\alpha^{\star}},\ \ \frac{\beta(\mathcal{E},\gamma)}{\beta^{\star}}\right\}.$
Proof.
See Appendix A.1. ∎
Theorem 1 offers a simple and novel expression of the regret which does not require researchers to specify $B$. As we show below, Theorem 1 will significantly simplify the optimization program. The key insight is that the adaptation regret for linear estimators and arbitrary designs leads to a (strictly) quasi-convex objective function, whose worst-case solution is the maximum between the variance and the bias regret. This provides a simple and intuitive trade-off between the two. Specifically, provided that the class of designs $\mathcal{D}$ is sufficiently flexible (outside boundary solutions), it follows that at the optimum we equalize
$\frac{\alpha}{\alpha^{\star}}\;=\;\frac{\beta}{\beta^{\star}}.$
That is, we allocate sample size to noisier treatment arms when the variance dominates, and select treatment arms with largest sensitivity when the bias dominates.
[…]
Figure 1: Each feasible design $(\mathcal{E},\Sigma,\gamma)$ maps to a point $\big(\alpha/\alpha^{\star},\ \beta/\beta^{\star}\big)$: the xx–axis is the variance ratio and the yy–axis is the worst–case bias ratio. The blue curve depicts the attainable frontier as we vary shrinkage $\gamma$ and precision $\Sigma$. Level sets of the objective $\mathcal{R}=\max\{\alpha/\alpha^{\star},\ \beta/\beta^{\star}\}$ are axis–aligned squares (the dotted inverted “L” shows the smallest such square touching the frontier). The minimizer outside boundary solutions is where the frontier meets the $45^{\circ}$ line (red dot).
Remark 3 (Comparison with previous literature on adaptation regret).
Previous literature has studied adaptation regret to combine an unbiased and a biased estimator for a single parameter $\theta\in\mathbb{R}$ under a fixed design (Armstrong et al., 2024; Tsybakov, 1998). There are three key distinctions from this literature. First (and foremost), we optimize over the design itself rather than taking it as given; regret is defined relative to an oracle that chooses both the estimator and the design ex ante. Second, we allow biases to arise across multiple parameters (rather than a single one), and do not require that researchers have access to an unbiased estimator for each component of $\theta$. Third, we study a linear class of estimators, instead of arbitrary nonlinear rules. [^10] In the univariate bounded-normal-mean problem, Donoho et al. (1990) show that linear estimators lose no more than 25% efficiency relative to the optimal nonlinear estimator.
As a consequence, the characterization in Theorem 1 differs from Armstrong et al. (2024), where regret is not quasi-convex in the bias radius and the optimal estimator solves a minimax problem via a least-favourable prior; see Appendix D. Their minimax solution for combining two estimators is not applicable here once we optimize over the choice of experiments, sample allocation and multiple estimators, since searching over nonlinear shrinkage rules becomes computationally demanding. By focusing on linear estimators, we obtain a tractable solution that also extends to common method-of-moments, see Section 4. ∎
Example 5 (Intuition with two parameters).
To build further intuition, it is useful to study the two-parameter model. Let $\theta=(\theta_{1},\theta_{2})^{\top}$, and mutually independent observational and experimental estimates $\tilde{\theta}^{\mathrm{obs}}_{j}-\theta_{j}\sim\mathcal{N}(b_{j},\ \sigma_{j}^{2}),\quad\tilde{\theta}^{\exp}_{j}-\theta_{j}\sim\mathcal{N}(0,\ v_{j}^{2}),\quad j=1,2.$ For simplicity $\sigma_{j}^{2},v_{j}^{2}$ are fixed. We may run one experiment due to budget constraints: $\mathcal{E}\in\{\{1\},\{2\}\}$. If we pick $j$, we estimate $\hat{\theta}_{j}=\gamma_{j}\,\tilde{\theta}^{\exp}_{j}+(1-\gamma_{j})\,\tilde{\theta}^{\mathrm{obs}}_{j},\quad\hat{\theta}_{-j}=\tilde{\theta}^{\mathrm{obs}}_{-j},$ and, to first order, $\tau(\theta)-\tau(\hat{\theta})=\omega_{1}(\theta_{1}-\hat{\theta}_{1})+\omega_{2}(\theta_{2}-\hat{\theta}_{2}).$ We can write
$\alpha^{\star}=\min_{k\in\{1,2\}}\left\{\ \omega_{-k}^{2}\sigma_{-k}^{2}\;+\;\omega_{k}^{2}\frac{\sigma_{k}^{2}\,v_{k}^{2}}{\sigma_{k}^{2}+v_{k}^{2}}\ \right\},\qquad\beta^{\star}=\min_{k\in\{1,2\}}\omega_{k}^{2}.$
Here, $\alpha^{\star}$ follows from the variance-only optimal shrinkage $\gamma_{k}=\sigma_{k}^{2}/(\sigma_{k}^{2}+v_{k}^{2})$ (i.e., the oracle choice when $B=0$); $\beta^{\star}$ equals the smallest between the two sensitivity parameters $\omega$. Both depend on the choice of the estimator and on the class of experiments.
We cannot achieve both $\alpha^{\star}$ and $\beta^{\star}$, since these require different choices of $\gamma$. Instead,
$\alpha(j,\gamma_{j})$ $\equiv\underbrace{\omega_{-j}^{2}\sigma_{-j}^{2}}_{\text{var obs estimate}}+\underbrace{\omega_{j}^{2}\!\big[(1-\gamma_{j})^{2}\sigma_{j}^{2}+\gamma_{j}^{2}v_{j}^{2}\big]}_{\text{var exp and obs estimate}},\quad\beta(j,\gamma_{j})$ $\equiv\big(\underbrace{|\omega_{-j}|}_{\begin{subarray}{c}\text{bias obs}\\ \text{estimate}/B\end{subarray}}+\underbrace{(1-\gamma_{j})|\omega_{j}|}_{\begin{subarray}{c}\text{bias obs minus}\\ \text{ bias exp}/B\end{subarray}}\big)^{2}.$
The very first step is optimizing over the estimator (assuming here for simplicity $\gamma$ can take any value). The following characterizes the regret optimal $\gamma^{\star}$ (see Appendix B.1).
[…]
At the boundaries, if the variance regret $\alpha/\alpha^{\star}$ is always smaller than the bias regret $\beta/\beta^{\star}$, the optimal weight is $\gamma_{j}^{\star}=1$ (i.e., use only the experimental estimate). If instead the variance regret always dominates, the optimal choice is the variance-optimal $\gamma_{j}^{\star}=\sigma_{j}^{2}/(\sigma_{j}^{2}+v_{j}^{2})$.
When the solution is an interior solution for $\gamma$, the optimal design minimizes both the bias and variance regret (now equalized), so that
$j^{\star}\in\arg\min_{j\in\{1,2\}}\underbrace{\left\{\omega_{-j}^{2}\sigma_{-j}^{2}+\omega_{j}^{2}\!\big[(1-\gamma_{j}^{\star})^{2}\sigma_{j}^{2}+(\gamma_{j}^{\star})^{2}v_{j}^{2}\big]\right\}}_{\text{overall variance at }\gamma_{j}^{\star}}=\arg\min_{j\in\{1,2\}}\underbrace{\left\{|\omega_{-j}|+|1-\gamma_{j}^{\star}||\omega_{j}|\right\}}_{\text{worst--case bias}/|B|\text{ at }\gamma_{j}^{\star}}.$
However, cases at the extreme boundaries can also occur. If the variance regret $\alpha/\alpha^{\star}$ or the bias regret $\beta/\beta^{\star}$ uniformly dominates the other for one of the experiments, the optimal design minimizes the dominating term. These are desiderable properties as the optimal design always prioritizes the dominating term. Appendix Table 5 provides a complete overview of all possible scenarios. Appendix B.2 provides a numerical example and Section 6 presents a more complex example that include optimally allocating sample size. ∎
3.2 Optimization
Next, we provide an explicit optimization routine that can be solved using off-the-shelf software. We consider observational estimates independent of the experimental ones (but possibly dependent with each other), and experimental estimates independent of each other. This occurs when conducting experiments on independent samples different from the observational sample as in our applications. More general dependences are possible but omitted.
Algorithm 1 Regret-optimal experimental design
1:Inputs: target functional weights $\omega$, variance parameters $(v_{j}^{2})$, observational covariance $\Sigma_{\mathrm{obs}}$, per-unit costs $(c_{j})$, total budget $n$, and feasible experiment set $\mathcal{X}$.
2:Define variance, bias, and sample sizes. For any $(x,\gamma)$ with $x\in\mathcal{X}$ and $0\leq\gamma_{j}\leq 1$, set $s_{j}=x_{j}\gamma_{j}$ and compute:
  • the sample allocation (8), and the definition of $\alpha(s)$ in (9);
  • the bias contribution $\beta(s)$ as defined below (9).
3:Compute oracle quantities. Obtain $\alpha^{\star}$ and $\beta^{\star}$ by solving the two optimization problems, i.e.,
  • minimize $\alpha(s)$ over $(x,\gamma,s)$ subject to $x\in\mathcal{X}$, $x_{j}\in\{0,1\}$, $s_{j}=x_{j}\gamma_{j}$;
  • minimize $\sum_{j}x_{j}|\omega_{j}|$ over $x\in\mathcal{X}$ with $x_{j}\in\{0,1\}$.
4:Solve the minimax MIQP. Solve (10)–(14) with a off-the-shelf solver.
5:Outputs. Let $(x^{\star},\gamma^{\star},s^{\star},t^{\star})$ denote an optimizer of (10)–(14). Report:
  • the regret-optimal experiment set $\mathcal{E}^{\star}=\{j:x_{j}^{\star}=1\}$ and shrinkage weights $\gamma^{\star}$;
  • the implied $s^{\star}$ and regret factor $t^{\star}=\max\{\alpha/\alpha^{\star},\beta/\beta^{\star}\}$;
  • the optimal sample sizes $n_{j}^{\star}$ obtained by plugging $s^{\star}$ into (8).
Notation and decision variables
For given $(\mathcal{E},\Sigma,\gamma)$, denote by $\Sigma_{\mathrm{obs}}$ the submatrix of $\Sigma$ corresponding to the variance–covariance matrix of the observational estimates (possibly non-diagonal), and let $\mathbb{V}(\tilde{\theta}_{j}^{\mathrm{exp}})=v_{j}^{2}/n_{j}$ be the variance of the experimental estimator for component $j$, where $n_{j}$ denotes the sample size allocated to that experiment. Let $x_{j}=1\{j\in\mathcal{E}\}\in\{0,1\}$ indicate whether the experiment identifies component $j$, and let $x\in\mathcal{X}$, where $\mathcal{X}$ denotes the constraint set of feasible experiments. We optimize over $(x,\gamma,n_{j})$ under a total budget constraint $\sum_{j\in\mathcal{E}}c_{j}n_{j}=n,$ where $c_{j}>0$ denotes a per-unit cost for experiment.
Variance, bias and optimal sample size
Under independence of experimental and observational estimators, the variance of $\tau(\theta)$ can be written as
$\sum_{j}\omega_{j}^{2}s_{j}^{2}\frac{v_{j}^{2}}{n_{j}}\;+\;\big(\omega\odot(1-s)\big)^{\top}\Sigma_{\mathrm{obs}}\big(\omega\odot(1-s)\big),\qquad s_{j}\equiv x_{j}\gamma_{j},$ (7)
where $\odot$ denotes the elementwise product. The sample sizes $n_{j}$ enter only through the first (experimental) component. By the standard Neyman allocation with variable costs (Gerber and Green, 2012), the minimizer over $(n_{j})$ for fixed $s$ is given by
$n_{j}^{\star}\;=\;n\,\frac{|\omega_{j}|v_{j}s_{j}/\sqrt{c_{j}}}{\sum_{k}|\omega_{k}|v_{k}s_{k}\sqrt{c_{k}}}.$ (8)
Substituting (8) into the experimental variance and with a slight abuse of notation, let
$\alpha(s)\;\equiv\;\frac{1}{n}\Big(\sum_{j}s_{j}|\omega_{j}|v_{j}\sqrt{c_{j}}\Big)^{2}+\big(\omega\odot(1-s)\big)^{\top}\Sigma_{\mathrm{obs}}\big(\omega\odot(1-s)\big),\quad\beta(s)\;\equiv\;\Big(\sum_{j}(1-s_{j})|\omega_{j}|\Big)^{2},$ (9)
the variance and bias contributions.
Oracle solutions
The next step is to compute $\alpha^{\star}$ and $\beta^{\star}$. These solve
$\alpha^{\star}=\min_{x,\gamma,s}\alpha(s),$ $\text{s.t. }0\leq\gamma_{j}\leq 1,\;\;x_{j}\in\{0,1\},\;\;x\in\mathcal{X},$
$0\leq s_{j}\leq\gamma_{j},\;\;s_{j}\leq x_{j},\;\;s_{j}\geq\gamma_{j}+x_{j}-1,\quad\forall j,$
$\sqrt{\beta^{\star}}=\min_{x^{\prime}}\sum_{j}(1-x_{j}^{\prime})|\omega_{j}|,$ $\text{s.t. }x^{\prime}\in\mathcal{X},\;\;x_{j}^{\prime}\in\{0,1\},\quad\forall j.$
The constraints for $s$ are standard inequalities enforcing $s_{j}=\gamma_{j}x_{j}$ when $x_{j}\in\{0,1\}$ and $0\leq\gamma_{j}\leq 1$. Because $\alpha(s)$ is convex quadratic and $\mathcal{X}$ typically admits a linear (or quadratic) description, the first problem is a mixed-integer convex quadratic program, while the second is a mixed-integer linear program. Both can be solved using off-the-shelf solvers.
Regret-optimal solution
We are now ready to solve for the regret-optimal design. By Theorem 1, we want to minimize $\max\{\alpha/\alpha^{\star},\beta/\beta^{\star}\}$. The corresponding optimization program can be written as
$\min_{x,\gamma,s,t}\quad$ $t$ (10)
s.t. $\alpha(s)\;\leq\;t\,\alpha^{\star},$ (11)
$\beta(s)\;\leq\;t\,\beta^{\star},$ (12)
$0\leq\gamma_{j}\leq 1,\quad x_{j}\in\{0,1\},\quad j=1,\dots,p,$ (13)
$0\leq s_{j}\leq\gamma_{j},\quad s_{j}\leq x_{j},\quad s_{j}\geq\gamma_{j}+x_{j}-1,\quad j=1,\dots,p,$ (14)
$x\in\mathcal{X}\quad\text{(constraint on feasible experiments).}$ (15)
The constraints (14) again enforce $s_{j}=x_{j}\gamma_{j}$ for $x_{j}\in\{0,1\}$ and $0\leq\gamma_{j}\leq 1$. The variable $t$ satisfies $t\geq\alpha/\alpha^{\star}$ and $t\geq\beta/\beta^{\star}$, so at the optimum we have $t=\max\{\alpha/\alpha^{\star},\beta/\beta^{\star}\}$ as desired. Because both $\alpha(s)$ and $\beta(s)$ are convex quadratic functions, the optimization problem (10)–(14) is a mixed-integer quadratic program (MIQP) with linear objective and convex quadratic constraints, and can be passed to standard MIQP solvers.
Remark 4 (“Bias-aware” Neyman allocation).
The optimal sample size solution in Equation (8) has the desirable feature that, conditional on the chosen experiment and shrinkage weights, it takes the form of a Neyman (variance-optimal) allocation. However, because both the experiment and the estimator are selected to trade off variance and worst-case bias, the resulting allocation generally differs from the standard Neyman allocation one would obtain by ignoring the sensitivity of the estimand to misspecification (and the fixed-cost constraints). It is therefore natural to refer to our sample size allocation as a (novel) “bias-aware” Neyman allocation, where the choice of the experiment and shrinkage weights trade-off bias and variance, and optimal sample size minimizes the variance conditional on these choices. ∎
4 General framework for moment selection
In many scenarios, we may expect that experiments only identify relevant moments in the data, and observational estimates are often based on model-implied moment conditions. We now extend the framework to generalized method of moments (GMM) estimators. In addition, we allow for a more general class of ambiguity sets and multivalued estimands.
Consider a vector of moment conditions $g(\theta)\in\mathbb{R}^{p}$ that stacks the “menu of moments” (observational and experimental), i.e. all moments potentially available for estimation. Let $\Lambda$ denote the Jacobian of the moment function with respect to $\theta$, evaluated at $\theta$ (i.e., $\Lambda:=\partial g(\theta)/\partial\theta^{\top}$). For simplicity, we will assume that $\Lambda$ is known (i.e., moments are linear in $\theta$). This simplifies exposition but it is not necessary in the local asymptotic framework we present in Section 5.3, where $\Lambda$ can be replaced by the Jacobian of $g(\theta)$ evaluated at the observational estimates. Let $\bar{g}_{\Sigma}$ denote the sample analogue with
$\mathbb{E}[\bar{g}_{\Sigma}]=b,\qquad\mathbb{V}(\bar{g}_{\Sigma})=\Sigma,$
where the vector $b$ captures the biases in the moments. The researcher may access only some of the entries of $\bar{g}$ by selecting a subset of experiments due to budget constraints.
General misspecification set.
Let $\mathcal{I}\subseteq\{1,\dots,p\}$ index all potential experimental moments researchers may select (unbiased by design) and $\mathcal{I}^{c}$ the observational ones (potentially misspecified). We posit bounded misspecification under arbitrary norms
[…]
where the ambiguity set is defined relative to an arbitrary norm $||\cdot||_{l}$. Examples include the $\ell_{1},\ell_{2},\ell_{\infty}$ norms as special cases, as well as norms that may weight observational biases differently, or may even assume no bias for some observational moments (see Example 7).
Estimator and design.
Given the empirical moments $\bar{g}_{\Sigma}$, we consider (asymptotically) linear estimators of the form
$\hat{\theta}(W,\Sigma)-\theta\;=\;\Gamma_{\Lambda}(W)\,\bar{g}_{\Sigma},\quad\Gamma_{\Lambda}(W)\equiv-(\Lambda^{\top}W\Lambda)^{-1}\Lambda^{\top}W$ (17)
where $W$ and $\Sigma$ are chosen by the researcher from a feasible set. The structure of the estimator follows a standard GMM asymptotic expansion (e.g. Andrews et al., 2020, 2017). [^11] The weighting matrix $W$ may be singular. The expansion $\hat{\theta}-\theta\;\dot{=}\;-(\Lambda^{\top}W\Lambda)^{-1}\Lambda^{\top}W\,\bar{g}_{\Sigma}$ holds whenever $\Lambda^{\top}W\Lambda$ is nonsingular, which can be forced through the set of constraints in Assumption 3.
Here, $W$ may have columns of zeros, which excludes the corresponding moments from the estimator. This occurs when researchers face a trade-off in which experiment to run, leading to different moments $\bar{g}$ observed in the experiment.
The matrix $\Sigma$ depends on the experimental design (e.g., sample-size allocations across experimental moments). The constraints on $W$ and $\Sigma$ effectively limit the class of experiments, sample size allocations and estimators the researcher may consider.
Assumption 3 (Constraint set).
Let $(W,\Sigma)\in\mathcal{D}^{\prime}$ for a set $\mathcal{D}^{\prime}$ so that $\Lambda^{\top}W\Lambda$ is invertible, $\Gamma_{\Lambda}(W)\Sigma\Gamma_{\Lambda}(W)^{\top}$ has uniformly bounded entries and is strictly positive definite.
Assumption 3 requires that the variance of the selected (reweighted) moments is uniformly bounded and strictly positive definite (even when $W$ does not select some moments). This ensures that any subset of moments the researcher chooses to use is nondegenerate. Note that for now we are abstracting from settings with incomplete models studied in Section 5.1.
Example 6 (GMM moment selection).
Let $\theta\in\mathbb{R}^{2}$ and
$g(\theta)=\big(g_{1}^{\text{obs}}(\theta),\,g_{2}^{\text{obs}}(\theta),\,g^{\text{exp}}_{1}(\theta),\,g^{\text{exp}}_{2}(\theta)\big)^{\top}\in\mathbb{R}^{4},$
where the first two components are observational and the last two are experimental. Suppose at most one experimental moment can be used. Selection is encoded by the GMM weight matrix $W\in\mathbb{R}^{4\times 4}$. A simple class of weighting matrices is
$W^{(1)}=\begin{pmatrix}w_{\mathrm{obs},1}&\rho&0&0\\ \rho&w_{\mathrm{obs},2}&0&0\\ 0&0&w_{\mathrm{exp},1}&0\\ 0&0&0&0\end{pmatrix}\qquad W^{(2)}=\begin{pmatrix}w_{\mathrm{obs},1}&\rho&0&0\\ \rho&w_{\mathrm{obs},2}&0&0\\ 0&0&0&0\\ 0&0&0&w_{\mathrm{exp},2}\end{pmatrix}$,
where $W^{(1)}$ selects the second experimental moment $g^{\text{exp}}_{1}$ and sets the weight on $g^{\text{exp}}_{2}$ to zero, while $W^{(2)}$ selects $g^{\text{exp}}_{2}$ and excludes $g^{\text{exp}}_{1}$. ∎
Target and MSE.
Let the estimand be potentially multivalued $\tau(\theta)\in\mathbb{R}^{q}$, where $\tau(\theta)=\Omega\theta$ with $\Omega\in\mathbb{R}^{q\times d}$ with $\Omega$ different from the zero matrix. Our objective is the sum of MSE for each entry of $\tau$, that is equal to
[…]
The key distinction now is that the variance and bias regret depend on $\Omega$ and $\Lambda$.
Definition 3 (Moment based variance and bias regret).
[…]
$\alpha_{\Omega}(W,\Sigma)\;\equiv\;\mathrm{Trace}\left(\Omega\Gamma_{\Lambda}(W)\,\Sigma\,\Gamma_{\Lambda}(W)^{\top}\Omega^{\top}\right),\qquad\beta_{l,\Omega}(W)\;\equiv\;\big\|[\Omega\Gamma_{\Lambda}(W)]_{.,\mathcal{I}^{c}}\big\|_{l,*}^{\,2}.$
Let $\alpha_{\Omega}^{\star},\beta_{l,\Omega}^{\star}$ their corresponding minimizers over $(W,\Sigma)\in\mathcal{D}^{\prime}$. The moment-based variance and bias regret are defined as $\alpha_{\Omega}/\alpha_{\Omega}^{\star},\beta_{l,\Omega}/\beta_{l,\Omega}^{\star}$ respectively.
The bias now depends on the interaction between the sensitivity $\Omega$, the moment Jacobian $\Lambda$ and the identity of which moments are observational and therefore potentially biased. For unidimensional $\tau(\theta)$ the bias corresponds to the squared dual norm of $\Omega\Gamma_{\Lambda}(W)$ with respect to $||\cdot||_{l}$. The weighting matrix $W$ determines how these moments influence $\hat{\theta}$.
As before, we consider the adaptation regret with respect to an oracle that knows $B$:
$\mathcal{R}_{l}(W,\Sigma)\;\equiv\;\sup_{B\geq 0}\ \frac{\sup_{b\in\mathcal{B}_{l}(B)}\mathrm{MSE}_{b}(W,\Sigma)}{\inf_{(W^{\prime},\Sigma^{\prime})\in\mathcal{D}^{\prime}}\sup_{b^{\prime}\in\mathcal{B}_{l}(B)}\ \mathrm{MSE}_{b^{\prime}}(W^{\prime},\Sigma^{\prime})}.$ (19)
The next theorem generalizes Theorem 1 to this more general framework.
Theorem 2 (Regret characterization).
Let Assumptions 2, 3 hold, $\mathcal{R}$ be as defined in Equation (19),with $\alpha,\beta,\alpha^{\star},\beta^{\star}$ as in Definition 3. Consider an estimator as in Equation (17). Then
$\mathcal{R}_{l}(W,\Sigma)\;=\;\max\!\left\{\,\frac{\alpha_{\Omega}(W,\Sigma)}{\alpha_{\Omega}^{\star}}\,,\ \frac{\beta_{l,\Omega}(W)}{\beta_{l,\Omega}^{\star}}\,\right\}.$ (20)
Proof.
See Appendix A.1. ∎
Theorem 2 shows how our framework generalizes to (i) choosing the moments and the estimator (via $W$); and (ii) arbitrary norms characterizing the ambiguity set. The key insight is that the regret is quasi-convex under (approximate) linearity in the moments relative to the dual norm of the ambiguity set.
This generalization encompasses, in addition to choosing which experiment to run, problems that involve whether to collect additional data (e.g., through surveys or data acquisition) since these can be directly incorporated as part of the moment conditions.
Example 7 (Weighted bias).
Let $\tau(\theta)\in\mathbb{R}$ and consider an ambiguity set of the form $\tilde{\mathcal{B}}(B)\;\equiv\;\Big\{b\in\mathbb{R}^{p}:|b_{j}|\leq k_{j}B\;\;\text{for all }j=1,\dots,p\Big\},$ for known rescaling factors $k$. Then the corresponding bias regret is $\tilde{\beta}/\tilde{\beta}^{\star}$ with $\tilde{\beta}(\mathcal{E},\gamma)\;\equiv\;(\sum_{j\notin\mathcal{E}}k_{j}|\omega_{j}|\;+\;\sum_{j\in\mathcal{E}}k_{j}|\omega_{j}|\big|1-\gamma_{j}\big|)^{2},\quad\tilde{\beta}^{*}\;\equiv\;\min_{\mathcal{E}\in\mathcal{S}}(\sum_{j\notin\mathcal{E}}k_{j}|\omega_{j}|)^{2},$ and the overall regret equals $\max\!\left\{\,\frac{\alpha(W,\Sigma)}{\alpha^{\star}}\,,\ \frac{\tilde{\beta}(W)}{\tilde{\beta}^{\star}}\,\right\}$. ∎
5 Confidence sets, incomplete models and extensions
5.1 Confidence sets and partially identified $\tau$
Next, we generalize our framework to minimizing confidence interval length and partially identified $\tau(\theta)$. For simplicity, we let $\tau(\theta)\in\mathbb{R}$. Researchers may not know $\tau(\cdot)$ exactly and only know functions $\bar{\tau},\underline{\tau}$ with
$\bar{\tau}(\theta)=\bar{\omega}^{\top}\theta,\qquad\underline{\tau}(\theta)=\underline{\omega}^{\top}\theta,\quad\tau(\theta)\in[\underline{\tau}(\theta),\bar{\tau}(\theta)]$
where linearity can be justified as a first-order linear approximation for $\bar{\tau}(\theta)$ and $\underline{\tau}(\theta)$ under the local asymptotics in Section 5.3 for smooth functions $\bar{\tau},\underline{\tau}$. [^12] When the envelope for $\tau(\theta)$ is non-smooth, we recommend replacing it with smooth upper and lower envelopes. In this case, following Section 5.3, $\bar{\omega}$ and $\underline{\omega}$ can be interpreted as the gradients of these smooth bounds evaluated at the observational estimates. We implicitly assume that at least one of $\bar{\omega}$ and $\underline{\omega}$ differ from the zero vector to avoid trivial solutions.
For a generic weight vector $\omega$ let $\alpha_{\omega}(W,\Sigma),\beta_{l,\omega}(W)$ as in Definition 3.
Confidence intervals under bias.
For a candidate design $(W,\Sigma)$ and a bias vector $b\in\mathbb{R}^{p}$, define two-sided $(1-\eta)$ lower and upper bounds for $\tau(\theta)$ by
$\ell_{b}(W,\Sigma)$ $\equiv\underline{\tau}(\hat{\theta}(W))\;-\;\underline{\omega}^{\top}\Gamma_{\Lambda}(W)b\;-\;z_{1-\eta/2}\,\sqrt{\alpha_{\underline{\omega}}(W,\Sigma)},$
$u_{b}(W,\Sigma)$ $\equiv\bar{\tau}(\hat{\theta}(W))\;-\;\bar{\omega}^{\top}\Gamma_{\Lambda}(W)b\;+\;z_{1-\eta/2}\,\sqrt{\alpha_{\bar{\omega}}(W,\Sigma)},$
where $z_{1-\eta/2}$ is the standard normal $(1-\eta/2)$ quantile and $\hat{\theta}(W)$ is the GMM estimator in (17). These bounds adjust for the bias $\omega^{\top}\Gamma_{\Lambda}(W)b$ in $\omega^{\top}\hat{\theta}(W)$ for each envelope separately.
As before, the Jacobian $\Lambda$ is known; when estimated see Section 5.3 and Remark 5.
An audience, indexed by a worst-case bias radius $B\geq 0$, forms a worst-case confidence interval (i.e., a bias-aware confidence interval a la Armstrong and Kolesár (2018)) by choosing the most conservative endpoints over $b$ in the ambiguity set $\mathcal{B}_{l}(B)$ in (16):
$L_{l,B}(W,\Sigma)~\equiv~\left[\inf_{b\in\mathcal{B}_{l}(B)}\ell_{b}(W,\Sigma),\;\sup_{b\in\mathcal{B}_{l}(B)}u_{b}(W,\Sigma)\right],$
and we denote by $|L_{l,B}(W,\Sigma)|$ its total length.
The researcher does not need to know the audience’s specific choice of $B$ and can simply report the point estimates and associated standard errors; the audience (indexed by $B$) can then form their own bias-aware confidence set from these statistics. The researcher would like to minimize the expected length of the confidence interval formed by the audience.
An oracle that knows $B$ minimizes the worst-case expected length of the audience’s confidence length relative to the size of the sharp identified set (worst-case under the audience’s prior $b\in\mathcal{B}_{l}(B)$). Its (ex-ante) expected loss is defined as
$\mathcal{L}_{l,B}(W,\Sigma)~\equiv~\sup_{b_{0}\in\mathcal{B}_{l}(B)}\Big\{\mathbb{E}_{W,\Sigma,b_{0}}\big[|L_{l,B}(W,\Sigma)|\big]-\big(\bar{\tau}(\theta)-\underline{\tau}(\theta)\big)\Big\}.$ (21)
Here $\mathcal{L}_{l,B}(W,\Sigma)$ measures how informative the expected confidence interval is, relative to the length of the identified set $\bar{\tau}(\theta)-\underline{\tau}(\theta)$, worst-case over $b_{0}$ in $\mathcal{B}_{l}(B)$. Here, $b_{0}$ determines the expectation of $\hat{\theta}(W)$, and ultimately the length of the bias-aware identified set.
As in previous sections, we define the corresponding regret as
$\tilde{\mathcal{R}}_{l}(W,\Sigma)~\equiv~\sup_{B\geq 0}\;\frac{\mathcal{L}_{l,B}(W,\Sigma)}{\inf_{(W^{\prime},\Sigma^{\prime})\in\mathcal{D}^{\prime}}\mathcal{L}_{l,B}(W^{\prime},\Sigma^{\prime})}\!.$
Example 8 (Confidence set in a complete model).
Consider a complete model with $\bar{\tau}(\theta)=\underline{\tau}(\theta)=\tau(\theta)$ and $\bar{\omega}=\underline{\omega}=\omega$. Then we can show
$\mathcal{L}_{l,B}(W,\Sigma)=2\,z_{1-\eta/2}\sqrt{\alpha_{\omega}(W,\Sigma)}+2\,B\,\sqrt{\beta_{l,\omega}(W)}.$
In particular, when the parameter is point identified, the excess confidence interval length depends only on the variance index $\alpha_{\omega}$ and the bias index $\beta_{l,\omega}$, corresponding to standard bias-aware confidence intervals (Armstrong and Kolesár, 2018).
Example 9 (Simple incomplete model).
[…]
t=0t=0t=1t=1t=2t=2t=3t=3 Design
Choose $(W,\Sigma)$
Reporting
Compute $\hat{\theta}(W)$
[…]
Report $\alpha_{\underline{\omega}}$, $\alpha_{\bar{\omega}}$
Audience
Observe $(W,\Sigma)$
Choose $B\in\mathbb{R}$
Form $L_{l,B}(W,\Sigma)$
Ex-post loss
Excess length
$|L_{l,B}(W,\Sigma)|-(\bar{\tau}(\theta)-\underline{\tau}(\theta))$
Figure 2: Timeline of design, reporting, audience confidence set, and loss.
Regret optimal design (and estimator).
Define
$A(W,\Sigma)\equiv z_{1-\eta/2}\big(\sqrt{\alpha_{\bar{\omega}}(W,\Sigma)}+\sqrt{\alpha_{\underline{\omega}}(W,\Sigma)}\big),\quad C_{l}(W)\equiv\sqrt{\beta_{l,\bar{\omega}}(W)}+\sqrt{\beta_{l,\underline{\omega}}(W)}+\sqrt{\beta_{l,\bar{\omega}-\underline{\omega}}(W)},$
and the corresponding oracle solutions
$A^{\star}\equiv\inf_{(W,\Sigma)\in\mathcal{D}^{\prime}}A(W,\Sigma),\qquad C_{l}^{\star}\equiv\inf_{(W,\Sigma)\in\mathcal{D}^{\prime}}C_{l}(W).$
The first component $A$ depends on the variance of the estimated parameters (and envelopes) and the second component $C_{l}$ depends on the bias. We now show how our results directly extend to this scenario.
Theorem 3 (Regret for confidence interval length).
Let Assumptions 2 and 3 hold and consider an estimator as in (17). Then, for any $(W,\Sigma)\in\mathcal{D}^{\prime}$,
$\tilde{\mathcal{R}}_{l}(W,\Sigma)=\max\Big\{\frac{A(W,\Sigma)}{A^{\star}},\;\frac{C_{l}(W)}{C_{l}^{\star}}\Big\}.$
Proof.
See Appendix A.2. ∎
Theorem 3 shows that the confidence-interval-length regret can be decomposed into a variance component $A(W,\Sigma)$ and a bias component $C_{l}(W)$, each divided by its smallest feasible value. This result mimics Theorem 2 allowing for partially identified models. It provides a ready-to-use expression for experimental design (and moment selection).
The following corollary shows that whenever the model is point identified the solutions that minimize the regret relative to confidence interval length coincides with the MSE optimal solution we derived in Section 4.
Corollary 1 (Equivalence of MSE and confidence-interval length optimal solutions).
Let Assumptions 2 and 3 hold. Let $\alpha,\beta_{l},\alpha^{\star},\beta_{l}^{\star}$ be as in Definition 3, and consider an estimator as in (17). Suppose $\bar{\omega}=\underline{\omega}=\omega$ (complete model). Then, for any $(W,\Sigma)$,
$\arg\min_{(W,\Sigma)\in\mathcal{D}^{\prime}}\mathcal{R}_{l}(W,\Sigma)~=~\arg\min_{(W,\Sigma)\in\mathcal{D}^{\prime}}\tilde{\mathcal{R}}_{l}(W,\Sigma),$
where $\mathcal{R}_{l}(W,\Sigma)$ denotes the MSE-based regret in Equation (19).
Proof.
See Appendix A.3. ∎
In summary, when $\tau(\theta)$ is point identified, minimizing the regret of the confidence interval length leads to the same optimal design as minimizing MSE-based regret. In partially identified models, Theorem 3 shows that the confidence interval length maintains the same bias-variance decomposition, a desirable property of the MSE solution in Section 4.
5.2 Optimization with partial knowledge on the bias bound
In this section we extend our results to settings where researchers have prior information about the bias upper bound. Following the framework in Section 4, suppose researchers know that $B\leq\bar{B}$ for a known upper bound $\bar{B}$, and consider an arbitrary (possibly weighted) norm $\|\cdot\|_{l}$ on the bias. The theorem below establishes the resulting characterization of the regret.
Theorem 4 (Regret characterization).
Consider the setting in Section 4. Let Assumptions 2 and 3 hold, and let $\mathcal{R}$ be as defined in Equation (19), with $\alpha_{\Omega},\beta_{l,\Omega},\alpha_{\Omega}^{\star},\beta_{l,\Omega}^{\star}$ as in Definition 3. Consider an estimator as in Equation (17). Then
$\sup_{B\in[0,\bar{B}]}\ \frac{\sup_{b\in\mathcal{B}_{l}(B)}\mathrm{MSE}_{b}(W,\Sigma)}{\inf_{(W^{\prime},\Sigma^{\prime})\in\mathcal{D}^{\prime}}\sup_{b^{\prime}\in\mathcal{B}_{l}(B)}\ \mathrm{MSE}_{b^{\prime}}(W^{\prime},\Sigma^{\prime})}\;=\;\max\!\left\{\,\frac{\alpha_{\Omega}(W,\Sigma)}{\alpha_{\Omega}^{\star}}\,,\ \frac{\alpha_{\Omega}(W,\Sigma)+\bar{B}^{2}\beta_{l,\Omega}(W)}{\min_{W^{\prime},\Sigma^{\prime}}\big\{\alpha_{\Omega}(W^{\prime},\Sigma^{\prime})+\bar{B}^{2}\beta_{l,\Omega}(W^{\prime})\big\}}\,\right\}.$ (22)
Proof.
See Appendix A.1. ∎
Theorem 4 shows that when an upper bound $\bar{B}$ on the bias is known, the worst-case proportional regret over all $B\in[0,\bar{B}]$ is again the maximum of two terms. The first term, $\alpha_{\Omega}(W,\Sigma)/\alpha_{\Omega}^{\star}$, is the familiar variance regret. The second term compares the MSE evaluated at the largest feasible bias level, $\bar{B}$, to the smallest achievable MSE at the same bias level across all feasible designs and weight matrices. The optimization problem is therefore analogous to our baseline case, but with the “bias” component of regret evaluated at $B=\bar{B}$ rather than in the limit as $B\to\infty$.
From an interpretation standpoint, this extension interpolates between two polar cases. When $\bar{B}$ is small, the second term is close to the variance ratio, and designs that minimize variance dominate. As $\bar{B}$ increases, the second term places more weight on robustness to misspecification, and the objective approaches the fully robust case studied in our baseline results. In practice, the theorem implies that any prior information that narrows $B$ to $0,\bar{B}]$ can be incorporated simply by replacing the worst-case bias component in our original criterion with the MSE at $B=\bar{B}$, while keeping the same quasi-convex structure.
5.3 Local asymptotics with non linear target
Next, we extend our framework for (i) nonlinear moment functions $g(\theta)$ used in Section 4 and (ii) a smooth, potentially nonlinear estimand $\tau(\theta)$. Although we focus on known $\tau(\cdot)$, similar reasoning applies to partially identified sets in Section 5.1, omitted for brevity.
Linearization of the moments
Assume the sample moments obey root-$n$ scaling, i.e., $\mathbb{V}(\bar{g}_{\Sigma})=O(n^{-1})$. Let $b=b_{n}$ be a sequence where we retain the same misspecification structure as in Section 4: the moment vector has mean $b=b_{n}$ with $(b_{n})_{j}=0$ for $j\in\mathcal{I}$ (experiment-based moments) and $(b_{n})_{\mathcal{I}^{c}}$ potentially nonzero. Our key assumption is a local condition on the moments bias of the form
$\|b_{n}\|^{2}\sqrt{n}\;\to\;0.$ (23)
This condition builds on the local asymptotic frameworks in Andrews et al. (2020) and Armstrong and Kolesár (2021), where $b_{n}||^{2}=1/n$ matching the rate of convergence of the variance. In our case, $b_{n}$ can grow faster, slower or at the same rate of the standard error (i.e., $b_{n}\propto n^{-\alpha},\alpha>1/4$), encompassing $b_{n}||^{2}=1/n$ as a special case.
Under standard smoothness conditions, a GMM estimator with weighting matrix $W$ satisfies
$\sqrt{n}\big(\hat{\theta}(W,\Sigma)-\theta\big)=-(\Lambda^{\top}W\Lambda)^{-1}\Lambda^{\top}W\,\sqrt{n}\big(\bar{g}_{\Sigma}-b_{n}\big)+o_{p}(1).$
The condition $\|b_{n}\|^{2}\sqrt{n}\to 0$ ensures that higher order terms are negligible under sufficient smoothness conditions, justifying the use of the linear expansion (17) for our design analysis. Because $b_{n}\to 0$, the Jacobian $\Lambda$ can be consistently estimated by evaluating the derivative of $g(\theta)$ at a preliminary observational estimate $\tilde{\theta}^{\mathrm{obs}}\rightarrow_{p}\theta$ (Andrews et al., 2020).
Linearization of the estimand.
Let $\tau:\mathbb{R}^{p}\to\mathbb{R}$ be twice continuously differentiable near $\theta$, with gradient $\omega(\theta):=\partial\tau(\theta)/\partial\theta$ and bounded Hessian. A Taylor expansion around $\theta$ yields
$\sqrt{n}\big(\tau(\hat{\theta})-\tau(\theta)\big)=\omega(\theta)^{\top}\,\Gamma_{\Lambda}(W)\,\sqrt{n}\big(\bar{g}_{\Sigma}-b_{n}\big)\;+\;o_{p}(1),$
because $\|\hat{\theta}-\theta\|=O_{p}(n^{-1/2}+\|b_{n}\|)$ and the local condition $\sqrt{n}\,\|b_{n}\|^{2}\to 0$ renders the quadratic remainder $o_{p}(1)$ after $\sqrt{n}$ scaling.
Consequently, the regret expressions derived for the linear case continue to apply up to $o_{p}(1)$, with $\omega$ interpreted as the gradient $\omega(\theta)$ (or any consistent plug-in $\omega(\tilde{\theta}^{\mathrm{obs}})$ from a preliminary estimator $\tilde{\theta}^{\mathrm{obs}}\to_{p}\theta$).
Remark 5 (Ex-post valid coverage with estimated $\Lambda$).
When some components of $\hat{\theta}$ are based on observational data and $\Lambda$ is estimated from the same observational sample used to select the design, the selection of the experiment and the construction of $\hat{\theta}$ both depend on the same data. Unless the observational sample is (asymptotically) much larger than the experimental sample, this case can lead to selection-induced distortions in the distribution of $\hat{\theta}$ and invalidate standard Gaussian confidence intervals (see, e.g., Andrews et al., 2024). A simple way to avoid this problem is to split the observational data into two independent subsamples (independent of the experimental samples): one used to estimate $\Lambda$ and choose the experiment at the design stage, and a second used, together with the experimental data to estimate the observational and experimental components of $\hat{\theta}$. Under this sample-splitting scheme the experimental and observational data used for inference are independent of the design choice, so standard CLT-based confidence intervals retain their nominal coverage. ∎
Remark 6 (Equivalent rescaling).
An equivalent formulation uses $\sqrt{n}$–rescaled coordinates (as in Example 4). Define the rescaled moment vector $\tilde{g}:=\sqrt{n}\,\bar{g}$ and the rescaled estimation error $\tilde{\theta}\equiv\sqrt{n}\,(\hat{\theta}-\theta)$. Then $\mathbb{V}(\tilde{g})=\tilde{\Sigma}=O(1),\mathbb{E}[\tilde{g}]=\tilde{b}_{n}\ \ \text{with}\ \ \tilde{b}_{n}\equiv\sqrt{n}\,b,$ so $\tilde{b}_{n}$ is the asymptotic bias and $\tilde{\Sigma}$ the asymptotic variance. Local misspecification in (23) is equivalent to $\frac{\|\tilde{b}_{n}\|^{2}}{\sqrt{n}}\;\to\;0,$ which coincides with notation in Example 4. ∎
6 Empirical applications
6.1 Experimental design for General Equilibrium effects
Because large-scale experimentation is often infeasible due to cost constraints (Muralidharan and Niehaus, 2017), researchers often combine experimental and observational data to extrapolate effects from small-scale experiments to learn general-equilibrium effects (e.g. Todd and Wolpin, 2006; Attanasio et al., 2012; Meghir et al., 2022). The goal of this subsection is to present a step-by-step practitioner’s guide with an example, and compare our regret-optimal design to a standard variance-minimizing (Neyman) benchmark.
Consider a researcher evaluating conditional cash transfers for sending children to school in rural Kenya. Due to budget constraints, the researcher can only randomize at a small scale (partial equilibrium), but ultimately wishes to predict GE effects. Researchers have access to preliminary (external) estimates from the Mexican PROGRESA experiment (Todd and Wolpin, 2006), although these preliminary estimates may lack external validity in Kenya.
Step 1: model description
The first step for a researcher is to describe how observational/external and experimental estimates in Kenya are jointly used for estimation. Here, we consider a stylized model of school choice following Bonhomme and Weidner (2022), Todd and Wolpin (2006).
Individual choices
Let $S\in\{0,1\}$ denote school attendance, $C$ consumption, $Y$ (pre-transfer) household income, $W$ the child’s potential wage, and $t$ the stipend when enrolled. Abstracting from covariates (we will introduce covariates in the estimation), utility is
$U(C,S,t,\varepsilon)=\xi_{1}\,C+\xi_{2}\,CS+(\xi_{2}-\xi_{3}-\xi_{1})\,tS+\xi_{4}\,S+S\varepsilon,\qquad\varepsilon\sim\mathcal{N}(0,1),$ (24)
with budget C=Y+W(1S)+tSC=Y+W(1-S)+tS. The parametrization $(\xi_{2}-\xi_{3}-\xi_{1})$ is without loss and simplifies expressions below. Enrollment satisfies $S=1\big\{U(Y+t,1,t,\varepsilon)>U(Y+W,0,0,0)\big\}.$ Letting $Z(Y,W,t)\equiv\xi_{3}W-\xi_{1}Y-\xi_{2}t-\xi_{4},$ we have $P(S=1\mid Y,W,t)=\Phi\!\big(-Z(Y,W,t)\big)$, with $\Phi(\cdot)$ denoting the standard Gaussian CDF.
Model for GE effects
We are interested in the effect of a small stipend $t$ to all eligible (poor) households in rural Kenya. GE feedback is allowed through income and wages which we write as functions of a common transfer $t$ to all eligible households in the population. Specifically for functions $y(t),w(t)$ and mean-zero idiosyncratic income and wage shocks $\varepsilon_{Yi},\varepsilon_{Wi}$, for a common transfer $t$, denote the potential individual income and wage [^13] Because we focus on the effect of a common transfer, we keep a scalar notation for $t$ for expositional convenience. Alternatively, here $y(t),w(t)$ can also be interpreted as exposure mappings (Aronow et al., 2017) and $t$ can be defined as a $n$ dimensional vector of transfers.
$Y_{i}(t)=y(t)+\varepsilon_{Yi},\hskip 18.49988ptW_{i}(t)=w(t)+\varepsilon_{Wi}.$
Let $y_{0}\equiv\partial_{t}y(t)|_{t=0}$ and $w_{0}\equiv\partial_{t}w(t)|_{t=0}$. Define $\phi_{0}\equiv\mathbb{E}\!\left[\phi\!\big(-Z\big(Y(0),W(0),0\big)\big)\right]$. Our estimand of interest is the marginal effect of an increase in a small transfer $t$ to all eligible households
$\mathbb{E}\!\left[\left.\frac{\partial\,\Pr(S{=}1\mid Y(t),W(t),t)}{\partial t}\right|_{t=0}\right]=\;\phi_{0}\cdot\big(\,\xi_{2}+\xi_{1}y_{0}-\xi_{3}w_{0}\,\big),$ (25)
which decomposes into the direct stipend effect $\phi_{0}\xi_{3}$ and the indirect GE effects via income and wages, $\phi_{0}\xi_{1}y_{0}$ and $\phi_{0}\xi_{1}w_{0}$. Under standard market clearing for labor supply and demand, we write [^14] This follows from the following: set $L_{s}(W,t)=L_{d}(W)$, the labor supply equals the labor demand, and differentiate at $t=0$: $\frac{\partial L_{s}}{\partial W}w_{0}+\frac{\partial L_{s}}{\partial t}=\frac{\partial L_{d}}{\partial W}w_{0}$. Suppose we have a constant number of hours worked for working child, denoted as $H$, and any child not at school is working. From the probit index, $\partial_{t}\mathbb{E}[S\,|\,\cdot]=\phi(-Z)\,\xi_{2}$ and $\partial_{W}\mathbb{E}[S\,|\,\cdot]=-\phi(-Z)\,\xi_{3}$ at $t=0$. Hence $\partial L_{s}/\partial t=-\phi_{0}\xi_{2}H$ and $\partial L_{s}/\partial W=\phi_{0}\xi_{3}H$. Dividing the equation by $H$ we obtain the desired result.
$w_{0}=\;\frac{\phi_{0}\,\xi_{2}}{\phi_{0}\,\xi_{3}-d},$ (26)
where $d$ denotes the slope of the demand curve divided by total hours of work. Wages in equilibrium depend both on the demand’s slope and the supply effect, captured through $\phi_{0}\xi_{2}$ (assuming that all non-students are employed for a constant number of hours).
Step 2: parametrization and transparency on biases
The following step is to define the parameters of interest, here defined as
$\theta_{1}\;\equiv\;\phi_{0}\,\xi_{1},\qquad\theta_{2}\;\equiv\;\phi_{0}\,\xi_{2},\qquad\theta_{3}\;\equiv\;-\phi_{0}\,\xi_{3},$ (27)
with corresponding estimand
$\tau(\theta)\;=\;\underbrace{\theta_{2}}_{\text{Direct effect}}\;+\;\underbrace{y_{0}}_{\text{Income multiplier}}\,\underbrace{\theta_{1}}_{\text{Income effect}}\;+\;\underbrace{w_{0}(\theta)\,\theta_{3}}_{\text{Wage effect}},\qquad w_{0}(\theta)\;=\;\frac{\theta_{2}}{-\theta_{3}-d},$ (28)
where the explicit form of $w_{0}(\theta)$ follows from (26), a smooth function of the parameters.
We treat $d$ as known for simplicity, using meta-analyses on demand elasticities, and $y_{0}=1.5$ based on experimental estimates in Kenya from Egger et al. (2022). [^15] Egger et al. (2022) report a total income/consumption multiplier $M_{\text{tot}}=\partial_{t}\mathbb{E}\big[Y_{\text{pre}}(t)+t\big]|_{t=0}=2.5$; therefore the target derivative of pre-transfer income is $y_{0}=\partial_{t}\mathbb{E}\big[Y_{\text{pre}}(t)\big]|_{t=0}=M_{\text{tot}}-1$. Specifically, $d=0.5\frac{1-S_{0}}{W_{0}}$, where $S_{0}$ and $W_{0}$ are baseline probability to go to school (assuming all children not in school go to work) and baseline wages calibrated to the pre-experimental data from PROGRESA and $0.5$ obtained from meta-studies in Espey and Thilmany (2000). [^16] One could also calibrate $S_{0},W_{0}$ to pre-experimental data in Kenya directly when this data is available.
Here, the primary concern is bias in $\tilde{\theta}^{\mathrm{obs}}$, abstracting from potential biases in $y_{0},d$. When biases of $(y_{0},d)$ are also considered first-order, researchers should let $y_{0},d$ be additional parameters in $\theta$, which is allowed in our framework. This difference illustrates a feature of our framework: the chosen parameterization ensures transparency about which bias concerns drive the design; these considerations can be collected in a pre-analysis plan.
Step 3: observational study estimates
Using experimental data from the large-scale Mexico’s PROGRESA program, we estimate $(\xi_{1},\xi_{2},\xi_{3},\xi_{4})$ through probit with standard controls (age, distance to school, eligibility, year, highest grade achieved) and use these estimates to obtain $\tilde{\theta}^{\mathrm{obs}}$. Estimation is conducted separately by gender; we will focus here on the effects on female students.
We pool treated and control observations to estimate the schooling effect $\xi_{4}$, controlling for the individual-level subsidy. In turn, we obtain observational estimates $\big(\tilde{\theta}_{1}^{\mathrm{obs}},\tilde{\theta}_{2}^{\mathrm{obs}},\tilde{\theta}_{3}^{\mathrm{obs}}\big)$ that map to the model parametrization. Standard errors are constructed using the Delta method with clustering at the village level. Following Section 5.3, we then construct $\omega=\frac{\partial\tau}{\partial\theta}\Big|_{\theta=\tilde{\theta}^{\mathrm{obs}}}$. Point estimates, estimated sensitivity $\omega$, and standard errors are reported below.
Parameter $\tilde{\theta}^{\mathrm{obs}}$ $\omega$ $\sqrt{\Sigma_{jj}^{\mathrm{obs}}}$
[…]
Observational estimates for female students, corresponding sensitivity parameter $\omega$, and standard errors.
$\Sigma^{\mathrm{obs}}\;=\;\begin{bmatrix}4.31&-11.31&5.57\\ -11.31&973.11&-126.16\\ 5.57&-126.16&1038.56\end{bmatrix}\times e^{-9}$
Variance–covariance matrix of observational estimates for female students.
Step 4: candidate experiments
In this example, we consider a researcher who can afford only small, partial-equilibrium experiments in Kenya, and examine three stylized experiments or combinations thereof:
  • $j=\{1\}$: Unconditional transfer (income shock). The researcher randomizes a small income shock to a small fraction of households (implying no general-equilibrium effects). Under a first-order (Taylor) approximation, the experiment identifies the average marginal effect of income on school attendance, $\theta_{1}\;=\;\mathbb{E}\left[\left.\frac{\partial}{\partial Y}\Pr\big(S{=}1\mid Y,W,t\big)\right|_{t=0}\right].$ The researcher optimizes over $\gamma_{1}$, while $\theta_{2}$ and $\theta_{3}$ are calibrated to PROGRESA.
  • $j=\{2\}$: Conditional cash transfer (stipend). The researcher randomizes a small stipend $t$, conditional on attending school, to a small fraction of households (with prices held fixed). Under a first-order approximation, this identifies the average marginal effect of the stipend on school attendance, $\theta_{2}\;=\;\mathbb{E}\left[\left.\frac{\partial}{\partial t}\Pr\big(S{=}1\mid Y,W,t\big)\right|_{t=0}\right],$ with variance $v_{2}^{2}/n$. The researcher optimizes over $\gamma_{2}$, while $\theta_{1}$ and $\theta_{3}$ are calibrated to the PROGRESA study.
  • $j=\{3\}$: Youth job program (wage shock). Last, consider a job program for teenagers, that hires teenagers with small randomized differential wages. The program identifies $\theta_{3}$, holding constant $\theta_{1}$ and $\theta_{2}$
The per-unit variance in the experiment is calibrated to the corresponding entry of $\sigma^{2}$ in observational study, appropriately multiplied by the observational study sample size.
We consider two scenarios. First, the researcher may run only one of the three programs. Second, the researcher may run combinations of two of the three programs allocating a fixed total sample size to each, but not all three simultaneously, reflecting higher fixed costs and institutional constraints. For instance, the job program requires a different implementation infrastructure than a cash-transfer program; at the same time, conducting conditional and unconditional cash transfers in the same experimental location is often infeasible due to fairness concerns (e.g. Wingfield et al., 2023), which raises costs for each cash-transfer arm.
Remark 7 (Additional constraints).
In some contexts additional constraints may apply. For instance, we may be able to only run either $\{1,2\}$ or $\{3\}$, or only run either $\{1\}$ or $\{2\}$ but not $\{3\}$. Researchers can directly include these constraints in our framework. ∎
Experiment type Identified parameters Choice Variables Calibrated standard errors Description
$j=\{1\}$ $\theta_{1}$ $\gamma_{1}$ $\frac{6.56e^{-5}}{\sqrt{n_{\mathrm{tot}}/n_{\mathrm{obs}}}}$ Unconditional transfer (income shock, UCT). Randomize an income shock to a small fraction of households.
$j=\{2\}$ $\theta_{2}$ $\gamma_{2}$ $\frac{9.86e^{-4}}{\sqrt{n_{\mathrm{tot}}/n_{\mathrm{obs}}}}$ Conditional cash transfer (education shock, CCT). Randomize a conditional transfer to a small fraction of households.
$j=\{3\}$ $\theta_{3}$ $\gamma_{3}$ $\frac{1.01e^{-3}}{\sqrt{n_{\mathrm{tot}}/n_{\mathrm{obs}}}}$ Job program (Wage shock). Teenager jobs program with (small) randomized wages differentials.
$\{j,j^{\prime}\}$ $(\theta_{j},\theta_{j^{\prime}})$ $\gamma_{j},\gamma_{j^{\prime}},n_{j}$ $\frac{\sqrt{\Sigma_{jj}}}{\sqrt{n_{j}/n_{\mathrm{obs}}}}$, $\frac{\sqrt{\Sigma_{j^{\prime}j^{\prime}}}}{\sqrt{(n_{\mathrm{tot}}-n_{j})/n_{\mathrm{obs}}}}$ Combination of two of the three possible experiments. Choose two treatment arms and allocate optimally sample size, with given total experimental sample size $n_{\mathrm{tot}}$.
Table 1: Design options, identified parameters, choice variables, and per-unit variances.
Step 5: experimental design
Figure 3: Optimal experimental sample sizes $n_{j}$ allocated to each arm (UCT, CCT, and Job program) as a function of the total experimental sample size $n_{\text{tot}}$. The left panel considers designs with only one treatment arm, while the right panel allows two treatment arms.
Figure 4: Variance, bias, and adaptive regret of the Neyman (variance–optimal) allocation relative to the proposed optimal design, as a function of the total experimental sample size $n_{\text{tot}}$. Left panel: one feasible experiment; right panel: two feasible experiments.
Finally, we design the experiment, with total number of experimental observations $n_{\mathrm{tot}}$; let $n_{\mathrm{obs}}$ denote instead the number of units in the original Mexican PROGRESA experiment.
Choice of the experiment and sample size
In Figure 3 we show the optimal sample size allocation as a function of the total experimental sample size, under the constraints that either one (left panel) or two (right panel) treatment arms can be selected. When the total sample size is small and only one arm can be chosen, the optimal design selects the UCT arm. Intuitively, the UCT strikes a good balance between sensitivity and precision of the experimental estimate when $n_{\text{tot}}$ is small.
When two treatment arms can be selected, the optimal design initially chooses both UCT and CCT for $n_{\mathrm{tot}}\leq 500$. In this region, almost all of the experimental sample is allocated to the CCT arm (which is more sensitive but noisier), with a smaller fraction allocated to the more precise UCT arm. For total sample sizes above five-hundred, the optimal design instead selects the job program as the main experiment, reflecting the high sensitivity of the estimand to the wage effect (Appendix Figure 12 reports the bias and variance regret of the design). With two experiments and large sample sizes, roughly 90% of the total sample is allocated to the job program and the remaining fraction to the UCT arm, using the precise UCT effect to help balance variance against bias.
In addition to sample allocation, our method optimizes the shrinkage weights of the parameters of interest. Appendix Figure 13 collects the estimated shrinkage weight across designs. We find that the shrinkage weight is above 0.7, and closer to one for most cases. This shows how the experimental variation gets significantly higher weight.
Main bias and variance regret comparisons
In Figure 4 we compare the proposed optimal design to the variance-optimal (Neyman) design in terms of both variance and bias. This comparison clarifies what is gained by our procedure relative to a benchmark that chooses the experiment, shrinkage weights, and sample size solely to minimize the variance of the estimated effect $\hat{\tau}$ (and which therefore achieves the smallest possible variance). We find that the variance ratio between the Neyman and optimal designs (the inverse of the variance regret) is close to one for all sample sizes and converges to one as the sample size grows. For experimental sample $n_{\mathrm{tot}}=1000$, the variance of the Neyman and our proposed procedure is only about $14\%$ smaller when choosing one treatment arm and $35\%$ when choosing two treatment arms. In contrast, the bias of the Neyman procedure is about 1.4 times larger for one treatment arm and about 4.5 times larger with two. As a result, the adaptation regret of the Neyman allocation is similarly 1.4 and 4.5 times larger than our proposed design. These comparisons illustrate the benefit of our method: its variance is only slightly larger than the Neyman variance, while the bias is more than four times smaller for two treatment arms.
6.2 Site selection for external validity
Next, we show how observational evidence can inform whom to recruit into an experiment. We consider the setting in Banerjee et al. (2024), who study how the expansion of microfinance affects village social networks in Karnataka, India. We focus on the first outcome reported by Banerjee et al. (2024) corresponding to the density of the network, i.e., the percentage of connections a random household in a village has relative to the village size.
Background and research question
In their observational analysis, the authors assemble panel data from 75 villages in Karnataka, 43 of which were exposed to microfinance. Because program rollout was not randomized, they estimate effects using Difference-in-Differences (DiD). Although DiD is informative, the lack of randomized assignment can bias estimates in the presence of selection (e.g., Ghanem et al., 2022). Therefore, the authors subsequently conducted an experimental evaluation in one metropolitan area, randomizing microfinance access across 104 urban neighborhoods. We only use observational variation for choosing the experimental design and estimator. We then use experimental variation from Banerjee et al. (2024) to validate our procedure by calibrating its worst-case bias.
In practice, implementation costs often depend on how geographically dispersed the study sites are due to coordination and survey costs. Using preliminary observational estimates from Banerjee et al. (2024) in Karnataka, we ask: Which area(s) in Karnataka should be prioritized for an experimental evaluation, and how many villages should be enrolled in each area? We focus on the experiment randomizing treatments in one or two areas.
Observational study
We partition the observational sample into four geographically contiguous areas that group nearby villages; each area contains 11–12 villages that were exposed to microfinance during the study period. Areas are heterogeneous in terms of their overall population size. Figure 5 maps these four areas and reports the corresponding DiD point estimates. Three of the four areas display negative estimated effects, with meaningful variation in magnitudes across areas. Table 2 summarizes, for each area, samples sizes, variances clustered at the village level, and population sizes ($\omega$). These area-specific DiD estimates and their variances serve as our primary observational inputs for experimental design. [^17] We use DiD estimates to mimic the estimator used by the authors. An analogous analysis can be conducted after empirical-Bayes shrinkage of area-level estimates, omitted for brevity. Using the 2011 population census, matched with the geographic boundaries of the areas obtained from merging village positions with the SHRUG platform (Asher et al., 2019), we compute $\omega$ as the population share of a given area.
Table 2: Area-level observational inputs for Karnataka villages. The outcome is network density (corresponding to the average share of connection of an individual to other individuals in a village). For each area, we report the numbers of treated ($n_{1}$) and untreated ($n_{0}$) villages, the pre-intervention variance of density pooled across arms ($v_{\text{pre}}^{2}$), the area-specific DiD estimate $\hat{\mu}$, and its sampling variance $\widehat{\sigma}^{2}$. Here $\omega$ denotes the relative population share obtained using the 2011 population census in India.
Area $n_{1}$ $n_{0}$ $v_{\text{pre}}^{2}$ $\hat{\mu}$ $\widehat{\sigma}^{2}$ $\omega$
11 $11$ 66 0.000 4570.000\,457 0.0187-0.0187 0.000 045 30.000\,045\,3 $0.033$
22 $12$ 99 0.001 750.001\,75 0.0377-0.0377 0.000 1400.000\,140 $0.766$
33 $11$ $12$ 0.001 380.001\,38 0.003 90-0.003\,90 0.000 1870.000\,187 $0.121$
44 $11$ 66 0.000 9320.000\,932 $0.0148$ 0.000 1300.000\,130 $0.078$
Figure 5: Observational villages and area-level DiD effects in Karnataka. The background heat map partitions the area of interest into the four contiguous zones used to compute area-specific DiD estimates of microfinance’s impact on network density; the color scale encodes the DiD value. Points mark village locations: circles denote villages exposed to microfinance and triangles denote unexposed villages in the observational sample (authors’ survey). Major cities (Bengaluru, Mysuru, Tumakuru) are labeled for orientation.
Experimental design
We consider a family of designs that (i) select $E\in\{1,2\}$ geographic areas in Karnataka from which to recruit experimental sites and (ii) assign $n_{1}$ villages to treatment and $n_{0}=n_{1}$ to control (so the total sample is $n=2n_{1}$); (iii) optimizes over shrinkage weights as in Algorithm 1. We examine a grid of values for $n_{1}$ ranging from 40 to 80 (with 52 corresponding to the size of the experiment in Banerjee et al. (2024)). For variance calibration, we take $v_{\mathrm{pre},a}^{2},a\in\{1,\cdots,4\}$ to be the pre-intervention, area-level variance of network density in Table 2. We assume that the variance of a single treated-control difference in the experiment is $2v_{\mathrm{pre},a}^{2}$.
Choice of the areas
Figure 6 reports the optimal sample size allocation across areas when either one or two areas can be selected. When only one area can be chosen, the method selects the area with the largest weight $\omega$, which in this application corresponds to the area containing Bengaluru (about 76% of the population). When two areas can be selected, the procedure chooses the two areas with the largest $\omega$ (Area 2 and Area 3, the latter containing Tumakuru), where the second area has weight $\omega\approx 12\%$. The method also optimally allocates the sample size across these areas, assigning roughly 10 to 20 villages in total (half of which treated) to Tumakuru and the remaining villages to the larger area containing Bengaluru. Figure 14 shows the optimal shrinkage weights: the shrinkage weight for the Bengaluru area is about 93%, reflecting an interior solution, while it is essentially one for Tumakuru. The resulting pattern is simple and intuitive: assign most of the sample size to the area with the largest weight, but, when possible, include a second area with substantial weight, letting the allocation of villages across areas balance robustness and variance.
Figure 6: Area-level allocations as total treated villages $n_{1}$ varies (from ten to fifty-two villages), for one or two areas included in the experiment.
Figure 7: Relative worst-case MSE across designs under a calibrated bias bound $B$, set to the largest absolute discrepancy between the experimental and observational ATE estimates in Banerjee et al. (2024). The y-axis reports $\mathrm{MSE}/\mathrm{MSE}^{\star}$, where $\mathrm{MSE}$ is the worst-case MSE over $\|b\|_{\infty}\leq B$ for a given design and $\mathrm{MSE}^{\star}$ is the (infeasible) oracle minimum that knows $B$ and optimally chooses the design. We compare our design (“regret optimal”) that does not use information about $B$ and minimizes the worst-case adaptation regret to the variance-minimizing Neyman allocation.
MSE comparisons
We compare two strategies: (i) our robust design, which jointly chooses the area(s), the sample allocation, and the shrinkage weights $\gamma^{\star}$ as described in Section 3; and (ii) the variance–optimal (Neyman) allocation. We evaluate each strategy using the worst-case MSE over bias vectors satisfying $b||_{\infty}\leq B$. Rather than treating $B$ as a free sensitivity parameter, we calibrate it using experimental evidence from Banerjee et al. (2024). Specifically, we set $B$ equal to the largest absolute discrepancy between the Hyderabad RCT ATE reported in the follow-up experiment of Banerjee et al. (2024) and the corresponding Karnataka DiD estimate in each area. Importantly, our design procedure does not use this calibrated bound $B$ as an input; it is used only for the ex-post MSE comparison.
Figure 7 plots regret (the ratio $\text{MSE}/\text{MSE}^{\star}$, $\text{MSE}^{\star}$ is the smallest achievable worst-case MSE for a bias upper bound at most $B$, and MSE is the worst case MSE for $b||_{\infty}\leq B$). As the number of treated villages $n_{1}$ increases, all designs improve; because the denominator also falls with $n_{1}$, the ratio need not be monotone. Across $E\in\{1,2\}$, the Neyman benchmark remains well above the oracle— with regret around 2.5 or more. On the other hand, our design tracks the oracle more closely, with the ratio closer to one and no larger than 1.2. Appendix Figure 15 collects similar comparisons for the worst-case regret over all possible values of $B$.
7 Implications for practice
This paper studies experimental design in the presence of observational evidence. A key challenge is that the bias of the observational estimators is unknown in practice. We adopt a minimax adaptation regret criterion that compares the mean-squared error (MSE) of a candidate design to that of an oracle that knows the worst-case bias. This reveals a fundamental trade-off between precision and robustness. The optimal design balances the design’s variance normalized by the smallest achievable variance (variance gap) and its worst-case bias normalized by the smallest attainable bias (bias gap). We propose a procedure that jointly determines: (i) how to combine observational and experimental evidence; (ii) how to allocate precision across experiments given budget constraints; and (iii) which treatment arm and/or sub-population to include in the experiment with fixed experimental costs.
In practice, the workflow is:
  • Define the estimand(s) of interest. Specify $\tau(\theta)$ for a known mapping $\tau$ and unknown parameters $\theta\in\mathbb{R}^{p}$. For example, $\tau(\theta)$ may represent a counterfactual, a general-equilibrium effect, or an average impact across locations. Adopt a parametrization in which some (but not necessarily all) components of $\theta$ can be learned experimentally; this clarifies what the experiment can identify. When such a parametrization is not readily available, Section 4 shows how our procedure applies to $\tau$ identified through arbitrary moment restrictions and Section 5.1 when it is partially identified.
  • Assemble informative observational evidence. Collect observational estimates $\tilde{\theta}^{\rm obs}$ and their covariance $\tilde{\Sigma}^{\rm obs}$. These serve as informative (but potentially biased) baselines. Such evidence may come from observational designs, structural estimates with pre-experimental data, or prior experiments conducted in different contexts.
  • Compute the sensitivity parameters. Using the observational baseline, compute the sensitivity weights $\omega\;=\;\frac{\partial\tau(\theta)}{\partial\theta}\big|_{\tilde{\theta}^{\rm obs}}$, which quantify how bias in each coordinate of $\theta$ propagates to $\tau(\theta)$. Large $|\omega_{j}|$ indicates greater payoff to learning the $j$th component.
  • Specify feasibility constraints and calibrate experimental variance. Enumerate the admissible design set, and the budget on sample size allocation. Calibrate per-unit experimental variances using pilot studies or historical data to form the set of feasible designs $\mathcal{D}$.
  • Run the method and decide. Optimize over the design and corresponding estimators. The output is a pre-analysis plan with the selected arm(s), sample sizes, and pre-specified combination rule ($\gamma$), and optimization is conducted via a mixed-integer quadratic program in Algorithm 1.
The applicability of our framework spans a wide range of settings in economics and beyond. Examples include estimating general equilibrium effects or structural models (Todd and Wolpin, 2006; Attanasio et al., 2012; Meghir et al., 2022; Kreindler et al., 2023; de Albuquerque et al., 2025); choosing among alternative treatment arms in factorial designs (Muralidharan et al., 2020; Bandiera et al., 2025); and deciding where to run the next experiment for external validity (Gechter et al., 2024; Olea et al., 2024). In industrial organization, applications include choices between demand-side and supply-side interventions (Bergquist and Dinerstein, 2020), the effects of information acquisition in markets (Allende et al., 2019; Larroucau et al., 2024), and decisions about which additional data source to acquire to improve statistical analysis (Allcott et al., 2025). Beyond economics, medical applications may include allocating sample size across subgroups and allocating doses across treatment arms (e.g., Porter et al., 2024; Morita et al., 2017; Manski, 2025).
Several open questions remain for future research. These include settings where researchers have well-specified priors about bias such as from Bayesian models (Gechter et al., 2024), or other modeling choices (Olea et al., 2024). This may be potentially incorporated through additional moment restrictions as in Section 4. Additional future extensions include sequential or adaptive experimental choices (e.g. Cesa-Bianchi et al., 2025).
Appendix A Proofs of main results
A.1 Proof of Theorems 1, 2, 4
[…]
Step 1 (worst-case MSE). By definition of dual norm,
[…]
Hence $\sup_{b\in\mathcal{B}_{l}(B)}\mathrm{MSE}_{b}(W,\Sigma)=\alpha(W,\Sigma)+B^{2}\beta_{l}(W).$
Step 2 (oracle envelope and basic properties). Let $\delta(t):=\inf_{W,\Sigma}\{\alpha(W,\Sigma)+t\beta_{l}(W)\}$. Recall that $\beta_{l}(W)$ and $\alpha(W,\Sigma)$ are bounded from above and $\alpha(W,\Sigma)>0$ by Assumption 3 (using the fact that $\Omega$ contains some non zero entry). Being the pointwise infimum of affine functions of $t$, $\delta$ is concave, nondecreasing, and finite for $t\geq 0$. Moreover,
$\delta(0)=\alpha^{\star},\qquad\lim_{t\to\infty}\frac{\delta(t)}{t}=\beta_{l}^{\star}.$
Step 3 (quasi-convexity and boundary maximization) Fix $(\alpha,\beta_{l})$. Consider
$\phi(t):=\frac{\alpha+t\beta_{l}}{\delta(t)},\qquad t\geq 0.$
Because the numerator is affine and the denominator is positive and concave in $t$, $\phi$ is quasi-convex on $0,\infty)$. Hence, on any compact interval $[0,T]$, $\max_{t\in[0,T]}\phi(t)$ is attained at the boundary $\{0,T\}$. Letting $T\to\infty$ and using the limits in Step 2,
$\sup_{t\geq 0}\phi(t)=\max\!\left\{\frac{\alpha}{\delta(0)},\ \lim_{t\to\infty}\frac{\alpha+t\beta_{l}}{\delta(t)}\right\}=\max\!\left\{\frac{\alpha}{\alpha^{\star}},\ \frac{\beta_{l}}{\beta_{l}^{\star}}\right\},$
with the stated conventions $0/0=1$. This proves Theorem 2.
Step 4 (worst-case with bound on the bias) From quasi-convexity, whenever $t\leq\bar{t}$, it follows that
$\sup_{t\in[0,\bar{t}]}\phi(t)=\max\!\left\{\frac{\alpha}{\delta(0)},\ \frac{\alpha+\bar{t}\beta_{l}}{\delta(\bar{t})}\right\},$
proving Theorem 4.
A.2 Proof of Theorem 3
Step 1 (closed form of $\mathcal{L}_{l,B}(W,\Sigma)$). Fix $(W,\Sigma)$ and $B\geq 0$. By definition,
$\ell_{b}(W,\Sigma)=\underline{\tau}(\hat{\theta}(W))-\underline{\omega}^{\top}\Gamma_{\Lambda}(W)b-z_{1-\eta/2}\sqrt{\alpha_{\underline{\omega}}(W,\Sigma)},$
so
$\inf_{\|b\|_{l}\leq B}\ell_{b}(W,\Sigma)=\underline{\tau}(\hat{\theta}(W))-z_{1-\eta/2}\sqrt{\alpha_{\underline{\omega}}(W,\Sigma)}-\sup_{\|b\|_{l}\leq B}\underline{\omega}^{\top}\Gamma_{\Lambda}(W)b.$
Since $b_{j}=0$ for $j\in\mathcal{I}$ and $\|b_{\mathcal{I}^{c}}\|_{l}\leq B$, the dual norm gives
$\sup_{\|b\|_{l}\leq B}\underline{\omega}^{\top}\Gamma_{\Lambda}(W)b=B\,\sqrt{\beta_{l,\underline{\omega}}(W)}.$
Hence
$\inf_{\|b\|_{l}\leq B}\ell_{b}(W,\Sigma)=\underline{\tau}(\hat{\theta}(W))-z_{1-\eta/2}\sqrt{\alpha_{\underline{\omega}}(W,\Sigma)}-B\,\sqrt{\beta_{l,\underline{\omega}}(W)}.$
Similarly, for the upper bound,
$\sup_{\|b\|_{l}\leq B}u_{b}(W,\Sigma)=\bar{\tau}(\hat{\theta}(W))+z_{1-\eta/2}\sqrt{\alpha_{\bar{\omega}}(W,\Sigma)}+B\,\sqrt{\beta_{l,\bar{\omega}}(W)}.$
Therefore the worst-case interval has length
$|L_{l,B}(W,\Sigma)|=\big(\bar{\tau}(\hat{\theta}(W))-\underline{\tau}(\hat{\theta}(W))\big)+z_{1-\eta/2}\big(\sqrt{\alpha_{\bar{\omega}}(W,\Sigma)}+\sqrt{\alpha_{\underline{\omega}}(W,\Sigma)}\big)+B\big(\sqrt{\beta_{l,\bar{\omega}}(W)}+\sqrt{\beta_{l,\underline{\omega}}(W)}\big).$
Let $b_{0}$ denote the true bias, with $\|b_{0}\|_{l}\leq B$. Under (17),
$\bar{\tau}(\hat{\theta}(W))-\underline{\tau}(\hat{\theta}(W))=(\bar{\omega}-\underline{\omega})^{\top}\theta+(\bar{\omega}-\underline{\omega})^{\top}\Gamma_{\Lambda}(W)\bar{g}_{\Sigma},$
so
$\mathbb{E}_{W,\Sigma,b_{0}}\big[\bar{\tau}(\hat{\theta}(W))-\underline{\tau}(\hat{\theta}(W))\big]=(\bar{\omega}-\underline{\omega})^{\top}\theta+(\bar{\omega}-\underline{\omega})^{\top}\Gamma_{\Lambda}(W)b_{0}.$
Subtracting the identified set length $\bar{\tau}(\theta)-\underline{\tau}(\theta)=(\bar{\omega}-\underline{\omega})^{\top}\theta$, we obtain
$\mathbb{E}_{W,\Sigma,b_{0}}\big[|L_{l,B}(W,\Sigma)|\big]-\big(\bar{\tau}(\theta)-\underline{\tau}(\theta)\big)=$
$(\bar{\omega}-\underline{\omega})^{\top}\Gamma_{\Lambda}(W)b_{0}+z_{1-\eta/2}\big(\sqrt{\alpha_{\bar{\omega}}(W,\Sigma)}+\sqrt{\alpha_{\underline{\omega}}(W,\Sigma)}\big)+B\big(\sqrt{\beta_{l,\bar{\omega}}(W)}+\sqrt{\beta_{l,\underline{\omega}}(W)}\big).$
Taking the worst case over $\|b_{0}\|_{l}\leq B$ implies
$\mathcal{L}_{l,B}(W,\Sigma)=z_{1-\eta/2}\big(\sqrt{\alpha_{\bar{\omega}}(W,\Sigma)}+\sqrt{\alpha_{\underline{\omega}}(W,\Sigma)}\big)+B\big(\sqrt{\beta_{l,\bar{\omega}}(W)}+\sqrt{\beta_{l,\underline{\omega}}(W)}+\sqrt{\beta_{l,\bar{\omega}-\underline{\omega}}(W)}\big),$
which yields the linear form $\mathcal{L}_{l,B}(W,\Sigma)=A(W,\Sigma)+B\,C_{l}(W)$.
Step 2 (oracle envelope and basic properties). Define
$\delta(t)~\equiv~\inf_{(W,\Sigma)\in\mathcal{D}^{\prime}}\{A(W,\Sigma)+t\,C_{l}(W)\},\qquad t\geq 0.$
By Assumption 3 (because either or both $\bar{\omega},\underline{\omega}$ are not zero vectors), $A(W,\Sigma)$ and $C_{l}(W)$ are bounded from above and $A(W,\Sigma)>0$ for all $(W,\Sigma)\in\mathcal{D}^{\prime}$, so $\delta(t)$ is finite and strictly positive for all $t\geq 0$. Being the pointwise infimum of affine functions of $t$, $\delta$ is concave and nondecreasing on $0,\infty)$. Moreover,
$\delta(0)=A^{\star},\qquad\lim_{t\to\infty}\frac{\delta(t)}{t}=C_{l}^{\star}.$
Step 3 (quasi-convexity and boundary maximization). Fix $(W,\Sigma)$ and set $A\equiv A(W,\Sigma)$ and $C\equiv C_{l}(W)$. For $t\geq 0$ define
$\phi(t)~\equiv~\frac{A+t\,C}{\delta(t)}.$
The numerator is affine in $t$ and the denominator is positive and concave, so $\phi$ is quasi-convex on $0,\infty)$. Hence, on any compact interval $[0,T]$, $\max_{t\in[0,T]}\phi(t)$ is attained at the boundary $\{0,T\}$. Letting $T\to\infty$ and using the limits in Step 2,
$\sup_{t\geq 0}\phi(t)=\max\left\{\frac{A}{\delta(0)},\;\lim_{t\to\infty}\frac{A+t\,C}{\delta(t)}\right\}=\max\left\{\frac{A(W,\Sigma)}{A^{\star}},\;\frac{C_{l}(W)}{C_{l}^{\star}}\right\}.$
By the definition of $\tilde{\mathcal{R}}_{l}(W,\Sigma)$ and the identity $\mathcal{L}_{l,B}(W,\Sigma)=A(W,\Sigma)+BC_{l}(W)$, this yields
$\tilde{\mathcal{R}}_{l}(W,\Sigma)=\sup_{B\geq 0}\frac{\mathcal{L}_{l,B}(W,\Sigma)}{\inf_{(W^{\prime},\Sigma^{\prime})\in\mathcal{D}^{\prime}}\mathcal{L}_{l,B}(W^{\prime},\Sigma^{\prime})}=\max\left\{\frac{A(W,\Sigma)}{A^{\star}},\;\frac{C_{l}(W)}{C_{l}^{\star}}\right\},$
which proves the theorem. ∎
A.3 Proof of Corollary 1
Whenever $\bar{\omega}=\underline{\omega}=\omega$, we can write $A(W,\Sigma)=\alpha_{\omega}(W,\Sigma)$ and $C_{l}(W)=\beta_{l,\omega}(W)$. From Theorem 3
[…]
Appendix B Examples
B.1 Two parameters examples in Equation (6)
By Theorem 1, for fixed $(\mathcal{E},\Sigma)$ the adaptation regret as a function of the shrinkage vector equals
$\mathcal{R}(\gamma)\;=\;\max\!\Big\{\alpha(\mathcal{E},\Sigma,\gamma)/\alpha^{\star},\;\beta(\mathcal{E},\gamma)/\beta^{\star}\Big\},$
with $\alpha^{\star},\beta^{\star}$ constants.
Step 1. With independence and two parameters,
$\alpha(\mathcal{E},\Sigma,\gamma)=C+\omega_{j}^{2}\!\Big(\gamma_{j}^{2}v_{j}^{2}+(1-\gamma_{j})^{2}\sigma_{j}^{2}\Big),$
where $C$ does not depend on $\gamma_{j}$. Hence $\alpha(j,\gamma_{j})\equiv\alpha(\mathcal{E},\Sigma,\gamma)$ is a strictly convex quadratic in $\gamma_{j}$ with unique minimizer at
$\gamma_{j}^{\mathrm{var}}=\frac{\sigma_{j}^{2}}{\sigma_{j}^{2}+v_{j}^{2}}.$
Moreover, $\alpha(j,\gamma_{j})$ is strictly increasing on $\big(\gamma_{j}^{\mathrm{var}},1\big)$ and strictly decreasing on $\big[0,\gamma_{j}^{\mathrm{var}}\big)$.
Write the worst–case bias component holding the other coordinate fixed as
$\beta(j,\gamma_{j})\;=\;\Big(A_{j}+|\omega_{j}|\cdot|1-\gamma_{j}|\Big)^{2},$
where $A_{j}\geq 0$ collects all terms not involving $\gamma_{j}$ (including the contribution from the other coordinate). On $[0,1]$, $|1-\gamma_{j}|=1-\gamma_{j}$, so $\beta(j,\gamma_{j})$ is strictly decreasing in $\gamma_{j}$ and convex with minimum at $\gamma_{j}=1$. Therefore, any minimizer satisfies
$\gamma_{j}^{\star}\in\big[\gamma_{j}^{\mathrm{var}},\,1\big].$
Step 2. On $\big(\gamma_{j}^{\mathrm{var}},1\big)$ the map $\gamma_{j}\mapsto\alpha(j,\gamma_{j})/\alpha^{\star}$ is strictly increasing, while $\gamma_{j}\mapsto\beta(j,\gamma_{j})/\beta^{\star}$ is strictly decreasing. Hence the function
$f(\gamma_{j})\equiv\max\!\Big\{\alpha(j,\gamma_{j})/\alpha^{\star},\ \beta(j,\gamma_{j})/\beta^{\star}\Big\},\qquad\gamma_{j}\in\big[\gamma_{j}^{\mathrm{var}},1\big],$
is minimized either (i) at a boundary point, or (ii) at the unique interior point where the two arguments are equal (by strict monotonicity, at most one intersection exists). Therefore,
  • If $\alpha(j,\gamma_{j})/\alpha^{\star}<\beta(j,\gamma_{j})/\beta^{\star}$ for all $\gamma_{j}\in\big(\gamma_{j}^{\mathrm{var}},1\big)$, then $f(\gamma_{j})=\beta(j,\gamma_{j})/\beta^{\star}$ on that interval. Since this term is strictly decreasing, the minimizer is the right boundary $\gamma_{j}^{\star}=1$.
  • If $\alpha(j,\gamma_{j})/\alpha^{\star}>\beta(j,\gamma_{j})/\beta^{\star}$ for all $\gamma_{j}\in\big(\gamma_{j}^{\mathrm{var}},1\big)$, then $f(\gamma_{j})=\alpha(j,\gamma_{j})/\alpha^{\star}$ on that interval. Since this term is strictly increasing there, the minimizer is the left boundary $\gamma_{j}^{\star}=\gamma_{j}^{\mathrm{var}}=\sigma_{j}^{2}/(\sigma_{j}^{2}+v_{j}^{2})$.
  • Otherwise, by the intermediate value theorem and strict monotonicity of the two curves, there exists a unique $\gamma_{j}\in\big(\gamma_{j}^{\mathrm{var}},1\big)$ such that $\frac{\alpha(j,\gamma_{j})}{\alpha^{\star}}\;=\;\frac{\beta(j,\gamma_{j})}{\beta^{\star}}.$
B.2 Numerical example
Here, we expand our discussion around Example 5.
Figure 8: Example for $j=2$. Top row: relative bias $\beta/\beta^{\star}$ (dashed) and relative variance $\alpha/\alpha^{\star}$ (solid), both evaluated at the optimal weight $\gamma_{2}^{\star}$. Bottom row: the optimal weight $\gamma_{2}^{\star}$. Columns vary, respectively, (a) $\omega_{2}$ with $v_{1}=0.5$, $v_{2}=1$, $\sigma_{1}=\sigma_{2}=1$, $\omega_{1}=1$; (b) $v_{2}$ with $\omega=(0.9,1)$, $v_{1}=1$, $\sigma=(1,1)$; (c) $\sigma_{2}$ with $\omega=(0.9,1)$, $v=(1,1)$, $\sigma_{1}=1$.
Figure 9: Example for $j=2$ (regret). Different colors correspond to the regret for choosing either experiment. Figure reports $\max\{\alpha(j,\gamma_{j}^{\star})/\alpha^{\star},\beta(j,\gamma_{j}^{\star})/\beta^{\star}\}$ for $j\in\{1,2\}$ as the x–axis parameter varies (columns: $\omega_{2}$, $v_{2}$, $\sigma_{2}$). The vertical dashed line marks indifference, where the two curves intersect; to its left/right, the optimal experiment is the one with lower max regret. Columns vary, respectively, (a) $\omega_{2}$ with $v_{2}=0.5$, $v_{1}=1$, $\sigma_{1}=\sigma_{2}=1$, $\omega_{1}=1$; (b) $v_{2}$ with $\omega=(0.9,1)$, $v_{1}=1$, $\sigma=(1,1)$; (c) $\sigma_{2}$ with $\omega=(0.9,1)$, $v=(1,1)$, $\sigma_{1}=1$.
Consider choosing $\gamma_{2}^{\star}$ when experiment $j=2$ is more precise as $j=1$, with $\sigma_{1}^{2}=\sigma_{2}^{2}=1$, $v_{1}=1$, $v_{2}=0.5$, and $\omega_{1}=1$. The optimal $\gamma_{2}^{\star}$ depends not only on the features of arm $2$ but also on those of the alternative arm 2-2, $(v_{-j}^{2},\sigma_{-j}^{2},\omega_{-j})$, through the oracle benchmarks $(\alpha^{\star},\beta^{\star})$. Figure 8 illustrates this by plotting $\alpha(2,\gamma_{2}^{\star})/\alpha^{\star}$ and $\beta(2,\gamma_{2}^{\star})/\beta^{\star}$ at the optimal weight (top row) and the corresponding $\gamma_{2}^{\star}$ (bottom row), as we vary $\omega_{2}$, $v_{2}$, and $\sigma_{2}$ column by column. Varying $\omega_{2}$ yields two key effects: both very small and very large $\omega_{2}$ push $\gamma_{2}^{\star}$ toward one, but for different reasons. When $\omega_{2}$ is small, the oracle bias declines faster than the experimental variance (since $\beta^{\star}=\omega_{2}$), while when $\omega_{2}$ is large, the oracle bias is $\beta^{\star}=\omega_{1}$ and $\omega_{2}$ affects only the feasible experiment’s bias. Consequently, both regret and the optimal estimator depend on the oracle’s choice of which experiment it would run.
In the interior region for $\omega_{2}$, the optimal weight first moves toward the variance-optimal solution (as the oracle bias falls) and then rises again toward a bias-preferred solution. When we vary the experimental and observational variances, we obtain the expected comparative statics: $\gamma_{2}^{\star}$ decreases with the experimental variance and increases with the observational variance. These patterns translate into the regret of each arm (Figure 9). For arm $j=2$, regret initially falls sharply as $\omega_{2}$ rises and then declines more gradually once the bias component is already small. For arm $j=1$, regret is at best slightly decreasing for small changes in $\omega_{2}$ (when variance effects dominate) but eventually increases with $\omega_{2}$, because the oracle increasingly favors arm $2$ and the opportunity cost of selecting arm $1$ rises. Table 4 summarizes these patterns. We now provide further details.
Choice of $\gamma^{\star}$
Figure 8 (top row) reports $\alpha(2,\gamma_{2}^{\star})/\alpha^{\star}$ and $\beta(2,\gamma_{2}^{\star})/\beta^{\star}$ evaluated at the optimal weight while varying, column by column, $\omega_{2}$, $v_{2}$, and $\sigma_{2}$. In the bottom row, we plot $\gamma_{2}^{\star}$.
Observation 1: small $\omega_{2}$ pushes $\gamma_{2}^{\star}$ toward one. The first (left) column sets $v_{1}=v_{2}/2$ (the first experiment is more precise) and $\omega_{1}=1$ (with $\sigma_{1}=\sigma_{2}=1$). For small $\omega_{2}$ (low sensitivity of the second coordinate to bias), the optimal choice is $\gamma_{2}^{\star}=1$. The reason is twofold. First, the bias ratio $\beta(2,\gamma)/\beta^{\star}=\big(|\omega_{1}|+|1-\gamma||\omega_{2}|\big)^{2}/\beta^{\star}$ is normalized by $\beta^{\star}=\min\{|\omega_{1}|,|\omega_{2}|\}^{2}=\omega_{2}^{2}$ when $\omega_{2}<\omega_{1}$; thus even a small observational component $|1-\gamma|>0$ makes the ratio increase substantially. Second, the variance ratio $\alpha(2,\gamma)/\alpha^{\star}$ is multiplied by $\omega_{2}^{2}$ in its $j=2$ contribution, so its dependence on $\gamma$ becomes negligible as $\omega_{2}\downarrow 0$. Together these forces make the bias ratio dominant and push $\gamma_{2}^{\star}$ to the boundary at $1$. The implication is that for parameters with low $\omega_{2}$, we typically shrink $\hat{\theta}_{2}$ more toward the experimental estimate $\tilde{\theta}_{2}^{\mathrm{exp}}$.
Observation 2: when $\omega_{2}$ is of the same order as $\omega_{1}$, an interior solution emerges. As $\omega_{2}$ increases toward $\omega_{1}$ (roughly $0.6\,\omega_{1}$ and above), the normalization ceases to penalize $j=2$ as harshly, and an interior solution appears: $\gamma_{2}^{\star}$ moves down from $1$ toward the variance–only weight, trading off bias and variance. The first kink in the bottom-left panel reflects this regime change, where $\gamma_{2}^{\star}$ is selected to equalize bias and variance.
Observation 3: for $\omega_{2}>1$, $\gamma_{2}^{\star}$ increases again. At $\omega_{2}=\omega_{1}$ we observe a second kink, since the worst–case bias normalizer switches to $\beta^{\star}=\omega_{1}^{2}$. This raises the bias cost of putting weight on the observational estimate; as $\omega_{2}$ increases above $1$, $\gamma_{2}^{\star}$ rises again. This pattern is driven by the oracle benchmark, which optimizes both the estimator and the design.
Observation 4: for $\omega_{2}\gtrsim 1.25$, $\gamma_{2}^{\star}$ increases slowly and stays interior. Around $\omega_{2}\approx 1.25$ we observe a third kink: the variance normalizer $\alpha^{\star}$ switches branches (from favoring $k=1$ to favoring $k=2$ in the variance–only comparison). The optimal $\gamma_{2}^{\star}$ remains interior but grows more slowly, because increases in $\gamma_{2}^{\star}$ now have a larger relative impact on $\alpha/\alpha^{\star}$. [^18] A further kink could occur if the solution transitioned into a pure variance–dominant regime (where the boundary weight is chosen), which does not arise for the range of $\omega_{2}$ shown here.
Observation 5: $\gamma_{2}^{\star}$ increases with experimental precision and decreases with observational precision The second and third columns vary $v_{2}$ and $\sigma_{2}$, respectively. In both, we fix $v_{1}=1$ and take $\omega_{1}=0.9\,\omega_{2}$ (so here $\omega_{1}=0.9$, $\omega_{2}=1$), with $\sigma_{1}=\sigma_{2}=1$ unless varied. Across $v_{2},\sigma_{2}\in[0.5,2]$, the solution is interior. As $v_{2}\downarrow 0.5$ or $\sigma_{2}\uparrow 2$, the optimal weight $\gamma_{2}^{\star}$ increases toward $1$, placing nearly all weight on the experimental estimate. Conversely, as $v_{2}$ becomes large or $\sigma_{2}$ becomes small, $\gamma_{2}^{\star}$ approaches its lower bound, the variance–minimizing weight $\sigma_{2}^{2}/(\sigma_{2}^{2}+v_{2}^{2})$. This is a typical (and desired) behavior of shrinkage estimators.
Table 3: Example: Regimes for $\gamma_{2}^{\star}$ as a function of $\omega_{2}$ for example in Appendix Figure 8 with $\sigma_{1}^{2}=\sigma_{2}^{2}=1,v_{1}=1,v_{2}=0.5,\omega_{1}=1$.
Regime $\beta^{\star}$ $\alpha^{\star}$ Solution regime ($\gamma_{2}^{\star}$) Trend of $\gamma_{2}^{\star}$
$\boldsymbol{\omega_{2}\ll\omega_{1}}$ $|\omega_{2}|^{2}$ $\omega_{2}^{2}\sigma_{2}^{2}\;+\;\omega_{1}^{2}\dfrac{\sigma_{1}^{2}v_{1}^{2}}{\sigma_{1}^{2}+v_{1}^{2}}$ Bias–dominant Constant $\gamma_{2}^{\star}=1$
$\boldsymbol{\omega_{2}<\omega_{1}}$ $|\omega_{2}|^{2}$ $\omega_{2}^{2}\sigma_{2}^{2}\;+\;\omega_{1}^{2}\dfrac{\sigma_{1}^{2}v_{1}^{2}}{\sigma_{1}^{2}+v_{1}^{2}}$ Interior Decrease in $\omega_{2}$
$\boldsymbol{\omega_{2}>\omega_{1}}$ $|\omega_{1}|^{2}$ $\omega_{1}^{2}\sigma_{1}^{2}\;+\;\omega_{2}^{2}\dfrac{\sigma_{2}^{2}v_{2}^{2}}{\sigma_{2}^{2}+v_{2}^{2}}$ Interior Increase in $\omega_{2}$
$\boldsymbol{\omega_{2}\gg\omega_{1}}$ $|\omega_{1}|^{2}$ $\omega_{1}^{2}\sigma_{1}^{2}\;+\;\omega_{2}^{2}\dfrac{\sigma_{2}^{2}v_{2}^{2}}{\sigma_{2}^{2}+v_{2}^{2}}$ Interior Slow increase in $\omega_{2}$
Choice of the design
Figure 10 compares the maximum regret of choosing experiment $j=1$ versus $j=2$ in three scenarios. The vertical dashed line marks the value where the two curves intersect and the designer is indifferent between $j=1$ and $j=2$. In each column, we keep the same parameterizations as in Figure 8.
Observation 1: The regret for $j=2$ decreases rapidly in $\omega_{2}$ and then stabilizes/slowly increases. In the first column, we vary $\omega_{2}$ with the first experiment $j=1$ having a smaller experimental variance, $v_{1}=v_{2}/2$. The panel shows that a small $\omega_{2}$ makes the regret of choosing experiment $2$ larger than that of choosing experiment $1$. This is because a small $\omega_{2}$ corresponds to a small bias from not choosing $j=2$. As $\omega_{2}$ increases, the max–regret curve for $j=2$ declines until it reaches $\omega_{2}\approx 1.25$; after this point, the regret curve for $j=2$ rises slowly, since a further increase in $\omega_{2}$ is associated with an increase in estimator variance.
Observation 2: The regret for $j=1$ decreases slowly and then increases rapidly in $\omega_{2}$. The regret curve for choosing $j=1$ as we vary $\omega_{2}$ (fixing $\omega_{1}=1$) first decreases slowly and then increases. The reason is that for $\omega_{2}<\omega_{1}$, $\gamma_{1}^{\star}$ is an interior solution and the oracle squared bias is $\beta^{\star}=\omega_{2}^{2}$. Therefore, an increase in $\omega_{2}^{2}$ raises the oracle’s bias. When $\omega_{2}>\omega_{1}$, $\gamma_{1}^{\star}$ becomes a boundary solution and $\beta^{\star}=\omega_{1}^{2}$. In this case, a larger $\omega_{2}$ increases the variance of the estimator while the oracle bias $\beta^{\star}$ remains constant in $\omega_{2}$. As a result, it becomes more attractive for the analyst to run the experiment with $j=2$.
Observation 3: The regret for $j=2$ is monotonically increasing in $v_{2}^{2}$, and the opposite holds for $j=1$. The second plot shows that the regret from choosing $j=2$ increases monotonically with the experimental variance $v_{2}^{2}$, while the regret from choosing $j=1$ decreases correspondingly. The kinks in the curves are driven by regime shifts in the oracle solutions $(\beta^{\star},\alpha^{\star})$.
Observation 4: The regret for $j=2$ is monotonically decreasing in $\sigma_{2}^{2}$. The third plot shows that the regret for $j=2$ decreases monotonically with the observational variance $\sigma_{2}^{2}$. This is expected, since a larger $\sigma_{2}^{2}$ makes choosing $j=2$ more appealing relative to relying on observation.
Observation 5: The regret for $j=1$ first decreases slowly and then increases rapidly in $\sigma_{2}^{2}$. As we vary $\sigma_{2}^{2}$, the regret for $j=1$ is initially (very) slowly decreasing. The reason is that a larger $\sigma_{2}^{2}$ increases the oracle variance $\alpha^{\star}$ faster than the variance of the estimator for $j=1$ (with $\gamma_{1}^{\star}\approx 1$). However, this behavior does not affect the optimal solution: for smaller values of $\sigma_{2}^{2}$, choosing $j=1$ remains preferable to choosing $j=2$. When $\sigma_{2}^{2}$ is larger than a tipping point, however, the regret of choosing $j=1$ increases rapidly with $\sigma_{2}^{2}$, making the first experiment no longer preferable to the second. This aligns with the intuition that, as observational variance grows, we should favor conducting the experiment for $j=2$.
In sum, these patterns show how our framework disentangles the competing forces across signal strength and observational/experimental noise, guiding the analyst to transparent design choices even in complex regimes. Table 4 summarizes the discussion.
Table 4: Example: Qualitative behavior of maximum regret as $\omega_{2}$, $v_{2}$, or $\sigma_{2}$ increase (rows) in Figure 10. Entries summarize the direction and relative speed of change in the maximum regret for choosing experiment $j=1$ or $j=2$.
Small regime ($\omega_{2}<\omega_{1}$) Large regime ($\omega_{2}>\omega_{1}$)
Trend regret $j=1$ regret $j=2$ regret $j=1$ regret $j=2$
$\uparrow~\omega_{2}$ slow decrease fast decrease fast increase slow increase
$\uparrow~v_{2}$ decrease increase decrease increase
$\uparrow~\sigma_{2}$ slow decrease decrease fast increase decrease
Figure 10: Regret comparisons. Top row: $\max\{\alpha(j,\gamma_{j}^{\star})/\alpha^{\star},\beta(j,\gamma_{j}^{\star})/\beta^{\star}\}$ for $j\in\{1,2\}$ as the x–axis parameter varies (columns: $\omega_{2}$, $v_{2}$, $\sigma_{2}$). The vertical dashed line marks indifference, where the two curves intersect; to its left/right, the optimal experiment is the one with lower max regret. Bottom row: optimal weights $\gamma_{j}^{\star}$ for $j=1,2$, which explain how sensitivity and precision interact to drive the design switch at $x^{\star}$. Columns vary, respectively, (a) $\omega_{2}$ with $v_{2}=0.5$, $v_{1}=1$, $\sigma_{1}=\sigma_{2}=1$, $\omega_{1}=1$; (b) $v_{2}$ with $\omega=(0.9,1)$, $v_{1}=1$, $\sigma=(1,1)$; (c) $\sigma_{2}$ with $\omega=(0.9,1)$, $v=(1,1)$, $\sigma_{1}=1$.
Table 5: Optimal shrinkage and experiment choice across different regimes in the presence of two parameters (Example 5) The first three regimes (row) corresponds to the case where the variance ratio $\alpha/\alpha^{\star}$ does not uniformly dominate $\beta/\beta^{\star}$ for all values of $\gamma_{j}$ and some $j\in\{1,2\}$ (which we view as leading cases). The other cases correspond to boundary solutions.
Case (condition) Optimal $\gamma_{j}^{\star}$ Optimal $j^{\star}$ solves Solution $(j^{\star},\gamma_{j}^{\star})$
Leading cases
Case 1 no bias/ variance dominance $\gamma_{j}^{\star}$ in the interior $\mathrm{arg}\min_{j}\left\{\omega_{-j}^{2}\sigma_{-j}^{2}+\omega_{j}^{2}\!\big[(1-\gamma_{j}^{\star})^{2}\sigma_{j}^{2}+(\gamma_{j}^{\star})^{2}v_{j}^{2}\big]\right\}$ $j^{\star}$ minimizes bias and variance
($\alpha(j,\gamma_{j})/\alpha^{\star}-\beta(j,\gamma_{j})/\beta^{\star}$ can flip sign) for $j\in\{1,2\}$ $\mathrm{arg}\min_{j}\left\{|\omega_{-j}|+|1-\gamma_{j}^{\star}||\omega_{j}|\right\}$ $\gamma_{j}^{\star}$ balances bias/variance
Case 2 $j=1$ is variance dominant $\gamma_{1}^{\star}=\frac{\sigma_{1}^{2}}{v_{1}^{2}+\sigma_{1}^{2}}$ $\mathrm{arg}\min_{j}\left\{\omega_{-j}^{2}\sigma_{-j}^{2}+\omega_{j}^{2}\!\big[(1-\gamma_{j}^{\star})^{2}\sigma_{j}^{2}+(\gamma_{j}^{\star})^{2}v_{j}^{2}\big]\right\}$ $j^{\star},\gamma_{1}^{\star}$ minimizes variance
$j=2$ has no bias/variance dominance $\gamma_{2}^{\star}$ is interior and $\gamma_{2}^{\star}$ balances bias/variance
Case 3 $j=1$ is bias dominant $\gamma_{1}^{\star}=1$ $\mathrm{arg}\min_{j}\left\{|\omega_{-j}|+|1-\gamma_{j}^{\star}||\omega_{j}|\right\}$ $j^{\star},\gamma_{1}^{\star}$ minimizes bias
$j=2$ has no bias/variance dominance $\gamma_{2}^{\star}$ is interior and $\gamma_{2}^{\star}$ balances bias/variance
Other (boundary) solutions
Case 4 Both $j\in\{1,2\}$ are variance-dominant $\gamma_{j}^{\star}=\frac{\sigma_{j}^{2}}{\sigma_{j}^{2}+v_{j}^{2}}$ $\mathrm{arg}\min_{j}\left\{\omega_{-j}^{2}\sigma_{-j}^{2}+\omega_{j}^{2}\frac{\sigma_{j}^{2}v_{j}^{2}}{\sigma_{j}^{2}+v_{j}^{2}}\right\}$ Minimizes variance
($\alpha/\alpha^{\star}>\beta/\beta^{\star}$ for all $\gamma$ and $j$) for $j\in\{1,2\}$
Case 5 Both $j\in\{1,2\}$ are bias-dominant $\gamma_{j}^{\star}=1$ $\mathrm{arg}\max_{j}|\omega_{j}|$ Minimizes bias
($\alpha/\alpha^{\star}<\beta/\beta^{\star}$ for all $\gamma$ and $j$) for $j\in\{1,2\}$
Case 6 bias dominant vs. variance dominant $\gamma_{1}^{\star}=1$ $\mathrm{arg}\min_{j}$ Minimizes bias $\beta/\beta^{\star}$ of $j=1$
($\alpha/\alpha^{\star}<\beta/\beta^{\star}$ for $j=1$ and vice versa for $j=2$) $\gamma_{2}^{\star}=\frac{\sigma_{2}^{2}}{\sigma_{2}^{2}+v_{2}^{2}}$ $\left\{\frac{|\omega_{2}|^{2}}{|\omega_{1}|^{2}}1\{j=1\}+(\omega_{1}^{2}\sigma_{1}^{2}+\omega_{2}^{2}\frac{\sigma_{2}^{2}v_{2}^{2}}{\sigma_{2}^{2}+v_{2}^{2}})1\{j=2\}\right\}$ vs. variance $\alpha/\alpha^{\star}$ of $j=2$
Appendix C Additional figures and tables
Figure 11: Share of experimental papers published in AEA journals also presenting experimental results in combination with observational estimates (either from a reduced form, or from a structural model or from both).
Figure 12: Experimental design for General Equilibrium Effects: Variance and bias regret of the proposed optimal design relative to the corresponding variance and bias oracles, as a function of total experimental sample size $n_{\text{tot}}$. The left panel considers the case with one feasible experiment and the right panel the case with two feasible experiments; values close to one indicate little loss relative to the oracle.
Figure 13: Experimental Design for General Equilibrium Effects: Optimal shrinkage weights $\gamma_{j}$ for each arm (UCT, CCT, and Training) as a function of the total experimental sample size $n_{\text{tot}}$. The left panel considers designs with one feasible experiment, while the right panel allows two feasible experiments, illustrating how the procedure gradually shifts weight from UCT toward the job program as sample size increases.
Figure 14: Site selection for external validity: Optimal shrinkage weight $\gamma_{j}^{\star}$ as a function of the treated villages in the experiment.
Figure 15: Site selection for external validity: Variance, bias, and adaptive regret of the Neyman (variance–optimal) allocation relative to the proposed optimal design, as a function of the total experimental sample size $n_{\text{tot}}$. Left panel: one feasible experiment; right panel: two feasible experiments.
Appendix D Review of adaptation regret with fixed design
This section reviews the definition of adaptation regret with a fixed design in Armstrong et al. (2024) and Tsybakov (1998) and illustrates distinctions from our framework.
Single parameter with a fixed design
Different from this paper, Armstrong et al. (2024) focus on a single parameter with two estimators under a fixed design where one of the two estimators is unbiased for the target estimand. In our notation, this corresponds to fixing a design $(E,\Sigma(E))$ with $E=\{1\}$, so there is an observational estimate $\tilde{\theta}^{\mathrm{obs}}\in\mathbb{R}$ and an experimental estimate $\tilde{\theta}^{\exp}\in\mathbb{R}$ satisfying
$\mathbb{E}[\tilde{\theta}^{\mathrm{obs}}]=\theta_{1}+b,\qquad\mathbb{E}[\tilde{\theta}^{\exp}]=\theta_{1},\quad\Delta\tilde{\theta}=\tilde{\theta}^{\mathrm{obs}}-\tilde{\theta}^{\exp}$
for an unknown scalar bias $b\in\mathbb{R}$. Without loss, standardize each estimator as well as the estimand $\theta_{1}$ by the variance of the difference $\Delta\tilde{\theta}$. With this normalization, the authors further assume
$\begin{pmatrix}\tilde{\theta}^{\exp}\\ \Delta\tilde{\theta}\end{pmatrix}\sim\mathcal{N}\!\left(\begin{pmatrix}\theta_{1}\\ b\end{pmatrix},\,\Sigma\right),\quad\Sigma=\begin{pmatrix}\sigma_{\exp}^{2}&\rho\,\sigma_{\exp}\\[4.0pt] \rho\,\sigma_{\exp}&1\end{pmatrix}$,
where $\sigma_{\exp}^{2}$ is the variance of $\tilde{\theta}^{\exp}$, and $\rho$ is the correlation between the experimental estimate and $\Delta\tilde{\theta}$. The question of interest in Armstrong et al. (2024) is how to optimally combine $\tilde{\theta}^{\mathrm{exp}}$ and $\Delta\tilde{\theta}$ in a single estimator.
Non linear estimator and corresponding adaptation regret
Using invariance arguments, Armstrong et al. (2024) study estimators of the form
$\hat{\tau}_{\delta}=\tilde{\theta}^{\exp}+\rho\,\sigma_{\exp}\,\big\{\delta(\Delta\tilde{\theta})-\Delta\tilde{\theta}\big\}$,
for some measurable function $\delta:\mathbb{R}\to\mathbb{R}$ to be optimized. The authors show that for this choice, the adaptation regret worst case for $|b|\leq B$, can be written as
$A(B,\delta)=\frac{\displaystyle\mathrm{MSE}_{B}(\delta)}{R^{*}(B)},\qquad R^{*}(B)=\sigma_{\exp}^{2}\Big[(1-\rho^{2})+\rho^{2}\,r_{\mathrm{BNM}}(B)\Big]$
where $r_{\mathrm{BNM}}(\cdot)$ denotes the B-minimax risk in the bounded normal-mean problem.
Properties of the objective function and solution
Because of the non-linearity of the estimator (and of the $r_{\mathrm{BNM}}$ function) $A(B,\delta)$ is not quasi-convex in $B^{2}$ for the chosen non-linear shrinkage rule. The optimal solution is obtained by solving a minimax problem by using a discrete approximation over the least-favorable prior over $B$ (Chamberlain, 2000).
This differs from our framework where quasi-convexity arises as we restrict the class of estimators to be linear. A formal argument is provided in the proof of Theorem 1. Such quasi-convexity simplifies estimation here, since the choice is not only for the estimator, but also for the class of designs and sample size allocations.